A Bibliography of ACM SIGACT-SIGMOD Symposia on Principles of Database Systems and SIGMOD

Conferences on Management of Data

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 2.27

Title word cross-reference

007 [636]. 1 [121]. 10 [165]. 2 [1191, 1328, 2196]. 3 [1369, 422]. 5 [165]. ^{∗} [407]. _{RS}^{∗} [1439]. ^{+} [1643, 1459, 1018]. ^{3} [747, 1490]. _{2} [514, 715]. k [1767, 2021, 2184, 2010, 1210, 2195, 1969, 1670]. L_{p} [2158]. N [129, 881]. ≠ [772]. O_{2} [262].

-anonymization [1969]. -ary [129]. -DB [1484]. -dimensional [881, 1191]. -histogram [2184]. -index [1750]. -medians [1767]. -nearest [1210]. -NN [1670]. -set [2021]. -tree [407]. -Trees [1643, 502, 1018, 1459]. -variable [1328].

.NET [1914].

/GraphLog [419].

10g [1912]. 13-15 [2292]. 16th [2240]. 1975 [2240]. 1978 [2243]. 1982 [2247]. 1983 [2249]. 1984 [2251]. 1986 [2256]. 1987 [2258]. 1989 [2262]. 1991 [2266]. 1992 [2268]. 1993 [2271]. 1994 [2273]. 1995 [2274]. 1996 [2276, 2277]. 1999 [2283, 2282].

2.0 [2038]. 2000 [1596, 635]. 2001 [2285, 2286]. 2002 [2288]. 2003 [2290, 2289]. 2004 [2291]. 2005 [2292]. 2006 [2293]. 2007 [2294]. 2008 [2295]. 2009 [2296]. 2010 [2127, 2297]. 2012 [2186]. 2013 [2217]. 23-26 [2249]. 2Tier [1476].

3 [2259, 1110, 2152, 1258]. 31st [2299]. 3NF [2004]. 3Tier [1476].

6.0 [547].

7.0 [1284].

’82 [2248]. ’83 [2250]. ’84 [2252]. ’85 [2253]. ’86 [2255, 2256]. ’87 [2257]. ’88 [2260]. ’89 [2261].

’90 [2264]. ’91 [2265]. ’92 [2269]. ’93 [2270, 2271, 833]. ’94 [2272, 2273]. ’95 [2274]. ’96 [2276]. ’97 [2279]. ’98 [2281]. ’99 [2282, 1386].

= [1651].

Abstract [119, 93, 7, 394, 432, 361, 781, 1117, 993, 1094, 911, 63, 483, 487, 1017, 1035, 1006, 1051, 1269, 908, 759, 863, 895, 1099, 804, 1038, 1028, 953, 809, 807, 878, 999, 1125, 1014, 1097, 219, 796, 492, 1092, 889, 1032, 467, 678, 885, 900, 1040, 1090, 563, 351, 1162, 477, 460, 1649, 624, 790, 2187, 802, 1088, 806, 892, 887, 1042, 1268, 756, 468, 80, 1216, 1130, 383, 470, 766, 413, 997, 764, 1005, 353, 668, 485, 574, 926, 445, 699, 903, 931]. abstract-object [1051]. abstraction [1896]. Abstracts [1807, 472]. Accelerating [1634]. acceleration [1604, 1632, 1792]. Access [591, 599, 414, 325, 621, 1966, 454, 407, 2231, 1027, 1261, 1480, 2002, 1304, 1286, 506, 168, 1355, 44, 213, 682, 1153, 1528, 1389, 1290, 830, 412, 1698, 633, 1263, 1786, 1886, 2226, 1692, 698, 1262, 2011]. accesses [2172, 165, 1978]. Accessing [993, 1068, 317, 292]. accounting [494]. Accumulated [1358]. accuracy [2058]. accurate [1028, 649]. ACDN [1704]. achieve [654]. Achieving [2008]. ACM [2248, 2250, 2252, 2253, 2255, 2257, 2260, 2261, 2264, 2265, 2269, 2270, 2272, 2275, 2274, 2276, 2279, 2281, 2283, 2285, 2288, 2290, 2289, 2291, 2292, 2293, 2271, 2284, 2262, 2266, 2282, 2287, 2263, 2280, 2300, 2277, 2299, 2295, 2298, 2254, 2296, 2278, 2286, 2273, 2268, 2297, 2256, 2097, 2217, 2244, 2245, 2266, 2243, 2186, 2246, 2241, 2242, 2127, 2251]. ACM-SIGMOD [2254, 2245, 2243, 2246, 2241, 2242, 2244, 2251]. Acquisition [13]. acquisitional [1795]. across [1741, 1359, 1764]. act [2221]. ACTA [382]. action [528, 456, 1830]. actions [1126]. activation [642]. Active [1234, 1033, 1983, 2083, 2102, 977, 974, 981, 540, 304, 1150, 892, 1413, 927, 1859]. activities [384]. activity [1082, 1409]. acyclic [2015, 83]. acyclicity [2233, 79, 447, 81]. ad [1940, 1137]. ad-hoc [1940]. ADABAS [562]. ADABAS/NATURAL [562]. Adaptable [1526, 1400]. Adapting [913, 1873, 2230]. Adaptive [1874, 785, 1841, 865, 1562, 1801, 886, 1448, 29, 1441, 1939, 1994, 1750, 1636, 1055, 32, 1353, 1620, 1530, 1804, 345, 351, 1498, 1624, 821]. Adding [804, 1412, 226]. additional [922]. address [506, 951, 428, 952]. administering [1075]. Admission [1196]. Advanced [532, 858, 493, 739, 980]. advantage [966]. after [913]. again [1610]. aganist [241]. agenda [837]. agent [1122]. agent-based [1122]. agents [2136]. Aggregate [1327, 1331, 784, 1223, 2074, 2167, 1557, 1541, 1626, 1980, 1458, 276, 1245, 2024, 1832, 1231, 1566, 2237]. aggregates [1765, 2070, 1977, 1505, 1543, 2055, 1334, 2164, 1545, 1116]. Aggregation [1318, 1001, 1723, 1899, 1521, 1779, 1584, 1351, 1118, 1150, 1618, 552, 886, 2165, 2209, 555, 1637, 1811]. aggregations [1915]. Agreement [6, 284]. ahead [570, 625]. aided [1602]. air [758]. AJAX [1491]. alarming [1933]. Alberto [2097, 2217, 2186, 2127]. Alexander [293]. Algebra [792, 637, 1247, 1321, 27, 318, 226, 236, 197, 1333, 1963, 171, 182, 620, 1060, 153, 1313].

Algebraic [49, 459, 1099]. Algebras [863, 643, 700, 1057, 2142]. Algorithm [591, 32, 59, 792, 1043, 1564, 2190, 1965, 901, 1195, 397, 2128, 113, 2149, 224, 2069, 1660, 192, 1367, 344, 691, 1472, 1793, 904, 1678, 191, 275, 1116]. algorithmic [1781]. Algorithmics [1623]. Algorithms [583, 1063, 54, 1729, 697, 623, 1204, 1310, 1117, 1547, 2081, 2036, 1899, 2012, 2019, 2080, 825, 1524, 1521, 996, 1970, 2153, 499, 1647, 203, 1229, 694, 2210, 404, 467, 398, 1955, 1300, 288, 2157, 2187, 592, 1668, 1455, 166, 1440, 1739, 285, 260, 886, 1186, 501, 581, 116, 392, 561, 1489]. Algres [974]. Allocating [1975]. allocation [394, 1561, 647, 498, 500]. Alpha [83]. Alpha-acyclic [83]. alphanumeric [1047]. AlphaSort [798]. already [1127]. Alternating [657, 264]. alternative [523, 1231]. alternatives [15, 180, 1242]. amalgamation [2218]. AMDB [1290]. among [1541, 2082, 803, 1223, 766, 353, 1536]. AMS [1082]. analog [747]. Analysis [1185, 36, 168, 753, 570, 674, 1211, 200, 527, 1966, 2083, 1035, 1788, 1244, 238, 551, 254, 1985, 523, 117, 169, 241, 1132, 2212, 461, 398, 33, 1333, 1957, 2192, 2197, 151, 460, 1888, 147, 1713, 215, 998]. Analytical [202]. Analyzing [2002]. Anatomy [1473]. ancestor [349]. ancestor-controlled [349]. Angeles [2248]. Ann [2246]. Annealing [776, 96]. Annotated [2088]. annotation [2205]. annotations [1642]. Annual [2259, 2258, 2249, 2251]. anonymity [2008, 1216]. anonymization [1969]. ANSI [860]. answer [1732]. Answering [1232, 1653, 2074, 47, 1106, 885, 1040, 110, 887, 1428, 1410, 1349, 1460, 1958, 2118, 1772, 2101, 1557, 1705, 2229]. Answers [1311, 1340, 297, 2141, 1241, 2010, 2120, 2159, 1862, 2140]. Any [792, 1609]. APEX [1636]. APPC [726]. append [616]. append-only [616]. Application [610, 1, 1610, 1476, 1189, 1697, 1006, 270, 728, 827, 36, 1258, 1252, 943, 1918, 845, 1721, 1270, 1360, 1776, 1495, 1481, 2090]. Application-Specific [610, 1495]. Applications [1333, 541, 1199, 1695, 1406, 1387, 1260, 1928, 819, 1576, 1810, 1925, 150, 484, 948, 720, 1704, 152, 2267, 39, 1955, 1293, 296, 1083, 1623, 970, 493, 739, 299, 986]. Applying [923, 911, 650, 1570]. Approach [780, 47, 4, 569, 687, 584, 781, 654, 490, 1727, 358, 1557, 1632, 1575, 607, 1500, 1443, 316, 671, 915, 467, 186, 2105, 1734, 542, 497, 1934, 14, 1392, 1694, 431, 677, 86, 1848, 1868]. Approaches [553, 1064, 724, 971]. Approximate [1318, 1733, 1666, 2020, 2197, 1249, 1862, 1334, 1410, 1349, 1460, 1819, 1799, 244, 1665, 1557, 1241, 1543, 1566, 1562, 1485]. approximated [770]. Approximating [1458, 2184, 2071, 2010]. Approximation [2081, 2080, 1898, 2036, 1977]. approximations [2206, 2146]. April [2252, 2264, 2246]. AQR [1498]. AQR-toolkit [1498]. Aqua [1410]. Arabic [347]. Araneus [1281]. Arbitrary [1025, 1986]. Arbitration [648]. Arbor [2246]. Architecture [131, 721, 962, 10, 1375, 1919, 111, 374, 136, 304, 1476, 709, 835, 496]. Architectures [515, 17, 1817, 495, 1079, 174, 959, 1563, 1157, 841, 944, 763, 1007, 1010, 611, 305, 306]. archiver [10]. archives [1457]. Archiving [1616, 660]. Area [56, 1796]. argument [470, 766, 353]. arguments [620]. Ariadne [1287]. Ariel [528]. ARIES [625, 763]. ARIES/CSA [763]. ARIES/IM [625]. arithmetic [1653, 1511]. arity [772, 326]. Arizona [2278, 2279]. arm [527]. Armstrong [55]. Array [1, 1994, 2048, 1116]. array-based [1116]. Arrays [606, 694, 1041, 1357, 1225, 198, 459, 500]. art [1072, 1954]. ary [129]. Asking [2013, 861]. ASP [1723]. Aspects [664, 486, 139, 35]. Assembly [1236, 457]. asset [960, 761]. assets [961]. assigning [303]. assignment [116]. assisted [342]. associated [1721]. Association [2259, 1320, 2258, 1030, 1151, 673, 1135, 996, 1081, 1136, 1325, 904, 1242, 1572, 1186, 994, 1181]. association-rule [1181]. associations [1183]. Associative [1, 557]. Assumption [60, 79, 77]. assumptions [925]. assurance [1924]. astronomy [1457]. asymmetric [910]. Asynchronous [522, 1990, 1396, 497, 1430]. Athens [2298]. Atlanta [2250]. Atlantic [2263]. Atomic [267, 667, 2151, 117, 922]. Atomicity [780, 1354, 781, 567, 765]. Attacks [1320]. Attribute [1182, 284, 421, 473, 523]. Attributes [1030, 1343, 1458, 303, 409, 1754, 288]. attrition [2222]. auditing [1967]. augmented [2237]. August [2242]. Aurora [1821]. Austin [2260, 2243, 2254]. author [1593]. author-X [1593]. Authority [1270]. auto [1496]. auto-configuration [1496]. AutoAdmin [1244]. Automata [1337, 2117, 1633, 279, 2086]. automated [1870, 740]. Automatic [1199, 1534, 1374, 1903, 175, 606, 1055, 1709, 130, 1595, 1894, 1463, 1850, 1495, 1428]. automatically [1237]. Automating [1691, 1464]. automation [235, 970]. autonomous [844]. Availability [88, 1938, 589, 1278, 586, 282, 839, 948, 545, 1439]. available [1478, 1393, 1394, 1909]. aviation [1581]. avoid [554]. Avoiding [624]. Award [2186, 2127, 2097, 2217]. aware [1540, 2212, 1857]. awareness [1951]. axes [2234]. axiomatic [186]. Axiomatization [143, 2137].

B [2016, 2151, 123, 1688, 1643, 881, 317, 310, 398, 502, 2112, 1459, 503, 501, 1018]. B* [57]. B*-Trees [57]. B-tree [881, 501, 1688, 398]. B-trees [2016, 2151, 123, 317, 310, 2112, 503]. B2B [1613]. B5 [575]. back [2063, 505]. backup [1685, 1422, 592]. backup/recovery [1685]. Backward [380, 1285]. backwater [1717]. bad [1479]. Bag [784, 1999]. bag-set [1999]. Bags [643]. Balance [125]. Balanced [1363, 74]. Balancing [1120, 920, 606, 805, 932, 1599]. Baltimore [2292]. bandwidth [939]. BANG [145]. Barbara [2285, 2286]. Base [899, 1194, 98, 596, 422, 660, 162, 258, 493, 399, 739]. Based [778, 599, 509, 788, 1185, 776, 870, 558, 575, 1182, 690, 687, 1023, 2179, 1598, 1555, 377, 1701, 476, 1418, 1122, 1744, 1921, 933, 2231, 1945, 1143, 1427, 389, 2101, 1736, 1177, 908, 1601, 1557, 491, 1488, 1854, 1057, 1193, 1079, 1220, 1286, 939, 1632, 1538, 843, 126, 718, 1674, 1241, 771, 1738, 316, 984, 1166, 463, 1082, 185, 626, 1492, 33, 2024, 2046, 1369, 1804, 711, 374, 1436, 1231, 2199, 846, 199, 830, 1163, 1288, 1499, 542, 2048, 1251, 1281, 1487, 1675, 498, 1007, 904, 1045, 525, 1089, 1556, 1060, 403, 14, 40, 617, 1240, 1005, 581, 1206, 434, 1622]. based [1461, 1848, 1868, 1871, 1791, 1116, 734, 1484, 1399, 1408]. Bases [49, 680, 1224, 314, 302, 440, 216, 2003, 681, 414, 478, 221, 93, 431]. basis [108]. basket [1368, 1133]. baskets [1135]. battlefield [731]. Bayan [347]. BC [2295]. BCNF [222, 359]. BDD [2224]. BDD/FC [2224]. be [770, 751, 286, 160]. beats [291]. beauty [2154]. bed [1493]. beef [1477]. Behavior [531, 382, 328]. behind [1725]. Beijing [2294]. being [1119, 408]. Belief [1319, 120, 2003]. benchmark [1112, 251, 638, 635, 636]. Benchmarking [164, 1462, 1493, 1693]. benchmarks [513]. Benefits [902, 443]. Bernoulli [2040, 1863]. BeSS [1084]. Best [1628, 1977, 987, 1849]. Best-effort [1628, 1849]. better [1039]. Between [648, 673, 1685, 1309, 908, 774, 1872, 1069, 1032, 228, 724, 544]. Beyond [1135, 753, 2164, 687, 2161, 1805, 972, 1466, 2112, 1376, 93]. bi [908, 1863, 1925]. bi-directional [908]. bi-level [1863]. biased [2019, 1426]. bidirectional [2189]. Bifocal [1046]. Big [861, 2182]. billion [799]. billion-record [799]. billions [1605]. Binary [129, 465]. binding [887]. bio [1950]. bio-diversity [1950]. biology [1955]. BIRCH [1015]. BIRN [1833]. BIRN-M [1833]. Bit [1511]. Bit-sliced [1511]. Bitmap [1243, 1338, 2112]. bitmaps [1341]. BLAS [1844]. block [1818, 1864]. block-level [1864]. Blocking [676, 584, 581]. bloom [1768]. BNR [710]. board [45]. boat [1248, 1330]. BODHI [1950]. Bonded [772]. boolean [2216, 909]. boosting [1517]. born [1610]. born-again [1610]. Boston [2244, 2251]. both [1775, 1643, 212, 359]. Bottom [1364, 291, 442]. Bottom-up [1364, 291, 442]. Bound [902, 1189, 1644, 443]. boundary [1373]. Bounded [441, 102, 895, 326, 2005, 2047, 2238, 137, 200, 926, 590]. bounded-arity [326]. bounded-variable [926]. boundedness [429, 252]. bounding [884]. Bounds [135, 175, 2109, 2098, 1978, 2017, 2158, 1191, 2035, 1755, 2113]. branching [1639]. breaches [1763]. breaking [1208, 2190, 518]. brick [840]. bridging [736]. Bringing [1385, 2063]. Broadcast [910, 1120, 1314]. brokering [929]. browser [1658]. browsing [510, 1297, 984, 1454, 1474]. BSR [939]. Bubba [196]. bubbles [1517]. bucket [460]. BUCKY [1112]. buffer [1053, 1203, 1561, 498, 1107]. Buffering [1976, 1856, 331, 691]. buffers [787]. build [1658]. builders [544]. Building [1918, 1141, 1332, 1818, 689, 547, 1929]. Built [1219]. Built-in [1219]. bulk [1103]. bursts [1851]. bushy [461, 1000]. business [1598, 1921, 1817, 1276, 728, 2154, 1925, 1464, 1696, 1465, 1579]. byte [165]. Byzantine [6].

C [1008, 1009, 335, 638, 966, 529]. C-logic [335]. CA [2269]. Cache [2016, 1796, 2150, 496, 1643, 2079, 1291, 1628, 1459, 1095, 1856, 1695]. Cache-and-query [1796]. Cache-oblivious [2016, 2150]. cache-obliviously [2079]. cached [1562]. Caches [1348, 368]. Caching [1233, 757, 658, 1022, 1697, 1271, 1695, 752, 1817, 1577, 495, 1703, 2106, 1604, 1059, 1620, 1816, 1696, 756, 399, 1694]. CAD [471, 1175, 150, 1803, 723]. CAD/CAM [150]. Cadlab [982]. CALANDA [990]. calculi [796, 2237]. Calculus [683, 69, 873, 802, 620, 155, 699, 392]. Calif [2289]. California [2274, 2285, 2290, 2240, 2286, 2268, 2248, 2257, 2275, 2245]. Call [634]. CAM [1937, 150]. CAMAS [1951]. Cambridge [2255]. Can [770, 286, 1667, 544]. Canada [2276, 2277, 2295, 2242, 2252]. candidate [1419]. Canonical [1896]. capabilities [1371]. Capability [1288, 1082]. capability-based [1082]. CapBasED [1082]. CapBasED-AMS [1082]. Capturing [1775, 2139]. cardinalities [1698]. cardinality [2129]. Carlo [1678]. Carnot [729, 955]. Cartesian [394, 624, 1000]. Case [558, 356, 888, 1869, 2111, 1157, 151, 2187, 198, 233, 1155, 1487]. Case-Based [558]. Catalog [1580]. Catching [1248]. categorical [110, 2194]. categorization [1239, 1903]. categorizing [1515]. CCUBE [1411]. cells [477]. centered [1416]. centric [2227]. Certain [2141, 2159, 2140, 2229]. certification [40]. Chain [134, 1771, 522, 2143, 539]. chaining [380]. challenge [507, 702, 946]. Challenges [1464, 760, 706, 1156, 2031, 1923, 1546, 1924, 704]. Change [1065, 648, 1091]. Changes [1322, 1802, 1269, 1403]. Changing [1193, 2176, 841]. Character [562]. Characteristics [1322, 605]. characterization [78, 222, 1786]. characterizations [2067, 539, 1585, 1537, 2025]. Characterizing [2147, 1655]. charging [1421]. Charting [2229]. chase [2076, 1442]. Chasing [1021]. Chebyshev [1890]. check [2226]. Checking [672, 762, 534, 280, 19, 2009, 1982, 1061]. checkpointing [415]. Chemical [569]. Chicago [2259, 2293]. Chickens [142]. CHIMERA [974]. China [2294]. choice [549]. chopping [608]. Chordality [48]. chromatic [465]. chronicle [889]. chunks [1233]. CIM [971]. circuit [45]. circumscription [77]. citizen [1951]. City [2263]. Class [794, 263, 339, 1002, 137, 896, 1039, 64]. class-division [896]. classes [1519, 678, 2238, 1846, 479, 70]. CLASSIC [274]. Classification [2205, 247, 2239, 1779, 1186]. classify [1515]. classifying [1519]. clauses [3, 157]. cleaning [1780, 1491, 1836]. Client [1362, 495, 568, 1024, 1252, 944, 763, 836, 496, 1095]. client-server [495, 568, 1024, 1252, 763, 1095]. Client-site [1362]. client/server [836, 496]. Clio [1595]. clips [747]. clocks [866]. cloning [1729]. closed [77, 2075]. Closest [1435]. Closure [1001, 792, 825, 158, 467, 248, 352, 903, 2026, 2086]. cloud [2182]. cluster [2041]. Clustered [308]. Clustering [599, 1382, 776, 882, 1676, 1877, 1310, 2008, 1199, 2081, 1308, 1517, 331, 433, 1974, 2080, 32, 2028, 1195, 2121, 1714, 409, 1306, 1804, 1879, 1472, 1808, 1426, 1678, 1360, 431, 560, 1848, 2203, 1015]. clustering-based [1848]. Clusters [1878, 1425, 1507, 1934]. CLV [238]. CMVF [1812]. co [2081, 654]. co-clustering [2081]. co-existence [654]. coalescing [1756]. CoBase [735]. CODASYL [36]. CODAYSL [44]. Codd [1504]. codes [237, 313, 1128, 1439]. cohabitation [724]. coherency [551]. COKO [1702]. Collaborative [2221]. Collection [782, 639, 818, 1055, 267, 667, 1139]. collections [1600, 1449, 2079]. Colony [2250]. Colorado [2266, 2265]. Colorful [1861]. column [1762]. Combinatorial [917, 918, 775, 334]. combined [654, 1443, 2099]. Combining [1038, 1444, 1999, 334, 987]. COMFORT [740]. Commerce [1215, 1317, 1257, 1580, 1393, 1378, 1394, 1473]. commercial [718, 948]. commit [1154, 922, 446, 559]. Commitment [256, 650, 548, 117, 114]. COMMIX [1705]. Common [5, 1220, 428, 832]. communicating [2001]. Communication [2236, 1572, 910, 114]. Communication-efficient [1572]. communities [1483]. compact [1688, 1506]. Comparing [694]. comparison [1343, 282, 12, 579, 1690, 410, 246]. comparisons [1653]. Compensation [617]. Compensation-based [617]. Competitive [1196]. Compilation [522, 1651, 2300, 1785]. compiler [214, 115]. Compiling [307, 245, 764]. complement [1734]. complements [1640]. Complete [1097, 144, 2161, 1531, 191, 275]. completed [67]. Completeness [680, 681, 2103]. Complex [788, 897, 792, 1230, 69, 1006, 933, 244, 1784, 210, 335, 433, 307, 1872, 1626, 489, 1222, 1503, 2168, 597, 457, 39, 1964, 333, 1090, 821, 1074, 743, 459, 1428]. complex-object [597]. Complexity [1232, 134, 1326, 664, 2084, 1227, 273, 1329, 1219, 794, 1230, 62, 2175, 612, 1539, 1772, 1665, 774, 2239, 2207, 596, 2039, 1757, 1226, 1518, 2168, 1988, 292, 2138, 2046, 2201, 2213, 1964, 1995, 1963, 2193, 151, 2145, 1088, 2099, 479, 1130, 1755, 2059, 2064, 190, 352, 926, 2038, 618]. component [962, 1170]. Components [1350, 212]. Composing [1797]. Composite [329, 538, 471, 1123, 512, 1953, 24]. Composition [1972, 2084]. compound [93]. comprehensive [1806]. compressed [1552, 2079, 2202, 2220, 1813, 1336]. Compressing [1684, 1884]. compression [1555, 684, 1747, 72]. compressor [1432]. computability [208]. computable [330]. computation [1364, 1543, 1858, 1513, 1503, 2024, 1640, 2197, 1345, 1507, 2015, 1633, 1334, 1545]. computations [2017, 1149]. Compute [792]. Computer [1224, 342, 1602, 719]. computer-assisted [342]. computers [754, 428]. Computing [2259, 2122, 1878, 2142, 2258, 104, 1970, 1740, 58, 1371, 701, 1965, 1505, 1996, 1533, 2181, 93, 2140]. Concept [753, 1492, 732]. Concepts [214, 140, 2093, 260]. Conceptual [48, 510, 845]. Concise [1748]. Concurrency [354, 874, 340, 317, 720, 374, 1098, 125, 692, 1356, 687, 62, 866, 15, 341, 377, 287, 1304, 450, 566, 201, 203, 615, 219, 621, 605, 625, 1314, 501, 127, 116, 496, 315, 66, 590]. Concurrent [223, 57, 1607, 398, 581, 1095]. condensed [1551]. condition [528, 160]. Conditional [1866]. conditions [405, 406, 1635, 927]. Conference [2259, 2275, 2290, 2291, 2244, 2271, 2245, 2284, 2266, 2243, 2282, 2287, 2263, 2280, 2277, 2240, 2246, 2254, 2278, 2241, 2247, 2286, 2242, 2273, 2268, 2251, 2256, 2262, 2258, 1952, 2300, 1393, 1394]. configuration [1496, 1155]. Configurations [1350]. Conflict [303]. Conflicts [1033, 1032]. confluence [531]. conforming [1825]. Congressional [1460]. conjecture [2224, 2123]. Conjunctive [1094, 1207, 1219, 1221, 1635, 1875, 2206, 578, 646, 2239, 1644, 2060, 289, 2200, 2098, 2000, 2170, 2173, 2140, 2229]. Conjunctive-Query [1221, 2200]. Connected [1878, 725]. connection [78, 544]. connection-trap-free [78]. Connections [48]. conscious [671, 1459]. Consensus [2120]. Conservative [639]. consistency [1254, 1661, 1958, 565, 2058, 1685, 1119, 2153, 1908, 1942, 443, 1062, 496, 1095]. Consistent [1311]. Constant [84, 102, 222, 1734, 2196, 286]. constant-time [102, 222]. constant-time-maintainability [286]. Constrained [354, 874, 1021, 361, 174, 482, 1328, 1183, 1404]. Constraint [995, 1200, 1318, 875, 762, 463, 402, 871, 868, 872, 1221, 534, 906, 1871, 869, 1411, 1585, 878, 999, 711, 1502, 1061, 609, 997]. Constraint-based [463, 1871]. constraint-like [1585]. Constraints [46, 1213, 778, 1205, 1309, 1209, 672, 175, 909, 588, 533, 766, 1109, 1775, 628, 300, 472, 2050, 598, 2072, 342, 130, 1523, 1134, 1746, 1942, 644, 408, 276, 1645, 1092, 2073, 678, 2174, 20, 1328, 400, 1989, 927, 764, 353, 2011]. constructing [1287]. construction [1250, 364, 1330, 1647]. constructivism [270]. constructs [440, 226, 133]. consumer [830, 1392]. Containing [1217, 179]. Containment [1209, 578, 1207, 1219, 1221, 1627, 1752, 1736, 1746, 1092, 2000, 2205, 1090, 1568, 2038]. Content [1579, 976, 1592, 1577, 1079, 1604, 1632, 1375, 1738, 1704, 1804, 1546, 1499, 1007]. content-based [1079, 1804, 1499, 1007]. contention [233]. Context [1957, 687, 36, 1164]. Context-Based [687]. Context-sensitive [1957]. Contexts [187]. contingency [2058]. continual [1937, 1505]. Continually [1674]. Continuous [1196, 2211, 616, 1655, 1289, 1823, 1451, 1773, 1940, 1624, 1892, 2052, 1115, 1801, 1010, 1876]. Continuously [1624, 1441, 1447]. Contorting [1794]. contradictions [263]. contribution [1984]. Control [1196, 902, 125, 692, 1356, 687, 62, 866, 15, 341, 195, 377, 287, 358, 1304, 1055, 551, 201, 340, 317, 203, 494, 615, 219, 197, 1153, 720, 443, 374, 605, 497, 1886, 1314, 501, 127, 116, 496, 315, 66, 240, 1289]. controllability [2030]. controlled [1966, 595, 349]. controller [1413]. controlling [259]. controls [494]. Convergence [65]. conversion [1214, 36]. cooperation [1628]. cooperative [418, 737]. Coordinating [1685]. Copies [594]. coping [2095]. Copy [941]. CORAL [741, 665]. CORBA [1417]. CORDS [1894]. core [1742]. cores [1970, 1996]. Cornell [1412]. correct [155]. correcting [313]. Correctness [1123, 1350, 50, 652, 915, 257, 343]. correlated [1505, 1188]. Correlating [1751]. Correlation [1878, 1598, 1832]. correlations [1135, 1894]. COSIMA [1588]. COSS [832]. Cost [1854, 1324, 1192, 755, 1060, 1240, 1206, 1101, 1667, 1028, 671, 922, 595, 1454, 1786, 2066, 114, 801, 1461, 240]. Cost-based [1854, 1060, 1240, 1206, 1461]. cost-controlled [595]. cost-effective [1454]. cost/performance [801]. costly [1149]. coterie [587]. COUGAR [1706]. count [1648, 2210, 1515, 482]. counterfactuals [596]. Counting [2018, 250, 1461, 1133, 2135, 2211, 103]. course [2045]. Cover [961]. Covering [1639, 1128]. covers [104]. CPU [165]. CQ [1282]. Crash [568, 52, 2045, 907]. crawlers [1377]. creating [623]. Creation [685, 1173]. credit [251]. crisis [1951, 837]. critical [1387]. critique [860, 858]. cross [1130]. crossroads [1159]. cryptographically [1966]. cryptography [1840]. CSA [763]. CSP [2231]. cube [1813, 1463, 1507, 1103, 1364]. CubeExplorer [1711]. Cubes [1236, 1503, 1711, 1036, 1128, 1100, 1105, 1490]. Cubetree [1103, 1413]. cubetrees [1231]. Curated [2062]. CURE [1195]. currency [2160, 1908, 1942]. current [1717]. curse [1208]. curtain [1725]. curves [2203]. customer [953, 1270, 1969]. customizable [2191]. Cut [1113]. CWA [2042]. CWA-solutions [2042]. cyclic [250]. cymbal [1390].

D [1750, 1110, 1369, 2196, 30]. D.C [2256]. Dallas [2284]. dangers [1029]. DARE [1597]. Darmstadt [131]. Data [2259, 2275, 2290, 2291, 206, 767, 591, 1820, 2161, 48, 49, 2244, 1027, 1480, 1860, 1517, 875, 2271, 1205, 495, 1158, 2245, 2284, 1194, 2266, 196, 2243, 1320, 2258, 2282, 585, 1742, 1267, 2287, 996, 1479, 1322, 2263, 771, 1996, 1331, 1125, 1913, 2280, 1073, 680, 1717, 776, 754, 217, 588, 2277, 1838, 1836, 424, 862, 1340, 2240, 1224, 1657, 1998, 2075, 2246, 1583, 1339, 90, 2254, 93, 218, 1378, 56, 674, 2278, 1420, 2241, 606, 2247, 2286, 59, 2242, 1236, 2273, 413, 957, 2268, 1398, 1573, 2251, 2256, 1821, 2179, 2221, 1332, 910, 1120, 1114, 1237, 2074]. data [2107, 1368, 1509, 1681, 1802, 314, 271, 448, 1148, 1199, 1456, 1547, 1731, 1947, 1887, 195, 2147, 220, 1527, 1400, 1901, 1655, 1782, 396, 1727, 1958, 2063, 2180, 2122, 1165, 1617, 1771, 1767, 1555, 1701, 2124, 1445, 483, 98, 507, 237, 1775, 1101, 1006, 1685, 2231, 1261, 1564, 1993, 274, 675, 1932, 1133, 844, 949, 1066, 1108, 1616, 2018, 389, 1933, 261, 706, 1576, 1931, 1381, 1780, 2182, 1091, 1591, 759, 1607, 1750, 1800, 1636, 1594, 2262, 1214, 2056, 722, 1853, 2019, 2080, 613, 1709, 1794, 684, 2031, 1733, 1898, 1658, 2044, 1587, 1538, 1370, 2001, 828, 1575]. data [1626, 506, 1897, 1763, 2095, 901, 2077, 2160, 1069, 2114, 2234, 1446, 1946, 1971, 1491, 1720, 1751, 1142, 1505, 1985, 607, 1585, 2082, 1970, 2017, 2121, 688, 316, 1654, 1172, 1136, 1711, 1036, 2227, 2152, 219, 2042, 1988, 1714, 681, 1128, 1100, 579, 1064, 2212, 2300, 484, 1353, 1483, 1873, 278, 889, 1306, 185, 438, 1397, 1492, 2055, 597, 33, 678, 1762, 1388, 1510, 545, 1781, 1279, 1729, 2061, 2014, 1389, 1962, 1995, 455, 1832, 1920, 1690, 1365, 2199, 1602, 199, 21, 1173, 1608, 1277, 830, 262, 1438, 1280, 1991, 2159, 2230, 1432, 136, 1439, 1498, 325, 2197, 193]. data [730, 85, 460, 821, 1531, 1487, 1888, 1151, 1777, 1747, 660, 1105, 2181, 1879, 1578, 1581, 2194, 147, 2133, 1115, 1238, 1629, 1270, 1569, 115, 72, 1801, 515, 1808, 1426, 1372, 1650, 1493, 1790, 1934, 1077, 1089, 1455, 1776, 258, 1263, 1437, 1609, 1954, 1103, 479, 2226, 1716, 1216, 1490, 170, 998, 2059, 1176, 179, 723, 1743, 1783, 399, 1383, 1692, 411, 1885, 1392, 1700, 725, 2209, 581, 1334, 359, 775, 1676, 1745, 1262, 561, 1842, 1871, 1294, 1015, 2023, 1969, 1876, 1670, 618]. data-aware [2212]. data-centric [2227]. Data-driven [1573, 2221, 2001]. data-intensive [195, 1709, 1487]. data-mining [901]. Data-valued [413]. data/knowledge [258]. Database [2248, 2250, 2252, 2253, 2255, 2257, 2260, 2261, 2264, 2265, 2269, 2270, 2272, 2274, 2276, 2279, 2281, 2283, 2285, 2288, 2289, 2292, 2293, 2294, 1553, 271, 701, 490, 1234, 18, 1019, 1402, 564, 685, 1265, 1110, 1228, 1180, 546, 513, 2300, 1033, 872, 580, 54, 550, 575, 610, 918, 2, 2295, 1, 55, 232, 125, 272, 704, 861, 6, 690, 362, 883, 1722, 553, 7, 1107, 94, 2297, 1202, 837, 144, 229, 940, 37, 923, 17, 851, 1870, 760, 531, 1726, 1697, 1901, 826, 1695, 1165, 207, 514, 214, 1387, 1292, 1950, 1175, 1260, 42, 1685, 600, 857, 1414]. database [1075, 2218, 1718, 1817, 1411, 844, 167, 280, 1577, 287, 1156, 1576, 187, 974, 102, 209, 1269, 176, 491, 1552, 1429, 614, 31, 307, 1286, 818, 1810, 1160, 1297, 1922, 1658, 702, 2249, 254, 36, 661, 2154, 1872, 1080, 703, 1011, 13, 375, 2148, 2177, 843, 959, 1706, 718, 231, 183, 540, 733, 1152, 512, 1167, 1246, 440, 2005, 1858, 1067, 777, 950, 1839, 1518, 1154, 1913, 426, 427, 1863, 1654, 107, 241, 397, 150, 111, 222, 796, 915, 173, 184, 492, 2212, 855, 1092, 132, 95, 1904, 185, 570, 511, 991, 871, 328, 1258, 423, 41, 346, 834, 347, 917]. database [2299, 20, 1272, 1957, 1816, 1438, 1076, 113, 2298, 2100, 542, 1696, 1906, 118, 281, 1840, 304, 943, 845, 740, 763, 1380, 864, 162, 1062, 1150, 379, 773, 748, 1256, 221, 1578, 344, 452, 691, 631, 1501, 82, 38, 1088, 806, 2296, 131, 233, 435, 488, 415, 230, 982, 1394, 741, 665, 256, 1691, 242, 1502, 133, 1495, 1437, 1609, 164, 172, 83, 1516, 849, 824, 1242, 43, 148, 2064, 1376, 355, 1812, 464, 14, 1481, 493, 25, 40, 927, 299, 1652, 986, 848, 1880, 658, 1295, 719, 1724, 698, 2045, 279, 713, 445, 1274, 395]. database [955, 590, 731, 390, 240, 101, 1095, 705, 1568, 1641, 1856, 1417, 1910, 2267]. database-driven [2218, 1577]. Database-friendly [1553]. database-like [1286]. database-service-provider [1654]. database-system [1724]. Databases [1003, 354, 874, 1311, 919, 509, 664, 921, 1196, 810, 992, 812, 897, 522, 859, 526, 1157, 914, 534, 569, 1217, 572, 1012, 519, 662, 1198, 61, 1023, 1302, 1020, 2092, 1818, 341, 548, 673, 1117, 1712, 1741, 1829, 471, 105, 582, 653, 120, 283, 2228, 2233, 154, 476, 1197, 1744, 602, 1102, 1087, 1035, 925, 76, 1718, 1417, 1316, 270, 949, 2062, 1772, 1374, 1477, 852, 1381, 908, 1451, 433, 188, 417, 2190, 27, 1259, 1220, 1055, 1435, 537, 1416, 106, 342, 1796, 1049, 269, 1379, 1099, 88, 1897, 1973, 822, 968, 1190, 1446, 1443, 809, 891]. databases [277, 44, 799, 1759, 489, 807, 878, 999, 1195, 462, 644, 853, 1078, 615, 235, 122, 463, 893, 1382, 1229, 579, 482, 292, 1096, 273, 1515, 707, 1319, 1212, 983, 1071, 1369, 736, 1530, 391, 186, 630, 2071, 2132, 374, 1834, 437, 737, 1424, 449, 189, 985, 2120, 110, 508, 651, 119, 1892, 1675, 708, 1367, 90, 1068, 1454, 1814, 321, 1086, 892, 1673, 559, 1360, 604, 2030, 552, 121, 9, 64, 295, 265, 370, 305, 1421, 1145, 1121, 954, 1615, 26, 306, 997, 1052, 338, 1671, 616, 1005, 456, 89, 35, 716, 116, 677, 988, 1174, 1485, 1470, 2140]. databases [86, 1482, 247, 1536, 191, 275, 705, 1015, 1401]. DataBlade [835]. DataBlitz [1387]. datacycle [111]. dataflow [195, 1823]. dataflows [1909]. Datalog [905, 1352, 539, 2123, 888, 376, 894, 206, 1990, 2231, 320, 300, 2101, 530, 269, 804, 1619, 2047, 2232, 208, 322, 533, 194, 161, 295, 291, 2090, 432, 772, 1035, 659, 770, 774, 372, 429, 356, 655, 900, 867, 475, 481]. datalog-based [2101]. DataMIME^{T M} [1935]. DataMine [973]. datasets [1852, 901, 1137, 1162, 1138, 1345]. dataspace [1992]. DataSplash [1283]. DataSpot [1297]. Daytona [1390]. DB [1027, 1484, 532, 535]. DB/DC [532]. DB2 [964, 1384, 1264, 1910, 1687, 1808, 836]. DBCache [1697, 1817]. DBIS [1400]. DBIS-toolkit [1400]. DBLearn [853]. DBMiner [1078]. DBMS [1478, 495, 1177, 331, 850, 1484, 745, 981, 235, 1245, 717, 1476, 1141, 1474, 260, 1176, 327, 714, 496]. DBMSs [504, 833]. DBSim [1076]. dbSwitch^{T M} [1922]. DBXplorer [1712]. DC [2271, 2270, 532, 2241]. DDB [723]. DDT [843]. deadlock [205, 255]. deadlock-free [205]. debit [251]. debugger [1702]. debugging [1290]. DEC [828]. decade [1260]. decaying [1765, 1899, 2070]. DECdtm [711]. Decidability [1234, 1209, 139, 252, 1772, 1099, 2030]. decidable [137]. Deciding [87, 2214, 1090, 1223, 1960, 186]. Decision [2036, 838, 958, 1158, 289, 839, 1330, 1693, 1039, 1240]. Declarative [1928, 281, 229, 2178, 1464]. Declustering [313, 1621, 523]. decomposable [2156]. decomposed [1670]. Decomposing [129, 53, 269]. Decomposition [212, 776, 28]. Decomposition-Based [776]. Decompositions [1303, 2032, 83]. DEDALE [1222]. Deductive [921, 522, 519, 456, 105, 1744, 602, 76, 188, 417, 462, 1319, 391, 186, 449, 466, 281, 119, 90, 488, 741, 665, 552, 849, 824, 370, 355, 464, 713, 247]. deep [1930, 461, 1848]. Default [517]. deferred [280, 1063]. defined [512, 212, 2226]. definedness [1960]. degradation [1031]. degrees [2003]. del [2248]. DELAB [217]. Delaunay [1160]. Delay [1194]. Delay-Sensitive [1194]. delayed [548]. delays [1206]. deletes [310]. Deleting [67]. deletion [2170, 2201]. deletions [1642]. Delivering [1278, 1546]. delivery [1400, 1704]. demand [937, 1561]. demo [1162]. demonstration [1820, 1593, 746]. demonstrators [1492]. Dense [878, 1934]. Dense-order [878]. densification [1913]. Density [1426, 1427]. density-based [1427]. Denver [2265, 2266]. dependence [494, 265]. Dependencies [4, 2077, 53, 1315, 588, 1021, 2092, 102, 2116, 104, 2152, 2042, 1894, 2201, 224, 1062, 1972, 82, 729]. dependency [1598, 36, 1538, 2004, 224]. dependency-based [1538]. dependent [2002]. deployment [1495]. depth [181]. derivation [193]. derived [1114]. Deriving [353, 2064]. describing [1614]. description [1820, 1104, 675]. descriptions [1748]. descriptors [1433]. Design [1194, 661, 1, 55, 572, 166, 553, 980, 2119, 1547, 1870, 220, 18, 938, 45, 187, 1243, 473, 614, 735, 827, 13, 12, 843, 2005, 150, 235, 1953, 484, 1492, 1602, 1041, 1795, 2145, 221, 344, 82, 1691, 121, 172, 709, 723, 29, 349, 101]. designers [163]. Designing [1478, 1457, 102, 1783]. designs [359]. desktop [1179]. Detecting [451]. detection [1509, 941, 1065, 1091, 470, 1052]. determinacy [1961]. determined [2033]. Determining [2172, 1984, 2160]. determinism [531, 574]. Deterministic [2021, 2188, 387, 388, 105, 2019, 439, 912, 2022, 464]. Developing [1275]. development [1702, 1931, 728, 743]. DEVise [1162, 1138]. diagnosing [1802]. Diagnosis [1990]. diagram [852]. Diamond [1853]. dichotomy [2060, 2201, 2237]. dictionaries [1360]. Diego [2257, 2269, 2290, 2289, 2268]. DIET [1940]. differences [1121]. different [2151]. different-sized [2151]. Differential [1989, 2135]. Digital [1273, 1224, 1457, 1692, 941, 24, 747]. Dimension [753, 1645, 2112, 1812]. Dimensional [1361, 1425, 1549, 1509, 1681, 1199, 1101, 1260, 1564, 1143, 1853, 1794, 1538, 1554, 881, 996, 1054, 1142, 1458, 1988, 1714, 1453, 1144, 455, 1191, 1565, 1336, 477, 1808, 1716, 2165, 2113, 1235]. Dimensionality [1212, 1549, 1208, 1388, 1530]. Dinur [2190]. Direct [9, 1389, 422, 508]. directed [1882]. direction [462]. directional [908]. directions [346, 1037, 1954]. directories [1323]. Directory [1348, 1481, 725]. DIRECTV [946]. disambiguation [779]. disc [323, 165]. disclosing [1216]. disclosure [1888]. Disco [1170]. discovering [1614, 1872]. Discovery [768, 1374, 1373, 777, 853, 1894, 973, 1934, 1716, 775, 988, 1174]. discrepancies [437]. discrepancy [1621]. discrete [1990, 328]. Discrimination [599]. discs [323]. discussion [42]. DISIMA [1501]. disjointness [289]. disjunct [234]. disjunction [301, 804]. disjunctions [1965, 816, 1740, 1042]. Disjunctive [664, 2231, 815, 380, 275]. Disk [601, 606, 394, 1643, 10, 694, 1715, 691, 1225, 527, 503, 368, 500]. Disks [1194, 910, 238, 1115, 198]. disorder [200]. disparate [1575]. Dissemination [1587]. Distance [1143, 1751, 1229, 1987, 1448]. Distance-based [1143]. distinct [1980, 2128]. distorted [693]. Distributed [2119, 1730, 536, 109, 550, 56, 6, 1633, 801, 249, 561, 1022, 993, 195, 504, 1938, 42, 287, 722, 2131, 1416, 805, 1396, 269, 1496, 644, 1154, 2210, 132, 1508, 696, 1620, 446, 695, 1439, 1498, 2211, 1578, 93, 1801, 1501, 38, 435, 415, 743, 497, 256, 1437, 255, 363, 1572, 305, 721, 40, 306, 658, 1170, 716, 116, 371, 2110, 2023]. Distributing [803, 1707]. Distribution [586, 1798, 33, 545, 211, 265, 1155, 535]. distributions [2184, 1776]. distributor [828]. diversification [2198]. diversity [1950]. divide [1974]. divide-and-merge [1974]. division [1963, 896]. DLFM [1467]. DNA [1606]. DNA-miner [1606]. do [959, 544, 928]. Document [1182, 510, 10, 174, 1433, 2025, 1739, 1753, 808, 988, 1174, 1274]. documents [1966, 1798, 2102, 1237, 1630, 1593, 1075, 2176, 941, 813, 1373, 1433, 1759, 1728, 1074, 1625, 1755, 1536]. DOM [2087]. domain [221]. domains [1220, 1312, 668, 618]. Dominated [810, 809]. DOMINO [1401, 1380]. Domino/Notes [1380]. DOODLE [537]. door [1930]. Doubly [693, 160]. down [2149, 291]. downward [2114]. dozen [2182]. driven [2221, 1921, 2218, 1577, 2001, 1496, 2013, 1082, 1842, 1573, 1699]. DS [1693]. DSQ [1443]. DSS [959]. DTD [1825]. DTD-conforming [1825]. DTDs [1959, 1523]. DTL [1297]. duality [2214]. duplicates [935, 2158]. during [451, 1253, 1031]. Dwarf [1682]. DWQ [1492]. DyDa [1607]. Dyn [794]. Dyn-FO [794]. DynaMat [1365]. Dynamic [1798, 1799, 1133, 919, 106, 1604, 929, 1228, 332, 138, 526, 1561, 794, 2195, 1236, 1679, 500, 2113, 1876, 1883, 1122, 2198, 1592, 1577, 1663, 783, 1375, 1806, 1587, 805, 1937, 2227, 1212, 23, 455, 1365, 557, 1918, 1879, 73, 2089, 1745, 561, 1409]. Dynamically [1508, 1778, 1632, 512].

e-business [1817, 1696, 1579]. E-commerce [1473, 1394]. E-salesperson [1588]. E-services [1614, 1725]. each [956]. early [1749]. earth [725, 506]. easier [521]. easy [1297, 1486]. ebXML [1686]. eCommerce [1479]. ECRC [991]. ed [1393]. Eddies [1441]. edit [1751]. edition [964]. Efendi [982]. Effective [2067, 1864, 433, 928, 331, 873, 1453, 1920, 208, 1056, 1879, 1454, 904, 1934, 1705, 1536]. effectiveness [777]. effects [1549]. Efficiency [362, 883, 2009, 72, 1724]. Efficient [866, 1213, 314, 1148, 396, 47, 360, 2206, 2176, 679, 1931, 1304, 1779, 1522, 1028, 1548, 149, 44, 1503, 1845, 1201, 2024, 1369, 457, 1423, 1252, 936, 1754, 554, 309, 1454, 1668, 1455, 1618, 1440, 1902, 1314, 59, 1885, 1689, 1536, 1631, 1545, 1637, 1670, 1559, 1818, 1869, 2111, 407, 589, 856, 301, 1338, 1780, 1844, 1431, 1981, 2019, 1794, 277, 1513, 524, 1195, 1550, 758, 2222, 2105, 1737, 2022, 1432, 1050, 1345, 1056, 2208, 625, 660, 2194, 2162, 1846, 1893, 824, 1572, 1812, 2209, 359, 1015]. Efficiently [1197, 1137, 1305, 148, 2011, 1114, 1824, 1036, 2014]. effort [1628, 1849]. eggs [142]. Eighteenth [2283]. Eighth [2261, 2296]. Electronic [1215, 1257, 558, 1317, 898, 1809, 1378]. elements [2128]. Eleventh [2269]. eliminate [584]. Eliminating [1149]. elimination [1811]. Embedded [53, 2116, 104, 1972, 452]. embeddings [1751]. embodied [224]. EMC [1389]. emerging [969]. enable [1604, 1464]. enabler [963]. Enabling [1917, 1577, 1712]. Encapsulation [365]. Encina [721]. encoding [1338, 1806]. encrypted [1654]. encryption [1887]. end [427]. end-user [427]. Energy [758]. enforcing [1153]. engine [1939, 1414, 1945, 1582, 1499, 824, 1700, 1485]. Engineering [541, 1070, 342, 751, 152]. engines [1278, 1466]. Enhanced [591, 1239, 1176, 1856]. enhancement [327]. Enhancing [915, 730, 1969]. Enough [1127, 1250, 302, 1908, 1904, 1861]. Ensuring [567, 765]. Enterprise [947, 841, 1262, 1932, 716, 1178, 963]. enterprise-wide [716, 1178]. entities [2130]. entity [2036, 343]. entity-relationship [343]. entropy [2129]. enumerating [1461]. Enumeration [2238]. envelope [1758]. envelopes [659]. Environment [271, 852, 722, 1179, 943, 743, 1403, 285, 260, 649, 1677, 934]. environments [910, 752, 18, 1122, 1607, 1923, 173, 1314, 731]. EOS [854]. ephemeral [669]. Epidemic [1117]. Epistemic [2078]. Epsilon [1564]. equality [2114]. Equi [181]. Equi-depth [181]. Equivalence [2116, 1999, 2115, 655, 1094, 530, 774, 2200, 492, 1092, 2048, 1627, 80]. Equivalences [1541, 1223]. equivalent [1403, 160]. Error [482, 2109, 313, 1683]. Error-constrained [482]. errors [480]. Estimating [1986, 1907, 1785, 2055, 1047, 2068, 787]. Estimation [670, 755, 1344, 865, 1302, 1681, 1864, 785, 1680, 1046, 1980, 1571, 619, 920, 1336, 345, 351, 1251, 2208, 2162, 1048, 2129, 649, 668, 1752, 1570, 2053]. estimators [1343, 2155, 236]. evaluate [1638]. Evaluating [1800, 1058, 265, 2066, 1674, 2193]. Evaluation [666, 52, 522, 682, 592, 519, 556, 1859, 65, 377, 1976, 1744, 925, 1882, 2085, 2176, 301, 1243, 451, 783, 91, 1522, 32, 2219, 1757, 332, 149, 482, 626, 2073, 1508, 1203, 669, 2132, 2173, 309, 452, 1846, 2099, 1141, 611, 166, 380, 285, 403, 1526, 442, 890, 444, 191, 275, 1631, 1235]. EVE [1403]. Event [540, 1990, 1598, 1177, 2168, 1082, 328, 2057, 1842]. event-based [1177]. event-driven [1082, 1842]. Every [266, 953]. everything [1069]. Evolution [2006, 154, 243]. Evolutionary [558]. Evolvable [1403]. Evolving [838, 1802, 2040, 1145]. Exact [1318, 1819]. example [2216]. examples [2147]. Exceeding [2121]. exchange [2074, 1958, 2161, 1828, 949, 2044, 1742, 2095, 1971, 1970, 1996, 2042, 1962, 1995, 1998, 2075, 1888]. exchanged [2063]. Exchanging [1777]. exclusive [1254]. Executing [1654]. execution [448, 237, 942, 1907, 1912, 929, 516, 876, 528, 1059, 1550, 789, 1353, 1201, 1556, 1461, 1689]. Executions [652, 1149]. Exercise [1324]. existence [654]. existential [194]. EXODUS [261, 568, 124, 133]. expansion [2238]. expansions [460]. expect [1667]. expectations [2121]. Expected [1324, 894, 1667, 481]. expensive [1672, 686, 1059, 815, 2052]. Experience [810, 603, 809, 2123, 1394]. Experiences [689, 1581, 1248, 115]. experiment [1179]. Experimental [52, 150, 258]. experiments [1722]. Expert [41]. Explicit [197]. exploit [1784]. Exploiting [1662, 331, 1585, 520, 1928]. exploration [1297, 1711, 1162, 1138, 1716, 998]. Exploratory [1183, 1404]. exploring [1591, 1936]. Exponential [792, 1644]. exponential-time [1644]. express [387, 388]. Expressibility [326, 324, 2140]. Expressing [1465, 1444]. expression [281, 1713]. Expressions [1335, 1207, 538, 1662, 1866, 1058, 1773, 2188, 236, 779, 2193, 1525, 182, 392]. Expressive [2124, 995, 800, 356, 888, 877, 1202, 2225, 1619, 184, 292, 2071, 21, 178, 322, 383, 355, 153, 485, 2237, 2038]. Expressiveness [2039, 1182, 539, 436, 892, 139]. expressivity [2026]. extend [1312]. Extended [1766, 1525, 650, 394, 432, 361, 781, 1117, 993, 1094, 911, 63, 483, 487, 1017, 1035, 1006, 761, 472, 908, 759, 863, 895, 1099, 804, 1038, 1028, 77, 807, 878, 999, 1125, 1014, 1097, 494, 2232, 796, 492, 1092, 889, 1032, 467, 678, 208, 885, 900, 1040, 1090, 351, 477, 343, 460, 1649, 624, 790, 2187, 802, 1088, 806, 892, 611, 887, 1042, 756, 468, 80, 1130, 383, 43, 470, 766, 413, 997, 764, 1005, 1295, 353, 668, 153, 485, 574, 926, 445, 699, 1470, 903, 931]. extender [505]. extendible [74]. Extending [280, 627, 381, 171, 1886, 486, 355]. extensibility [1246]. Extensible [336, 525, 854, 1491, 1500, 1940, 1293, 1437, 1440, 835]. Extensible/rule [525]. extension [1900, 136, 1497]. extensions [831, 655, 1362]. extent [1637]. External [22, 676, 757, 1204, 942, 1978, 2227, 1253, 1040, 756, 363, 2163, 2165, 2195]. Extracting [1782, 450, 1238, 1237, 1600, 1433]. extraction [1818, 1709, 2215, 12, 1619, 1399, 957, 1705]. Extrema [921]. Eye [1714].

F [290]. F-logic [290]. FAcADE [1934]. face [1267]. facilities [813, 682]. facility [508, 1408]. FACT [1488]. factoring [903]. Factorizing [1784]. failure [75]. Failures [1023]. family [436, 721]. far [1442]. FARMER [1852]. Fast [1310, 582, 1087, 1342, 822, 1858, 1647, 1586, 1955, 2133, 1524, 901, 150, 1472, 1678, 1934, 16, 1987]. Fast-Start [1586]. FastMap} [901]. FastSort [363]. Fault [938, 1010, 1599, 59, 1586, 559, 587, 1909, 954]. Fault-Tolerant [59, 1010, 1599, 559, 1909, 954]. FC [2224]. FD [84, 81]. FD-acyclicity [81]. FDM [44]. Feasibility [672]. Feasible [1776]. feature [1803]. Features [532, 437, 717, 836, 987]. featuring [1495]. Federated [546, 855, 1687, 982, 677]. federation [1157]. feedback [785, 1804, 163]. fetches [787]. few [1986, 1978]. FIFO [2156]. Fifteenth [2276]. Fifth [2255, 2293]. File [594, 940, 1685, 145, 211, 311, 460, 944, 200, 987, 801, 500]. FileNet [1274]. files [394, 811, 12, 213, 682, 697, 695, 416, 24, 308]. filling [2203]. filter [2012, 2069]. Filtering [875, 1524]. filters [1874, 1768, 2052, 1801, 1841]. finally [841]. finance [1955]. financial [1842]. find [949, 1551, 36, 1774]. find-statements [36]. Finding [1425, 659, 2041, 1449, 2010, 2174, 2023, 1852, 1853, 2158, 2022, 288, 2163]. Fine [817, 2239, 1886, 411]. Fine-grained [817, 1886, 411]. fingerprinting [1739]. Finite [928, 2047, 436, 2100, 2030, 81, 787]. Finitely [807]. finiteness [790]. firm [1152]. First [2288, 1003, 1228, 324, 2238, 75, 2140]. First-Order [1003, 1228, 324, 2238, 75, 2140]. fits [959]. Five [150]. Fixed [670, 1532, 326, 796, 266]. fixed-point [326]. Fixed-Precision [670]. fixed-size [1532]. Fixpoint [657, 3, 2143, 227, 264]. Flag [2226]. flat [182]. flavor [79, 1687]. flaws [1052]. Flexible [1510, 498, 1847, 554, 660, 765]. FleXPath [1847]. flocks [1181]. Florida [2247]. Flow [2012, 575, 2069]. FlowBack [1285]. flows [720]. FO [794]. FOIES [895]. footprint [1376]. forest [1475]. form [1630, 944, 121]. Formal [1788, 175, 130, 257, 1111, 2215, 843, 1888]. formalism [425]. formalization [270, 378]. formats [949]. Forms [639, 1727, 288]. formulas [155, 247]. foruth [1390]. foruth-generation [1390]. forward [391]. foundation [66]. Foundations [2144, 2212, 2031, 2189]. Founded [666, 2232, 268, 225, 555]. four [285]. Fourier [2220]. Fourteenth [2274]. Fourth [2253, 2292]. Fractal [1643, 753]. Fractals [312]. fragment [1627]. fragments [2051, 1960, 2038]. Framework [1184, 1322, 519, 1681, 1802, 1444, 1968, 2101, 382, 629, 1943, 2215, 372, 1129, 1064, 1153, 398, 2191, 208, 449, 1590, 246, 963]. France [2291]. Francisco [2258]. Free [1280, 78, 205, 1056, 1420]. freely [401]. freely-reorderable [401]. frequencies [2210]. frequent [1551, 1778, 2214, 1419, 2213, 2105, 1328, 2022, 1404, 1493, 1868]. fresh [1717]. freshness [1429]. friendly [1553, 420, 1927]. frontier [2229]. FTW [1987]. Full [53, 535, 1847, 1945, 1965, 1740, 1042]. full-text [1847, 1945]. Fully [502, 1607]. Function [478]. Functional [571, 4, 1315, 796, 601, 569, 2092, 490, 102, 876, 104, 1894, 2201, 1964, 224, 180, 82]. functionality [654, 1495]. Functions [683, 784, 206, 2198, 1245, 1533, 2237]. Fundamental [1004, 238]. Future [1037, 1382, 94]. Fuzzy [1180, 1780, 1038].

G [419]. gains [498]. game [1537]. Games [918, 917, 322, 1277]. gamma [79, 254]. gamma-acyclicity [79]. garbage [818, 1055, 267, 667, 1139]. Garlic [1687, 1085]. GEA [1713]. gene [1713]. General [1350, 1675, 628, 301, 2101, 629, 881, 1002, 2159, 16, 225]. general-purpose [16]. generalization [1181]. Generalized [2032, 2094, 1425, 1058, 231, 1098, 1675, 475, 1186, 248]. Generalizing [1216, 1135]. generated [1632]. Generating [1559, 1021, 799, 1130]. Generation [1184, 175, 1604, 702, 130, 1390, 2267, 963, 1920, 1253, 969, 1256]. generation0 [1419]. generator [1701, 124, 1399]. generators [908]. generic [2069, 1760, 1408]. Genericity [1016, 1017]. Genesis [38]. GENESYS [856]. Genome [992, 891]. geo [1141]. geo-spatial [1141]. Geographic [1011, 703]. geometric [2091]. geometrical [1097]. geometry [1369]. geometry-based [1369]. GeoMiner [1172]. Georgia [2250]. Get [2155, 1720, 1291]. getting [1742]. gigabyte [1715]. Gigascope [1708, 1810]. GIOSS [777]. GIS [746, 1690, 1502, 986]. given [1972]. Global [1542, 650, 2221, 644, 704, 2023]. globally [993]. Glue [488, 661]. Glue-Nail [488, 661]. go [1608, 2181]. Goal [1053, 1496]. goal-driven [1496]. Goal-oriented [1053]. goals [2182]. Going [1718, 1466]. golden [1570]. good [1538, 1479, 1908, 733]. gossip [2024]. gossip-based [2024]. grained [817, 1886, 411]. Grammar [180, 539, 316]. grammar-based [316]. Grammar-like [180]. Grammars [1182]. granularities [1006]. granularity [1756]. Graph [2183, 676, 1868, 390, 2228, 2233, 2124, 2171, 1750, 2190, 4, 364, 733, 426, 427, 203, 2230, 2006, 1673, 2068, 1623]. graph-oriented [733, 426, 427]. graph-structured [2124, 1750, 1673]. Graph-theoretic [390, 4]. graphical [350, 420, 852, 156, 422, 344]. graphics [1858]. GraphLog [419, 425]. Graphs [48, 772, 2235, 401]. Greece [2298]. Greedy [549, 2153]. GREO [718]. grid [1564, 145, 213, 1838, 697, 1815]. GridDB [1815]. ground [705]. Group [2259, 2258, 1363, 1460, 201, 1520]. group-by [1460]. groups [1854, 1852]. groupware [340]. growing [803]. Guaranteeing [650]. guarantees [2151, 1683, 1656, 1115]. guards [1537]. GUI [1067]. guidance [608]. guide [1809]. guided [418].

H [557]. H-trees [557]. habitat [1950]. ’half [358, 358, 2182]. ’half-and-half [358]. half-dozen [2182]. Handling [107, 424, 862, 1470, 277, 1646, 338]. handwritten [898]. Hangul [562]. happening [1384]. Hard [357, 276, 1462]. hardness [2036, 2032]. Hardware [1792, 718, 16]. harvesting [2130]. hash [789, 647, 1044, 1660, 416, 74, 645, 904, 308]. hash-based [904]. hash-joins [1044]. hashes [2082]. Hashing [215, 594, 2220, 23, 22, 24, 2150, 73, 695]. haystack [1519]. HD [1714]. HD-Eye [1714]. healthcare [1947]. heap [267, 667]. Heavy [1853, 2109, 1988, 2220, 2110]. Height [1363]. Height-Balanced [1363]. helps [1119]. heterogeneity [1096]. heterogeneous [504, 1165, 1359, 1220, 985, 1218, 435, 1263, 979]. heuristics [334]. hidden [1515, 1849]. hiding [1031]. Hierarchical [1681, 1517, 2082, 1647, 316, 1988, 1879, 147, 349, 1853]. hierarchically [1065]. hierarchies [263, 339, 326]. hierarchy [3, 278, 1861, 1186]. High [1131, 1260, 253, 948, 545, 1422, 233, 1425, 1549, 1509, 1681, 1199, 1938, 1101, 589, 1564, 1143, 1278, 282, 1708, 1794, 1538, 615, 111, 1714, 1144, 1279, 1565, 113, 1439, 625, 1275, 1293, 1716, 493, 35, 820, 1605]. high-availability [282, 1439]. High-dimensional [1260, 1101, 1564, 1143, 1538, 1144, 1565, 1716]. high-level [493, 35]. High-performance [1131, 113, 1293]. high-scale [1279]. higher [2126, 290]. higher-order [2126, 290]. Highly [1909, 1095, 1817, 1478, 1923, 1578]. HiLog [468]. hippocratically [1947]. histogram [1540, 1250, 1538, 1647, 2184, 920, 1336]. Histograms [1977, 1332, 1542, 1251, 181, 1048, 1679]. Historical [572, 1860, 27, 1884, 1618, 319]. Histories [690]. hitters [2109, 1988, 2220, 2110, 1853]. Hoare [2087]. hoc [1940, 1137]. hold [706]. Holistic [1669, 1843, 2058]. HOMER [1487]. homogeneous [205]. Honey [1376]. Hoo [1929]. horizontal [1870, 1915]. horizontally [14]. Horn [3, 277, 157, 662]. Hosting [1914]. Hot [1031, 1778, 445]. Hotel [2248, 2250, 2244, 2245, 2254]. humming [1920, 1830, 1758]. Hy [734]. Hybrid [219]. HYDRO [435]. Hygraph [734]. Hygraph-based [734]. Hypergraph [933, 1125]. hyperlinks [1239]. HyperStorM [1075]. Hypertext [1381, 1239, 1333]. Hypertree [1303, 1537, 2032]. Hypothetical [320, 1129].

I/O [1034, 2222]. I/O-efficient [2222]. IBM [717]. Iceberg [1364, 1503, 1507]. Iceberg-cube [1507]. icebergs [2023]. icons [978]. IDEA [1296, 998]. identification [2036]. identifiers [492]. identify [1308]. Identifying [1851, 1427, 2214, 2105]. identity [294, 459]. IDLOG [355]. if [1244]. ignorance [441]. II [2223, 505]. Illinois [2259]. IM [625]. Image [1494, 976, 1601, 1804, 1163, 1499, 1367, 1501, 1812]. images [1469]. iMAP [1872]. impact [1734, 1583, 315]. implement [63]. Implementation [628, 610, 773, 115, 572, 665, 154, 210, 614, 735, 1632, 661, 1686, 1524, 122, 1041, 151, 1141, 131, 133, 64, 1060, 980]. Implemented [2]. Implementing [602, 367, 1386, 1719, 1036, 260, 1880, 907, 1641, 1129, 401]. implication [1133, 299]. implications [15, 958, 827, 1099, 461, 224, 1242]. implicit [2197]. implied [2002]. imprecise [1800]. IMPRESS [843]. improve [565, 1429]. Improved [1093, 1048, 1426]. Improving [1544, 1397, 587, 1724, 2077, 1241, 1904]. In-memory [1392]. in-network [1979]. INA [421]. incidence [2142]. incidents [1933]. Including [588, 452]. Inclusion [4]. Incompatible [1354]. Incomplete [484, 779, 1340, 914, 2159, 1529, 2118, 1772, 1756, 277, 122, 189, 1998, 2030, 89]. inconsistency [590]. Inconsistent [1311, 1772]. Incorporating [2072, 278, 1536]. Increasing [354, 874, 922, 590]. Incremental [1882, 935, 1229, 2132, 1879, 808, 444, 1881, 705, 1965, 667, 1892, 1103, 1430]. incrementally [663]. indefinite [618]. Independence [49, 209, 1001, 1538, 375, 753]. Independence-reducible [209]. independent [360, 90, 486, 479]. Index [642, 1475, 2048, 1020, 1596, 1260, 1532, 1243, 1244, 1284, 1544, 1750, 1636, 547, 1144, 1528, 1565, 557, 625, 1893, 1468, 1511, 787, 1745]. Index-based [2048]. Indexability [1361, 2196, 2113]. indexed [1867, 2202]. indexes [1784, 1639, 1905, 455, 1690, 623, 1093, 2112, 2156, 1700, 1881, 2113, 1758]. Indexing [2107, 1469, 1361, 1890, 968, 1956, 678, 1388, 1346, 1211, 1447, 1189, 1368, 1770, 898, 2111, 1143, 1452, 1943, 901, 1554, 1612, 1769, 1132, 758, 1453, 1530, 455, 1191, 1944, 1454, 2112, 896, 1812, 1145, 1891, 2164, 1868]. Indiana [2297]. Indianapolis [2297]. indicator [1906]. indices [1145, 1095]. indivisibility [2196]. Induction [509, 1519]. Inductive [322]. industrial [1807]. industry [732, 727]. inequalities [2000]. inequality [472]. inexpensive [198]. Inference [300, 472, 1320, 1339, 494, 162, 272, 1313, 2090]. Inferred [517]. Infinite [880, 424, 862, 1875, 483, 188, 297, 157, 295].

Influence [1436]. InfoHarness [979]. Infomaster [1168]. Information [1741, 1901, 762, 984, 2138, 580, 1224, 749, 1924, 517, 648, 1529, 1818, 1917, 993, 1727, 750, 11, 2118, 1418, 1122, 1950, 769, 1359, 1065, 2155, 1884, 1761, 1756, 1038, 2215, 2103, 1168, 34, 1619, 277, 122, 1494, 243, 1203, 39, 1333, 1389, 2004, 189, 1336, 1407, 1998, 2159, 1399, 1252, 477, 730, 841, 1888, 704, 960, 1042, 1618, 2030, 172, 1216, 979, 725, 658, 1170, 89, 1622, 1705, 2130, 712, 29, 275]. Information-theoretic [1901, 1727, 2004]. Informix [547, 965, 1470, 1497]. InfoSleuth [1122, 1169, 955]. infrastructure [1686, 1936, 1944]. INGRES [627]. Inherent [1326]. inheritance [339, 331, 290, 466, 459]. initial [1206]. ink [898]. Inktomi [1278]. innovation [1919]. input [363, 352]. input/output [352]. inputs [1300]. ins [1069]. insert [1927]. insert-friendly [1927]. inserts [310]. Instance [47, 108, 283, 360, 2002]. instance-dependent [2002]. Instances [58, 2064]. instant [649]. instruction [1856]. instructions [1641]. Instrumental [743]. Integrated [532, 1494, 1949, 39, 1162, 1138, 1274]. Integrating [1870, 389, 1500, 1042, 1242, 750, 962, 1932, 1417, 1930, 1848]. Integration [1220, 1809, 1300, 971, 1122, 1775, 253, 1613, 1431, 2044, 1984, 1925, 1582, 1168, 1064, 1353, 1873, 1905, 1657, 985, 1618, 732, 1579, 1398, 1705, 712, 1871]. Integrity [46, 672, 175, 19, 534, 280, 598, 342, 130, 1523, 644, 228, 1061, 927, 89]. intelligence [1925]. Intelligent [712, 280, 1490]. intensional [1035, 1777, 2237]. intensive [195, 1709, 1822, 1498, 1487, 649]. inter [520, 730]. inter-operability [730]. inter-operation [520]. interacting [1474]. Interaction [1309, 1203]. Interactive [517, 1084, 1078, 830, 1716, 1490, 998, 1848, 973]. InterBase [738, 710]. Interchange [1164]. Interconnection [546]. interdependencies [727]. Interest [2259, 2258]. interesting [1852]. interface [940, 350, 420, 418, 1708, 1416, 34, 422, 737, 1163, 1815, 35]. interfaces [1142, 427, 1764, 1930, 1848, 1849]. interfering [448]. intermediate [184]. internal [1057]. internals [748]. International [2275, 2290, 2291, 2244, 2271, 2245, 2284, 2266, 2243, 2282, 2287, 2263, 2280, 2277, 2246, 2254, 2278, 2241, 2247, 2286, 2242, 2273, 2268, 2251, 2256, 2262, 2300, 725, 2240]. Internet [1287, 1402, 1471, 1451, 1941, 1822, 1124, 1546]. Internet-scale [1822]. Interoperability [726, 728, 1167, 437, 633]. interoperable [1080, 1501]. interpolation [452]. interpretation [106]. interpreted [1102]. interpreter [259]. Interpreting [699]. interrelationship [142]. interrupted [1423]. INtersect [732]. Intersection [585]. intertask [729]. interval [1806, 1897, 455, 1151, 1005]. interval-based [1005]. Invariants [1198, 1188]. invention [330]. inverses [2043]. inverted [1905, 808]. Inverting [1997]. investigation [1490, 1107]. invisible [2037]. Invited [305, 306, 279, 2111, 2093, 1985, 1807, 2189]. Involving [54, 1006, 492]. IPSOFACTO [1832]. IRIS [1494, 423]. IrisNet [1822]. ISIS [34]. Island [2296]. islands [2153]. isn’t [1861]. isolation [860, 1975]. Isomorphism [1228]. Issues [10, 707, 526, 724, 760, 701, 1617, 204, 613, 1011, 889, 1781, 1378, 892, 1394, 1274, 1482]. Italia [1398]. ITASCA [536]. items [673, 1778, 2022]. Itemset [1184, 1133, 1776]. itemsets [2214, 2105]. iterated [266]. ITQS [1949].

Jaguar [1412]. Japanese [562]. Java [1259, 1262]. javax.XXL [1489]. jobs [1305]. Join [1349, 1301, 942, 1980, 53, 670, 755, 933, 1564, 1539, 1733, 1749, 1434, 1046, 1229, 404, 480, 1140, 224, 1646, 1660, 1300, 2187, 1045, 166, 1430, 285, 656, 1448, 177, 334, 668, 1000, 1752, 890]. Join-distinct [1980]. join-order [1000]. Joining [1897, 2061]. Joins [54, 1307, 791, 1669, 280, 1347, 1977, 1666, 1351, 789, 1963, 647, 793, 1044, 1754, 624, 2015, 645, 2104, 1902, 2066, 403, 649, 1792, 369, 1689]. joints [679]. Jose [2275, 2274, 2249, 2240]. July [2296]. June [2259, 2269, 2276, 2281, 2283, 2288, 2290, 2289, 2291, 2292, 2293, 2294, 2244, 2262, 2243, 2282, 2287, 2280, 2300, 2277, 2295, 2298, 2296, 2241, 2247, 2268, 2297, 2251]. Jungle [1414]. just [1590, 1481].

K-d [30]. Kabiria} [510]. Kaleidoscope [418]. Kalman [1841]. Kanellakis [1008, 1009]. Kanji [562]. KBMS [980]. keep [1946]. kernel [131]. Key [594, 312, 2073, 73]. keyed [1094]. Keynote [506, 951, 952, 428]. keys [2151, 1532, 24]. keyword [1712, 1728, 2010]. kind [145]. Kleisli [1298]. KNN [1794]. Knocking [1930]. know [953, 242]. Knowledge [899, 1651, 509, 768, 5, 517, 314, 993, 491, 302, 13, 596, 117, 2003, 853, 1078, 132, 243, 422, 162, 379, 221, 258, 14, 493, 739, 980, 2130]. knowledge-base [162]. knowledge-based [14]. knowledge-theoretic [117]. Knowledgebase [593]. Korean [562]. KRISYS [980]. KSR1 [719]. Kweelt [1590].

lab [1296]. Labeling [1663, 1854]. labels [1927]. LaMansion [2254]. lambda [796]. Language [788, 575, 437, 577, 572, 786, 1202, 144, 294, 2179, 271, 98, 421, 2027, 869, 1066, 1374, 261, 187, 1837, 156, 537, 259, 843, 1390, 290, 846, 737, 1041, 466, 272, 8, 464, 179, 493, 35]. Languages [995, 1102, 1318, 1167, 897, 868, 872, 641, 784, 1182, 1217, 782, 639, 229, 387, 388, 490, 2049, 2124, 507, 1535, 1614, 1057, 1431, 2143, 873, 876, 2146, 1619, 489, 1097, 796, 436, 402, 871, 1964, 178, 2230, 542, 43, 355, 1005, 434]. Large [591, 1600, 314, 673, 1400, 1901, 600, 342, 1264, 2017, 1195, 235, 893, 1382, 579, 404, 1865, 983, 1530, 667, 1137, 1162, 1138, 1139, 2197, 1345, 623, 1454, 1455, 1954, 527, 260, 1812, 1421, 994, 177, 334, 1676, 1470, 1015]. large-scale [2197]. larger [2153]. lasting [1583]. Latent [1211]. laws [1434]. Layered [585, 1375]. layout [1808]. Lazy [1859, 696, 1119]. LDAP [1348, 1481]. LDL [179]. LDL1 [98]. Learning [2216, 524, 1698, 1488, 1575, 1125, 2149, 1462]. Least [1324, 1667, 266]. leaves [460]. Left [461, 309]. left- [309]. Left-deep [461]. legacy [724]. lenses [2027]. LEO [1698]. Lessons [1155]. level [253, 1864, 1863, 1078, 494, 573, 1693, 493, 430, 35, 66, 366]. levels [860, 1975]. Leveraging [960]. LexEQUAL [1948]. LH [695, 1439]. Libraries [1224]. library [1489]. life [425, 603]. light [989]. Lightweight [1077]. like [1286, 1585, 180]. limitations [182]. Limited [1301, 1355, 1040, 1249]. Limiting [1763]. limits [130]. line [1020, 1289, 1828, 579, 1424, 1422, 821, 1146, 617, 1018]. Linear [1213, 1358, 409, 997, 1202, 2049, 602, 2085, 320, 1452, 1134, 2188, 1092, 2184, 1333, 2135, 453, 309, 2099, 73, 252, 160, 695]. Linearising [298]. linearity [1099]. linearly [618]. Linguistic [2189]. linkages [1536]. Liquid [1932]. List [788]. List-Based [788]. lists [1905, 808]. LLO [466]. Load [1938, 358, 606, 805, 1599]. load-balancing [1599]. Loading [675, 697]. loads [1423]. Local [50, 2200, 2087, 644, 56, 2221, 1427, 2153, 1739]. Locality [1360]. Locality-preserving [1360]. Locally [1530]. Location [1791, 1634]. Location-based [1791]. lock [551, 255, 581]. locked [68]. Locking [778, 919, 556, 581, 584, 548, 123, 358, 112, 205, 223, 554, 1556]. Locks [354, 361, 874, 1556]. LockX [1824]. LOF [1427]. log [570, 1422]. logarithmic [2239]. logging [109, 669, 1312, 625]. Logic [1003, 270, 1185, 97, 362, 883, 877, 1230, 63, 135, 98, 543, 1993, 2176, 472, 335, 375, 2029, 458, 197, 290, 337, 178, 381, 333, 151, 232, 266, 230, 268, 447, 64, 385, 357, 383, 139, 179, 470, 766, 442, 225, 264, 353, 249, 378, 2086]. Logical [1312, 170, 61, 101, 105, 1537, 21, 1422, 1126, 651, 1846, 716, 86]. logics [1104, 2067, 2152]. LOGRES [744]. logspace [2214]. long [1197, 384, 1677]. long-running [384]. look [1720, 1725]. looking [1332]. loops [542]. loosely [866]. LORE [1077]. loss [565]. Lossless [1659]. Lots [1605]. Lotus [1380, 945]. low [253, 1707]. low-level [253]. low-power [1707]. Lowell [1835]. Lower [1978, 1189, 1644, 2017, 2035, 2113]. LRU [691, 787]. LRU-K [691]. LyriC} [869].

M [1833]. MA [2251, 2255]. Machiavelli [272]. Machine [2035, 844, 254, 1575, 1125, 563, 798]. machine-learning [1575]. Machinery [2259, 2258]. machines [17, 287]. macro [1982]. made [1486]. Madison [2288, 2287]. Magic [63, 622, 657, 405, 406, 393, 103, 903, 146, 1744, 773, 376, 1060]. Magic-sets [622, 773]. magnitude [966]. MAIDS [1933]. Main [1532, 582, 1387, 1794, 1528, 1293, 1459, 1635]. Main-memory [1532, 1387, 1293]. mainstream [1385]. maintainability [222, 286]. maintainable [102]. Maintaining [1767, 1765, 2040, 663, 1062, 1114, 2202]. Maintenance [120, 175, 526, 1463, 1105, 1148, 1607, 1594, 1063, 1147, 130, 84, 935, 889, 1865, 1173, 936, 1558, 1146, 1061, 1403, 1430, 202, 1881, 934]. Making [800, 391, 1459, 1406, 1648, 1715, 581]. Management [2259, 2275, 2290, 2291, 2244, 2271, 2245, 2284, 564, 2266, 2243, 2031, 2258, 2282, 2287, 2263, 2280, 2277, 2240, 1224, 2246, 1, 2254, 56, 2278, 2241, 2247, 1356, 2286, 59, 2242, 2273, 2268, 2251, 2256, 1821, 2179, 910, 314, 471, 760, 2144, 1165, 1938, 237, 857, 1860, 386, 1053, 844, 2182, 473, 31, 2262, 613, 2044, 566, 1265, 990, 1171, 1248, 1479, 183, 1496, 1656, 2300, 1082, 1179, 1838, 1841, 511, 1203, 1279, 1285, 347, 711, 1962, 1365, 1602, 199, 422, 1253, 1608, 446, 1407, 136, 1580, 1056, 304, 1281, 1760, 1583, 625, 1072, 1578, 218, 515, 1378]. management [258, 1486, 25, 739, 1885, 1392, 848, 1415, 1274, 1178, 731, 1568, 1409]. manager [220, 1387, 847, 989, 1467, 1007, 723, 1820]. Managing [1947, 45, 1264, 1096, 1032, 1033, 243, 795, 430, 1383, 600, 267, 747, 2059, 411]. Manhattan [2133]. manipulating [1049]. manipulation [422, 223, 359]. Map [577]. Map-Oriented [577]. mapped [820]. Mapping [1781, 118, 2063, 2144, 1359, 2065, 1595, 1699]. mapping-driven [1699]. mappings [2147, 2096, 1594, 1997, 2043, 1962, 2075, 2094, 1972, 2064, 1573]. maps [454, 1921]. March [2248, 2250, 2253, 2255, 2257, 2260, 2261]. MARELA [564]. marginal [477, 498]. Marina [2248]. market [1368, 1135, 1133]. Markov [2143, 2125]. Marrying [2054]. marshals [1537]. Maryland [2292]. mass [1854]. Massachusetts [2244]. masses [1027]. massive [1555, 1564, 966]. massively [838, 1079, 1007]. Match [583, 594, 1780, 211, 1675]. matches [1872]. Matching [1358, 1819, 1669, 822, 607, 397, 1764, 1762, 1948, 1675, 988, 1174, 1842]. materialization [169, 478, 1450]. Materialized [1232, 1596, 1558, 1061, 1560, 913, 1064, 936, 202, 1826]. Materializing [1732]. Matrix [564]. Max [2198]. Max-Sum [2198]. maximal [2213]. Maximally [2061]. maximization [2129]. maximize [72]. Maximizing [2170]. maximum [458]. May [2265, 2270, 2272, 2275, 2274, 2279, 2283, 2285, 2244, 2271, 2245, 2284, 2262, 2266, 2243, 2258, 2249, 2263, 2240, 2299, 2246, 2254, 2278, 2286, 2273, 2256]. MDM [1165]. meaning [163]. Meaningful [1091]. measures [1766, 1612, 1503]. Measuring [880, 1322]. Mechanism [52, 158, 16]. mechanisms [941, 2136, 1397, 715]. Media [1273, 1196, 575, 282, 771, 830, 1010, 604]. medians [1767, 1249]. mediated [936]. mediation [1418, 984, 1288]. mediator [1022, 1164, 1833, 1631]. mediators [1287, 1214, 1371]. medical [730, 221]. medium [1388]. medium-dimensionality [1388]. MEDUSA [844, 1938]. Meet [1224, 2152]. meeting [2249, 2251]. meets [1276]. memoriam [1008, 1009]. Memory [1253, 1, 1204, 582, 1655, 1771, 1387, 1994, 1532, 1794, 165, 1978, 1528, 113, 1249, 864, 795, 1293, 1459, 1635, 656, 2163, 2165, 2195, 1392]. memory-resident [113]. memoryless [494]. men [1377]. Mendelzon [2097, 2217, 2186, 2127]. MENTOR [1178]. mentorlite [1496]. menu [418]. menu-guided [418]. Merge [1391, 1974, 893, 1646, 1045]. merge/purge [893]. Mergeable [2185]. messages [413]. meta [1124]. meta-searching [1124]. metadata [1962]. metamessages [187]. metanoia [1576]. metasearch [1536]. Metatheory [861]. Method [348, 685, 526, 1368, 911, 407, 1592, 1977, 1290, 412, 621, 573, 466, 1031, 625, 660, 763, 1675, 1426, 649, 1745, 1015]. methodologies [18]. methodology [1974, 450]. Methods [640, 534, 627, 1304, 12, 168, 1985, 250, 1059, 1382, 325, 466, 151, 554, 103, 1752, 390]. Metric [1192, 1343, 1143]. METU [850, 1080]. mice [1377]. Michigan [2246, 2246]. microarray [1852]. Microsoft [1914, 1596, 1445, 1395, 1261, 1284, 1391, 1393]. Microsoft.com [1279]. Mid [1694, 1201, 1816]. mid-query [1201]. Mid-tier [1694, 1816]. Middle [1696, 1817, 1926, 1522]. Middle-tier [1696, 1817]. middle-ware [1522]. middleware [1400, 1521, 1498, 1495, 1437, 1526]. Migration [797, 1931, 686, 1729, 724, 411, 1876]. migration/cohabitation [724]. min [2082]. min-hashes [2082]. Mind [1359]. miner [1606]. Minimal [48, 1672, 1320, 1732, 1769, 2174, 114]. minimalism [76]. Minimality [662]. minimax [2136]. Minimization [1574, 2166, 1771, 578]. Minimizing [240, 1668]. Minimum [458, 884]. Mining [673, 767, 1658, 1030, 1419, 1519, 994, 1677, 851, 1199, 1456, 1547, 1901, 1197, 1006, 1933, 1381, 1981, 1763, 901, 1956, 996, 1377, 1720, 1125, 1078, 1073, 1172, 1136, 1606, 1325, 1494, 2213, 1280, 1991, 1581, 1270, 1183, 1404, 1426, 904, 1493, 1455, 1776, 1420, 1242, 1572, 1421, 1186, 1457, 1181, 1294]. Minneapolis [2272, 2273]. Minnesota [2273]. minute [165]. Miramar [2245]. Miramar-Sheraton [2245]. Miro [714]. mirrored [694]. mirroring [1031]. mirrors [693]. missing [2139]. mitigation [1951]. MIX [1418]. Mixed [467, 2115]. Mixed-approach [467]. MLPQ [1502]. MLPQ/GIS [1502]. MLR [573]. MLS [1319]. MMDB [499]. MMSNP [2231]. MN [2272]. Mobile [1346, 752, 1809, 754]. mobility [1412, 1376]. MOCHA [1495, 1437]. Model [1921, 1192, 564, 585, 558, 2, 1026, 541, 1598, 271, 1271, 108, 283, 360, 1165, 1555, 543, 1101, 274, 261, 210, 1899, 28, 566, 1446, 231, 876, 2013, 426, 427, 1654, 494, 484, 278, 889, 185, 438, 597, 33, 39, 257, 21, 262, 2100, 118, 343, 1760, 1487, 115, 266, 153, 116, 86]. model-based [1555, 1487]. Model-driven [1921, 2013]. model-theoretic [86]. Modeling [263, 1185, 1565, 333, 445, 731, 11, 771, 217, 638, 845, 170, 202]. modelling [105, 389]. Models [15, 48, 1617, 1614, 1985, 1916, 2118, 507, 1374, 1594, 1028, 1571, 688, 316, 678, 328, 2159, 1111, 858, 2006, 255, 385, 2035, 1752]. modes [163]. modification [1020]. Modified [23]. modifies [310]. modifying [612]. Modular [341, 376, 447, 745, 305, 306]. Modularity [715]. module [2169]. modules [677]. MOL [577]. Monadic [1619, 2047]. monadically [2226]. Monica [2245]. Monitoring [1598, 1569, 1730, 1708, 1937, 1282, 2091]. monitors [702, 1610]. monotone [2198, 2214]. Monotonic [552, 2211]. monotonicity [300]. Montage [835]. Monte [1678]. Montréal [2276, 2277]. mosaic [984]. mosaic-based [984]. MOSAICO [845]. most [2155, 1778]. mostly [840]. motion [1891]. movement [527]. MovINg [1401, 1680, 1446, 1649, 1447, 1783, 1891, 1482]. MTCache [1942]. much [1250]. Multi [791, 585, 1973, 1054, 575, 1336, 633, 1808, 503, 366, 1247, 1853, 1514, 929, 1142, 523, 1167, 1458, 1988, 1550, 1453, 991, 455, 1520, 1566, 573, 460, 1558, 221, 309, 1115, 604, 1440, 1210, 1448, 430, 1457, 66, 890, 1826]. multi-attribute [523]. multi-bucket [460]. multi-database [1167]. Multi-dimensional [1054, 1336, 1808, 1853, 1142, 1458, 1988, 1453, 455]. Multi-disk [503]. multi-domain [221]. multi-join [890]. Multi-Layered [585]. Multi-level [366, 573, 430, 66]. multi-linear [309]. Multi-Media [575, 604]. multi-parametric [1550]. multi-query [1514, 1558]. multi-resolution [1566]. multi-similarity [1247]. multi-stage [1448]. Multi-step [791, 1210]. Multi-structural [1973]. multi-terabyte [1457]. multi-user [929]. Multi-vendor [633]. multi-way [1520]. multi-XQuery [1826]. multi-zone [1115]. multiattribute [308]. multicast [1269]. Multidatabase [204, 727, 386, 738, 932, 720, 846, 567, 1062, 732, 765]. multidatabases [605]. Multidimensional [1189, 1292, 1540, 1304, 1621, 1233, 32, 1528, 1041, 2048, 181, 24, 74, 1679, 1334, 1116]. multigraph [1981]. Multikey [30, 599]. multilevel [377, 438, 374, 162, 70]. multilevel-secure [374]. multilexical [1948]. Multimedia [31, 1180, 1086, 1321, 1087, 769, 938, 1013, 1079, 1416, 901, 968, 707, 830, 1007, 1615, 1294]. MultiMediaMiner [1294]. Multiobjective [1512]. Multiple [594, 668, 750, 1165, 1006, 1594, 1147, 1766, 1038, 1078, 789, 19, 409, 1153, 1300, 508, 624, 234, 486, 1638, 369, 1235]. multiple-data [1165]. multiple-disjunct [234]. Multiple-Key [594]. multiple-level [1078]. multiple-relation [234]. multiprocessor [17, 285]. multiset [239]. multisets [2040]. MultiSync [1413]. multitransputer [743]. multivalued [82]. multiversion [341, 448, 203, 615, 325, 412, 127]. multiversioning [565]. MultiView [1083]. Multiway [1307]. music [1955]. musical [172]. mutually [1254]. MVS [1389]. MYRIAD [855].

Nagano [1277]. Nail [488, 661, 192]. naive [2219]. names [421, 1762, 620]. NAOS [975]. NASA [506]. Nashville [2264]. native [1608, 410, 1827]. natural [737, 266, 562]. NAUDA [737]. Navigation [580, 1289, 703, 2025, 1603]. NC [786]. Near [2089, 1621, 2041, 1769]. near-minimal [1769]. Near-optimal [2089, 1621]. Nearest [2204, 2223, 879, 1681, 1770, 1101, 1144, 1436, 1210]. Nearest-neighbor [2204]. nearly [1420]. necessary [160]. need [1214, 1396]. needed [378]. needle [1519]. Needs [792, 1011]. Negation [75, 662, 206, 98, 320, 1541, 227, 268, 376, 468, 385, 264]. negotiation [1142]. neighbor [2204, 2223, 1681, 1770, 1101, 1144, 1436, 879, 1210]. neighborhood [2174]. neighbors [2041]. Nested [491, 112, 1315, 319, 541, 318, 2115, 364, 317, 99, 226, 140, 748, 182, 121, 2237, 2086]. Network [2, 760, 1899, 1708, 1810, 1706, 1323, 1620, 1704, 1832, 1578, 1268, 1979, 725, 1877]. networked [1834, 93]. networking [2178]. Networks [56, 1131, 1883, 1884, 1940, 859, 1707, 1795, 1918, 218, 2089]. neuroscience [1833]. Next [1256, 1260, 2182, 1839, 2267, 1693, 2057]. next-generation [2267]. NF [121]. NiagaraCQ [1451]. NICE [529]. NICE-C [529]. NICE-OS [529]. nimble [1582]. Ninth [2264, 2297]. Nissim [2190]. NJ [2263]. NN [1670]. no [922]. node [1927]. NoDoSE [1237, 1405]. noisy [1934, 1677]. nomadic [701]. Non [387, 105, 2049, 439, 912, 652, 662, 448, 962, 339, 251, 205, 2211, 1403, 265, 464, 68]. Non-deterministic [387, 105, 439, 912, 464]. non-equivalent [1403]. Non-Horn [662]. non-interfering [448]. Non-linear [2049]. non-monotonic [2211]. non-relational [962]. Non-Serializable [652]. non-stop [251]. non-strict [339]. non-two-phase [205]. non-uniform [265]. non-well-locked [68]. nonblocking [581]. nondeterminism [385]. Nondeterministic [388, 588]. nonlinear [1770, 298]. Nonrecursive [1230, 530, 659, 774, 622, 1964]. Nonstop [839, 586]. Normal [639, 1630, 1727, 288, 121]. normalization [2004, 1920]. Normalizing [914]. notes [945, 1380]. nothing [1020, 844, 285, 658]. novel [813, 415, 1812]. NP [2032]. NP-hardness [2032]. NT [947, 1389]. null [191]. nulls [2095]. number [803]. Numeric [1030, 1887, 1081]. Numerical [1753].

O

[2217, 2186, 2127, 514, 2097, 1384, 1034, 715]. O-efficient [2222]. O-O [1384]. Object [294, 271, 207, 800, 1156, 685, 897, 173, 526, 831, 797, 943, 836, 1077, 690, 792, 529, 1023, 654, 653, 514, 154, 476, 1788, 854, 857, 1051, 1075, 1411, 167, 389, 1112, 187, 833, 974, 1177, 331, 908, 433, 614, 629, 818, 1055, 537, 850, 1049, 168, 873, 1190, 1484, 745, 876, 183, 540, 733, 847, 989, 489, 426, 427, 216, 615, 292, 492, 484, 1245, 597, 736, 152, 414, 478, 337, 630, 346, 834, 963, 1385, 1424, 595, 262, 333, 1163, 466, 1139, 832, 821, 651, 845, 1649, 1007, 802, 515, 255, 1083, 327, 954, 723]. object [739, 1052, 431, 560, 713, 1415, 820, 731]. Object-Based [690, 1023, 476, 908, 1163]. Object-Oriented [800, 526, 207, 173, 831, 943, 836, 654, 653, 514, 154, 1075, 1411, 389, 187, 833, 974, 1177, 331, 433, 629, 537, 850, 1049, 1190, 1484, 745, 183, 540, 292, 736, 152, 337, 630, 346, 834, 595, 262, 333, 466, 651, 845, 255, 1083, 327, 739, 1052, 713, 1415]. Object-relational [1156, 1112, 1245, 963, 1385]. Objectivity [535]. Objectivity/DB [535]. Objects [474, 788, 967, 963, 1346, 882, 471, 69, 600, 1878, 274, 869, 244, 210, 335, 433, 473, 1680, 188, 450, 1446, 512, 2082, 409, 457, 290, 329, 1803, 1090, 1050, 93, 587, 486, 1447, 430, 1891, 459, 1482, 1877, 349, 1637, 1401, 273]. ObjectStore [631]. oblivious [2016, 2150]. obliviously [2079]. observable [531]. observing [506]. occurrence [2162]. ODBMS [536]. odds [1991]. ODE [271, 350, 420]. OdeView [350, 420]. ODMG [833, 1484, 1408]. ODMG-93 [833]. ODMG-based [1484, 1408]. Odyssey [1504]. office [11, 174]. OFL [876]. OLAP [1158, 1128, 1100, 1645, 1563, 1620, 1737, 1813, 1490, 1121, 1735]. Old [648]. OLE [1027]. OLTP [1397, 1420, 1107]. Olympic [1277]. OminiSearch [1592]. On-line [1424, 1146, 1018, 1020, 1289, 1828, 1422, 821, 617]. One [79, 959, 159, 1720, 1904, 1861, 1249, 2112, 2113, 29]. one-dimensional [2113]. One-sided [159]. onion [1452]. OnLine [547, 1611, 1118, 1325, 1865, 1468, 1842, 1780, 939, 965, 1513, 1351, 1711, 1034, 1345, 1664, 1851]. only [448, 421, 1720, 119, 554, 1275, 368, 616]. Ontario [2252]. ontological [2232]. ontologies [2101, 1900, 1700]. Ontology [2231]. Ontology-based [2231]. OODB [487, 571, 689, 578, 745, 557, 896, 485]. OODBMS [977, 817, 987]. OODBs [916, 682]. opaque [1762]. Open [857, 745, 711, 632, 1317, 1122, 689, 1051, 1718, 840, 2075, 748, 1263]. open-source [1718]. Opening [2093]. operability [730]. operating [38, 743]. Operation [123, 520]. operational [1925]. Operations [585, 912, 610, 57, 1535, 439, 607, 1858, 239, 1422, 164, 1641, 1856]. Operator [1979, 1771, 1948, 2237]. operators [182]. optical [10, 238, 323]. OPTICS [1308]. Optimal [1361, 2108, 875, 759, 2131, 1521, 1543, 1769, 211, 882, 757, 1210, 2110, 1869, 2109, 1669, 1621, 2155, 2136, 239, 1306, 1201, 2128, 647, 2187, 1793, 756, 1130, 2066, 2089]. optimality [394, 920, 369, 2203]. optimisation [2090]. Optimism [550]. optimistic [866, 1330, 408, 446]. Optimization [1855, 646, 1187, 924, 1324, 783, 99, 909, 1328, 595, 2052, 234, 452, 1307, 1516, 43, 177, 334, 1022, 1875, 1662, 1066, 1452, 1665, 942, 1159, 1557, 1552, 1667, 629, 318, 1514, 2065, 1355, 126, 1584, 516, 1028, 1585, 1002, 2013, 688, 463, 1857, 96, 461, 1542, 1201, 449, 2192, 900, 1041, 180, 1895, 2134, 1558, 1512, 525, 373, 1896, 1786, 1440, 1831, 1060, 128, 1004, 1526, 1240, 442, 1000, 1622, 1341, 1826, 1235]. optimization-based [1557]. optimizations [150, 1183]. Optimized [1030, 996, 1081]. optimizer [1702, 689, 124, 1785, 1461, 603]. optimizers [1057, 1193]. Optimizing [1013, 811, 1560, 815, 1528, 2135, 194, 161, 1643, 686, 213, 404, 542]. OR-DBMS [1176]. OR-objects [273]. OR-predicates [1845]. OR-Sets [641]. oracle [1586, 1271, 1695, 1912, 827, 946, 1388, 1690, 830]. Oracle8 [1265, 1272]. Oracle8i [1469].

Order [1003, 1887, 966, 634, 650, 73, 2126, 1564, 2070, 324, 1228, 878, 2020, 688, 796, 2238, 290, 1345, 75, 1004, 997, 1000, 2140]. Order-of-magnitude [966]. ordered [1556, 1652, 618]. Ordering [650, 1308, 1874, 2012, 2069, 192, 190, 2011]. ORDPATHs [1927]. Oregon [2262, 2253]. ORES [848]. organization [1231, 1407, 749, 8, 200, 1103, 1413, 801]. organizations [460]. Organizing [384, 1483]. Oriented [800, 526, 577, 654, 653, 207, 514, 154, 1075, 1411, 1053, 167, 389, 187, 833, 974, 1177, 331, 433, 629, 537, 1925, 850, 1049, 168, 1190, 1484, 745, 183, 540, 733, 440, 426, 427, 173, 292, 736, 152, 337, 630, 346, 834, 831, 1291, 595, 262, 333, 466, 651, 943, 845, 836, 255, 1083, 327, 723, 739, 1052, 713, 1415, 395]. Orlando [2247]. Orthogonal [1584, 2165]. OSAM*.KBMS [739]. other [63, 956, 1955, 1249]. o’Ticks [1605]. out-of-order [2070]. outer [933]. outerjoin [1896]. Outerjoins [816, 1042, 401]. Outlier [1509]. outliers [1427, 1455]. output [1289, 2104, 363, 352]. outs [1069]. Overbound [453]. overheads [1831]. Overtaking [57]. Overview [1187, 601, 964, 2003, 1836, 969].

p [992]. P2P [1940]. P2P-DIET [1940]. PA [2282, 2261]. pack [2033]. package [2207]. packed [1994, 9]. packed-memory [1994]. page [817, 1604, 2006, 691, 787]. PageRank [2068]. pages [1782]. pair [1435]. Pan [2157]. Pan-private [2157]. panacea [504]. Panel [1834, 18, 42, 949, 1477, 31, 967, 956, 1475, 1717, 1608, 1610, 1256, 1476, 544, 1609, 25, 1255]. paper [760]. Papyrus [746]. paradigm [389, 201, 1014, 371]. Paradise [986]. paradox [2152]. Parallel [134, 826, 924, 91, 1912, 547, 950, 576, 610, 2173, 794, 519, 1186, 719, 890, 1020, 838, 964, 2236, 237, 1087, 844, 1079, 805, 372, 516, 1028, 1054, 77, 1136, 789, 1245, 1438, 697, 1280, 1007, 743, 1691, 166, 285, 1909, 656, 886, 1470, 371]. Parallelism [287, 1123, 586, 365, 520, 966, 444]. parallelization [302]. Parallelizing [475]. parameter [410]. parameterized [2239, 1518, 1671]. parameters [1786]. parametric [1550]. Parametricity [1016, 1017]. Paris [2291, 1008, 1009]. parity [1031]. Park [2244]. parsing [1849]. part [1722]. Partial [762, 1128, 594, 1664, 1532, 2155, 211, 460, 587]. Partial-Match [594]. Partial-sum [1128]. Partially [645, 1531]. particle [237]. Partition [818, 51, 189, 1045]. Partitioned [1139, 88, 256]. Partitioning [14, 1870, 344, 479, 413]. pass [1249]. past [1382, 94]. Pasta [422]. Pasta- [422]. paste [1113]. patchwork [1402]. patent [1298]. Path [1205, 1309, 1335, 756, 757, 1109, 2235, 2124, 1058, 1636, 1746, 779, 1639, 2193, 1525]. Pattern [509, 607, 988, 1574, 1669, 1674, 2046, 2174, 1944, 1493, 1668, 775, 1676]. Pattern-Based [509, 2046]. patterns [2171, 1197, 1551, 1355, 1419, 887, 1891, 1677]. PBIR [1601]. PC [943, 1507]. pebble [322]. pebbles [2037]. Peer [1971, 1943, 2044, 1940, 1620, 1781, 2089, 1885]. peer-to-peer [1943, 1620, 1781, 2089]. PeerDB [1814]. peering [1814]. Pegasus [709]. penalty [1031]. Pennsylvania [2283]. percentage [1915]. perception [1601]. perception-based [1601]. perfect [22, 2009]. Performance [377, 2151, 238, 551, 1024, 669, 460, 674, 611, 327, 556, 501, 127, 1026, 15, 1131, 565, 1387, 600, 253, 1517, 287, 1544, 1643, 323, 1708, 825, 254, 1110, 12, 32, 523, 169, 241, 615, 398, 545, 1076, 113, 412, 740, 1093, 1293, 1493, 285, 403, 560, 801, 820, 907, 1605, 1856]. periodicities [1851]. Persistence [736]. Persistent [529, 819, 614, 847, 989, 502, 1050, 430]. personal [1809, 1814]. personalizable [1407]. personalization [1466]. personalized [1282]. perspective [653, 1267, 2212, 95, 1657, 1991, 1380]. Perspectives [119]. perturbation [2104]. pessimistic [338]. Petabyte [1379]. PetaCube [1682]. pharmaceutical [732]. phase [1519, 205]. PhD [2300]. Philadelphia [2261, 2283, 2282]. Phoenix [1406]. phrase [1819]. physical [1870, 8, 1691]. physics [237, 1955]. pictorial [9]. Picture [1161]. pipelined [1874, 2012, 718, 789, 647, 166]. Pipelining [1514]. PIPES [1936]. PIPs [1686]. pivoting [475]. PIX [1819]. Placement [1194, 606, 196, 939, 814, 1438, 1979]. plan [1831, 1876]. planar [2222]. planning [484]. plans [1559, 783, 332, 1201, 1130, 1461]. PLASTIC [1831]. platform [1261, 1710, 1760, 1262, 1926]. platforms [840]. Plaza [2244]. PODS [2274, 2276, 2283, 2285, 2288, 2289, 2292, 2294, 2217, 2186, 2300, 2127, 2248, 2250, 2252, 2253, 2255, 2257, 2260, 2261, 2264, 2265, 2269, 2270, 2272, 2274, 2276, 2279, 2281, 2097]. PODS’08 [2295]. PODS’09 [2296]. PODS’10 [2297]. PODS’11 [2298]. PODS’12 [2299]. PODS’13 [2300]. Point [1005, 326, 266]. PointCast [1268]. pointer [403]. pointer-based [403]. points [1308, 407]. Policies [919, 1147, 818, 551, 1500, 1153]. policy [939, 1923]. Polyhedra [875]. polymorphic [272, 802, 1313]. Polynomial [905, 417, 2051, 370, 359, 432, 1996, 224, 2099, 298]. Polynomial-time [370]. polynomials [1890]. popularity [2006]. populated [1018]. population [147]. Port [1270]. portable [1246]. portals [1377]. Portland [2253, 2262]. positions [1447]. Positive [2126, 578]. Positivism [76]. Possibilities [182]. possible [100, 2219]. Power [995, 768, 356, 888, 2225, 146, 700, 2190, 1434, 2082, 1619, 2153, 184, 292, 21, 178, 322, 1707, 1629, 293, 355, 434, 153, 485, 2237].

PowerBookmarks [1407]. powered [1298]. Powers [877, 383]. Powerset [792, 226, 2237]. Practical [1968, 684, 814, 345, 288, 1443, 1560, 1970, 999, 391, 1920, 286]. practicality [1254, 920]. practically [1869]. Practice [810, 809, 751, 215, 2267]. Pragmatics [25]. precedence [116]. precedence-assignment [116]. Precision [670, 1562]. PREDATOR [1412, 1176]. predeclared [127]. Predicate [686, 397, 814]. Predicates [921, 1219, 933, 1539, 1672, 659, 1784, 458, 1789, 686, 1845, 815, 2071, 1048, 1902, 1545]. predicted [1893]. predicting [1076]. Prediction [1891, 1026]. Predictive [805, 571]. preemptible [645]. PREFER [1550]. preferences [1444]. prefetching [1855, 1544, 1643, 684]. prefixes [2049]. preliminary [402, 21, 775, 392]. presence [1959, 2116, 342, 1523, 1929, 1355, 1047, 2134, 1062]. presentation [1321, 1295]. presentations [1807]. preservation [2004]. preserving [1456, 1547, 1887, 2175, 1517, 1761, 1763, 1985, 1759, 2168, 1360, 73]. Presumptions [1354]. pretty [751]. Prevent [1320]. price [1883]. pricing [2199]. PRIMA [512]. prime [288]. primitive [294]. primitives [2225]. Principles [2248, 2250, 2252, 2253, 2255, 2257, 2260, 2261, 2264, 2265, 2269, 2270, 2272, 2274, 2276, 2279, 2281, 2283, 2285, 2288, 2289, 2292, 2293, 2294, 1194, 1992, 2300, 2299, 2295, 2298, 2296, 2297, 1156, 2267, 709, 1722]. printed [45]. priority [454, 2222, 1869]. Privacy [1456, 1731, 2058, 2007, 1969, 1547, 1968, 2190, 2169, 1761, 1763, 2078, 1985, 2136, 2168, 2137, 2191, 2135, 2009, 2104]. Privacy-enhancing [1969]. Privacy-preserving [1456, 1985, 2168]. private [1741, 2157]. proactive [1269]. Probabilistic [49, 1211, 1800, 2072, 2117, 2060, 2031, 2143, 1571, 2055, 2061, 2071, 2120, 2059]. probabilities [2122]. probability [2142]. probably [521]. Probe [1515]. probes [1986, 2013]. probing [1672]. Problem [1234, 1358, 1329, 1219, 1644, 881, 1067, 777, 1738, 893, 2000, 2128, 193, 605]. Problems [924, 992, 1317, 2012, 2207, 891, 1833, 2158, 2069, 1037, 357, 2156, 1255, 299]. Procedural [229, 268]. procedure [289, 399]. procedures [619, 241, 452, 148]. Proceedings [2248, 2250, 2252, 2253, 2255, 2257, 2260, 2261, 2264, 2265, 2269, 2270, 2272, 2274, 2276, 2279, 2281, 2299, 2298, 2251, 2244, 2245, 2266, 2249, 2300, 2246, 2241, 2242, 2297, 2259, 2275, 2283, 2285, 2288, 2290, 2289, 2291, 2292, 2293, 2294, 2271, 2284, 2262, 2258, 2282, 2287, 2263, 2280, 2277, 2295, 2254, 2296, 2278, 2247, 2286, 2273, 2268, 2256]. Process [195, 1356, 1556, 45, 1952, 2212]. processes [1598, 2154, 1556]. processibility [269]. Processing [1626, 1773, 1134, 241, 276, 588, 610, 2014, 1307, 396, 1441, 1799, 1939, 1938, 838, 2236, 679, 856, 791, 301, 1823, 1488, 1844, 417, 1276, 1794, 109, 1733, 805, 1024, 1911, 372, 1152, 946, 365, 513, 524, 1154, 1789, 688, 336, 1611, 107, 2168, 947, 273, 1873, 1245, 1845, 1687, 1279, 1936, 1300, 1754, 1357, 1895, 1892, 410, 631, 1225, 743, 592, 611, 1664, 258, 2030, 2035, 823, 656, 721, 1448, 842, 617, 1979, 980, 719, 2057, 392, 249, 371, 1699, 1489]. processor [718, 647, 1795]. processors [1043, 1858, 803, 1040]. producing [1749, 80]. product [394]. production [531, 259, 260, 395]. products [1156, 956, 717, 1072, 624, 1130, 1000]. program [36, 1957, 266, 370, 988]. Programmable [2028]. Programming [872, 133, 877, 637, 529, 490, 571, 602, 270, 389, 187, 458, 1067, 1161, 226, 871, 736, 337, 97, 178, 381, 333, 542, 1760, 232, 272, 383, 493, 378]. Programs [657, 362, 883, 1230, 63, 135, 543, 300, 472, 530, 774, 122, 197, 900, 624, 268, 376, 447, 385, 161, 357, 475, 470, 766, 442, 225, 264, 353, 249, 903]. progress [1907, 1906]. Progressive [1566, 1499, 1895, 1793]. progressively [1638]. project [851, 742, 1593, 1412, 844, 1161, 1085, 1083, 1169]. projected [1310, 1425]. projection [212]. Projections [54, 58, 1553, 2153, 80]. projective [1678]. Prolog [37, 44]. Proof [296]. Proof-tree [296]. propagation [135, 1316, 1642, 480, 2170, 2201]. Properties [92, 48, 1019, 517, 639, 87, 2118, 78, 1873, 2203]. proposal [1124]. proposed [1049]. propositional [596]. prospective [1096]. prospector [1079]. protect [477]. protection [1743]. protein [1298]. protocol [615, 570, 205, 446, 559, 1556]. Protocols [778, 902, 556, 584, 548, 377, 1316, 117, 443, 864, 587]. prototype [1600, 1164, 738, 744, 1597, 981, 853, 1172, 1606, 855, 1173, 985, 740, 1294, 1489]. prototyping [845]. Provenance [2167, 2169, 2034, 2166, 2088]. prover [228]. provide [1216]. Providence [2296]. provider [1654]. Providing [1286, 1039, 589, 1285, 830]. proximity [2010]. Proxy [1632, 1563]. Proxy-based [1632]. Proxy-server [1563]. pruning [2041, 1183]. PSA [1270]. pseudorandom [2007]. Ptool [847, 989]. public [1718, 1809, 1936, 1692]. publish [1269, 1524, 1880]. publish/subscribe [1269, 1524, 1880]. publishing [1882, 1805, 2039, 1985, 1797, 1295]. pull [1120]. pure [485]. purge [893]. purpose [16]. push [1120, 1267]. Pushing [609]. Puzzles [5]. pyramid [396, 1208]. pyramid-technique [1208].

QBI [978]. QBIC [976]. QC [1737].

QC-trees [1737]. QCluster [1804]. QUAD [30]. QUAD-trees [30]. quadratic [2214]. Quadtree [1690]. qualitative [579]. quality [1517, 1658, 2077, 1656, 1397, 1836]. quantification [1547]. quantified [2216]. quantifier [301]. quantile [1513]. quantiles [2019, 2020, 1249, 2110]. quantitative [994]. Quark [1273]. Quasi [2043]. Quasi-inverses [2043]. Quasilinear [392]. Quebec [2277]. queriable [1715, 1747]. Queries [521, 1232, 134, 1200, 1335, 1192, 1327, 1326, 930, 683, 1180, 1207, 1303, 909, 588, 870, 1340, 1219, 2033, 1339, 1217, 1223, 1025, 861, 1012, 519, 1189, 894, 928, 1202, 1109, 2102, 2216, 1460, 772, 1559, 1653, 2074, 105, 1574, 2167, 1655, 396, 1875, 2235, 2124, 2206, 483, 1104, 2126, 933, 1788, 1343, 301, 244, 578, 1338, 176, 1452, 1672, 646, 942, 1013, 1557, 1784, 1907, 1621, 1703, 2239, 1800, 1644, 1732, 2116, 1680, 1058, 629, 1220, 1286, 1541, 1999, 2056, 811, 1435, 91, 324, 2060, 2141, 2115, 1233, 1626, 269, 539, 1106, 375, 2148, 1522]. queries [2088, 372, 99, 1674, 1937, 1054, 1002, 1134, 1560, 2098, 149, 1746, 1129, 1518, 1222, 1458, 1789, 107, 241, 686, 1128, 1100, 276, 1550, 184, 330, 1900, 1092, 404, 1873, 2000, 1845, 1510, 1144, 1639, 2238, 815, 2014, 2071, 1137, 1436, 2173, 1328, 1957, 595, 1566, 2192, 885, 1040, 1090, 2120, 2135, 1906, 2048, 1624, 1707, 110, 85, 1111, 1944, 1892, 1649, 406, 453, 2052, 234, 75, 452, 1404, 1801, 795, 1793, 1088, 806, 1790, 1089, 887, 194, 1668, 1039, 1188, 2104, 879, 479, 2226, 64, 1516, 80, 295, 265, 1638, 1961, 1753, 481, 1599, 139, 248, 1633, 997]. queries [1240, 177, 334, 1671, 616, 2045, 252, 926, 1851, 1462, 2164, 1605, 2140, 890, 1631, 1428, 1568, 1791, 1235, 1876, 931]. Query [1003, 1022, 1311, 47, 995, 1318, 1209, 852, 583, 1159, 1187, 924, 666, 1552, 1324, 788, 419, 1190, 1207, 1355, 1911, 516, 1227, 1746, 1759, 897, 1910, 1059, 909, 96, 2073, 868, 610, 1221, 2199, 641, 351, 865, 1357, 1182, 1217, 1337, 572, 631, 674, 882, 1372, 401, 786, 1181, 639, 1341, 1830, 1702, 294, 1859, 1410, 1349, 65, 1094, 1958, 976, 1441, 1799, 1939, 2049, 1976, 2118, 2236, 1744, 1535, 925, 689, 421, 1855, 856, 1662, 1866, 1066, 1651, 1772, 2101, 1736, 261, 418, 1903, 1837, 3, 1359, 1665, 451, 785, 1298, 1488, 1451, 297, 417, 813]. query [1431, 1667, 318, 783, 734, 156, 106, 1514, 2142, 929, 1796, 2143, 805, 932, 326, 289, 2200, 873, 1024, 126, 1028, 876, 1241, 1585, 1002, 1757, 1226, 332, 365, 524, 489, 688, 1097, 336, 619, 1611, 1764, 1930, 463, 796, 482, 1949, 1785, 1857, 273, 461, 920, 1353, 626, 1508, 1203, 1687, 1201, 402, 186, 1964, 2132, 2205, 1920, 449, 655, 900, 1040, 1163, 1041, 2230, 1498, 180, 1795, 151, 1895, 978, 1362, 2134, 508, 1558, 1093, 410, 1512, 1225, 525, 373, 1862, 1664, 258, 1786, 1886, 2030, 479, 1440, 1590, 2035, 43, 1831, 823, 128, 1526, 617, 1979]. query [764, 1885, 1005, 1206, 1622, 1461, 603, 1705, 2229, 1570, 1848, 2113, 29, 1871, 191, 275, 1849, 1758, 2090, 2038, 1489]. Query-based [2199]. query-by-humming [1920]. Query-Optimal [882]. Query-preserving [1759]. query-reachability [655]. Querying [1726, 2180, 2228, 2171, 1941, 1720, 1331, 1323, 630, 400, 1531, 379, 1506, 1198, 61, 100, 1529, 1847, 339, 869, 852, 1824, 863, 1416, 1375, 1889, 607, 1443, 1014, 1940, 1834, 1162, 1138, 1629, 1650, 1827, 1755, 1652, 1745, 618]. Quest [851, 2154]. questions [2013]. Queue [912, 439, 205]. queues [367, 2222]. quick [1586]. Quickly [799, 1774]. QuickStore [820, 907]. quiescing [623]. quotes [1605]. QURSED [1650]. QXtract [1818].

R [1869, 407, 679, 1110, 753, 576, 1258, 1690, 884, 1307, 9, 1026]. R-tree [1869, 1869, 1690, 1026]. R-trees [679, 753, 576, 884, 1307, 9]. R/3 [1110, 1258]. RAID [198, 972]. RAID5 [694, 1031]. Rainbow [1826, 1699]. Random [1250, 1034, 1345, 416, 1553, 1978, 1865, 2006]. Randomized [2017, 2210, 404]. randsom [1347]. Range [1361, 909, 1100, 674, 1025, 1189, 2111, 1343, 1621, 1543, 1647, 1458, 2222, 2194, 1048, 1638, 2165, 2195, 2196, 2113, 1545]. range-sum [1638]. Rank [1857, 1779, 2066]. Rank-aware [1857]. ranked [1965, 1728, 1550, 1700]. ranks [2210]. Rapid [1000, 1464, 943, 845]. rare [1519]. RasDaMan [1292]. Rate [1622]. Rate-based [1622]. rates [1055]. ratio [939, 1991]. rational [608, 997]. Ray [2248]. Rdb [1271, 946, 604]. Rdb/VMS [604]. RDBMS [958, 840, 1735]. RDF [2029, 2230]. re [1201, 1408]. re-optimization [1201]. re-structuring [1408]. REACH [977]. reachability [655]. read [112, 840, 554, 1275]. read-mostly [840]. read-only [554, 1275]. read-write [112]. reads [527]. Real [958, 598, 1276, 671, 553, 556, 923, 548, 425, 613, 1110, 1152, 1458, 1833, 408, 1473, 1603, 795, 1589, 603, 1605, 646, 2000]. Real-Time [553, 556, 598, 671, 923, 548, 1152, 408, 1603, 795, 1589]. real-world [613, 1833]. Realization [538]. Realizing [597]. Rearranging [72]. reasoners [675]. Reasoning [2050, 812, 1315, 517, 1993, 210, 382, 2087, 462, 2232, 1319, 2046, 290]. rebalancing [465]. rebuild [694, 1468]. rebuilds [1031]. receivers [911]. reclamation [167]. reclustering [821]. recommendation [2207]. Reconciling [1575, 2044]. reconstructed [699]. Record [1373, 1271, 799]. Record-boundary [1373]. records [1534, 1277, 554]. recoverable [367]. Recovery [17, 778, 50, 52, 864, 692, 1356, 1023, 2063, 1685, 282, 568, 140, 570, 720, 1285, 667, 1098, 1586, 113, 621, 573, 1252, 1312, 1787, 763, 315, 366, 907]. Rectangle [2209]. Rectangle-efficient [2209]. rectangles [407, 884]. recursion [700, 320, 425, 156, 158, 311, 90, 447]. Recursions [522, 137, 159, 245, 298]. Recursive [1326, 680, 519, 894, 530, 774, 318, 91, 324, 1106, 149, 107, 681, 595, 373, 157, 481, 252, 247, 160]. recycling [1831]. Red [840]. redefinitions [913]. redo [1787]. reducible [209]. Reducing [778, 932, 1831]. reduction [1549, 1681, 1212, 1530, 1812]. Redundancy [321, 107, 2004, 533, 2196]. redundant [451, 1149, 198]. redux [2123]. Reed [1439]. Referential [1126, 280]. refinement [1289, 1573]. Reflective [637]. reformulation [1885]. regions [863]. registries [1686]. Regular [1109, 1983, 1335, 1207, 538, 2235, 1659, 2146, 1746, 2188]. regularization [2197]. Related [5, 2158]. Relation [129, 2, 79, 491, 234, 620]. Relational [1215, 2178, 995, 2027, 1187, 297, 564, 1228, 328, 1, 534, 637, 1712, 1829, 1094, 654, 1727, 962, 476, 1102, 1828, 280, 1112, 1156, 1805, 27, 1259, 106, 259, 1897, 231, 1759, 1913, 226, 853, 1078, 2227, 212, 236, 276, 482, 1245, 278, 438, 143, 736, 834, 963, 1385, 737, 1608, 189, 1640, 1963, 1815, 118, 85, 343, 171, 717, 406, 773, 221, 1068, 82, 121, 479, 83, 265, 1242, 148, 14, 1743, 248, 994, 1652, 1880, 155, 70, 153, 1313, 35, 699, 395, 392, 2237, 101, 191, 275, 1568]. Relations [880, 51, 129, 55, 620, 1600, 318, 967, 71, 250, 884, 157, 319]. Relationship [2104, 538, 77, 343, 444]. relationships [314, 1006, 2130]. Relative [780, 781, 2103, 755, 132, 1991]. Relaxed [1908, 1520, 1363, 125, 842, 1942, 765]. Relaxing [780, 60, 781]. release [2058]. relevance [2102, 2172, 1804]. relevant [956, 393]. Reliability [42, 930, 1227, 1226, 931]. Reliable [386, 62, 1919]. reload [499]. remote [592]. reorderable [401]. reordering [1002]. reorderings [933]. reorganization [1424, 1018]. replacement [1500, 691]. Replica [902, 497]. Replicated [722, 59, 1117, 1316, 759, 1449, 88, 374, 441, 199, 559, 587, 338]. Replication [1254, 1119, 1798, 827, 828, 953, 829, 1029, 1391, 754, 443, 327, 954, 2089, 561]. report [402, 831, 21, 392, 1835]. reporting [2194, 1650, 2195]. repositories [993, 1013, 2145, 1783]. Repository [1070, 1395, 1077]. Representability [349]. representable [807]. representation [100, 483, 1551, 491]. Representations [641]. Representative [47, 1025, 58]. Representing [1529, 180, 343]. requests [367]. Requirements [1923, 1655, 958, 613, 703]. rescue [1990]. Research [956, 891, 992, 346, 1482, 1718, 2182, 1011, 1067, 1475, 95, 511, 1037, 1072, 1954, 970, 725]. researchers [2154, 1609]. resident [113, 1305]. resilience [922]. resistant [1046]. resolution [303, 1566, 255]. Resolving [228]. resource [929, 1822, 1054, 1467, 1841]. resource-intensive [1822]. respect [102]. respond [1937]. response [759, 932, 1464]. responsive [1474]. restarts [37]. Restricted [58]. restriction [212]. restrictions [292]. restructuring [1014]. restructurings [492]. result [226, 920]. Results [1234, 680, 2036, 1903, 1749, 681, 480, 1620, 1664, 258, 2165, 127, 252, 775]. resumption [1423]. RETE [591, 440]. retention [551, 1270]. Rethinking [1952]. RETINA [1603]. retinal [1494]. Retrieval [323, 594, 30, 510, 1601, 759, 312, 1548, 469, 1203, 211, 1804, 1333, 563, 1163, 1499, 1367, 979, 649, 1700, 658, 808, 35]. retrieval-intensive [649]. Retrieving [750]. Reusing [1188, 612]. revealed [1830]. Revealing [1761]. Reverse [2095, 1436]. review [41, 1589]. reviewing [1952]. Revised [899]. revision [120, 596]. revisited [1527, 1053, 301, 2076, 1977, 2077, 99, 329, 2196]. Revisiting [1154]. revolution [1839]. Rewrite [2177, 525]. Rewriting [1104, 1335, 1327, 179, 1983, 1651, 1746, 1372, 1886, 1961, 1871]. Rhode [2296]. rich [751]. RIDL* [342]. right [2013, 453, 309, 1255]. right- [309]. right-linear [453]. Rights [1743]. rigorous [2191, 2105]. ring [2132]. Ripple [1351, 1660]. RISC [1043, 798]. risk [1991]. RoadRunner [1709]. Robbers [1537]. Robust [1780, 1895, 1406, 407, 1557]. ROLAP [1231]. Role [728, 1840, 1476, 1609]. roles [486]. ROLEX [1828]. roll [1430]. Rondo [1760]. ROSE [150]. RosettaNet [1686]. Rotating [1]. rough [1853]. routing [1738, 1498]. Rs [607]. Rs-operations [607]. Rufus [749]. Rule [1057, 528, 1033, 2179, 1114, 653, 642, 389, 1193, 1852, 91, 805, 126, 440, 165, 397, 1325, 197, 973, 1519, 1466, 525, 1242, 1181, 434, 371, 1570, 160]. rule-based [2179, 389, 1057, 1193, 126, 434]. rule-processing [371]. rulebases [612]. Rules [1234, 1030, 640, 673, 531, 1135, 1133, 2002, 1177, 1193, 2177, 996, 1081, 1136, 1464, 303, 1032, 143, 180, 1126, 1151, 1150, 309, 1183, 904, 1465, 1572, 1186, 179, 994, 399, 395, 444]. run [1253]. Running [2117, 384]. Runtime [1914, 2172].

S [2152]. S-3 [2152]. S3 [1175]. Safe

[1200, 919, 1050]. Safety

[683, 157, 295, 68, 155, 186, 208, 2048, 1581]. Sagas [141]. SAGRES [1483]. salesperson [1588]. SAMOS [981]. sample [1799, 2155, 1904]. sampled [2208]. samplers [2158]. samples [1460, 2040, 1865]. Sampling [755, 768, 865, 1681, 2108, 1250, 1347, 1864, 2131, 1046, 1241, 1910, 619, 1863, 1565, 345, 351, 1345, 416, 1426, 1461, 1570]. sampling-based [1241]. San [2257, 2269, 2275, 2274, 2290, 2289, 2258, 2249, 2240, 2268]. sandwiches [1306]. Sangam [1594]. Santa [2285, 2245, 2286]. SAP [1110, 1258]. Satisfaction [1221]. Satisfiability [2102, 2114, 1959, 655, 790]. say [1908]. saying [1127]. scalability [839]. Scalable [1136, 2015, 1817, 1451, 1604, 937, 1560, 847, 1920, 1439, 1660, 1892, 1266, 1880, 801].

Scalar [683, 1043, 1245]. Scale [591, 1400, 1941, 1822, 1279, 2197]. scaleable [1141]. scaleless [454]. Scaling [1342, 658]. scans [787]. Scheduling [924, 1771, 1054, 1034, 671, 729]. Schema [1962, 2145, 2147, 2096, 2063, 2144, 154, 1882, 1997, 2043, 2065, 2146, 1764, 1595, 1762, 2075, 2094, 1238, 1403, 2064, 1689, 1573]. schema-directed [1882]. schema-mapping [2065]. schema-mappings [2094]. Schemas [92, 800, 685, 87, 348, 1094, 1872, 1575, 2046]. schemata [212]. schematic [437]. schematically [1218]. Scheme [599, 60, 589, 1338, 996, 1863, 290, 2022, 24, 1808, 415, 1812, 338]. Schemes [1189, 360, 1165, 78, 102, 209, 1132, 222, 213, 1191, 83, 40, 101]. Science [1224]. sciences [725]. Scientific [613, 1783, 1616, 511, 775, 988]. Scottsdale [2299]. scrambling [1206]. Screen [575]. SDSS [1692]. Seamless [546]. Search [1361, 1549, 1681, 1712, 582, 2111, 454, 1101, 1087, 1175, 1414, 1945, 1278, 1779, 2148, 2177, 2082, 1728, 696, 1369, 2010, 1098, 1803, 803, 557, 465, 9, 1987, 1210, 979, 2196, 1700, 1170, 1851, 1485, 1670]. Searching [1213, 1342, 676, 757, 2204, 2223, 1592, 2079, 1124, 1212, 756, 1623]. Seattle [2281, 2280]. Second [2250, 2289, 688, 963]. Second-order [688]. Secondary [2112, 312]. Secretaries [558]. secrets [1946]. Secure [1889, 1152, 1246, 1919, 1824, 1923, 438, 374]. Securing [1593]. Security [1966, 162, 1954, 1889, 1840, 228, 1609, 1052]. seeded [793]. seek [1291]. Segment [455, 579]. segmentation [1534, 1850]. Selection [1329, 1596, 1799, 135, 1338, 1644, 818, 1738, 1904, 1558, 1635]. Selections [54, 609, 649, 1792, 1341]. Selectively [650]. Selectivity [1302, 1680, 1571, 670, 755, 1344, 1681, 1343, 1866, 785, 1434, 1047, 1336, 345, 1251, 1048]. Self [1332, 1301, 1483, 821, 1055, 259, 1438, 1437, 1266]. Self-adaptive [821, 1055]. self-controlling [259]. self-extensible [1437]. Self-Join [1301]. Self-organizing [1483]. Self-tuning [1332, 1438, 1266]. Semantic [48, 2233, 925, 1613, 900, 641, 2134, 1211, 255, 479, 1555, 1122, 1855, 1703, 1872, 34, 1833, 1096, 185, 449, 1737, 1813, 985, 118, 115, 128, 2180, 2192]. Semantics [154, 666, 51, 664, 199, 892, 648, 877, 687, 377, 543, 1788, 244, 331, 1999, 2115, 2025, 2232, 1781, 189, 651, 1256, 518, 268, 383, 225, 555, 246]. semantics-based [377]. SEMCOG [1163]. Semi [1055, 574, 1237, 1099, 1595]. semi-algebraic [1099]. Semi-automatic [1055, 1595]. semi-automatically [1237]. Semi-determinism [574]. semi-linearity [1099]. Semigroup [373]. Semint [985]. semirings [2034, 2205]. Semistructured [1108, 1205, 1340, 1339, 1237, 1591, 1370, 1510, 1238, 1372, 1650, 1077]. sense [428]. sensing [2220]. Sensitive [1194, 1957, 477, 2091]. sensitivity [1786]. sensor [1884, 1796, 1822, 1707, 1795, 731]. Sentinel [1177]. Separability [875]. separable [245]. separation [1140]. Sequence [823, 2202, 1516]. sequenced [607]. Sequences [867, 1606, 2125, 1137]. sequencing [1305]. Sequential [619, 527, 1677]. SEQUOIA [635]. SERF [1408]. SERFing [1486]. Serializability [780, 650, 781, 257]. Serializable [652]. Serialization [203, 364]. Series [1342, 990, 822, 1674, 1612, 1530, 1920, 1955, 1675, 1089, 1605]. Server [1596, 1284, 529, 1914, 495, 817, 1079, 568, 1024, 829, 1391, 1563, 1272, 1816, 830, 1252, 944, 763, 1393, 836, 747, 496, 1470, 1095, 632, 1692]. servers [1697, 938, 1610, 1721, 1010, 1266]. Service [1925, 1723, 1929, 1656, 1654, 1940, 1115]. Service-oriented [1925]. services [1932, 1614, 1822, 2001, 2084, 1710, 2227, 1725, 832, 1481, 1917, 1919, 1953, 1916, 1918, 1928]. services- [1481]. session [18, 1480, 1478, 1477, 31, 967, 956, 1479, 1475, 1608, 1610, 1476, 1609, 1481, 25, 1482]. Set [440, 784, 395, 911, 1744, 1999, 1773, 2021, 1548, 239, 436, 682, 1328, 223, 1963, 1754, 1404, 1902, 179]. Set-oriented [440, 395]. set-valued [1754]. Sets [98, 588, 641, 562, 100, 673, 1901, 63, 1099, 2017, 622, 330, 1436, 1803, 97, 178, 1748, 773, 1455, 376, 1716, 434, 225, 1676, 2023]. setting [1562]. settings [2042]. Seventh [2260, 2295]. Shape [2091]. shapes [469]. Share [1946]. Shared [354, 874, 1020, 844, 2069, 864, 2052, 1954, 285, 1556, 656, 658]. shared-nothing [844, 285, 658]. sharing [361, 1741, 1917, 817, 1483, 1389, 1407, 1050, 730, 802]. Sheraton [2245]. shiftable [1920]. Shifting [1342]. Shore [987]. Shoring [819]. should [953, 242]. shrinkage [1904]. Shrinking [1366, 1682]. shrunk [1376]. sided [159]. sideways [65]. SIG [2279]. SIG-SIGMOD-SIGART [2279]. SIGACT [2250, 2252, 2253, 2255, 2257, 2260, 2261, 2264, 2265, 2269, 2270, 2272, 2274, 2276, 2281, 2283, 2285, 2288, 2289, 2292, 2293, 2294, 2299, 2295, 2298, 2296, 2297]. SIGACT-SIGMOD [2250, 2252, 2253, 2255]. SIGACT-SIGMOD-SIGART [2257, 2260, 2261, 2264, 2269, 2270, 2272, 2274, 2276]. SIGACT-SIGMOD-SOGART [2265]. SIGART [2257, 2260, 2261, 2264, 2269, 2270, 2272, 2274, 2276, 2279, 2281, 2283, 2285, 2288, 2289, 2292, 2293, 2294, 2299, 2295, 2298, 2296, 2297]. SIGMOD [2250, 2252, 2253, 2255, 2257, 2260, 2261, 2264, 2265, 2269, 2270, 2272, 2275, 2274, 2276, 2279, 2281, 2283, 2285, 2288, 2290, 2289, 2291, 2292, 2293, 2294, 2271, 2284, 2262, 2266, 2282, 2287, 2263, 2280, 2300, 2277, 2299, 2295, 2298, 2254, 2296, 2278, 2286, 2273, 2268, 2297, 2256, 760, 2245, 2243, 2249, 2300, 2246, 2241, 2242, 2244, 2251]. SIGMOD-SIGACT-SIGART [2283, 2285, 2288, 2289, 2292, 2293, 2294, 2299, 2295, 2298, 2296, 2297]. SIGMOD/PODS [2300]. SIGMOD/PODS’13 [2300]. Signature [599, 12, 688, 682]. significant [2105]. SIM [628, 185]. SIMD [1641]. similar [1548, 469]. similarities [1121, 1851]. Similarity [2106, 1192, 870, 1225, 1089, 1247, 1368, 1549, 1087, 1175, 1564, 1220, 1286, 1779, 1674, 2082, 1612, 1900, 1212, 1369, 1803, 1367, 1987, 1902, 1676]. Similarity-Based [870, 1089, 1674]. Simple [134, 608, 2235, 421, 164, 64, 357, 2164]. simpler [2022]. simplification [143]. Simplifying [2146]. SIMS [750]. Simulatable [1967]. Simulated [776, 96]. simulation [1076]. Simultaneous [1235, 1116]. SINA [1892]. Singapore [1270]. Single [716, 302, 2146, 363]. single-input [363]. single-output [363]. single-type [2146]. site [1171, 1362, 1275, 1393, 1486, 1248]. sites [1577, 1709, 1850, 1487]. Sixteenth [2279]. Sixth [2257, 2294]. Size [899, 2098, 1140, 865, 894, 1532, 1732, 2056, 959, 1046, 619, 480, 920, 351, 265, 481, 649, 1107, 668, 1752]. sized [2151]. Sizes [1301, 470, 766, 353]. skeletons [1506]. sketches [2183, 2157, 2007, 2133]. Sketching [2056, 2220]. Skew [1646, 1046]. skew-resistant [1046]. Sky [1457, 1692]. skyline [2122, 2222, 1793]. skylines [2163]. Skyserver [1692]. SL/B5 [575]. Sleepers [752]. sliced [1511]. slides [1394, 1393]. sliding [1875, 2108, 2022, 2053]. Sloan [1457, 1692]. Small [55]. smart [1588, 1291]. smart-cache [1291]. SMP [840]. Snakes [1306]. Snowball [1600]. SOCQET [1813]. Soft [1744, 923, 1894, 2073]. soft-key [2073]. Software [1723, 173, 736, 743]. SOGART [2265]. Solomon [1439]. solution [2058, 1594, 1604, 881, 1560, 1029, 1279]. solutions [1970, 2042, 605]. Solving [299, 1833, 80]. Some [583, 517, 70, 12, 751, 1755, 775]. sometimes [1119]. SONAR [1081]. sort [1043, 1646, 798, 363]. sort-merge [1646]. sorter [718]. sorting [1131, 1805, 1978, 1253]. sound [1531, 191, 275]. SoundCompass [1920]. source [1718, 1984, 1873, 1628, 1848]. sources [1287, 750, 1359, 942, 1575, 1531, 1437, 1622, 2130]. sovereign [1917]. Space [2109, 1981, 2019, 895, 1513, 1988, 2208, 2194, 2162, 792, 442, 1101, 2239, 323, 939, 703, 2202, 1769, 239, 213, 1345, 1056, 1061, 725, 2203, 240]. Space- [2019]. Space-bounded [895]. Space-efficient [1513, 2208, 2194, 2162]. Space-optimal [2109]. Spaces [1192, 1425, 1143, 461, 410]. Spanners [2215, 2183]. SPARQL [2180, 2029, 2193]. sparse [1364, 1334, 2220]. sparsely [1018]. sparsely-populated [1018]. sparsification [2183]. SPARTAN [1555]. Spatial [454, 875, 583, 585, 1434, 909, 1602, 793, 1044, 674, 882, 1307, 1012, 1198, 1202, 1302, 1445, 679, 856, 791, 1435, 1898, 1099, 703, 168, 1222, 1172, 463, 1229, 1369, 1140, 1499, 1300, 321, 410, 806, 1045, 1141, 1934, 9, 1792, 2209, 1877, 1791, 1690]. spatial-merge [1045]. Spatially [1899]. Spatially-decaying [1899]. spatio [1890, 1680, 1892, 1671]. spatio-temporal [1890, 1680, 1892, 1671]. speaking [1588]. Special [2259, 2258, 703]. specialized [237]. Specific [610, 123, 1495]. Specification [122, 7, 144, 1928, 843, 540, 230]. specifications [1661, 297, 832, 349]. specifying [382, 688]. spectra [1854]. Spectral [1768]. speed [1422]. speeds [1843]. SPIFFI [937]. SPIRE [1499]. spirit [1386]. spots [445]. Spreadsheets [1735]. spruce [1608]. SQL [1596, 1914, 860, 253, 1386, 1127, 586, 1284, 1684, 1907, 2116, 1708, 1912, 1806, 967, 839, 1911, 99, 1002, 251, 1908, 1654, 1391, 2152, 1948, 1816, 1149, 969, 633, 1472, 1263, 1295, 1428]. SQL-99 [1386]. SQL3 [831]. SQLEM [1472]. SQLJ [1259].

Square [2250]. SR [1144]. SR-tree [1144]. Stable [239, 385, 267, 667]. stage [1448]. Staggered [769]. stagnant [1717]. Standard [1615, 833, 969, 221]. standards [504]. Stanford [1820, 1124]. star [1689]. starburst [336, 505, 525]. Start [1586]. STARTS [1124]. State [1072, 1954]. statements [36]. States [2286]. Static [2083, 1875, 1035, 2192, 1052, 272, 338]. Statistical [564, 1764, 236, 1673, 2190, 2003, 915, 2055, 2137, 477, 1121]. statistically [2105]. statistics [1662, 1866, 1864, 1241, 1543, 1345, 2208, 2157, 73]. StatiX [1648]. status [831]. step [791, 391, 740, 1210]. steps [2236, 1634]. STHoles [1540]. Stochastic [1115, 265, 431]. stock [233]. stop [251]. Storage [1301, 167, 1387, 600, 1084, 1051, 1576, 28, 972, 1034, 1231, 945, 1293, 611, 1786, 1413, 635]. store [854, 706, 1139, 430, 820]. STORED [1370]. StorHouse [1576]. Storing [1370, 1652]. StorM [1500]. strange [63, 967]. strategies [752, 1621, 523, 524, 169, 1306, 595, 369]. strategy [302, 461, 1188]. Stratification [906, 1744, 266, 376]. stratified [120, 380]. Stream [1820, 1789, 1821, 1617, 1771, 1767, 1874, 1939, 1938, 1860, 1765, 1810, 2020, 1841, 2181, 1979, 1820]. streaming [1843, 1674, 1790, 1625, 1622]. StreaMon [1939]. streams [923, 1802, 1655, 1875, 1976, 1855, 2018, 1933, 2056, 1981, 2019, 2070, 2131, 1733, 1626, 1956, 1773, 1980, 1720, 1751, 1505, 1717, 1988, 2055, 2158, 1936, 2211, 1624, 2208, 2133, 1801, 2068, 2209, 1842, 1876]. Street [1155]. strengthen [544]. Strict [690, 339]. strikes [505]. String [1535, 2016, 2079]. Strings [812, 2202, 1453]. STRIP [1114]. STRIPES [1893]. Striping [1194, 769]. strong [2109]. Structural [2025, 449, 906, 1174, 274, 1750, 1973, 518, 1485, 1881]. Structure [569, 1847, 1901, 1308, 396, 331, 382, 1658, 2021, 696, 1144, 1905, 1737, 311, 1566, 1850, 1439, 581, 1868]. structure-based [1868]. Structured [1205, 1182, 1237, 1726, 1782, 2124, 1075, 1534, 1065, 1091, 1750, 813, 2014, 1748, 1629, 1673, 587, 485]. Structures [1049, 125, 674, 1102, 1260, 600, 1532, 2060, 1446, 2047, 579, 2238, 1565, 333, 343, 1218, 147, 70, 1745]. structuring [1408]. STRUDEL [1171, 1248]. studies [1157, 1155]. Study [356, 888, 2231, 825, 579, 158, 2004, 638, 884, 1493, 233, 907]. Studying [784, 15]. sub [2184, 1201, 2208]. sub-linear [2184]. sub-optimal [1201]. sub-sampled [2208]. subgoals [192, 190, 2011]. subgraphs [2183, 2213]. submodular [2198]. subpopulation [2056]. subpopulation-size [2056]. subqueries [1584]. subquery [1811]. subscribe [1269, 1524, 1880]. Subsequence [1358, 822, 1920, 1675, 1842]. subset [1986, 149]. subsets [1748]. subspace [1681, 1199]. Substring [1344, 1554, 2162]. Subsystem [50]. Subtyping [487, 916]. suite [1580]. SuLQ [1968]. sum [1128, 1638, 2198]. summaries [2185, 1513]. summarization [1813, 1879]. summary [1750, 1241, 1543, 1867, 1737, 1463, 193, 1105, 8, 2164, 1428]. sums [1986]. SUPER [742, 1043, 1940]. super-peer [1940]. SuperSQL [1295]. Support [1942, 550, 653, 838, 958, 1260, 1084, 642, 908, 1158, 1805, 1594, 839, 1142, 1867, 173, 414, 647, 171, 221, 1086, 1693, 743, 1474, 1039, 604, 486, 1240, 459]. Supporting [1147, 174, 610, 562, 761, 1672, 156, 512, 1064, 39, 1137, 311, 587, 148, 954, 739, 1568]. Suppressing [477]. Survey [583, 534, 858, 1457]. suspects [1955]. Sybase [829, 632]. symmetric [827]. Symposium [2248, 2250, 2252, 2253, 2255, 2257, 2260, 2261, 2264, 2265, 2269, 2270, 2272, 2274, 2276, 2279, 2281, 2283, 2285, 2288, 2289, 2292, 2293, 2294, 2300, 2295, 2297, 2299, 2298, 2296]. synchronisation [216]. Synchronization [687, 341, 1628, 16]. synchronized [866].

Synchronizing [1429]. synopses

[1349, 1538, 1683, 1673]. Syntax [928, 1849]. synthesis [1651]. synthesized [2084]. synthetic [799]. System [550, 601, 610, 2, 532, 6, 562, 1821, 1410, 923, 1114, 1600, 1829, 701, 1287, 120, 976, 1555, 1938, 514, 214, 1292, 237, 1070, 761, 1051, 253, 386, 856, 1411, 710, 1597, 974, 1298, 1488, 1451, 1703, 1844, 473, 1824, 614, 975, 734, 1160, 661, 1080, 506, 990, 2148, 1171, 1248, 937, 1081, 1433, 183, 733, 1168, 512, 1496, 34, 1809, 365, 1222, 1913, 984, 1166, 853, 1172, 1606, 150, 1550, 1494, 1949, 1082, 973, 1353, 185, 243, 570, 1473, 511, 991, 983, 152, 1258, 423, 834, 347, 1920, 1365, 846, 1603, 113, 985, 1163, 1407, 304, 1281, 1951, 845, 945]. system [773, 631, 1501, 38, 1827, 1293, 131, 435, 488, 743, 982, 741, 665, 1502, 133, 1420, 1495, 1437, 544, 732, 128, 979, 1589, 739, 1652, 986, 848, 1880, 1724, 713, 1415, 731, 29, 1294, 1699, 1409, 1273, 419, 715]. Systems [2248, 2250, 2252, 2253, 2255, 2257, 2260, 2261, 2264, 2265, 2269, 2270, 2272, 2274, 2276, 2279, 2281, 2283, 2285, 2288, 2289, 2292, 2293, 2294, 1187, 1180, 546, 2300, 2295, 517, 690, 553, 2297, 1990, 2083, 37, 1022, 1731, 195, 1123, 654, 826, 1617, 1771, 207, 11, 1175, 769, 42, 1685, 1084, 857, 2218, 339, 738, 167, 1156, 1552, 1943, 1984, 1265, 1264, 1011, 340, 1524, 1038, 2177, 1152, 440, 950, 622, 1154, 1992, 397, 751, 2227, 111, 1034, 1838, 328, 1781, 41, 1285, 346, 2267, 39, 441, 2299, 422, 1608, 1438, 1561, 2298, 1498, 573, 1252, 730, 567, 864, 162, 1062]. systems [1256, 724, 1578, 704, 2296, 415, 743, 592, 497, 255, 1954, 1516, 1242, 305, 142, 148, 260, 40, 927, 306, 399, 68, 1885, 987, 725, 1880, 658, 445, 395, 590, 1095, 705, 765, 1568].

table [2058, 494, 477, 1698]. table-level [494]. Tables [1014, 1555, 277, 2152, 1463, 1850, 623, 1105, 8, 994, 1428]. tagging [1805]. tail [447]. Tailored [657]. tailoring [1074]. taking [1693]. talk [2093, 1394, 305, 306, 279, 1393]. talk-slides [1394, 1393]. TAM [1409]. tape [1305]. tape-resident [1305]. Tapes [706]. target [2042, 1783]. Task [729]. TAX [1900]. taxis [115]. taxonomy [26]. teach [708]. teaching [1724]. technical [1924, 837]. Technique [776, 757, 454, 1208, 1452, 32, 469, 8, 756]. Techniques [784, 1020, 1066, 942, 282, 1898, 13, 1073, 1059, 626, 455, 199, 1646, 1041, 1345, 1357, 1454, 410, 373, 1886, 1722, 1004, 334, 560]. technologies [1157, 717, 1721]. technology [1289, 1075, 10, 827, 1267, 972, 391, 1385, 1141, 1083, 970, 955, 1830]. Telecom [1398]. telecommunication [720]. telecommunications [760, 218]. telegraph [1599]. TelegraphCQ [1823]. Template [1166, 1701]. Template-based [1166, 1701]. templates [887, 293]. Temporal [1003, 188, 672, 1756, 424, 862, 1150, 927, 677, 483, 925, 1006, 2176, 1890, 1680, 417, 71, 231, 597, 20, 1892, 170, 1526, 1671, 848, 1005, 1497, 1545]. Tennessee [2264]. tension [228]. Tenth [2265]. terabyte [1457]. teradata [838, 1007]. Term [599, 563]. terminals [1809]. Termination [1234, 470, 531, 2094]. terminological [13]. terminology [221]. terms [179]. TerraServer [1445]. tertiary [706, 1034]. Test [2186, 85, 2127, 2097, 2217, 1493]. Test-of-Time [2186, 2217]. Test-of-Time-Award [2127, 2097]. testbed [974, 258]. Testing [1006, 528, 2184, 208, 224, 288]. tests [2114, 2234]. Texas [2243, 2260, 2284, 2254]. TeXQuery [1945]. TeXQuery-based [1945]. Text [809, 810, 1237, 1600, 1818, 1726, 1847, 2175, 1534, 1945, 1374, 942, 863, 777, 1904, 347, 808, 1536]. text-preserving [2175]. textual [1220, 1286]. texture [33]. their [1594, 779, 296]. Theorem [1189, 228]. theorems [296]. theoretic [1901, 1727, 4, 1537, 117, 2004, 86, 390]. theoretical [1096, 1657, 66]. theoretician [2100]. theoreticians [279]. theories [380]. Theory [176, 810, 918, 2181, 648, 692, 2065, 809, 751, 2212, 917, 2100, 1787, 806, 1360, 1776, 544, 142, 94, 279, 390]. there [1377]. these [1254]. they’re [841]. Things [953]. think [2221]. Thinksheet [1074]. Third [2252, 702, 969]. third-generation [969]. Thirteenth [2272]. thirtieth [2298]. Thor [1050]. thought [521]. Thoughts [95, 751]. Three [378, 600]. Three-valued [378]. Throughput [1196, 111]. Throughput-Competitive [1196]. tickets [1809]. Tie [518]. Tie-breaking [518]. tier [1817, 1816, 1696, 1694]. Tight [902, 2158, 1191, 1925, 2017, 443]. TIMBER [1827]. Time [905, 11, 1342, 2070, 1612, 2186, 553, 556, 2127, 1671, 923, 432, 548, 2097, 2217, 2085, 102, 759, 1644, 417, 598, 1765, 2019, 990, 932, 822, 2051, 1674, 1152, 771, 1996, 84, 165, 2188, 408, 222, 671, 276, 239, 1785, 2184, 1530, 1137, 1920, 1603, 1955, 224, 171, 1675, 795, 2099, 1089, 1061, 1987, 298, 370, 1589, 26, 286, 359, 1605, 29, 240]. time-based [771]. Time-decaying [2070, 1765]. time-efficient [2019]. Time-parameterized [1671]. Time-Series [1342, 822, 1920, 1675]. time-space [240]. Timestamp [40, 127]. TimesTen [1694]. TIP [1497]. Token [411]. tolerance [587]. Tolerant [59, 938, 1010, 559, 1599, 1909, 954]. Tolerating [590]. too [2058, 706, 1442]. Tool [53, 1237, 1596, 1165, 342, 1491, 1496, 277, 688, 2025, 1595, 1832, 1290, 1076, 1487, 1074]. toolbox [843, 2100]. toolkit [1400, 1142, 1282, 1498, 1713]. Tools [429, 1953, 356, 888, 1901, 18, 539]. top [1730, 1672, 2010, 2149, 2195, 291]. top- [2010, 2195]. top-down [2149, 291]. top-k [1730, 1672]. Topological [884, 1012, 1198]. Toronto [2242]. TOSS [1900]. Totality [906, 518]. ToXgene [1701]. TP} [702, 1610]. TPC [1110, 638, 966, 1693]. TPC-C [638, 966]. TPC-D [1110]. TPC-DS [1693]. track [1807]. Tracking [1301, 1778, 2210, 1401, 2110]. tractability [2153, 2094, 2145, 2229]. Tractable [1303, 2005, 489, 643, 897, 2101, 2032]. tradeoffs [495, 238, 1024]. trades [1605]. trading [165, 233, 1061]. traditional [901]. traffic [1471, 1832, 1603, 1275]. Trait [473]. trajectories [1770, 1890, 1893]. transacted [1919]. Transaction [37, 183, 1867, 216, 114, 541, 1023, 584, 386, 382, 109, 566, 1152, 946, 513, 251, 140, 671, 947, 1279, 39, 711, 199, 446, 171, 567, 858, 592, 230, 611, 721, 842, 163, 68]. Transactional [92, 1350, 1269, 1356, 87, 1467, 1556, 1409]. Transactions [46, 780, 1019, 1352, 1916, 650, 448, 781, 761, 614, 384, 1396, 112, 364, 1975, 317, 67, 140, 143, 554, 748, 608, 40, 411, 1392, 765]. Transducers [1215, 2178, 2037, 2039, 1982, 867]. Transducing [2125]. Transform [2220]. transformation [2037, 622, 542, 296, 957]. transformation-based [542]. transformations [387, 388, 2175, 1931, 1193, 2200, 593, 2149, 1797, 2189, 370, 1825]. Transforming [362, 883]. transforms [1758]. transient [554, 1240]. transient-views [1240]. transition [307, 837]. Transitive [792, 314, 825, 2234, 158, 467, 248, 352, 2026, 2086]. Translating [54]. Translation [683, 144, 1806, 984, 155]. Transparent [1816, 1252]. transport [1809, 1949]. TRANSSOFT [743]. transversals [1125]. trap [78]. traversal [311]. Tree [2092, 129, 1228, 1846, 1574, 1869, 2067, 1688, 2117, 932, 2037, 1751, 1330, 2146, 2153, 626, 398, 2174, 2014, 1690, 803, 1566, 1982, 1944, 1629, 74, 296, 1668, 80, 1623, 1026, 1745, 2086, 1869, 407, 881, 1144, 501]. tree-based [626]. tree-edit [1751]. tree-structured [2014, 1629]. tree-walking [2037]. Tree-width [2092]. Trees [1643, 1363, 57, 2054, 30, 2016, 2151, 123, 1993, 679, 2036, 1663, 753, 317, 1475, 461, 310, 576, 1528, 1098, 1737, 1520, 793, 557, 465, 2112, 884, 1307, 1459, 401, 9, 503, 1462, 30, 502, 1018]. treewidth [2005, 2047, 2098]. trend [1716]. TREX [1825]. Trial [2230]. triangles [2018]. trichotomy [2235]. trie [898, 2202]. trigger [1149]. triggers [627, 307, 384]. triumph [84]. troubleshooting [1722]. truth [1462]. TSIMMIS [984, 1166, 1288]. Tucson [2279, 2278]. tumble [1548]. Tuning [606, 1332, 1438, 740, 1155, 1722, 1266, 715]. Tuple [1021]. Tuple-Generating [1021]. tuples [706, 451, 2139]. Turbo [1421]. Turbo-charging [1421]. turning [1284]. Tutorial [767, 972, 810, 992, 872, 918, 1721, 782, 1257, 2111, 1260, 1480, 1478, 949, 1156, 1259, 2029, 1479, 1720, 1985, 1611, 2138, 871, 1258, 2189, 1481, 1482]. Twelfth [2270]. Twentieth [2285]. Twenty [2288, 2289, 2292, 2293, 2294, 2295, 2297, 2296]. twenty-eighth [2296]. Twenty-Fifth [2293]. Twenty-First [2288]. Twenty-Fourth [2292]. Twenty-Ninth [2297]. Twenty-Second [2289]. Twenty-Seventh [2295]. Twenty-Sixth [2294]. twig [1669, 1845]. Twin [213]. Two [1020, 1361, 1993, 1554, 517, 27, 2012, 996, 2082, 1519, 205, 477, 2165]. Two-Dimensional [1361, 1554, 996, 477, 2165]. two-phase [1519]. Two-variable [1993]. Type [339, 1309, 1339, 1313, 2090, 1433, 2146, 1982, 272]. typechecking [1527]. Typed [1217, 796]. Types [7, 782, 639, 1775, 174, 219, 184, 2134, 1176]. Typing [1755].

U [1035]. U-Datalog [1035]. Ubiquitous [1266]. UDAFs [1843]. UDB [1264]. UDDI [1686]. ugly [1479]. UI [1921]. ultra [1478]. unaggregated [2056]. unambiguity [1983]. unbiased [2155]. uncertain [2107, 2122, 2080, 2121, 2140]. uncertainty [2204, 2223, 462, 1533]. Uncoupling [465]. Undecidability [1234, 252]. Understanding [2148, 2129, 1849, 1573]. Unfounded [225]. Unified [692, 1681, 607, 1153, 834, 698]. unifies [1940]. uniform [449, 265]. Uniformity [753]. Unifying [82, 316]. UniSQL [834]. UniSQL/X [834]. United [2286]. Universal [1910, 2, 790, 60, 79, 1261, 1272, 215]. Universally [2136]. University [2246]. UNIX [838, 1389]. unknown [1891]. unnesting [1190]. unnormalized [153]. unstructured [1066, 2089]. Untyped [330]. updatable [2027]. Update [1316, 144, 229, 940, 923, 2225, 911, 204, 2050, 908, 1773, 1980, 953, 1366, 1282, 1031, 258, 35, 246]. Updates [1228, 54, 1363, 360, 2198, 1049, 596, 375, 1911, 19, 696, 1520, 281, 623, 232, 1103, 808]. Updating [283, 476, 1567, 122, 1734, 465, 86]. Upgrading [1320]. Upper [902, 443]. Upsizing [944]. USA [2293, 2282, 2297, 2287, 2280, 2300, 2299, 2278]. usage [1274]. USD [511]. Use [962, 1864, 460]. Useful [6]. User [1291, 420, 1416, 929, 427, 95, 35]. user-centered [1416]. user-friendly [420]. User-oriented [1291]. users [1272]. Using [1232, 448, 548, 654, 565, 908, 1348, 1327, 990, 753, 1803, 1850, 1218, 1307, 594, 656, 866, 1559, 1653, 1917, 1104, 367, 1075, 1855, 679, 1239, 785, 10, 1220, 1286, 2155, 1806, 1233, 1106, 13, 313, 1434, 996, 1751, 1571, 2013, 1560, 1809, 1858, 1746, 1769, 1128, 1064, 1904, 1841, 1397, 570, 1804, 1690, 1565, 595, 1336, 885, 1040, 1439, 793, 1422, 749, 228, 1558, 625, 726, 1472, 1225, 182, 887, 1506, 2129, 73, 9, 14, 470, 1633, 1240, 1652, 1700, 1880, 1334, 729, 1341, 1428, 1826, 1641, 1811]. Utility [1324, 1244, 1922, 2137]. Utilization [310, 323]. Utilizing [588, 538, 328].

V [747]. valid [543]. Validating [1625]. validity [1883]. value [1062]. valued [1754, 413, 378]. Values [1230, 1527, 2139, 303, 1762, 1964, 1562, 485, 191]. Vancouver [2295]. Variable [1001, 1993, 1328, 926]. variables [2051]. Variance [2053, 1767]. variant [1093]. variants [2032]. Various [664]. vectors [1803]. vehicle [703]. vendor [633]. Verifiable [1019]. Verification [2218, 2001, 2227, 7, 644, 20]. verifying [2216, 1661]. VERSANT [954]. Version [471, 1405, 341, 240]. versioning [554]. Versions [1395]. Versus [1003, 49, 323, 428]. Vertical [344, 1915, 1870, 1421]. vertically [1670]. very [1524, 111, 1865, 2197, 623, 1470, 1015]. Via [2, 2008, 2147, 2218, 684, 1228, 1779, 2220, 1519, 2157, 2007, 1404, 1198]. video [939, 937, 1561, 1454, 747]. video-on-demand [937, 1561]. View [1736, 889, 1329, 54, 934, 1148, 1596, 1651, 1644, 1063, 1147, 126, 1738, 169, 1365, 1640, 1734, 1558, 1146, 1061, 1083, 1403, 1430, 202, 716]. View-based [1736]. View-Selection [1329, 1644]. viewed [120]. Views [1232, 1299, 1327, 54, 1961, 1236, 538, 474, 1559, 1653, 476, 1104, 79, 2027, 1642, 1659, 1644, 1732, 27, 2169, 1106, 1889, 1560, 1867, 1746, 935, 663, 913, 2170, 1231, 885, 1797, 936, 2033, 802, 399, 1240, 1826]. virtual [1064, 303, 656, 1589, 413, 668]. VisDB [983]. ViST [1745]. Visual [1702, 419, 580, 577, 572, 425, 537, 1714, 1832, 1163, 1162, 1138, 1716, 698]. visualization [1084, 734, 1160, 901, 996, 1071]. visualizing [983, 748]. VLDB [946]. VLSI [473, 723]. VMS [604]. vocabulary [1359]. VODAK [748]. Volcano [365]. voting [138]. voxelized [1803]. VQBD [1591]. vs [432, 905, 76, 461, 2004, 997, 1005]. vs. [751]. walking [2037, 1629, 2086]. walkthrough [1589]. Wall [1155]. WALRUS [1367]. ware [1522]. warehouse [1445, 1471, 1607, 840, 1306, 1397, 1173, 1366, 1423, 1105, 1146, 957]. warehouses [1148, 1492, 1365]. warehousing [1576, 1158, 828, 1069, 1398, 934]. Warping [1758, 1987]. WASA2 [1415]. Washington [2270, 2271, 2280, 2241, 2256, 2281]. Waterloo [2252]. watermarking [1829, 1759]. Wave [1145]. Wave-indices [1145]. Wavelet [1683, 1251, 2202]. Wavelet-based [1251]. wavelets [1766, 1334]. way [1520, 1462]. ways [63]. Weak [71, 108, 283, 360, 2196]. wealth [2154]. Web [1695, 2179, 1697, 1577, 1709, 2093, 1632, 2084, 1710, 1619, 1764, 1515, 1583, 1633, 1705, 1917, 1695, 1782, 2180, 1417, 1928, 1592, 1296, 1488, 1449, 1276, 1286, 1297, 1604, 1375, 2001, 1373, 1171, 1248, 1937, 1377, 1443, 1919, 1930, 1953, 1157, 1291, 1850, 2192, 1407, 1916, 1399, 1037, 1281, 1111, 1487, 1918, 1610, 2006, 1275, 1068, 1569, 1486, 1383, 1170, 1504, 2130, 955, 1848, 1849]. Web-based [1281]. Web-CAM [1937]. Web-site [1248]. WebLogic [1932, 1926]. websites [1474, 1506]. WebSources [1631]. websphere [1580]. WebView [1450]. week [2249]. weight [989]. Well [666, 2232, 268, 68, 225, 555, 1960]. well-definedness [1960]. Well-Founded [666, 2232, 225, 555]. what-if [1244]. Where [1224, 1477, 2181]. while [1761]. WHIPS [1173]. WI [2287]. Wide [1417, 1037, 1068, 1170, 955, 1632, 1796, 716, 1178]. width [2092, 1537]. wild [1583]. wildcards [1047]. Will [751, 1609]. Window [1358, 882, 1025, 396, 1811, 1366]. Window-Accumulated [1358]. windows [1875, 1767, 2108, 2022, 1675, 2053, 947]. WinMagic [1811]. Winnowing [1739]. wireless [1707]. Wisconsin [2288]. within [743]. without [1332, 565, 1220, 1419, 257, 223, 623, 1256, 581]. witness [908, 1774]. wizard [1284]. won’t [949, 1256]. words [2054]. work [959, 720, 1466, 1256]. workaholics [752]. workbench [1178]. Workflow [1352, 1656, 970, 1496, 1285, 1072, 1415, 1178]. Workflows [1185, 2221, 1928, 2002]. working [1255]. workload [1540]. workload-aware [1540]. workloads [1684, 1107]. workshop [41]. workspaces [1395]. workstations [1131, 93]. World [558, 580, 958, 613, 1632, 1110, 77, 1833, 884, 1417, 1037, 1068, 1170, 955]. worlds [100, 2075]. WORM [323]. Worst [2111, 151, 2187, 1869]. Worst-case [2111, 151, 2187, 1869]. wrapper [1399]. wrappers [1166]. wrapping [1431]. Write [368, 448, 112, 570, 625]. write-ahead [570, 625]. Write-only [368, 448]. writes [587]. writesets [127]. WSDL [1929]. WSQ [1443]. WSQ/DSQ [1443].

X [1593, 834]. XBase [1715]. XCache [1703]. XDM [532]. XJML [1399]. XJML-based [1399]. XL [1710]. Xmas [1293]. XmdvTool [1716]. XMill [1432]. XML [1966, 1299, 1529, 1798, 1859, 1983, 2083, 2119, 1726, 1527, 2096, 1819, 1847, 2175, 1630, 1661, 1727, 1958, 1976, 1701, 2118, 1418, 1593, 1828, 1882, 1993, 2176, 1860, 1945, 1669, 1477, 1926, 2050, 1837, 1805, 1703, 1824, 1431, 1636, 1663, 2072, 2117, 2141, 2037, 1523, 1889, 2039, 1522, 2088, 1648, 1433, 1751, 2146, 1759, 1666, 1728, 1738, 1861, 1845, 2046, 1608, 2149, 1797, 1432, 1982, 2145, 1777, 1747, 1525, 1569, 1927, 1827, 1673, 1862, 1590, 1625, 1755, 2059, 1567, 1652, 1700, 1880, 1504, 1745, 1881, 1871, 1826, 1825]. XML-based [1418]. XPath [1959, 2085, 1844, 2114, 2234, 2051, 1757, 1634, 1789, 2025, 1627, 2099, 1790, 2026, 2038, 2086]. XPRESS [1747]. XPRS [520]. XPS [965]. XQuery [1926, 1837, 1719, 1806, 1929, 1846, 1699, 1826, 1964, 1960]. XRANK [1728]. XSB [849, 824]. XSeq [1944]. XSL [1797]. XTRACT [1433]. XXL [1700].

year [837]. Yoo [1929]. Yoo-Hoo [1929]. York [2300].

zone [1115]. Zoo [1179].

References

Lin:1976:DRA

[1] C. S. Lin, D. C. P. Smith, and J. M. Smith. The design of a rotating associative array memory for a relational database management application. ACM SIGACT-SIGMOD Symposium on Principles of Database Systems, 1(1):??, March 1976. Also published in/as: Proceedings of the First Conference on Very Large Databases, Morgan Kaufman pubs. (Los Altos CA), Kerr (ed.), 1975, pp. 453–455.

Kuck:1982:URD

[2] S. M. Kuck and Y. Sagiv. A universal relation database system implemented via the network model. In ACM [2248], page ?? ISBN 0-89791-070-2. LCCN QA76.9.D3 A33 1982.

Chandra:1983:HCF

[3] A. K. Chandra and D. Harel. Horn clauses and the fixpoint query hierarchy. In ACM [2250], page ?? ISBN 0-89791-097-4. LCCN QA76.9.D3 A15 1983. ACM order number 475830.

Cosmadakis:1984:FID

[4] S. S. Cosmadakis and P. C. Kanellakis. Functional and inclusion dependencies: a graph-theoretic approach. In ACM [2252], page ?? ISBN 0-89791-128-8.

LCCN QA76.9.D3 A15 1984. ACM order number 475840.

Lehman:1984:KCK

[5] D. Lehman. Knowledge, common knowledge, and related puzzles. In ACM [2252], page ?? ISBN 0-89791-128-8. LCCN QA76.9.D3 A15 1984. ACM order number 475840.

Pitelli:1984:BAU

[6] F. Pitelli, H. García-Molina, and S. Davidson. Is Byzantine agreement useful in a distributed database system. In ACM [2252], page ?? ISBN 0-89791-128-8. LCCN QA76.9.D3 A15 1984. Also published in/as: to appear in ACM Transactions on Database Systems 1985.

Stemple:1984:SVA

[7] D. Stemple and T. Sheard. Specification and verification of abstract database types. In ACM [2252], page ?? ISBN 0-89791-128-8. LCCN QA76.9.D3 A15 1984. ACM order number 475840.

Ozsoyoglu:1985:LPO

[8] Gultekin Ozsoyoglu, Z. Meral Ozsoyoglu, and Francisco Mata. A language and a physical organization technique for summary tables. In Navathe [2254], pages 3–16. ISBN 0-89791-160-

1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p3-ozsoyoglu/p3-ozsoyoglu.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/318898/p3-ozsoyoglu/

.

Roussopoulos:1985:DSS

[9] Nick Roussopoulos and Daniel Leifker. Direct spatial search on pictorial databases using packed R-trees. In Navathe [2254], pages 17–31. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1 .A87. URL http://www.

acm.org/pubs/articles/proceedings/
mod/318898/p17-roussopoulos/p17-
roussopoulos.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/318898/p17-roussopoulos/.

Christodoulakis:1985:IAD

[10] S. Christodoulakis. Issues in the architecture of a document archiver using optical disk technology. In Navathe [2254], pages 34–50. ISBN 0-89791-160-

1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p34-christodoulakis/p34-christodoulakis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/318898/
p34-christodoulakis/.

Barbic:1985:TMO

[11] F. Barbic and B. Pernici. Time modeling in office information systems. In Navathe [2254], pages 51–62. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1 .A87. URL http://www.

acm.org/pubs/articles/proceedings/
mod/318898/p51-barbic/p51-barbic.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/318898/
p51-barbic/.

Faloutsos:1985:SFD

[12] Chris Faloutsos. Signature files: design and performance comparison of some signature extraction methods. In Navathe [2254], pages 63–82. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1 .A87. URL http://www.

acm.org/pubs/articles/proceedings/

                                        mod/318898/p63-faloutsos/p63-faloutsos.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/318898/
                                        p63-faloutsos/.

Eick:1985:ATK

[13] Christoph F. Eick and Peter C. Lockemann. Acquisition of terminological knowledge using database design techniques. In Navathe [2254], pages 84–94. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/318898/
                                        p84-eick/p84-eick.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/318898/p84-eick/.

Shin:1985:PRD

[14] D. G. Shin and K. B. Irani. Partitioning a relational database horizontally using a knowledge-based approach. In Navathe [2254], pages 95–105. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/318898/
                                        p95-shin/p95-shin.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/318898/p95-shin/.

Agrawal:1985:MSC

[15] Rakesh Agrawal, Michael J. Carey, and Miron Livny. Models for studying concurrency control performance: alternatives and implications. In Navathe [2254], pages 108–121. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/318898/
                                        p108-agrawal/p108-agrawal.pdf;
                                        http://www.acm.org/pubs/citations/

proceedings/mod/318898/p108-agrawal/

.

Robinson:1985:FGP

[16] John T. Robinson. A fast generalpurpose hardware synchronization mechanism. In Navathe [2254], pages 122–130. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p122-robinson/p122-robinson.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/318898/p122-robinson/

.

Agrawal:1985:RAM

[17] Rakesh Agrawal and David J. DeWitt. Recovery architectures for multiprocessor database machines. In Navathe [2254], pages 131–145. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985;

QA1 .A87. URL http://www.acm.
org/pubs/articles/proceedings/mod/
318898/p131-agrawal/p131-agrawal.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/318898/
p131-agrawal/.

Batini:1985:DDM

[18] Carlo Batini, Stefano Ceri, Al Hershey, George Gardarin, and David Reiner. Database design: methodologies, tools, and environments (panel session). In Navathe [2254], pages 148–150. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p148-batini/p148-batini.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/318898/p148-batini/.

Hsu:1985:ICM

[19] Arding Hsu and Tomasz Imielinski. Integrity checking for multiple updates. In Navathe [2254], pages 152–168. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/citations/proceedings/mod/
                                        318898/p152-hsu/.

Kung:1985:VDT

[20] C. H. Kung. On verification of database temporal constraints. In Navathe [2254], pages 169–179. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985;

                                        QA1 .A87. URL http://www.acm.
                                        org/pubs/articles/proceedings/mod/
                                        318898/p169-kung/p169-kung.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/318898/p169-kung/

.

Kuper:1985:EPL

[21] Gabriel M. Kuper and Moshe Y. Vardi. On the expressive power of the logical data model: preliminary report. In Navathe [2254], pages 180–187. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/318898/
                                        p180-kuper/p180-kuper.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/318898/p180-kuper/.

Larson:1985:EPH

[22] Per-Ake Larson and M. V. Ramakrishna. External perfect hashing. In Navathe [2254], pages 190–200. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/318898/

p190-larson/p190-larson.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/318898/p190-larson/.

Kawagoe:1985:MDH

[23] Kyoji Kawagoe. Modified dynamic hashing. In Navathe [2254], pages 201–213. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p201-kawagoe/p201-kawagoe.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/318898/p201-kawagoe/

.

Otoo:1985:MDH

[24] Ekow J. Otoo. A multidimensional digital hashing scheme for files with composite keys. In Navathe [2254], pages 214–229. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p214-otoo/p214-otoo.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/318898/p214-otoo/.

Sibley:1985:PDM

[25] Edgar H. Sibley, Matthias Jarke, Cecil S. McMinn, John Murray, Randall Rustin, and Ken Sloan. Pragmatics of database management (panel session). In Navathe [2254], pages 232–234. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p232-sibley/p232-sibley.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/318898/p232-sibley/.

Snodgrass:1985:TTD

[26] Richard Snodgrass and Ilsoo Ahn. A taxonomy of time databases. In Navathe [2254], pages 236–246. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1 .A87. URL http://www.

                                        acm.org/pubs/articles/proceedings/
                                        mod/318898/p236-snodgrass/p236-
                                        snodgrass.pdf;   http://www.acm.
                                        org/pubs/citations/proceedings/
                                        mod/318898/p236-snodgrass/.

Clifford:1985:AHR

[27] James Clifford and Abdullah Uz Tansel. On an algebra for historical relational databases: two views. In Navathe [2254], pages 247–265. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/318898/
                                        p247-clifford/p247-clifford.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/318898/p247-clifford/

.

Copeland:1985:DSM

[28] George P. Copeland and Setrag N. Khoshafian. A decomposition storage model. In Navathe [2254], pages 268–279. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/318898/
                                        p268-copeland/p268-copeland.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/318898/p268-copeland/

.

Yu:1985:AIS

[29] C. T. Yu and C. H. Chen. Adaptive information system design: one query at a time. In Navathe [2254], pages 280–290. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p280-yu/p280-yu.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/318898/p280-yu/.

Beckley:1985:MRK

[30] D. A. Beckley, M. W. Evens, and V. K. Raman. Multikey retrieval from K-d trees and QUADtrees. In Navathe [2254], pages 291–301. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p291-beckley/p291-beckley.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/318898/p291-beckley/

.

Christodoulakis:1985:MDM

[31] Starvos Christodoulakis, D. Badal, A. Cardenas, P. Mantey, F. Tompa, and D. Tsichritzis. Multimedia database management (panel session). In Navathe [2254], pages 304–305. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p304-christodoulakis/p304-christodoulakis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/318898/
p304-christodoulakis/.

Fushimi:1985:APE

[32] Shinya Fushimi, Masaru Kitsuregawa, Masaya Nakayama, Hidehiko Tanaka, and Tohru Moto-oka. Algorithm and performance evaluation of adaptive multidimensional clustering technique. In Navathe [2254], pages 308–318. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/318898/
                                        p308-fushimi/p308-fushimi.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/318898/p308-fushimi/

.

Kamel:1985:MDD

[33] Nabil Kamel and Roger King. A model of data distribution based on texture analysis. In Navathe [2254], pages 319–325. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/318898/
                                        p319-kamel/p319-kamel.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/318898/p319-kamel/.

Goldman:1985:IIS

[34] Kenneth J. Goldman, Sally A. Goldman, Paris C. Kanellakis, and Stanley B. Zdonik. ISIS: interface for a semantic information system. In Navathe [2254], pages 328–342. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/318898/
                                        p328-goldman/p328-goldman.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/318898/p328-goldman/

.

Vossen:1985:HLU

[35] Gottfried Vossen and Volkert Brosda. A high-level user interface for update and retrieval in relational databases– language aspects. In Navathe [2254], pages 343–353. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p343-vossen/p343-vossen.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/318898/p343-vossen/.

Demo:1985:ACD

[36] G. Barbara Demo and Sukhamay Kundu. Analysis of the context dependency of CODASYL find-statements with application to a database program conversion. In Navathe [2254], pages 354–361. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p354-demo/p354-demo.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/318898/p354-demo/.

Acharya:1985:TRP

[37] Shridhar Acharya and Gael Buckley. Transaction restarts in Prolog database systems. In Navathe [2254], pages 364–373. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p364-acharya/p364-acharya.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/318898/p364-acharya/

.

Page:1985:GDD

[38] Thomas W. Page, Matthew J. Weinstein, and Gerald J. Popek. Genesis: a distributed database operating system. In Navathe [2254], pages 374–387. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p374-page/p374-page.pdf;   http:

                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/318898/p374-page/.

Klahold:1985:TMS

[39] P. Klahold, G. Schlageter, R. Unland, and W. Wilkes. A transaction model supporting complex applications in integrated information systems. In Navathe [2254], pages 388–401. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/citations/proceedings/mod/
                                        318898/p388-klahold/.

Sinha:1985:TBC

[40] Mukul K. Sinha, P. D. Nandikar, and S. L. Mehndiratta. Timestamp based certification schemes for transactions in distributed database systems. In Navathe [2254], pages 402–411. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/318898/
                                        p402-sinha/p402-sinha.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/318898/p402-sinha/.

Kerschberg:1985:EDS

[41] Larry Kerschberg, Michael Brodie, Charles Kellogg, D. Stott Parker, Gio Wiederhold, and Carlo Zaniolo. Expert database systems (workshop review). In Navathe [2254], pages 414–417. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/318898/
                                        p414-kerschberg/p414-kerschberg.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/318898/
                                        p414-kerschberg/.

Bhargava:1985:RDD

[42] Bharat Bhargava. Reliability in distributed database systems (panel discussion). In Navathe [2254], pages 420–422. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p420-bhargava/p420-bhargava.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/318898/p420-bhargava/

.

Sellis:1985:OED

[43] Timos K. Sellis and Leonard Shapiro. Optimization of extended database query languages. In Navathe [2254], pages 424–436. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p424-sellis/p424-sellis.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/318898/p424-sellis/.

Gray:1985:EPA

[44] P. M. D. Gray. Efficient Prolog access to CODAYSL and FDM databases. In Navathe [2254], pages 437–443. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/318898/
p437-gray/p437-gray.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/318898/p437-gray/.

Blain:1985:MPC

[45] Tomas Blain, Michael Dohler, Ralph Michaelis, and Emran Qureshi. Managing the printed circuit board design process. In Navathe [2254], pages 447–456. ISBN 0-89791-160-1.

LCCN QA 76.9 D3 I59 1985; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/318898/
                                        p447-blain/p447-blain.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/318898/p447-blain/.

Abiteboul:1985:TIC

[46] S. Abiteboul and V. Vianu. Transactions and integrity constraints. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Atzeni:1985:EQA

[47] P. Atzeni and E. P. F. Chan. Efficient query answering in the representative instance approach. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Ausiello:1985:CPG

[48] G. Ausiello and A. D’Atri. Chordality properties on graphs and minimal conceptual connections in semantic data models. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Bancilhon:1985:AVP

[49] F. Bancilhon and M. Spyratos. Algebraic versus probabilistic independence in data bases. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Casanova:1985:CLR

[50] M. A. Casanova, A. V. Moura, and L. Tucherman. On the correctness of a local recovery subsystem. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Cosmadakis:1985:PSR

[51] S. S. Cosmadakis, P. C. Kanellakis, and N. Spyratos. Partition semantics for relations. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Garcia-Molina:1985:EEC

[52] H. García-Molina and J. Kent. An experimental evaluation of crash recovery mechanism. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Gyssens:1985:EJD

[53] Marc Gyssens. Embedded join dependencies as a tool for decomposing full join dependencies. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Keller:1985:ATV

[54] A. M. Keller. Algorithms for translating view updates to database updates for views involving selections, projections, and joins. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Mannila:1985:SAR

[55] H. Mannila and K-J. Raiha. Small Armstrong relations for database design. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Page:1985:DDM

[56] T. W. Page, Jr. and G. J. Popek. Distributed data management in local area networks. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Sagiv:1985:COB

[57] Yehoshua Sagiv. Concurrent operations on B*-trees with overtaking. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Sagiv:1985:CRP

[58] Yehoshua Sagiv. On computing restricted projections of representative instances. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Skeen:1985:EFT

[59] D. Skeen, F. Cristian, and A. ElAbbadi. An efficient fault-tolerant algorithm for replicated data management. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Stein:1985:RUS

[60] J. Stein and D. Maier. Relaxing the universal scheme assumption. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Vardi:1985:QLD

[61] Moshe Vardi. Querying logical databases. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Yannakakis:1985:CRC

[62] Mihalis Yannakakis and C. H. Papadimitriou. The complexity of reliable concurrency control. In ACM [2253], page ?? ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Bancilhon:1986:MSO

[63] François Bancilhon, David Maier, Yehoshua Sagiv, and Jeffrey D.

Ullman. Magic sets and other strange ways to implement logic programs (extended abstract). In ACM [2255], pages 1–15. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

1986. URL http://www.acm.org/
pubs/articles/proceedings/pods/
6012/p1-bancilhon/p1-bancilhon.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/6012/

p1-bancilhon/. ACM order number 475860.

Sacca:1986:ISC

[64] Domenico Saccà and Carlo Zaniolo. On the implementation of a simple class of logic queries for databases. In ACM [2255], pages 16–23. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

1986. URL http://www.acm.org/
pubs/articles/proceedings/pods/
6012/p16-sacca/p16-sacca.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/6012/p16-sacca/

. ACM order number 475860.

Afrati:1986:CSQ

[65] Foto Afrati, Christos Papadimitriou, George Papageorgiou, Athena Roussou, Yehoshua Sagiv, and Jeffrey D. Ullman. Convergence of sideways query evaluation. In ACM [2255], pages 24–30. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

1986. URL http://www.acm.org/
pubs/articles/proceedings/pods/
6012/p24-afrati/p24-afrati.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/6012/p24-afrati/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/6012/

p24-afrati/. ACM order number 475860.

Weikum:1986:TFM

[66] Gerhard Weikum. A theoretical foundation of multi-level concurrency control. In ACM [2255], pages 31–43. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

                                        1986. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        6012/p31-weikum/p31-weikum.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/6012/p31-weikum/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/6012/

p31-weikum/. ACM order number

475860.

Hadzilacos:1986:DCT

[67] Thanasis Hadzilacos and Mihalis Yannakakis. Deleting completed transactions. In ACM [2255], pages 43–

46. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296 1986. URL http://

                                        www.acm.org/pubs/articles/proceedings/
                                        pods/6012/p43-hadzilacos/p43-hadzilacos.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/6012/
                                        p43-hadzilacos/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/6012/p43-hadzilacos/. ACM

order number 475860.

Su:1986:SNW

[68] Jianwen Su. Safety of non-well-locked transaction systems. In ACM [2255], pages 47–52. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

                                        1986. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        6012/p47-su/p47-su.pdf;   http:/
                                        /www.acm.org/pubs/citations/proceedings/
                                        pods/6012/p47-su/;   http://www.
                                        acm.org:80/pubs/citations/proceedings/
                                        pods/6012/p47-su/. ACM order num-

ber 475860.

Bancilhon:1986:CCO

[69] François Bancilhon and Setrag Khoshafian. A calculus for complex objects. In ACM [2255], pages 53–60. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

1986. URL http://www.acm.org/
pubs/articles/proceedings/pods/
6012/p53-bancilhon/p53-bancilhon.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/6012/
p53-bancilhon/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/6012/p53-bancilhon/. ACM

order number 475860.

VanGucht:1986:SCM

[70] Dirk Van Gucht and Patrick C. Fischer. Some classes of multilevel relational structures. In ACM [2255], pages 60–69. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

1986. URL http://www.acm.org/
pubs/articles/proceedings/pods/
6012/p60-van_gucht/p60-van_gucht.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/6012/
p60-van_gucht/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/6012/p60-van_gucht/. ACM

order number 475860.

Gadia:1986:WTR

[71] Shashi K. Gadia. Weak temporal relations. In ACM [2255], pages 70–

77. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296 1986. URL http://

www.acm.org/pubs/articles/proceedings/
pods/6012/p70-gadia/p70-gadia.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/6012/p70-gadia/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/6012/

p70-gadia/. ACM order number 475860.

Olken:1986:RDM

[72] Frank Olken and Doron Rotem. Rearranging data to maximize the efficiency of compression. In ACM [2255], pages 78–90. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

1986. URL http://www.acm.org/
pubs/articles/proceedings/pods/
6012/p78-olken/p78-olken.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/6012/p78-olken/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/6012/

p78-olken/. ACM order number 475860.

Robinson:1986:OPL

[73] John T. Robinson. Order preserving linear hashing using dynamic key statistics. In ACM [2255], pages 91–99. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

1986. URL http://www.acm.org/
pubs/articles/proceedings/pods/
6012/p91-robinson/p91-robinson.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/6012/
p91-robinson/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/6012/p91-robinson/. ACM or-

der number 475860.

Otoo:1986:BME

[74] Ekow J. Otoo. Balanced multidimensional extendible hash tree. In ACM [2255], pages 100–113. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

1986. URL http://www.acm.org/
pubs/articles/proceedings/pods/
6012/p100-otoo/p100-otoo.pdf;
http://www.acm.org/pubs/citations/

proceedings/pods/6012/p100-otoo/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/6012/

p100-otoo/. ACM order number 475860.

Naqvi:1986:NFF

[75] Shamim A. Naqvi. Negation as failure for first-order queries. In ACM [2255], pages 114–122. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

1986. URL http://www.acm.org/
pubs/articles/proceedings/pods/
6012/p114-naqvi/p114-naqvi.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/6012/p114-naqvi/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/6012/

p114-naqvi/. ACM order number 475860.

Bidoit:1986:PVM

[76] Nicole Bidoit and Richard Hull. Positivism vs. minimalism in deductive databases. In ACM [2255], pages 123–132. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

1986. URL http://www.acm.org/
pubs/articles/proceedings/pods/
6012/p123-bidoit/p123-bidoit.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/6012/p123-bidoit/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/6012/

p123-bidoit/. ACM order number 475860.

Gelfond:1986:ECW

[77] M. Gelfond, H. Przymusinska, and T. Przymusinski. The extended closed world assumption and its relationship to parallel circumscription. In ACM [2255], pages 133–139. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

1986. URL http://www.acm.org/
pubs/articles/proceedings/pods/
6012/p133-gelfond/p133-gelfond.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/6012/
p133-gelfond/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/6012/p133-gelfond/. ACM or-

der number 475860.

Chan:1986:PCC

[78] E. P. F. Chan and Paolo Atzeni. On the properties and characterization of connection-trap-free schemes. In ACM [2255], pages 140–147. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

1986. URL http://www.acm.org/
pubs/articles/proceedings/pods/
6012/p140-chan/p140-chan.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/6012/p140-chan/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/6012/

p140-chan/. ACM order number 475860.

Biskup:1986:OFA

[79] H. Biskup and L. Schnetgoke. One flavor assumption and gamma-acyclicity for universal relation views. In ACM [2255], pages 148–159. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

1986. URL http://www.acm.org/
pubs/articles/proceedings/pods/
6012/p148-biskup/p148-biskup.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/6012/p148-biskup/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/6012/

p148-biskup/. ACM order number 475860.

Sagiv:1986:ESQ

[80] Yehoshua Sagiv and Oded Shmueli.

The equivalence of solving queries and producing tree projections (extended abstract). In ACM [2255], pages 160–172. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

1986. URL http://www.acm.org/
pubs/articles/proceedings/pods/
6012/p160-sagiv/p160-sagiv.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/6012/p160-sagiv/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/6012/

p160-sagiv/. ACM order number 475860.

Sagiv:1986:FFA

[81] Yehoshua Sagiv. On finite FDacyclicity. In ACM [2255], pages 173–

182. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296 1986. URL http://

www.acm.org/pubs/articles/proceedings/
pods/6012/p173-sagiv/p173-sagiv.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/6012/
p173-sagiv/;   http://www.acm.org:
80/pubs/citations/proceedings/pods/
6012/p173-sagiv/. ACM order num-

ber 475860.

Ozsoyoglu:1986:UFM

[82] Meral Ozsoyoglu and Li Yan Yuan. Unifying functional and multivalued dependencies for relational database design. In ACM [2255], pages 183–

190. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296 1986. URL http://

www.acm.org/pubs/articles/proceedings/
pods/6012/p183-ozsoyoglu/p183-ozsoyoglu.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/6012/
p183-ozsoyoglu/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/6012/p183-ozsoyoglu/. ACM

order number 475860.

Ruland:1986:AAD

[83] Detlev Ruland and Dietmar Seipel. Alpha-acyclic decompositions of relational database schemes. In ACM [2255], pages 191–201. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

                                        1986. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        6012/p191-ruland/p191-ruland.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/6012/p191-ruland/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/6012/

p191-ruland/. ACM order number 475860.

Graham:1986:CTM

[84] Marc H. Graham and Ke Wang. Constant time maintenance or the triumph of the FD. In ACM [2255], pages 202–216. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

                                        1986. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        6012/p202-graham/p202-graham.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/6012/p202-graham/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/6012/

p202-graham/. ACM order number 475860.

Mannila:1986:TDR

[85] Heikki Mannila and Kari Jouko Raiha. Test data for relational queries. In ACM [2255], pages 217–223. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

                                        1986. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        6012/p217-mannila/p217-mannila.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/6012/
                                        p217-mannila/;   http://www.acm.

org:80/pubs/citations/proceedings/
pods/6012/p217-mannila/. ACM or-

der number 475860.

Wilkins:1986:MTA

[86] Marianne Winslett Wilkins. A model-theoretic approach to updating logical databases. In ACM [2255], pages 224–234. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

1986. URL http://www.acm.org/
pubs/articles/proceedings/pods/
6012/p224-wilkins/p224-wilkins.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/6012/
p224-wilkins/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/6012/p224-wilkins/. ACM or-

der number 475860.

Abiteboul:1986:DPT

[87] Serge Abiteboul and Victor Vianu. Deciding properties of transactional schemas. In ACM [2255], pages 235–

239. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296 1986. URL http://

www.acm.org/pubs/articles/proceedings/
pods/6012/p235-abiteboul/p235-abiteboul.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/6012/
p235-abiteboul/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/6012/p235-abiteboul/. ACM

order number 475860.

ElAbbadi:1986:APR

[88] Amr El Abbadi and Sam Toueg. Availability in partitioned replicated databases. In ACM [2255], pages 240–251. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

1986. URL http://www.acm.org/
pubs/articles/proceedings/pods/
6012/p240-el_abbadi/p240-el_abbadi.

                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/6012/
                                        p240-el_abbadi/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/6012/p240-el_abbadi/. ACM

order number 475860.

Vardi:1986:IDI

[89] Moshe Vardi. On the integrity of databases with incomplete information. In ACM [2255], pages 252–266. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

                                        1986. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        6012/p252-vardi/p252-vardi.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/6012/p252-vardi/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/6012/

p252-vardi/. ACM order number

475860.

Naughton:1986:DIR

[90] Jeff Naughton. Data independent recursion in deductive databases. In ACM [2255], pages 267–279. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296

                                        1986. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        6012/p267-naughton/p267-naughton.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/6012/
                                        p267-naughton/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/6012/p267-naughton/. ACM

order number 475860.

Cosmadakis:1986:PER

[91] S. Cosmadakis and P. Kanellakis. Parallel evaluation of recursive rule queries. In ACM [2255], pages 280–

293. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296 1986. URL http://

                                        www.acm.org/pubs/articles/proceedings/

pods/6012/p280-cosmadakis/p280-
cosmadakis.pdf;   http://www.acm.
org/pubs/citations/proceedings/
pods/6012/p280-cosmadakis/;   http:
//www.acm.org:80/pubs/citations/
proceedings/pods/6012/p280-cosmadakis/

. ACM order number 475860.

Abiteboul:1986:PTS

[92] S. Abiteboul and V. Vianu. Properties of transactional schemas. In ACM [2255], page ?? ISBN 0-89791-179-2. LCCN QA76.9.D3 A296 1986. ACM order number 475860.

Neff:1987:DBC

[93] R. K. Neff. Data bases, compound objects, and networked workstations: Beyond distributed computing (Abstract). In Dayal and Traiger [2258], page 1. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p1-neff/p1-neff.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/38713/p1-neff/. ACM order

number 472870.

Ullman:1987:DTP

[94] J. D. Ullman. Database theory — past and future. In ACM [2257], pages 1–10. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p1-ullman/p1-ullman.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/28659/p1-ullman/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/28659/

p1-ullman/. ACM Order Number 475870.

Ingenthron:1987:TDR

[95] Kurt Ingenthron. Thoughts on database research: a user perspective. In Dayal and Traiger [2258], page 2. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/38713/
                                        p2-ingenthron/p2-ingenthron.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/38713/p2-ingenthron/

. ACM order number 472870.

Ioannidis:1987:QOS

[96] Yannis E. Ioannidis and Eugene Wong. Query optimization by simulated annealing. In Dayal and Traiger [2258], pages 9–22. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/38713/
                                        p9-ioannidis/p9-ioannidis.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/38713/p9-ioannidis/

. ACM order number 472870.

Kuper:1987:LPS

[97] G. M. Kuper. Logic programming with sets. In ACM [2257], pages 11–20. ISBN 0-89791-223-3. LCCN QA76.9.D3A296 1987. URL http://

                                        www.acm.org/pubs/articles/proceedings/
                                        pods/28659/p11-kuper/p11-kuper.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/28659/
                                        p11-kuper/;   http://www.acm.org:
                                        80/pubs/citations/proceedings/pods/
                                        28659/p11-kuper/. ACM Order Num-

ber 475870.

Beeri:1987:SNL

[98] C. Beeri, S. Naqvi, R. Ramakrishnan, O. Shmueli, and S. Tsur.

Sets and negation in a logic data base language (LDL1). In ACM [2257], pages 21–37. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p21-beeri/p21-beeri.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/28659/p21-beeri/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/28659/

p21-beeri/. ACM Order Number 475870.

Ganski:1987:ONS

[99] Richard A. Ganski and Harry K. T. Wong. Optimization of nested SQL queries revisited. In Dayal and Traiger [2258], pages 23–33. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p23-ganski/p23-ganski.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/38713/p23-ganski/. ACM order

number 472870.

Abiteboul:1987:RQS

[100] Serge Abiteboul, Paris Kanellakis, and Gosta Grahne. On the representation and querying of sets of possible worlds. In Dayal and Traiger [2258], pages 34–48. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p34-abiteboul/p34-abiteboul.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/38713/p34-abiteboul/

. ACM order number 472870.

Yuan:1987:LDR

[101] L. Y. Yuan and Z. M. Ozsoyoglu. Logical design of relational database schemes. In ACM [2257], pages 38–

47. ISBN 0-89791-223-3. LCCN QA76.9.D3A296 1987. URL http://

                                        www.acm.org/pubs/articles/proceedings/
                                        pods/28659/p38-yuan/p38-yuan.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/28659/p38-yuan/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/28659/

p38-yuan/. ACM Order Number

475870.

Chan:1987:DDS

[102] E. P. F. Chan and H. J. Hernandez. On designing database schemes bounded or constant-time maintainable with respect to functional dependencies. In ACM [2257], pages 48–57. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        28659/p48-chan/p48-chan.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        pods/28659/p48-chan/;   http://
                                        www.acm.org:80/pubs/citations/proceedings/
                                        pods/28659/p48-chan/. ACM Order

Number 475870.

Sacca:1987:MCM

[103] Domenico Sacca and Carlo Zaniolo. Magic counting methods. In Dayal and Traiger [2258], pages 49–59. ISBN 0-89791-236-5. LCCN QA 76.9 D3

                                        P76 1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/38713/
                                        p49-sacca/p49-sacca.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/38713/p49-sacca/. ACM order

number 472870.

Gottlob:1987:CCE

[104] G. Gottlob. Computing covers for embedded functional dependencies. In ACM [2257], pages 58–69. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p58-gottlob/p58-gottlob.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/28659/p58-gottlob/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/28659/

p58-gottlob/. ACM Order Number 475870.

Aly:1987:NDM

[105] Hussien Aly and Z. Meral Ozsoyoglu. Non-deterministic modelling of logical queries in deductive databases. In Dayal and Traiger [2258], pages 60–72. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p60-aly/p60-aly.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/38713/p60-aly/. ACM order

number 472870.

DAtri:1987:DQI

[106] A. D’Atri, P. Di Felice, and M. Moscarini. Dynamic query interpretation in relational databases. In ACM [2257], pages 70–78. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p70-d_atri/p70-d_atri.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/28659/p70-d_atri/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/28659/

p70-d_atri/. ACM Order Number 475870.

Han:1987:HRP

[107] Jiawei Han and Lawrence J. Henschen. Handling redundancy in the processing of recursive database queries. In Dayal and Traiger [2258], pages 73–81. ISBN 0-89791-236-5. LCCN QA 76.9 D3

P76 1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p73-han/p73-han.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/38713/p73-han/. ACM order

number 472870.

Atzeni:1987:NBW

[108] P. Atzeni and M. C. De Bernardis. A new basis for the weak instance model. In ACM [2257], pages 79–86. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p79-atzeni/p79-atzeni.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/28659/p79-atzeni/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/28659/

p79-atzeni/. ACM Order Number 475870.

Daniels:1987:DLT

[109] Dean S. Daniels, Alfred Z. Spector, and Dean S. Thompson. Distributed logging for transaction processing. In Dayal and Traiger [2258], pages 82–96. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p82-daniels/p82-daniels.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/38713/p82-daniels/. ACM or-

der number 472870.

Malvestuto:1987:AQC

[110] F. M. Malvestuto. Answering queries in categorical databases. In ACM [2257], pages 87–96. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p87-malvestuto/p87-malvestuto.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/28659/
p87-malvestuto/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/28659/p87-malvestuto/. ACM

Order Number 475870.

Herman:1987:DAV

[111] Gary Herman, K. C. Lee, and Abel Weinrib. The datacycle architecture for very high throughput database systems. In Dayal and Traiger [2258], pages 97–103. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p97-herman/p97-herman.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/38713/p97-herman/. ACM order

number 472870.

Fekete:1987:NTR

[112] A. Fekete, N. Lynch, M. Merrit, and W. Weihl. Nested transactions and read-write locking. In ACM [2257], pages 97–111. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p97-fekete/p97-fekete.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/28659/p97-fekete/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/28659/

p97-fekete/. ACM Order Number 475870.

Lehman:1987:RAH

[113] Tobin J. Lehman and Michael J. Carey. A recovery algorithm for a high-performance memory-resident database system. In Dayal and Traiger [2258], pages 104–117. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/38713/
                                        p104-lehman/p104-lehman.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/38713/p104-lehman/. ACM or-

der number 472870.

Segall:1987:TCM

[114] A. Segall and O. Wolfson. Transaction commitment at minimal communication cost. In ACM [2257], pages 112–118. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        28659/p112-segall/p112-segall.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/28659/p112-segall/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/28659/

p112-segall/. ACM Order Number

475870.

Nixon:1987:ICS

[115] Brian Nixon, Lawrence Chung, John Mylopoulos, David Lauzon, Alex Borgida, and M. Stanley. Implementation of a compiler for a semantic data model: Experiences with taxis. In Dayal and Traiger [2258], pages 118–131. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/38713/
                                        p118-nixon/p118-nixon.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/38713/p118-nixon/. ACM order

number 472870.

Wang:1987:PAM

[116] C. P. Wang and V. O. K. Li.

The precedence-assignment model for distributed databases concurrency control algorithms. In ACM [2257], pages 119–128. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p119-wang/p119-wang.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/28659/p119-wang/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/28659/

p119-wang/. ACM Order Number 475870.

Hadzilacos:1987:KTA

[117] V. Hadzilacos. A knowledgetheoretic analysis of atomic commitment protocols. In ACM [2257], pages 129–134. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p129-hadzilacos/p129-hadzilacos.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/28659/
p129-hadzilacos/;   http://www.
acm.org:80/pubs/citations/proceedings/
pods/28659/p129-hadzilacos/. ACM

Order Number 475870.

Lyngbaek:1987:MSD

[118] Peter Lyngbaek and Victor Vianu. Mapping a semantic database model to the relational model. In Dayal and Traiger [2258], pages 132–142. ISBN 0-89791-236-5. LCCN QA 76.9 D3

P76 1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p132-lyngbaek/p132-lyngbaek.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/38713/p132-lyngbaek/

. ACM order number 472870.

Minker:1987:PDD

[119] J. Minker. Perspectives in deductive databases (Abstract only). In ACM [2257], pages 135–136 (or 135–135??). ISBN 0-89791-223-3. LCCN QA76.9.D3A296

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        28659/p135-minker/p135-minker.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/28659/p135-minker/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/28659/

p135-minker/. ACM Order Number

475870.

Apt:1987:MSD

[120] K. Apt and J. M. Pugin. Maintenance of stratified databases viewed as a belief revision system. In ACM [2257], pages 136–145. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        28659/p136-apt/p136-apt.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        pods/28659/p136-apt/;   http://
                                        www.acm.org:80/pubs/citations/proceedings/
                                        pods/28659/p136-apt/. ACM Order

Number 475870.

Roth:1987:DRD

[121] Mark A. Roth and Henry F. Korth. The design of 1NF relational databases into nested normal form. In Dayal and Traiger [2258], pages 143–159. ISBN 0-89791-236-5. LCCN QA 76.9 D3

                                        P76 1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/38713/
                                        p143-roth/p143-roth.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/38713/p143-roth/. ACM order

number 472870.

Hegner:1987:SIP

[122] S. Hegner. Specification and implementation of programs for updating incomplete information databases. In ACM [2257], pages 146–158. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p146-hegner/p146-hegner.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/28659/p146-hegner/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/28659/

p146-hegner/. ACM Order Number 475870.

Biliris:1987:OSL

[123] A. Biliris. Operation specific locking in B-trees. In ACM [2257], pages 159–169. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p159-biliris/p159-biliris.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/28659/
p159-biliris/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/28659/p159-biliris/. ACM

Order Number 475870.

Graefe:1987:EOG

[124] Goetz Graefe and David J. De-Witt. The EXODUS optimizer generator. In Dayal and Traiger [2258], pages 160–172. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p160-graefe/p160-graefe.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/38713/p160-graefe/. ACM or-

der number 472870.

Nurmi:1987:CCD

[125] O. Nurmi, E. Soisalon-Soininen, and D. Wood. Concurrency control in database structures with relaxed balance. In ACM [2257], pages 170–176. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        28659/p170-nurmi/p170-nurmi.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/28659/p170-nurmi/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/28659/

p170-nurmi/. ACM Order Number

475870.

Freytag:1987:RBV

[126] Johann Christoph Freytag. A rule-based view of query optimization. In Dayal and Traiger [2258], pages 173–180. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/38713/
                                        p173-freytag/p173-freytag.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/38713/p173-freytag/

. ACM order number 472870.

Sun:1987:PRM

[127] R. Sun and G. Thomas. Performance results on multiversion timestamp concurrency control with predeclared writesets. In ACM [2257], pages 177–184. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        28659/p177-sun/p177-sun.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        pods/28659/p177-sun/;   http://
                                        www.acm.org:80/pubs/citations/proceedings/

pods/28659/p177-sun/. ACM Order

Number 475870.

Shenoy:1987:SSQ

[128] Sreekumar T. Shenoy and Z. Meral Ozsoyoglu. A system for semantic query optimization. In Dayal and Traiger [2258], pages 181–195. ISBN 0-89791-236-5. LCCN QA 76.9 D3

P76 1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p181-shenoy/p181-shenoy.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/38713/p181-shenoy/. ACM or-

der number 472870.

Dechter:1987:DAR

[129] R. Dechter. Decomposing an N-ary relation into a tree of binary relations. In ACM [2257], pages 185–189. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p185-dechter/p185-dechter.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/28659/

p185-dechter/. ACM Order Number 475870.

Delgrande:1987:FLA

[130] J. P. Delgrande. Formal limits on the automatic generation and maintenance of integrity constraints. In ACM [2257], pages 190–196. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p190-delgrande/p190-delgrande.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/28659/
p190-delgrande/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/28659/p190-delgrande/. ACM

Order Number 475870.

Paul:1987:AID

[131] H. B. Paul, H. J. Schek, and M. H. Scholl. Architecture and implementation of the Darmstadt database kernel system. In Dayal and Traiger [2258], pages 196–207. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/38713/
                                        p196-paul/p196-paul.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/38713/p196-paul/. ACM order

number 472870.

Imielinski:1987:RKD

[132] T. Imielinski. Relative knowledge in a distributed database. In ACM [2257], pages 197–209. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        28659/p197-imielinski/p197-imielinski.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/28659/
                                        p197-imielinski/;   http://www.
                                        acm.org:80/pubs/citations/proceedings/
                                        pods/28659/p197-imielinski/. ACM

Order Number 475870.

Richardson:1987:PCD

[133] Joel E. Richardson and Michael J. Carey. Programming constructs for database system implementation in EXODUS. In Dayal and Traiger [2258], pages 208–219. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/38713/
                                        p208-richardson/p208-richardson.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/38713/
                                        p208-richardson/. ACM order num-

ber 472870.

Afrati:1987:PCS

[134] F. Afrati and C. Papadimitriou. The parallel complexity of simple chain queries. In ACM [2257], pages 210–213. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p210-afrati/p210-afrati.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/28659/p210-afrati/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/28659/

p210-afrati/. ACM Order Number 475870.

Beeri:1987:BPS

[135] C. Beeri, P. Kanellakis, F. Bancilhon, and R. Ramakrishnan. Bounds on the propagation of selection into logic programs. In ACM [2257], pages 214–226. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p214-beeri/p214-beeri.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/28659/p214-beeri/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/28659/

p214-beeri/. ACM Order Number 475870.

Lindsay:1987:DME

[136] Bruce Lindsay, John McPherson, and Hamid Pirahesh. A data management extension architecture. In Dayal and Traiger [2258], pages 220–226. ISBN 0-89791-236-5. LCCN QA 76.9 D3

P76 1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p220-lindsay/p220-lindsay.pdf;
http://www.acm.org/pubs/citations/

proceedings/mod/38713/p220-lindsay/

. ACM order number 472870.

Naughton:1987:DCB

[137] J. F. Naughton and Y. Sagiv. A decidable class of bounded recursions. In ACM [2257], pages 227–236. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p227-naughton/p227-naughton.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/28659/
p227-naughton/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/28659/p227-naughton/. ACM

Order Number 475870.

Jajodia:1987:DV

[138] Sushil Jajodia and David Mutchler. Dynamic voting. In Dayal and Traiger [2258], pages 227–238. ISBN 0-89791-236-5. LCCN QA 76.9 D3

P76 1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p227-jajodia/p227-jajodia.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/38713/p227-jajodia/

. ACM order number 472870.

Shmueli:1987:DEA

[139] O. Shmueli. Decidability and expressiveness aspects of logic queries. In ACM [2257], pages 237–249. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p237-shmueli/p237-shmueli.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/28659/
p237-shmueli/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/28659/p237-shmueli/. ACM

Order Number 475870.

Haerder:1987:CTR

[140] Theo Haerder and Kurt Rothermel. Concepts for transaction recovery in nested transactions. In Dayal and Traiger [2258], pages 239–248. ISBN 0-89791-236-5. LCCN QA 76.9 D3

P76 1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p239-haerder/p239-haerder.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/38713/p239-haerder/

. ACM order number 472870.

Garcia-Molina:1987:S

[141] Hector Garcia-Molina and Kenneth Salem. Sagas. In Dayal and Traiger [2258], pages 249–259. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p249-garcia-molina/p249-garcia-
molina.pdf;   http://www.acm.org/
pubs/citations/proceedings/mod/
38713/p249-garcia-molina/. ACM

order number 472870.

Selinger:1987:CEI

[142] P. Selinger. Chickens and eggs — the interrelationship of systems and theory. In ACM [2257], pages 250–253. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p250-selinger/p250-selinger.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/28659/
p250-selinger/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/28659/p250-selinger/. ACM

Order Number 475870.

Karabeg:1987:ASR

[143] A. Karabeg, D. Karabeg, K. Papakonstantinou, and V. Vianu. Axiomatization and simplification rules for relational transactions. In ACM [2257], pages 254–259. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p254-karabeg/p254-karabeg.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/28659/
p254-karabeg/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/28659/p254-karabeg/. ACM

Order Number 475870.

Abiteboul:1987:TLC

[144] S. Abiteboul and V. Vianu. A translation language complete for database update and specification. In ACM [2257], pages 260–268. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p260-abiteboul/p260-abiteboul.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/28659/
p260-abiteboul/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/28659/p260-abiteboul/. ACM

Order Number 475870.

Freeston:1987:BFN

[145] Michael Freeston. The BANG file: a new kind of grid file. In Dayal and Traiger [2258], pages 260–269. ISBN 0-89791-236-5. LCCN QA 76.9 D3

P76 1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p260-freeston/p260-freeston.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/38713/p260-freeston/

. ACM order number 472870.

Beeri:1987:PM

[146] C. Beeri and R. Ramakrishnan. On the power of magic. In ACM [2257], pages 269–284. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p269-beeri/p269-beeri.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/28659/p269-beeri/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/28659/

p269-beeri/. ACM Order Number 475870.

Nelson:1987:PAH

[147] Randal C. Nelson and Hanan Samet. A population analysis for hierarchical data structures. In Dayal and Traiger [2258], pages 270–277. ISBN 0-89791-236-5. LCCN QA 76.9 D3

P76 1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p270-nelson/p270-nelson.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/38713/p270-nelson/. ACM or-

der number 472870.

Sellis:1987:ESP

[148] Timos K. Sellis. Efficiently supporting procedures in relational database systems. In Dayal and Traiger [2258], pages 278–291. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p278-sellis/p278-sellis.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/38713/p278-sellis/. ACM or-

der number 472870.

Grahne:1987:EES

[149] G. Grahne, S. Sippu, and E. Soisalon-Soininen. Efficient evaluation for a subset of recursive queries. In ACM [2257], pages 284–293. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p284-grahne/p284-grahne.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/28659/p284-grahne/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/28659/

p284-grahne/. ACM Order Number 475870.

Hardwick:1987:WRF

[150] Martin Hardwick. Why ROSE is fast: Five optimizations in the design of an experimental database system for CAD/CAM applications. In Dayal and Traiger [2258], pages 292–298. ISBN 0-89791-236-5. LCCN QA 76.9 D3

P76 1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p292-hardwick/p292-hardwick.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/38713/p292-hardwick/

. ACM order number 472870.

Marchetti-Spaccamella:1987:WCC

[151] A. Marchetti-Spaccamella, A. Pelaggi, and D. Sacca. Worst-case complexity analysis of methods for logic query implementation. In ACM [2257], pages 294–301. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p294-marchetti-spaccamella/
p294-marchetti-spaccamella.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/28659/p294-marchetti-
spaccamella/;   http://www.acm.
org:80/pubs/citations/proceedings/

pods/28659/p294-marchetti-spaccamella/

. ACM Order Number 475870.

Kemper:1987:OOS

[152] Alfons Kemper, Peter C. Lockemann, and Mechtild Wallrath. An objectoriented system for engineering applications. In Dayal and Traiger [2258], pages 299–310. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p299-kemper/p299-kemper.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/38713/p299-kemper/. ACM or-

der number 472870.

VanGucht:1987:EPE

[153] D. Van Gucht. On the expressive power of the extended relational algebra for the unnormalized relational model. In ACM [2257], pages 302–312. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p302-van_gucht/p302-van_gucht.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/28659/
p302-van_gucht/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/28659/p302-van_gucht/. ACM

Order Number 475870.

Banerjee:1987:SIS

[154] Jay Banerjee, Won Kim, Hyoung-Joo Kim, and Henry F. Korth. Semantics and implementation of schema evolution in object-oriented databases. In Dayal and Traiger [2258], pages 311–322. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p311-banerjee/p311-banerjee.pdf;

                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/38713/p311-banerjee/

. ACM order number 472870.

VanGelder:1987:SCT

[155] A. Van Gelder and R. Topor. Safety and correct translation of relational calculus formulas. In ACM [2257], pages 313–327. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        28659/p313-van_gelder/p313-van_
                                        gelder.pdf;   http://www.acm.org/
                                        pubs/citations/proceedings/pods/
                                        28659/p313-van_gelder/. ACM Or-

der Number 475870.

Cruz:1987:GQL

[156] Isabel F. Cruz, Alberto O. Mendelzon, and Peter T. Wood. A graphical query language supporting recursion. In Dayal and Traiger [2258], pages 323–330. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/38713/
                                        p323-cruz/p323-cruz.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/38713/p323-cruz/. ACM order

number 472870.

Ramakrishnan:1987:SRH

[157] R. Ramakrishnan, F. Bancilhon, and A. Silberschatz. Safety of recursive Horn clauses with infinite relations. In ACM [2257], pages 328–339. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        28659/p328-ramakrishnan/p328-ramakrishnan.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/28659/
                                        p328-ramakrishnan/;   http://www.

acm.org:80/pubs/citations/proceedings/
pods/28659/p328-ramakrishnan/. ACM

Order Number 475870.

Jagadish:1987:STC

[158] H. V. Jagadish, Rakesh Agrawal, and Linda Ness. A study of transitive closure as a recursion mechanism. In Dayal and Traiger [2258], pages 331–344. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p331-jagadish/p331-jagadish.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/38713/p331-jagadish/

. ACM order number 472870.

Naughton:1987:OSR

[159] J. F. Naughton. One-sided recursions. In ACM [2257], pages 340–348. ISBN 0-89791-223-3. LCCN QA76.9.D3A296

1987. URL http://www.acm.org/
pubs/articles/proceedings/pods/
28659/p340-naughton/p340-naughton.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/28659/
p340-naughton/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/28659/p340-naughton/. ACM

Order Number 475870.

Zhang:1987:NCD

[160] Weining Zhang and C. T. Yu. A necessary condition for a doubly recursive rule to be equivalent to a linear recursive rule. In Dayal and Traiger [2258], pages 345–356. ISBN 0-89791-236-5. LCCN QA 76.9 D3

P76 1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p345-zhang/p345-zhang.pdf;   http:
//www.acm.org/pubs/citations/proceedings/

mod/38713/p345-zhang/. ACM order

number 472870.

Sagiv:1987:ODP

[161] Y. Sagiv. Optimizing datalog programs. In ACM [2257], pages 349–

362. ISBN 0-89791-223-3. LCCN QA76.9.D3A296 1987. URL http://

www.acm.org/pubs/articles/proceedings/
pods/28659/p349-sagiv/p349-sagiv.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/28659/
p349-sagiv/;   http://www.acm.org:
80/pubs/citations/proceedings/pods/
28659/p349-sagiv/. ACM Order

Number 475870.

Morgenstern:1987:SIM

[162] Matthew Morgenstern. Security and inference in multilevel database and knowledge-base systems. In Dayal and Traiger [2258], pages 357–373. ISBN 0-89791-236-5. LCCN QA 76.9 D3

P76 1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p357-morgenstern/p357-morgenstern.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/38713/
p357-morgenstern/. ACM order num-

ber 472870.

Stemple:1987:MMF

[163] David Stemple, Subhasish Mazumdar, and Tim Sheard. On the modes and meaning of feedback to transaction designers. In Dayal and Traiger [2258], pages 374–386. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p374-stemple/p374-stemple.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/38713/p374-stemple/

. ACM order number 472870.

Rubenstein:1987:BSD

[164] W. B. Rubenstein, M. S. Kubicar, and R. G. G. Cattell. Benchmarking simple database operations. In Dayal and Traiger [2258], pages 387–394. ISBN 0-89791-236-5. LCCN QA 76.9 D3

P76 1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p387-rubenstein/p387-rubenstein.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/38713/
p387-rubenstein/. ACM order num-

ber 472870.

Gray:1987:MRT

[165] Jim Gray and Franco Putzolu. The 5 minute rule for trading memory for disc accesses and the 10 byte rule for trading memory for CPU time. In Dayal and Traiger [2258], pages 395–398. ISBN 0-89791-236-5. LCCN QA 76.9 D3

P76 1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p395-gray/p395-gray.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/38713/p395-gray/. ACM order

number 472870.

Richardson:1987:DEP

[166] James P. Richardson, Hongjun Lu, and Krishna Mikkilineni. Design and evaluation of parallel pipelined join algorithms. In Dayal and Traiger [2258], pages 399–409. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/
p399-richardson/p399-richardson.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/38713/
p399-richardson/. ACM order num-

ber 472870.

Butler:1987:SRO

[167] Margaret H. Butler. Storage reclamation in object oriented database systems. In Dayal and Traiger [2258], pages 410–425. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/38713/
                                        p410-butler/p410-butler.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/38713/p410-butler/. ACM or-

der number 472870.

Faloutsos:1987:AOO

[168] Christos Faloutsos, Timos Sellis, and Nick Roussopoulos. Analysis of object oriented spatial access methods. In Dayal and Traiger [2258], pages 426–439. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/38713/
                                        p426-faloutsos/p426-faloutsos.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/38713/p426-faloutsos/

. ACM order number 472870.

Hanson:1987:PAV

[169] Eric N. Hanson. A performance analysis of view materialization strategies. In Dayal and Traiger [2258], pages 440–453. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76

                                        1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/38713/
                                        p440-hanson/p440-hanson.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/38713/p440-hanson/. ACM or-

der number 472870.

Segev:1987:LMT

[170] Arie Segev and Arie Shoshani. Logical modeling of temporal data. In Dayal and Traiger [2258], pages 454–466. ISBN 0-89791-236-5. LCCN QA 76.9

D3 P76 1987. URL http://www.acm.
org/pubs/articles/proceedings/mod/
38713/p454-segev/p454-segev.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/38713/p454-segev/

. ACM order number 472870.

McKenzie:1987:ERA

[171] Edwin McKenzie and Richard Snodgrass. Extending the relational algebra to support transaction time. In Dayal and Traiger [2258], pages 467–478. ISBN 0-89791-236-5. LCCN QA 76.9

D3 P76 1987. URL http://www.acm.
org/pubs/articles/proceedings/mod/
38713/p467-mckenzie/p467-mckenzie.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/38713/

p467-mckenzie/. ACM order number 472870.

Rubenstein:1987:DDM

[172] W. Bradley Rubenstein. A database design for musical information. In Dayal and Traiger [2258], pages 479– 490. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76 1987. URL http://www.

acm.org/pubs/articles/proceedings/
mod/38713/p479-rubenstein/p479-
rubenstein.pdf;   http://www.acm.
org/pubs/citations/proceedings/
mod/38713/p479-rubenstein/. ACM

order number 472870.

Hudson:1987:OOD

[173] Scott E. Hudson and Roger King. Object-oriented database support for software environments. In Dayal and Traiger [2258], pages 491–503. ISBN 0-89791-236-5. LCCN QA 76.9 D3

P76 1987. URL http://www.acm.org/
pubs/articles/proceedings/mod/38713/

                                        p491-hudson/p491-hudson.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/38713/p491-hudson/. ACM or-

der number 472870.

Croft:1987:SOD

[174] W. B. Croft and D. W. Stemple. Supporting office document architectures with constrained types. In Dayal and Traiger [2258], pages 504–509. ISBN 0-89791-236-5. LCCN QA 76.9 D3

                                        P76 1987. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/38713/
                                        p504-croft/p504-croft.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/38713/p504-croft/. ACM order

number 472870.

Delgrande:1987:FBA

[175] J. P. Delgrande. Formal bounds on automatic generation and maintenance of integrity constraints. In ACM [2257], page ?? ISBN 0-89791-223-3. LCCN QA76.9.D3A296 1987. ACM Order Number 475870.

Chandra:1988:TDQ

[176] Ashok K. Chandra. Theory of database queries. In ACM [2260], pages 1–

9. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15 1988. URL http://

                                        www.acm.org/pubs/articles/proceedings/
                                        pods/308386/p1-chandra/p1-chandra.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/308386/
                                        p1-chandra/.

Swami:1988:OLJ

[177] Arun Swami and Anoop Gupta. Optimization of large join queries. In ACM [2259], pages 8–17. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/50202/
                                        p8-swami/p8-swami.pdf;   http://

www.acm.org/pubs/citations/proceedings/
mod/50202/p8-swami/. ACM order

number 472880.

Kuper:1988:EPL

[178] Gabriel M. Kuper. On the expressive power of logic programming languages with sets. In ACM [2260], pages 10–14. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p10-kuper/p10-kuper.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/308386/p10-kuper/

.

Shmueli:1988:RRC

[179] Oded Shmueli, Shalom Tsur, and Carlo Zaniolo. Rewriting of rules containing set terms in a logic data language LDL. In ACM [2260], pages 15–28. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p15-shmueli/p15-shmueli.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/308386/
p15-shmueli/.

Lohman:1988:GLF

[180] Guy M. Lohman. Grammar-like functional rules for representing query optimization alternatives. In ACM [2259], pages 18–27. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p18-lohman/p18-lohman.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/50202/p18-lohman/. ACM order

number 472880.

Muralikrishna:1988:EDM

[181] M. Muralikrishna and David J. De-Witt. Equi-depth multidimensional histograms. In ACM [2259], pages

28–36.  ISBN ????  LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/50202/
                                        p28-muralikrishna/p28-muralikrishna.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/50202/
                                        p28-muralikrishna/. ACM order

number 472880.

Paredaens:1988:PLU

[182] Jan Paredaens and Dirk Van Gucht. Possibilities and limitations of using flat operators in nested algebra expressions. In ACM [2260], pages 29–

38. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15 1988. URL http://

                                        www.acm.org/pubs/articles/proceedings/
                                        pods/308386/p29-paredaens/p29-paredaens.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/308386/
                                        p29-paredaens/.

Garza:1988:TMO

[183] Jorge F. Garza and Won Kim. Transaction management in an object-oriented database system. In ACM [2259], pages 37–45. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/50202/
                                        p37-garza/p37-garza.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/50202/p37-garza/. ACM order

number 472880.

Hull:1988:EPD

[184] Richard Hull and Jianwen Su. On the expressive power of database queries with intermediate types. In ACM [2260], pages 39–51. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p39-hull/p39-hull.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/308386/p39-hull/

.

Jagannathan:1988:SDS

[185] D. Jagannathan, B. L. Fritchman, R. L. Guck, J. P. Thompson, and D. M. Tolbert. SIM: a database system based on the semantic data model. In ACM [2259], pages 46–55. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p46-jagannathan/p46-jagannathan.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/50202/
p46-jagannathan/. ACM order num-

ber 472880.

Kifer:1988:AAD

[186] Michael Kifer, Raghu Ramakrishnan, and Avi Silberschatz. An axiomatic approach to deciding query safety in deductive databases. In ACM [2260], pages 52–60. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p52-kifer/p52-kifer.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/308386/p52-kifer/

.

Caruso:1988:CMO

[187] Michael Caruso and Edward Sciore. Contexts and metamessages in objectoriented database programming language design. In ACM [2259], pages 56–65. ISBN ???? LCCN

????  URL http://www.acm.org/

                                        pubs/articles/proceedings/mod/50202/
                                        p56-caruso/p56-caruso.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/50202/p56-caruso/. ACM order

number 472880.

Chomicki:1988:TDD

[188] Jan Chomicki and Tomasz Imieliński. Temporal deductive databases and infinite objects. In ACM [2260], pages 61–73. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15 1988. URL http://

                                        www.acm.org/pubs/articles/proceedings/
                                        pods/308386/p61-chomicki/p61-chomicki.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/308386/
                                        p61-chomicki/.

Laurent:1988:PSI

[189] D. Laurent and N. Spyratis. Partition semantics for incomplete information in relational databases. In ACM [2259], pages 66–73. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/50202/
                                        p66-laurent/p66-laurent.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/50202/p66-laurent/. ACM or-

der number 472880.

Ullman:1988:COS

[190] Jeffrey D. Ullman and Moshe Y. Vardi. The complexity of ordering subgoals. In ACM [2260], pages 74–81. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

                                        1988. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        308386/p74-ullman/p74-ullman.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/308386/p74-ullman/

.

Yuan:1988:SCQ

[191] Li Yan Yuan and Ding-An Chiang. A sound and complete query evaluation algorithm for relational databases with null values. In ACM [2259], pages 74–81. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p74-yuan/p74-yuan.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/50202/p74-yuan/. ACM order

number 472880.

Morris:1988:AOS

[192] Katherine A. Morris. An algorithm for ordering subgoals in NAIL? In ACM [2260], pages 82–88. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p82-morris/p82-morris.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/308386/p82-morris/

.

Malvestuto:1988:DPS

[193] F. M. Malvestuto. The derivation problem of summary data. In ACM [2259], pages 82–89. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p82-malvestuto/p82-malvestuto.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/50202/p82-malvestuto/

. ACM order number 472880.

Ramakrishnan:1988:OED

[194] Raghu Ramakrishnan, Catriel Beeri, and Ravi Krishnamurthy. Optimizing existential datalog queries. In ACM [2260], pages 89–102. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/

pubs/articles/proceedings/pods/
308386/p89-ramakrishnan/p89-ramakrishnan.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/308386/
p89-ramakrishnan/.

Alexander:1988:PDC

[195] W. Alexander and G. Copeland. Process and dataflow control in distributed data-intensive systems. In ACM [2259], pages 90–98. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p90-alexander/p90-alexander.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/50202/p90-alexander/

. ACM order number 472880.

Copeland:1988:DPB

[196] George Copeland, William Alexander, Ellen Boughter, and Tom Keller. Data placement in Bubba. In ACM [2259], pages 99–108. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p99-copeland/p99-copeland.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/50202/p99-copeland/

. ACM order number 472880.

Imielinski:1988:ECL

[197] Tomasz Imielinski and Shamim Naqvi. Explicit control of logic programs through rule algebra. In ACM [2260], pages 103–116. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p103-imielinski/p103-imielinski.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/308386/
p103-imielinski/.

Patterson:1988:CRA

[198] David A. Patterson, Garth Gibson, and Randy H. Katz. A case for redundant arrays of inexpensive disks (RAID). In ACM [2259], pages

109–116.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p109-patterson/p109-patterson.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/50202/p109-patterson/

. ACM order number 472880.

Kumar:1988:SBT

[199] Akhil Kumar and Michael Stonebraker. Semantics based transaction management techniques for replicated data. In ACM [2259], pages

117–125.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p117-kumar/p117-kumar.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/50202/p117-kumar/. ACM order

number 472880.

Ramakrishna:1988:ABD

[200] M. V. Ramakrishna and P. Mukhopadhyay. Analysis of bounded disorder file organization. In ACM [2260], pages 117–125. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p117-ramakrishna/p117-ramakrishna.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/308386/
p117-ramakrishna/.

ElAbbadi:1988:GPC

[201] Amr El Abbadi and Sam Toueg. The group paradigm for concurrency control. In ACM [2259], pages

126–134.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p126-el_abbadi/p126-el_abbadi.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/50202/p126-el_abbadi/

. ACM order number 472880.

Srivastava:1988:AMM

[202] Jaideep Srivastava and Doron Rotem. Analytical modeling of materialized view maintenance. In ACM [2260], pages 126–134. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p126-srivastava/p126-srivastava.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/308386/
p126-srivastava/.

Hadzilacos:1988:SGA

[203] Thanasis Hadzilacos. Serialization graph algorithms for multiversion concurrency control. In ACM [2260], pages 135–141. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p135-hadzilacos/p135-hadzilacos.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/308386/
p135-hadzilacos/.

Breitbart:1988:MUI

[204] Yuri Breitbart and Avi Silberschatz. Multidatabase update issues. In ACM [2259], pages

135–142.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p135-breitbart/p135-breitbart.pdf;
 http://www.acm.org/pubs/citations/

proceedings/mod/50202/p135-breitbart/

. ACM order number 472880.

Kelter:1988:QPD

[205] Udo Kelter. The queue protocol: a deadlock-free, homogeneous, nontwo-phase locking protocol. In ACM [2260], pages 142–151. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p142-kelter/p142-kelter.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/308386/
p142-kelter/.

Abiteboul:1988:DFD

[206] Serge Abiteboul and Richard Hull. Data functions, datalog and negation. In ACM [2259], pages 143–

153.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p143-abiteboul/p143-abiteboul.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/50202/p143-abiteboul/

. ACM order number 472880.

Banciihon:1988:OOD

[207] François Banciihon. Object-oriented database systems. In ACM [2260], pages 152–162. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p152-banciihon/p152-banciihon.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/308386/
p152-banciihon/.

Krishnamurthy:1988:FTS

[208] Ravi Krishnamurthy, Raghu Ramakrishnan, and Oded Shmueli. A framework for testing safety and effective computability of extended datalog. In ACM [2259], pages

154–163.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p154-krishnamurthy/p154-krishnamurthy.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/50202/
p154-krishnamurthy/. ACM order

number 472880.

Chan:1988:IRD

[209] Edward P. F. Chan and Hector J. Hernandez. Independence-reducible database schemes. In ACM [2260], pages 163–173. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p163-chan/p163-chan.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/308386/p163-chan/

.

Chen:1988:IMR

[210] Qiming Chen and Georges Gardarin. An implementation model for reasoning with complex objects. In ACM [2259], pages

164–172.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p164-chen/p164-chen.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/50202/p164-chen/. ACM order

number 472880.

Kim:1988:OFD

[211] Myoung Ho Kim and Sakti Pramanik. Optimal file distribution for partial match retrieval. In ACM [2259], pages 173–182. ISBN ???? LCCN

????  URL http://www.acm.org/

pubs/articles/proceedings/mod/50202/
p173-kim/p173-kim.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/50202/p173-kim/. ACM order

number 472880.

Hegner:1988:DRS

[212] Stephen J. Hegner. Decomposition of relational schemata into components defined by both projection and restriction. In ACM [2260], pages 174–183. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p174-hegner/p174-hegner.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/308386/
p174-hegner/.

Hutflesz:1988:TGF

[213] Andreas Hutflesz, Hans-Werner Six, and Peter Widmayer. Twin grid files: space optimizing access schemes. In ACM [2259], pages 183–190. ISBN ????

LCCN ???? URL http://www.acm.
org/pubs/articles/proceedings/mod/
50202/p183-hutflesz/p183-hutflesz.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/50202/

p183-hutflesz/. ACM order number 472880.

Batory:1988:CDS

[214] D. S. Batory. Concepts for a database system compiler. In ACM [2260], pages 184–192. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p184-batory/p184-batory.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/308386/
p184-batory/.

Ramakrishna:1988:HPA

[215] M. V. Ramakrishna. Hashing practice: analysis of hashing and universal hashing. In ACM [2259], pages 191–199. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/50202/
                                        p191-ramakrishna/p191-ramakrishna.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/50202/
                                        p191-ramakrishna/. ACM order num-

ber 472880.

Hadzilacos:1988:TSO

[216] Thanasis Hadzilacos and Vassos Hadzilacos. Transaction synchronisation in object bases. In ACM [2260], pages 193–200. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

                                        1988. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        308386/p193-hadzilacos/p193-hadzilacos.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/308386/
                                        p193-hadzilacos/.

Ioannidis:1988:DMD

[217] Yannis E. Ioannidis and Miron Livny. Data modeling in DELAB. In ACM [2259], page 200. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/50202/
                                        p200-ioannidis/p200-ioannidis.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/50202/p200-ioannidis/

. ACM order number 472880.

Ono:1988:DMT

[218] Kiyoshi Ono, Mikio Aoyama, and Hiroshi Fujimoto. Data management of telecommunications networks. In ACM [2259], page 201. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p201-ono/p201-ono.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/50202/p201-ono/. ACM order

number 472880.

Herlihy:1988:HCC

[219] Maurice P. Herlihy and William E. Weihl. Hybrid concurrency control for abstract data types. In ACM [2260], pages 201–210. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p201-herlihy/p201-herlihy.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/308386/
p201-herlihy/.

Alho:1988:DDM

[220] Kari Alho, Hannu Peltonen, Martti Mäntylä, and Rejio Sulonen. A design data manager. In ACM [2259], page 202. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p202-alho/p202-alho.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/50202/p202-alho/. ACM order

number 472880.

Naeymi-Rad:1988:RDD

[221] Frank Naeymi-Rad, Lowell Carmony, David Trace, Christine Georgakis, and Max Harry Weil. A relational database design in support of standard medical terminology in multidomain knowledge bases. In ACM [2259], page 203. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p203-naeymi-rad/p203-naeymi-rad.
pdf;   http://www.acm.org/pubs/

                                        citations/proceedings/mod/50202/
                                        p203-naeymi-rad/. ACM order num-

ber 472880.

Hernandez:1988:CCT

[222] Héctor J. Hernández and Edward P. F. Chan. A characterization of constant-time maintainability for BCNF database schemes. In ACM [2259], pages 209–217. ISBN ????

                                        LCCN ???? URL http://www.acm.
                                        org/pubs/articles/proceedings/mod/
                                        50202/p209-hernandez/p209-hernandez.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/50202/
                                        p209-hernandez/. ACM order number

472880.

Lanin:1988:CSM

[223] Vladimir Lanin and Dennis Shasha. Concurrent set manipulation without locking. In ACM [2260], pages 211–220. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

                                        1988. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        308386/p211-lanin/p211-lanin.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/308386/p211-lanin/

.

Leuchner:1988:PTA

[224] J. Leuchner, L. Miller, and G. Slutzki. A polynomial time algorithm for testing implications of a join dependency and embodied functional dependencies. In ACM [2259], pages

218–224.  ISBN ????  LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/50202/
                                        p218-leuchner/p218-leuchner.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/50202/p218-leuchner/

. ACM order number 472880.

VanGelder:1988:USW

[225] Allen Van Gelder, Kenneth Ross, and John S. Schlipf. Unfounded sets and well-founded semantics for general logic programs. In ACM [2260], pages 221–230. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p221-van_gelder/p221-van_
gelder.pdf;   http://www.acm.org/
pubs/citations/proceedings/pods/
308386/p221-van_gelder/.

Gyssens:1988:PAR

[226] Marc Gyssens and Dirk van Gucht. The powerset algebra as a result of adding programming constructs to the nested relational algebra. In ACM [2259], pages 225–232. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p225-gyssens/p225-gyssens.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/50202/p225-gyssens/

. ACM order number 472880.

Kolaitis:1988:WNF

[227] Phokion G. Kolaitis and Christos H. Papadimitriou. Why not negation by fixpoint? In ACM [2260], pages 231–239. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p231-kolaitis/p231-kolaitis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/308386/
p231-kolaitis/.

Mazumdar:1988:RTB

[228] Subhasish Mazumdar, David Stemple, and Tim Sheard. Resolving the tension between integrity and security using a theorem prover. In ACM [2259], pages 233–242. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p233-mazumdar/p233-mazumdar.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/50202/p233-mazumdar/

. ACM order number 472880.

Abiteboul:1988:PDD

[229] Serge Abiteboul and Victor Vianu. Procedural and declarative database update languages. In ACM [2260], pages 240–250. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p240-abiteboul/p240-abiteboul.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/308386/
p240-abiteboul/.

Qian:1988:TLD

[230] Xiaolei Qian and Richard Waldinger. A transaction logic for database specification. In ACM [2259], pages

243–250.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p243-qian/p243-qian.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/50202/p243-qian/. ACM order

number 472880.

Gadia:1988:GMR

[231] Shashi K. Gadia and Chuen-Sing Yeung. A generalized model for a relational temporal database. In ACM [2259], pages 251–259. ISBN ????

LCCN ???? URL http://www.acm.
org/pubs/articles/proceedings/mod/
50202/p251-gadia/p251-gadia.pdf;

http://www.acm.org/pubs/citations/
proceedings/mod/50202/p251-gadia/

. ACM order number 472880.

Naqvi:1988:DUL

[232] Shamim Naqvi and Ravi Krishnamurthy. Database updates in logic programming. In ACM [2260], pages 251–262. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p251-naqvi/p251-naqvi.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/308386/p251-naqvi/

.

Peinl:1988:HCS

[233] Peter Peinl, Andreas Reuter, and Harald Sammer. High contention in a stock trading database: a case study. In ACM [2259], pages

260–268.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p260-peinl/p260-peinl.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/50202/p260-peinl/. ACM order

number 472880.

Muralikrishna:1988:OMR

[234] M. Muralikrishna and David J. De-Witt. Optimization of multiple-relation multiple-disjunct queries. In ACM [2260], pages 263–275. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p263-muralikrishna/p263-
muralikrishna.pdf;   http://www.
acm.org/pubs/citations/proceedings/
pods/308386/p263-muralikrishna/.

Haynie:1988:DLD

[235] M. Haynie. A DBMS for large design automation databases. In ACM [2259], pages 269–276. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p269-haynie/p269-haynie.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/50202/p269-haynie/. ACM or-

der number 472880.

Hou:1988:SER

[236] Wen-Chi Hou, Gultekin Ozsoyoglu, and Baldeo K. Taneja. Statistical estimators for relational algebra expressions. In ACM [2260], pages 276–287. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p276-hou/p276-hou.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/308386/p276-hou/

.

Bell:1988:SDM

[237] Jean L. Bell. A specialized data management system for parallel execution of particle physics codes. In ACM [2259], pages 277–285. ISBN ????

LCCN ???? URL http://www.acm.
org/pubs/articles/proceedings/mod/
50202/p277-bell/p277-bell.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/50202/p277-bell/

. ACM order number 472880.

Christodoulakis:1988:PAF

[238] Stavros Christodoulakis and Daniel Alexander Ford. Performance analysis and fundamental performance tradeoffs for CLV optical disks. In ACM [2259], pages 286–294. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p286-christodoulakis/p286-christodoulakis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/50202/
p286-christodoulakis/. ACM order

number 472880.

Huang:1988:SSM

[239] Bing-Chao Huang and Michael A. Langston. Stable set and multiset operations in optimal time and space. In ACM [2260], pages 288–293. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p288-huang/p288-huang.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/308386/p288-huang/

.

Yu:1988:MTS

[240] Lin Yu and Daniel J. Rosenkrantz. Minimizing time-space cost for database version control. In ACM [2260], pages 294–301. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p294-yu/p294-yu.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
pods/308386/p294-yu/.

Hanson:1988:PQA

[241] Eric N. Hanson. Processing queries aganist database procedures: a performance analysis. In ACM [2259], pages 295–302. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p295-hanson/p295-hanson.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/50202/p295-hanson/. ACM or-

der number 472880.

Reiter:1988:WSD

[242] Raymond Reiter. What should a database know? In ACM [2260], pages 302–304. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

                                        1988. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        308386/p302-reiter/p302-reiter.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/308386/
                                        p302-reiter/.

Jarke:1988:MKA

[243] Matthias Jarke and Thomas Rose. Managing knowledge about information system evolution. In ACM [2259], pages 303–311. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/50202/
                                        p303-jarke/p303-jarke.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/50202/p303-jarke/. ACM order

number 472880.

Buneman:1988:SCO

[244] Peter Buneman, Susan Davidson, and Aaron Watters. A semantics for complex objects and approximate queries. In ACM [2260], pages 305–314. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

                                        1988. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        308386/p305-buneman/p305-buneman.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/308386/
                                        p305-buneman/.

Naughton:1988:CSR

[245] Jeffrey F. Naughton. Compiling separable recursions. In ACM [2259], pages 312–319. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/50202/

p312-naughton/p312-naughton.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/50202/p312-naughton/

. ACM order number 472880.

Winslett:1988:FCU

[246] Marianne Winslett. A framework for comparison of update semantics. In ACM [2260], pages 315–324. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p315-winslett/p315-winslett.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/308386/
p315-winslett/.

Youn:1988:CRF

[247] Cheong Youn, Lawrence J. Henschen, and Jiawei Han. Classification of recursive formulas in deductive databases. In ACM [2259], pages 320–328. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p320-youn/p320-youn.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/50202/p320-youn/. ACM order

number 472880.

Sippu:1988:GTC

[248] Seppo Sippu and Eljas Soisalon-Soininen. A generalized transitive closure for relational queries. In ACM [2260], pages 325–332. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p325-sippu/p325-sippu.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/308386/p325-sippu/

.

Wolfson:1988:DPL

[249] Ouri Wolfson and Avi Silberschatz. Distributed processing of logic programs. In ACM [2259], pages

329–336.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p329-wolfson/p329-wolfson.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/50202/p329-wolfson/

. ACM order number 472880.

Haddad:1988:CMC

[250] Ramsey W. Haddad and Jeffrey F. Naughton. Counting methods for cyclic relations. In ACM [2260], pages 333–340. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/
308386/p333-haddad/p333-haddad.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/308386/
p333-haddad/.

Group:1988:BNS

[251] Tandem Performance Group. A benchmark of non-stop SQL on the debit credit transaction. In ACM [2259], pages 337–341. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/citations/proceedings/mod/
50202/p337-tandem_performance_group/

. ACM order number 472880.

Vardi:1988:DUR

[252] Moshe Y. Vardi. Decidability and undecidability results for boundedness of linear recursive queries. In ACM [2260], pages 341–351. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15

1988. URL http://www.acm.org/
pubs/articles/proceedings/pods/

308386/p341-vardi/p341-vardi.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/308386/p341-vardi/

.

Borr:1988:HPS

[253] A. Borr. High performance SQL through low-level system integration. In ACM [2259], pages 342–

349.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p342-borr/p342-borr.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/50202/p342-borr/. ACM order

number 472880.

DeWitt:1988:PAG

[254] D. J. DeWitt, S. Ghandeharizadeh, and D. Schneider. A performance analysis of the gamma database machine. In ACM [2259], pages

350–360.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p350-dewitt/p350-dewitt.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/50202/p350-dewitt/. ACM or-

der number 472880.

Roesler:1988:SLM

[255] M. Roesler and W. A. Burkhard. Semantic lock models in object-oriented distributed systems and deadlock resolution. In ACM [2259], pages

361–370.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p361-roesler/p361-roesler.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/50202/p361-roesler/

. ACM order number 472880.

Ramarao:1988:CPD

[256] K. V. S. Ramarao. Commitment in a partitioned distributed database. In ACM [2259], pages

371–378.  ISBN ????  LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/50202/
                                        p371-ramarao/p371-ramarao.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/50202/p371-ramarao/

. ACM order number 472880.

Korth:1988:FMC

[257] H. K. Korth and G. Speegle. Formal model of correctness without serializability. In ACM [2259], pages

379–386.  ISBN ????  LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/50202/
                                        p379-korth/p379-korth.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/50202/p379-korth/. ACM order

number 472880.

Ramnarayan:1988:DKB

[258] R. Ramnarayan and H. Lu. A data/ knowledge base management testbed and experimental results on data/ knowledge base query and update processing. In ACM [2259], pages

387–395.  ISBN ????  LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/50202/
                                        p387-ramnarayan/p387-ramnarayan.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/50202/
                                        p387-ramnarayan/. ACM order num-

ber 472880.

Delcambre:1988:SCI

[259] L. M. L. Delcambre and J. N. Etheredge. A self-controlling interpreter for the relational production language. In ACM [2259], pages

396–403.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p396-delcambre/p396-delcambre.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/50202/p396-delcambre/

. ACM order number 472880.

Sellis:1988:ILP

[260] T. Sellis, C. C. Lin, and L. Raschid. Implementing large production systems in a DBMS environment: concepts and algorithms. In ACM [2259], pages 404–423. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p404-sellis/p404-sellis.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/50202/p404-sellis/. ACM or-

der number 472880.

Carey:1988:DMQ

[261] Michael J. Carey and David J. De-Witt. A data model and query language for EXODUS. In ACM [2259], pages 413–423. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/50202/
p413-carey/p413-carey.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/50202/p413-carey/. ACM order

number 472880.

Lecluse:1988:OOD

[262] C. Lecluse, P. Richard, and F. Velez. O_{2}, an object-oriented data model. In ACM [2259], pages 424–433. ISBN ????

LCCN ???? URL http://www.acm.
org/pubs/articles/proceedings/mod/
50202/p424-lecluse/p424-lecluse.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/50202/

p424-lecluse/. ACM order number

472880.

Borgida:1988:MCH

[263] A. Borgida. Modeling class hierarchies with contradictions. In ACM [2259], pages 434–443. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/50202/
                                        p434-borgida/p434-borgida.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/50202/p434-borgida/

. ACM order number 472880.

VanGelder:1989:AFL

[264] A. Van Gelder. The alternating fixpoint of logic programs with negation. In ACM [2261], pages 1–10. ISBN 0-89791-308-6. LCCN QA 76.9 D3

                                        A26 1989. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        73721/p1-van_gelder/p1-van_gelder.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/73721/
                                        p1-van_gelder/.

Salza:1989:ESQ

[265] Silvio Salza and Mario Terranova. Evaluating the size of queries on relational databases with non-uniform distribution and stochastic dependence. In Clifford et al. [2262], pages 8–14. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p8-salza/p8-salza.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/67544/p8-salza/.

Przymusinski:1989:ELP

[266] T. C. Przymusinski. Every logic program has a natural stratification and an iterated least fixed point model.

In ACM [2261], pages 11–21. ISBN 0-89791-308-6. LCCN QA 76.9 D3

A26 1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p11-przymusinski/p11-przymusinski.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/73721/
p11-przymusinski/;   http://www.
acm.org:80/pubs/citations/proceedings/
pods/73721/p11-przymusinski/.

Kolodner:1989:AGC

[267] Elliot Kolodner, Barbara Liskov, and William Weihl. Atomic garbage collection: managing a stable heap. In Clifford et al. [2262], pages

15–25.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p15-kolodner/p15-kolodner.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/67544/p15-kolodner/

.

Ross:1989:PSW

[268] K. A. Ross. A procedural semantics for well founded negation in logic programs. In ACM [2261], pages 22–33. ISBN 0-89791-308-

6. LCCN QA 76.9 D3 A26

1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p22-ross/p22-ross.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
pods/73721/p22-ross/;   http://
www.acm.org:80/pubs/citations/proceedings/
pods/73721/p22-ross/.

Dong:1989:DPD

[269] Guozhu Dong. On distributed processibility of datalog queries by decomposing databases. In Clifford et al. [2262], pages 26–35. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p26-dong/p26-dong.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/67544/p26-dong/.

Bry:1989:LPC

[270] F. Bry. Logic programming as constructivism: a formalization and its application to databases. In ACM [2261], pages 34–50. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

                                        1989. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        73721/p34-bry/p34-bry.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        pods/73721/p34-bry/;   http://www.
                                        acm.org:80/pubs/citations/proceedings/
                                        pods/73721/p34-bry/.

Agrawal:1989:OOD

[271] R. Agrawal and N. H. Gehani. ODE (Object Database and Environment): the language and the data model. In Clifford et al. [2262], pages 36–45. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p36-agrawal/p36-agrawal.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/67544/p36-agrawal/.

Ohori:1989:DPM

[272] Atsushi Ohori, Peter Buneman, and Val Breazu-Tannen. Database programming in Machiavelli — a polymorphic language with static type inference. In Clifford et al. [2262], pages 46–57. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p46-ohori/p46-ohori.pdf;   http:

//www.acm.org/pubs/citations/proceedings/
mod/67544/p46-ohori/.

Imielinski:1989:CQP

[273] T. Imielinski and K. Vadaparty. Complexity of query processing in databases with OR-objects. In ACM [2261], pages 51–65. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p51-imielinski/p51-imielinski.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/73721/
p51-imielinski/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/73721/p51-imielinski/.

Borgida:1989:CSD

[274] Alexander Borgida, Ronald J. Brachman, Deborah L. McGuinness, and Lori Alperin Resnick. CLASSIC: a structural data model for objects. In Clifford et al. [2262], pages 58–67. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p58-borgida/p58-borgida.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/67544/p58-borgida/.

Yuan:1989:SCQ

[275] L. Y. Yuan and D.-A. Chiang. A sound and complete query evaluation algorithm for relational databases with disjunctive information. In ACM [2261], pages 66–74. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p66-yuan/p66-yuan.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
pods/73721/p66-yuan/;   http://

                                        www.acm.org:80/pubs/citations/proceedings/
                                        pods/73721/p66-yuan/.

Hou:1989:PAR

[276] Wen-Chi Hou, Gultekin Ozsoyoglu, and Baldeo K. Taneja. Processing aggregate relational queries with hard time constraints. In Clifford et al. [2262], pages 68–77. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p68-hou/p68-hou.pdf;   http://www.
                                        acm.org/pubs/citations/proceedings/
                                        mod/67544/p68-hou/.

Grahne:1989:HTE

[277] G. Grahne. Horn tables — an efficient tool for handling incomplete information in databases. In ACM [2261], pages 75–82. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

                                        1989. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        73721/p75-grahne/p75-grahne.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/73721/p75-grahne/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/73721/
                                        p75-grahne/.

Jagadish:1989:IHR

[278] H. V. Jagadish. Incorporating hierarchy in a relational model of data. In Clifford et al. [2262], pages 78–87. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p78-jagadish/p78-jagadish.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/67544/p78-jagadish/

.

Vardi:1989:ITA

[279] M. Y. Vardi. Invited talk: automata theory for database theoreticians. In ACM [2261], pages 83–92. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p83-vardi/p83-vardi.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/73721/p83-vardi/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/73721/
p83-vardi/.

Cammarata:1989:ERD

[280] Stephanie Cammarata, Prasadram Ramachandra, and Darrell Shane. Extending a relational database with deferred referential integrity checking and intelligent joins. In Clifford et al. [2262], pages 88–97. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p88-cammarata/p88-cammarata.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/67544/p88-cammarata/

.

Manchanda:1989:DED

[281] S. Manchanda. Declarative expression of deductive database updates. In ACM [2261], pages 93–100. ISBN 0-89791-308-6. LCCN QA 76.9 D3

A26 1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p93-manchanda/p93-manchanda.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/73721/
p93-manchanda/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/73721/p93-manchanda/.

Copeland:1989:CHA

[282] George Copeland and Tom Keller. A comparison of high-availability media recovery techniques. In Clifford et al. [2262], pages 98–

109.  ISBN ????  LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p98-copeland/p98-copeland.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/67544/p98-copeland/

.

Atzeni:1989:UDW

[283] P. Atzeni and R. Torlone. Updating databases in the weak instance model. In ACM [2261], pages 101–109. ISBN 0-89791-308-6. LCCN QA 76.9 D3

                                        A26 1989. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        73721/p101-atzeni/p101-atzeni.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/73721/p101-atzeni/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/73721/
                                        p101-atzeni/.

Tay:1989:AA

[284] Y. C. Tay. Attribute agreement. In ACM [2261], pages 110–119. ISBN 0-89791-308-6. LCCN QA 76.9 D3

                                        A26 1989. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        73721/p110-tay/p110-tay.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        pods/73721/p110-tay/;   http://
                                        www.acm.org:80/pubs/citations/proceedings/
                                        pods/73721/p110-tay/.

Schneider:1989:PEF

[285] Donovan A. Schneider and David J. DeWitt. A performance evaluation of four parallel join algorithms in a shared-nothing multiprocessor environment. In Clifford et al. [2262], pages 110–121. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p110-schneider/p110-schneider.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/67544/p110-schneider/

.

Wang:1989:CCT

[286] K. Wang. Can constant-time-maintainability be more practical? In ACM [2261], pages 120–127. ISBN 0-89791-308-6. LCCN QA 76.9 D3

A26 1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p120-wang/p120-wang.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/73721/p120-wang/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/73721/
p120-wang/.

Carey:1989:PCC

[287] Michael J. Carey and Miron Livny. Parallelism and concurrency control performance in distributed database machines. In Clifford et al. [2262], pages 122–133. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p122-carey/p122-carey.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/67544/p122-carey/.

Mannila:1989:PAF

[288] H. Mannila and K.-J. Raiha. Practical algorithms for finding prime attributes and testing normal forms. In ACM [2261], pages 128–133. ISBN 0-89791-308-6. LCCN QA 76.9 D3

A26 1989. URL http://www.acm.org/

                                        pubs/articles/proceedings/pods/
                                        73721/p128-mannila/p128-mannila.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/73721/
                                        p128-mannila/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/73721/p128-mannila/.

Elkan:1989:DPC

[289] C. Elkan. A decision procedure for conjunctive query disjointness. In ACM [2261], pages 134–139. ISBN 0-89791-308-6. LCCN QA 76.9 D3

                                        A26 1989. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        73721/p134-elkan/p134-elkan.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/73721/p134-elkan/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/73721/
                                        p134-elkan/.

Kifer:1989:FLH

[290] Michael Kifer and Georg Lausen. Flogic: a higher-order language for reasoning about objects, inheritance, and scheme. In Clifford et al. [2262], pages 134–146. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p134-kifer/p134-kifer.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/67544/p134-kifer/.

Ullman:1989:BBT

[291] J. D. Ullman. Bottom-up beats top-down for datalog. In ACM [2261], pages 140–149. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

                                        1989. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        73721/p140-ullman/p140-ullman.pdf;
                                         http://www.acm.org/pubs/citations/

proceedings/pods/73721/p140-ullman/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/73721/
p140-ullman/.

Hull:1989:AOO

[292] Richard Hull and Jianwen Su. On accessing object-oriented databases: expressive power, complexity, and restrictions. In Clifford et al. [2262], pages 147–158. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p147-hull/p147-hull.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/67544/p147-hull/.

Seki:1989:PAT

[293] H. Seki. On the power of Alexander templates. In ACM [2261], pages 150–159. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p150-seki/p150-seki.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/73721/p150-seki/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/73721/
p150-seki/.

Abiteboul:1989:OIQ

[294] Serge Abiteboul and Paris C. Kanellakis. Object identity as a query language primitive. In Clifford et al. [2262], pages 159–173. ISBN ????

LCCN ???? URL http://www.acm.
org/pubs/articles/proceedings/mod/
67544/p159-abiteboul/p159-abiteboul.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/67544/
p159-abiteboul/.

Sagiv:1989:SDQ

[295] Y. Sagiv and M. Y. Vardi. Safety of datalog queries over infinite databases. In ACM [2261], pages 160–171. ISBN 0-89791-308-6. LCCN QA 76.9 D3

                                        A26 1989. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        73721/p160-sagiv/p160-sagiv.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/73721/p160-sagiv/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/73721/
                                        p160-sagiv/.

Ramakrishnan:1989:PTT

[296] R. Ramakrishnan, Y. Sagiv, J. D. Ullman, and M. Y. Vardi. Prooftree transformation theorems and their applications. In ACM [2261], pages 172–181. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

                                        1989. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        73721/p172-ramakrishnan/p172-ramakrishnan.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/73721/
                                        p172-ramakrishnan/;   http://www.
                                        acm.org:80/pubs/citations/proceedings/
                                        pods/73721/p172-ramakrishnan/.

Chomicki:1989:RSI

[297] Jan Chomicki and Tomasz Imieliński. Relational specifications of infinite query answers. In Clifford et al. [2262], pages 174–183. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p174-chomicki/p174-chomicki.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/67544/p174-chomicki/

.

Saraiya:1989:LNR

[298] Y. P. Saraiya. Linearising nonlinear recursions in polynomial time. In ACM [2261], pages 182–189. ISBN 0-89791-308-6. LCCN QA 76.9 D3

A26 1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p182-saraiya/p182-saraiya.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/73721/
p182-saraiya/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/73721/p182-saraiya/.

Sun:1989:SIP

[299] Xian-He Sun, Nabil Kamel, and Lionel M. Ni. Solving implication problems in database applications. In Clifford et al. [2262], pages 185–192. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p185-sun/p185-sun.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/67544/p185-sun/.

Brodsky:1989:IMC

[300] A. Brodsky and Y. Sagiv. Inference of monotonicity constraints in datalog programs. In ACM [2261], pages 190–199. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p190-brodsky/p190-brodsky.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/73721/
p190-brodsky/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/73721/p190-brodsky/.

Bry:1989:TEE

[301] François Bry. Towards an efficient evaluation of general queries: quantifier and disjunction processing revisited. In Clifford et al. [2262], pages 193–204. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p193-bry/p193-bry.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/67544/p193-bry/.

Cohen:1989:WSP

[302] S. Cohen and O. Wolfson. Why a single parallelization strategy is not enough in knowledge bases. In ACM [2261], pages 200–216. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

                                        1989. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        73721/p200-cohen/p200-cohen.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/73721/p200-cohen/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/73721/
                                        p200-cohen/.

Ioannidis:1989:CRR

[303] Yannis E. Ioannidis and Timos K. Sellis. Conflict resolution of rules assigning values to virtual attributes. In Clifford et al. [2262], pages

205–214.  ISBN ????  LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p205-ioannidis/p205-ioannidis.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/67544/p205-ioannidis/

.

McCarthy:1989:AAD

[304] Dennis McCarthy and Umeshwar Dayal. The architecture of an active database management system. In Clifford et al. [2262], pages

215–224.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p215-mccarthy/p215-mccarthy.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/67544/p215-mccarthy/

.

Sector:1989:ITM

[305] A. Z. Sector. Invited talk: modular architectures for distributed and databases systems. In ACM [2261], pages 217–224. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

1989. URL http://www.acm.org:80/
pubs/citations/proceedings/pods/
73721/p217-sector/.

Spector:1989:ITM

[306] A. Z. Spector. Invited talk: modular architectures for distributed and databases systems. In ACM [2261], pages 217–224. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p217-spector/p217-spector.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/73721/
p217-spector/.

Cohen:1989:CCD

[307] D. Cohen. Compiling complex database transition triggers. In Clifford et al. [2262], pages 225–

234.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p225-cohen/p225-cohen.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/67544/p225-cohen/.

Rotem:1989:CMH

[308] D. Rotem. Clustered multiattribute hash files. In ACM [2261], pages 225–234. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

                                        1989. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        73721/p225-rotem/p225-rotem.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/73721/p225-rotem/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/73721/
                                        p225-rotem/.

Naughton:1989:EER

[309] J. F. Naughton, R. Ramakrishnan, Y. Sagiv, and J. D. Ullman. Efficient evaluation of right-, left-, and multilinear rules. In Clifford et al. [2262], pages 235–242. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p235-naughton/p235-naughton.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/67544/p235-naughton/

.

Johnson:1989:UBT

[310] T. Johnson and D. Shasha. Utilization of B-trees with inserts, deletes and modifies. In ACM [2261], pages 235–246. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

                                        1989. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        73721/p235-johnson/p235-johnson.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/73721/
                                        p235-johnson/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/73721/p235-johnson/.

Larson:1989:FSS

[311] P.-A. Larson and V. Deshpande. A file structure supporting traversal recursion. In Clifford et al. [2262], pages 243–252. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p243-larson/p243-larson.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/67544/p243-larson/.

Faloutsos:1989:FSK

[312] C. Faloutsos and S. Roseman. Fractals for secondary key retrieval. In ACM [2261], pages 247–252. ISBN 0-89791-308-6. LCCN QA 76.9 D3

A26 1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p247-faloutsos/p247-faloutsos.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/73721/
p247-faloutsos/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/73721/p247-faloutsos/.

Faloutsos:1989:DUE

[313] C. Faloutsos and D. Metaxas. Declustering using error correcting codes. In ACM [2261], pages 253–258. ISBN 0-89791-308-6. LCCN QA 76.9 D3

A26 1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p253-faloutsos/p253-faloutsos.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/73721/
p253-faloutsos/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/73721/p253-faloutsos/.

Agrawal:1989:EMT

[314] R. Agrawal, A. Borgida, and H. V. Jagadish. Efficient management of transitive relationships in large data and knowledge bases. In Clifford et al. [2262], pages 253–

262.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p253-agrawal/p253-agrawal.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/67544/p253-agrawal/

.

Weihl:1989:IRC

[315] W. E. Weihl. The impact of recovery on concurrency control. In ACM [2261], pages 259–269. ISBN 0-89791-308-6. LCCN QA 76.9 D3

A26 1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p259-weihl/p259-weihl.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/73721/p259-weihl/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/73721/
p259-weihl/.

Gyssens:1989:GBA

[316] M. Gyssens, J. Paredaens, and D. van Gucht. A grammar-based approach towards unifying hierarchical data models. In Clifford et al. [2262], pages 263–272. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p263-gyssens/p263-gyssens.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/67544/p263-gyssens/

.

Fu:1989:CCN

[317] A. Fu and T. Kameda. Concurrency control of nested transactions accessing B-trees. In ACM [2261], pages 270–285. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p270-fu/p270-fu.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
pods/73721/p270-fu/;   http://www.
acm.org:80/pubs/citations/proceedings/
pods/73721/p270-fu/.

Colby:1989:RAQ

[318] Latha S. Colby. A recursive algebra and query optimization for nested relations. In Clifford et al. [2262], pages 273–283. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p273-colby/p273-colby.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/67544/p273-colby/.

Tansel:1989:NHR

[319] A. U. Tansel and L. Garnett. Nested historical relations. In Clifford et al. [2262], pages 284–

294.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p284-tansel/p284-tansel.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/67544/p284-tansel/.

Bonner:1989:HDN

[320] A. J. Bonner. Hypothetical datalog negation and linear recursion. In ACM [2261], pages 286–300. ISBN 0-89791-308-6. LCCN QA 76.9 D3

A26 1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p286-bonner/p286-bonner.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/73721/p286-bonner/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/73721/
p286-bonner/.

Orenstein:1989:RSD

[321] J. A. Orenstein. Redundancy in spatial databases. In Clifford et al. [2262], pages 295–305. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p295-orenstein/p295-orenstein.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/67544/p295-orenstein/

.

Lakshmanan:1989:IPG

[322] V. S. Lakshmanan and A. O. Mendelzon. Inductive pebble games and the expressive power of datalog. In ACM [2261], pages 301–310. ISBN 0-89791-308-6. LCCN QA 76.9 D3

                                        A26 1989. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        73721/p301-lakshmanan/p301-lakshmanan.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/73721/
                                        p301-lakshmanan/;   http://www.
                                        acm.org:80/pubs/citations/proceedings/
                                        pods/73721/p301-lakshmanan/.

Christodoulakis:1989:RPV

[323] Stavros Christodoulakis and Daniel Alexander Ford. Retrieval performance versus disc space utilization on WORM optical discs. In Clifford et al. [2262], pages 306–314. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p306-christodoulakis/p306-christodoulakis.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/67544/
                                        p306-christodoulakis/.

Cosmadakis:1989:FOE

[324] S. S. Cosmadakis. On the first-order expressibility of recursive queries. In ACM [2261], pages 311–323. ISBN 0-89791-308-6. LCCN QA 76.9 D3

A26 1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p311-cosmadakis/p311-cosmadakis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/73721/
p311-cosmadakis/;   http://www.
acm.org:80/pubs/citations/proceedings/
pods/73721/p311-cosmadakis/.

Lomet:1989:AMM

[325] David Lomet and Betty Salzberg. Access methods for multiversion data. In Clifford et al. [2262], pages

315–324.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p315-lomet/p315-lomet.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/67544/p315-lomet/.

Dublish:1989:EBA

[326] P. Dublish and S. N. Maheshwari. Expressibility of bounded-arity fixedpoint query hierarchies. In ACM [2261], pages 324–335. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p324-dublish/p324-dublish.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/73721/
p324-dublish/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/73721/p324-dublish/.

Shekita:1989:PET

[327] Eugene J. Shekita and Michael J. Carey. Performance enhancement through replication in an objectoriented DBMS. In Clifford et al. [2262], pages 325–336. ISBN ????

LCCN ???? URL http://www.acm.

                                        org/pubs/articles/proceedings/mod/
                                        67544/p325-shekita/p325-shekita.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/67544/
                                        p325-shekita/.

Kedem:1989:RDB

[328] Z. M. Kedem and A. Tuzhilin. Relational database behavior: utilizing relational discrete event systems and models. In ACM [2261], pages 336–346. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

                                        1989. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        73721/p336-kedem/p336-kedem.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/73721/p336-kedem/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/73721/
                                        p336-kedem/.

Kim:1989:COR

[329] Won Kim, Elisa Bertino, and Jorge F. Garza. Composite objects revisited. In Clifford et al. [2262], pages

337–347.  ISBN ????  LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p337-kim/p337-kim.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/67544/p337-kim/.

Hull:1989:USI

[330] R. Hull and J. Su. Untyped sets, invention, and computable queries. In ACM [2261], pages 347–359. ISBN 0-89791-308-6. LCCN QA 76.9 D3

                                        A26 1989. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        73721/p347-hull/p347-hull.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/pods/73721/p347-hull/

;   http://www.acm.org:80/pubs/
citations/proceedings/pods/73721/
p347-hull/.

Chang:1989:EIS

[331] E. E. Chang and R. H. Katz. Exploiting inheritance and structure semantics for effective clustering and buffering in an object-oriented DBMS. In Clifford et al. [2262], pages

348–357.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p348-chang/p348-chang.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/67544/p348-chang/.

Graefe:1989:DQE

[332] G. Graefe and K. Ward. Dynamic query evaluation plans. In Clifford et al. [2262], pages 358–

366.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p358-graefe/p358-graefe.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/67544/p358-graefe/.

Lecluse:1989:MCS

[333] C. Lécluse and P. Richard. Modeling complex structures in objectoriented logic programming. In ACM [2261], pages 360–368. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p360-lecluse/p360-lecluse.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/73721/
p360-lecluse/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/73721/p360-lecluse/.

Swami:1989:OLJ

[334] A. Swami. Optimization of large join queries: combining heuristics and combinatorial techniques. In Clifford et al. [2262], pages 367–

376.  ISBN ????  LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p367-swami/p367-swami.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/67544/p367-swami/.

Chen:1989:CLC

[335] W. Chen and D. S. Warren. Clogic of complex objects. In ACM [2261], pages 369–378. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

                                        1989. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        73721/p369-chen/p369-chen.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/pods/73721/p369-chen/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/73721/
                                        p369-chen/.

Haas:1989:EQP

[336] L. M. Haas, J. C. Freytag, G. M. Lohman, and H. Pirahesh. Extensible query processing in starburst. In Clifford et al. [2262], pages 377–388. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p377-haas/p377-haas.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/67544/p377-haas/.

Kifer:1989:LOO

[337] M. Kifer and J. Wu. A logic for object-oriented logic programming. In ACM [2261], pages 379–393. ISBN 0-89791-308-6. LCCN QA 76.9 D3

A26 1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p379-kifer/p379-kifer.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/73721/p379-kifer/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/73721/
p379-kifer/.

Tang:1989:SPS

[338] T. Tang and N. Natarajan. A static pessimistic scheme for handling replicated databases. In Clifford et al. [2262], pages 389–

398.  ISBN ????  LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p389-tang/p389-tang.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/67544/p389-tang/.

Borgida:1989:TSQ

[339] A. Borgida. Type systems for querying class hierarchies with nonstrict inheritance. In ACM [2261], pages 394–400. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26

1989. URL http://www.acm.org/
pubs/articles/proceedings/pods/
73721/p394-borgida/p394-borgida.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/73721/
p394-borgida/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/73721/p394-borgida/.

Ellis:1989:CCG

[340] C. A. Ellis and S. J. Gibbs. Concurrency control in groupware systems. In Clifford et al. [2262], pages

399–407.  ISBN ????  LCCN

????  URL http://www.acm.org/

                                        pubs/articles/proceedings/mod/67544/
                                        p399-ellis/p399-ellis.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/67544/p399-ellis/.

Agrawal:1989:MSM

[341] D. Agrawal and S. Sengupta. Modular synchronization in multiversion databases: version control and concurrency control. In Clifford et al. [2262], pages 408–417. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p408-agrawal/p408-agrawal.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/67544/p408-agrawal/

.

DeTroyer:1989:RTC

[342] O. De Troyer. RIDL*: a tool for the computer-assisted engineering of large databases in the presence of integrity constraints. In Clifford et al. [2262], pages 418–429. ISBN ???? LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p418-de_troyer/p418-de_troyer.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/67544/p418-de_troyer/

.

Markowitz:1989:CRE

[343] Victor M. Markowitz and Arie Shoshani. On the correctness of representing extended entity-relationship structures in the relational model. In Clifford et al. [2262], pages

430–439.  ISBN ????  LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/67544/
                                        p430-markowitz/p430-markowitz.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/67544/p430-markowitz/

.

Navathe:1989:VPD

[344] Shamkant B. Navathe and Mingyoung Ra. Vertical partitioning for database design: a graphical algorithm. In Clifford et al. [2262], pages 440–450. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/67544/
p440-navathe/p440-navathe.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/67544/p440-navathe/

.

Lipton:1990:PSE

[345] Richard J. Lipton, Jeffrey F. Naughton, and Donovan A. Schneider. Practical selectivity estimation through adaptive sampling. In Garcia-Molina and Jagadish [2263], pages 1–11. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p1-lipton/p1-lipton.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/93597/p1-lipton/.

Kim:1990:RDO

[346] Won Kim. Research directions in object-oriented database systems. In ACM [2264], pages 1–

15. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p1-kim/p1-kim.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
pods/298514/p1-kim/. A few papers

from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

King:1990:BAT

[347] Roger King and Ali Morfeq. Bayan: an Arabic text database management system. In Garcia-Molina and Jagadish [2263], pages 12–23. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/93597/
                                        p12-king/p12-king.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/93597/p12-king/.

Abiteboul:1990:MS

[348] Serge Abiteboul, Paris C. Kanellakis, and Emmanuel Waller. Method schemas. In ACM [2264], pages 16–27. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

                                        1990. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        298514/p16-abiteboul/p16-abiteboul.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/298514/

p16-abiteboul/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Yu:1990:RDO

[349] Lin Yu and Daniel J. Rosenkrantz. Representability of design objects by ancestor-controlled hierarchical specifications. In ACM [2264], pages 28–39. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

                                        1990. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        298514/p28-yu/p28-yu.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/

pods/298514/p28-yu/. A few papers

from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Agrawal:1990:OGI

[350] R. Agrawal, N. H. Gehani, and J. Srinivasan. OdeView: the graphical interface to Ode. In Garcia-Molina and Jagadish [2263], pages 34–43. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p34-agrawal/p34-agrawal.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/93597/p34-agrawal/.

Lipton:1990:QSE

[351] Richard J. Lipton and Jeffrey F. Naughton. Query size estimation by adaptive sampling (extended abstract). In ACM [2264], pages 40–

46. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p40-lipton/p40-lipton.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/298514/p40-lipton/

. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Ullman:1990:IOC

[352] Jeffrey D. Ullman and Mihalis Yannakakis. The input/output complexity of transitive closure. In Garcia-Molina and Jagadish [2263], pages 44–53. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p44-ullman/p44-ullman.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/93597/p44-ullman/.

VanGelder:1990:DCA

[353] Allen Van Gelder. Deriving constraints among argument sizes in logic programs (extended abstract). In ACM [2264], pages 47–60. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3

A26 1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p47-van_gelder/p47-van_gelder.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/
p47-van_gelder/. A few papers from

this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Agrawal:1990:CSL

[354] D. Agrawal and A. El Abbadi. Constrained shared locks for increasing concurrency in databases. In ACM [2264], pages 53–63. ISBN 0-89791-352-

3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26 1990. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Shen:1990:IEE

[355] Yeh-Heng Shen. IDLOG: extending the expressive power of deductive database languages. In Garcia-Molina and Jagadish [2263], pages 54–63. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p54-shen/p54-shen.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/93597/p54-shen/.

Kolaitis:1990:EPD

[356] Phokion G. Kolaitis and Moshe Y. Vardi. On the expressive power of Datalog: Tools and a case study. In ACM [2264], pages 61–71. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3

A26 1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p61-kolaitis/p61-kolaitis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/

p61-kolaitis/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Saraiya:1990:HPS

[357] Yatin P. Saraiya. Hard problems for simple logic programs. In Garcia-Molina and Jagadish [2263], pages 64–73. ISBN 0-89791-365-

5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p64-saraiya/p64-saraiya.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/93597/p64-saraiya/.

Carey:1990:LCL

[358] Michael J. Carey, Sanjay Krishnamurthi, and Miron Livny. Load control for locking: the ’half-and-half’ approach. In ACM [2264], pages 72–84. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

                                        1990. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        298514/p72-carey/p72-carey.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/298514/p72-carey/

. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Wang:1990:PTD

[359] Ke Wang. Polynomial time designs toward both BCNF and efficient data manipulation. In Garcia-Molina and Jagadish [2263], pages 74–83. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/93597/
                                        p74-wang/p74-wang.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/93597/p74-wang/.

Atzeni:1990:EUI

[360] Paolo Atzeni and Riccardo Torlone. Efficient updates to independent schemes in the weak instance model. In Garcia-Molina and Jagadish [2263], pages 84–93. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/93597/
                                        p84-atzeni/p84-atzeni.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/93597/p84-atzeni/.

Agrawal:1990:LCS

[361] Divyakant Agrawal and Amr El Abbadi. Locks with constrained sharing (extended abstract). In ACM [2264], pages 85–93. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3

A26 1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p85-agrawal/p85-agrawal.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/

p85-agrawal/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Saraiya:1990:ETD

[362] Y. P. Saraiya. On the efficiency of transforming database logic programs. In ACM [2264], pages 87–109. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26 1990. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Salzberg:1990:FDS

[363] Betty Salzberg, Alex Tsukerman, Jim Gray, Michael Stuewart, Susan Uren, and Bonnie Vaughan. FastSort: a distributed single-input single-output external sort. In Garcia-Molina and Jagadish [2263], pages 94–101. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p94-salzberg/p94-salzberg.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/93597/p94-salzberg/

.

Fekete:1990:SGC

[364] Alan Fekete, Nancy Lynch, and William E. Weihl. A serialization graph construction for nested transactions. In ACM [2264], pages 94–

108. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p94-fekete/p94-fekete.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/298514/p94-fekete/

. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Graefe:1990:EPV

[365] Goetz Graefe. Encapsulation of parallelism in the Volcano query processing system. In Garcia-Molina and Jagadish [2263], pages 102–111. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p102-graefe/p102-graefe.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/93597/p102-graefe/.

Weikum:1990:MLR

[366] Gerhard Weikum, Christof Hasse, Peter Broessler, and Peter Muth. Multilevel recovery. In ACM [2264], pages 109–123. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p109-weikum/p109-weikum.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/

p109-weikum/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Bernstein:1990:IRR

[367] Philip A. Bernstein, Meichun Hsu, and Bruce Mann. Implementing recoverable requests using queues. In Garcia-Molina and Jagadish [2263], pages 112–122. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p112-bernstein/p112-bernstein.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/93597/p112-bernstein/

.

Solworth:1990:WOD

[368] Jon A. Solworth and Cyril U. Orji. Write-only disk caches. In Garcia-Molina and Jagadish [2263], pages 123–132. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p123-solworth/p123-solworth.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/93597/p123-solworth/

.

Tay:1990:OSM

[369] Y. C. Tay. On the optimality of strategies for multiple joins. In ACM [2264], pages 124–131. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3

A26 1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p124-tay/p124-tay.pdf;
http://www.acm.org/pubs/citations/

proceedings/pods/298514/p124-tay/

. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Saraiya:1990:PTP

[370] Yatin P. Saraiya. Polynomial-time program transformations in deductive databases. In ACM [2264], pages 132–144. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p132-saraiya/p132-saraiya.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/

p132-saraiya/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Wolfson:1990:NPP

[371] Ouri Wolfson and Aya Ozeri. A new paradigm for parallel and distributed rule-processing. In Garcia-Molina and Jagadish [2263], pages 133–142. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p133-wolfson/p133-wolfson.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/93597/p133-wolfson/

.

Ganguly:1990:FPP

[372] Sumit Ganguly, Avi Silberschatz, and Shalom Tsur. A framework for the parallel processing of Datalog queries. In Garcia-Molina and Jagadish [2263], pages 143–152. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p143-ganguly/p143-ganguly.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/93597/p143-ganguly/

.

Plambeck:1990:STR

[373] Thane Plambeck. Semigroup techniques in recursive query optimization. In ACM [2264], pages 145–

153. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p145-plambeck/p145-plambeck.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/

p145-plambeck/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Kogan:1990:CCM

[374] Boris Kogan and S. Jajodia. Concurrency control in multilevel-secure databases based on replicated architecture. In Garcia-Molina and Jagadish [2263], pages 153–162. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p153-kogan/p153-kogan.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/93597/p153-kogan/.

Elkan:1990:ILD

[375] Charles Elkan. Independence of logic database queries and updates. In ACM [2264], pages 154–160. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3

                                        A26 1990. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        298514/p154-elkan/p154-elkan.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/298514/p154-elkan/

. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Ross:1990:MSM

[376] Kenneth A. Ross. Modular stratification and magic sets for datalog programs with negation. In ACM [2264], pages 161–171. ISBN 0-89791-352-

3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

                                        1990. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        298514/p161-ross/p161-ross.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/298514/p161-ross/

. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Badrinath:1990:PES

[377] B. R. Badrinath and Krithi Ramamritham. Performance evaluation of semantics-based multilevel concurrency control protocols. In Garcia-Molina and Jagadish [2263], pages 163–172. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/93597/
                                        p163-badrinath/p163-badrinath.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/93597/p163-badrinath/

.

You:1990:TVF

[378] Jia-Huai You and Li Yan Yuan. Threevalued formalization of logic programming: is it needed? In ACM [2264], pages 172–182. ISBN 0-89791-352-

3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p172-you/p172-you.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/298514/p172-you/

. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Motro:1990:QDK

[379] Amihai Motro and Qiuhui Yuan. Querying database knowledge. In Garcia-Molina and Jagadish [2263], pages 173–183. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p173-motro/p173-motro.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/93597/p173-motro/.

Royer:1990:BCE

[380] Véronique Royer. Backward chaining evaluation in stratified disjunctive theories. In ACM [2264], pages 183–195. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p183-royer/p183-royer.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/298514/p183-royer/

. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Laenens:1990:ELP

[381] Els Laenens, Domenico Sacca, and Dirk Vermeir. Extending logic programming. In Garcia-Molina and Jagadish [2263], pages 184–193. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/93597/
                                        p184-laenens/p184-laenens.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/93597/p184-laenens/

.

Chrysanthis:1990:AFS

[382] Panayiotis K. Chrysanthis and Krithi Ramamritham. ACTA: a framework for specifying and reasoning about transaction structure and behavior. In Garcia-Molina and Jagadish [2263], pages 194–203. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/93597/
                                        p194-chrysanthis/p194-chrysanthis.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/93597/
                                        p194-chrysanthis/.

Schlipf:1990:EPL

[383] John S. Schlipf. The expressive powers of the logic programming semantics (extended abstract). In ACM [2264], pages 196–204. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3

                                        A26 1990. URL http://www.acm.org/

pubs/articles/proceedings/pods/
298514/p196-schlipf/p196-schlipf.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/

p196-schlipf/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Dayal:1990:OLR

[384] Umeshwar Dayal, Meichun Hsu, and Rivka Ladin. Organizing long-running activities with triggers and transactions. In Garcia-Molina and Jagadish [2263], pages 204–214. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p204-dayal/p204-dayal.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/93597/p204-dayal/.

Sacca:1990:SMN

[385] Domenico Sacca and Carlo Zaniolo. Stable models and nondeterminism in logic programs with negation. In ACM [2264], pages 205–217. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3

A26 1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p205-sacca/p205-sacca.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/298514/p205-sacca/

. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Breitbart:1990:RTM

[386] Yuri Breitbart, Avi Silberschatz, and Glenn R. Thompson. Reliable transaction management in a multidatabase system. In Garcia-Molina and Jagadish [2263], pages 215–224. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p215-breitbart/p215-breitbart.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/93597/p215-breitbart/

.

Abiteboul:1990:NDL

[387] Serge Abiteboul, Eric Simon, and Victor Vianu. Non-deterministic languages to express deterministic transformations. In ACM [2264], pages 218–229. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p218-abiteboul/p218-abiteboul.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/
p218-abiteboul/. A few papers from

this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Abiteboul:1990:NLE

[388] S. Abiteboul, E. Simon, and V. Vianu. Nondeterministic languages to express deterministic transformations. In ACM [2264], pages 218–229. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26 1990. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Cacace:1990:IOO

[389] F. Cacace, S. Ceri, S. Crespi-Reghizzi, L. Tanca, and R. Zicari. Integrating object-oriented data modelling with a rule-based programming paradigm. In Garcia-Molina and Jagadish [2263], pages 225–236. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p225-cacace/p225-cacace.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/93597/p225-cacace/.

Yannakakis:1990:GTM

[390] Mihalis Yannakakis. Graph-theoretic methods in database theory. In ACM [2264], pages 230–242. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3

A26 1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p230-yannakakis/p230-yannakakis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/
p230-yannakakis/. A few papers from

this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Kiernan:1990:MDD

[391] G. Kiernan, C. de Maindreville, and E. Simon. Making deductive databases a practical technology: a step forward. In Garcia-Molina and Jagadish [2263], pages 237–246. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/

                                        p237-kiernan/p237-kiernan.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/93597/p237-kiernan/

.

Willard:1990:QAP

[392] Dan E. Willard. Quasilinear algorithms for processing relational calculus expressions (preliminary report). In ACM [2264], pages 243–257. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3

                                        A26 1990. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        298514/p243-willard/p243-willard.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/298514/

p243-willard/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Mumick:1990:MR

[393] I. S. Mumick, S. J. Finkelstein, Hamid Pirahesh, and Raghu Ramakrishnan. Magic is relevant. In Garcia-Molina and Jagadish [2263], pages 247–258. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/93597/
                                        p247-mumick/p247-mumick.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/93597/p247-mumick/.

Abdel-Ghaffar:1990:ODA

[394] Khaled A. S. Abdel-Ghaffar and Amr El Abbadi. On the optimality of disk allocation for Cartesian product files (extended abstract). In ACM [2264], pages 258–264. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3

A26 1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p258-abdel-ghaffar/p258-
abdel-ghaffar.pdf;   http://www.
acm.org/pubs/citations/proceedings/
pods/298514/p258-abdel-ghaffar/.

A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Widom:1990:SOP

[395] Jennifer Widom and S. J. Finkelstein. Set-oriented production rules in relational database systems. In Garcia-Molina and Jagadish [2263], pages 259–270. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p259-widom/p259-widom.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/93597/p259-widom/.

Aref:1990:EPW

[396] Walid G. Aref and Hanan Samet. Efficient processing of window queries in the pyramid data structure. In ACM [2264], pages 265–272. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3

A26 1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p265-aref/p265-aref.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/298514/p265-aref/

. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Hanson:1990:PMA

[397] Eric N. Hanson, Moez Chaabouni, Chang-Ho Kim, and Yu-Wang Wang. A predicate matching algorithm for database rule systems. In Garcia-Molina and Jagadish [2263], pages 271–280. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p271-hanson/p271-hanson.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/93597/p271-hanson/.

Johnson:1990:FPA

[398] Theodore Johnson and Dennis Shasha. A framework for the performance analysis of concurrent B-tree algorithms. In ACM [2264], pages 273–

287. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p273-johnson/p273-johnson.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/

p273-johnson/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Stonebraker:1990:RPC

[399] Michael Stonebraker, Anant Jhingran, Jeffrey Goh, and Spyros Potamianos. On rules, procedure, caching and views in data base systems. In Garcia-Molina and Jagadish [2263], pages 281–290. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/

p281-stonebraker/p281-stonebraker.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/93597/
p281-stonebraker/.

Lassez:1990:QC

[400] Jean-Louis Lassez. Querying constraints. In ACM [2264], pages 288–

298. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p288-lassez/p288-lassez.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/

p288-lassez/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Rosenthal:1990:QGI

[401] Arnon Rosenthal and Cesar Galindo-Legaria. Query graphs, implementing trees, and freely-reorderable outerjoins. In Garcia-Molina and Jagadish [2263], pages 291–299. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p291-rosenthal/p291-rosenthal.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/93597/p291-rosenthal/

.

Kanellakis:1990:CQL

[402] Paris C. Kanellakis, Gabriel M. Kuper, and Peter Z. Revesz. Constraint query languages (preliminary report). In ACM [2264], pages 299–

313. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p299-kanellakis/p299-kanellakis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/
p299-kanellakis/. A few papers from

this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Shekita:1990:PEP

[403] Eugene J. Shekita and Michael J. Carey. A performance evaluation of pointer-based joins. In Garcia-Molina and Jagadish [2263], pages 300–311. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p300-shekita/p300-shekita.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/93597/p300-shekita/

.

Ioannidis:1990:RAO

[404] Y. E. Ioannidis and Younkyung Kang. Randomized algorithms for optimizing large join queries. In Garcia-Molina and Jagadish [2263], pages 312–321. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p312-ioannidis/p312-ioannidis.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/93597/p312-ioannidis/

.

Mumick:1990:MC

[405] Inderpal Singh Mumick, Sheldon J.

Finkelstein, Hamid Pirahesh, and Raghu Ramakrishnan. Magic conditions. In ACM [2264], pages 314– 330. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p314-mumick/p314-mumick.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/

p314-mumick/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Mumick:1990:MCR

[406] I. S. Mumick, S. J. Finkelstein, H. Pirahesh, Ramakrishnan, and R. Magic conditions (relational queries). In ACM [2264], pages 314–380. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26 1990. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Beckmann:1990:RTE

[407] Norbert Beckmann, Hans-Peter Kriegel, Ralf Schneider, and Bernhard Seeger. The R^{∗}-tree: an efficient and robust access method for points and rectangles. In Garcia-Molina and Jagadish [2263], pages 322–331. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p322-beckmann/p322-beckmann.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/93597/p322-beckmann/

.

Haritsa:1990:BOA

[408] Jayant R. Haritsa, Michael J. Carey, and Miron Livny. On being optimistic about real-time constraints. In ACM [2264], pages 331–343. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3

                                        A26 1990. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        298514/p331-haritsa/p331-haritsa.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/298514/

p331-haritsa/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Jagadish:1990:LCO

[409] H. V. Jagadish. Linear clustering of objects with multiple attributes. In Garcia-Molina and Jagadish [2263], pages 332–342. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/93597/
                                        p332-jagadish/p332-jagadish.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/93597/p332-jagadish/

.

Orenstein:1990:CSQ

[410] Jack Orenstein. A comparison of spatial query processing techniques for native and parameter spaces. In Garcia-Molina and Jagadish [2263], pages 343–352. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/93597/

p343-orenstein/p343-orenstein.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/93597/p343-orenstein/

.

Tam:1990:TTM

[411] Va-On Tam and Meichun Hsu. Token transactions: managing fine-grained migration of data. In ACM [2264], pages 344–356. ISBN 0-89791-352-

3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p344-tam/p344-tam.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/298514/p344-tam/

. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Lomet:1990:PMA

[412] David Lomet and Betty Salzberg. The performance of a multiversion access method. In Garcia-Molina and Jagadish [2263], pages 353–363. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p353-lomet/p353-lomet.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/93597/p353-lomet/.

Soparkar:1990:DVP

[413] Nandit Soparkar and Abraham Silberschatz. Data-valued partitioning and virtual messages (extended abstract). In ACM [2264], pages 357–

364. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p357-soparkar/p357-soparkar.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/

p357-soparkar/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Kemper:1990:ASO

[414] Alfons Kemper and Guido Moerkotte. Access support in object bases. In Garcia-Molina and Jagadish [2263], pages 364–374. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p364-kemper/p364-kemper.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/93597/p364-kemper/.

Pilarski:1990:NCS

[415] Slawomir Pilarski and Tiko Kameda. A novel checkpointing scheme for distributed database systems. In ACM [2264], pages 368–378. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3

A26 1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p368-pilarski/p368-pilarski.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/

p368-pilarski/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Olken:1990:RSH

[416] Frank Olken, Doron Rotem, and Ping Xu. Random sampling from hash files. In Garcia-Molina and Jagadish [2263], pages 375–386. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p375-olken/p375-olken.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/93597/p375-olken/.

Chomicki:1990:PTQ

[417] Jan Chomicki. Polynomial time query processing in temporal deductive databases. In ACM [2264], pages 379–391. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p379-chomicki/p379-chomicki.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/

p379-chomicki/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Cha:1990:KCM

[418] Sang K. Cha and Gio Wiederhold. Kaleidoscope: a cooperative menuguided query interface. In Garcia-Molina and Jagadish [2263], page 387. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p387-cha/p387-cha.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/93597/p387-cha/.

Consens:1990:GGV

[419] Mariano P. Consens and Alberto O. Mendelzon. The G+/GraphLog Visual Query System. In Garcia-Molina and Jagadish [2263], page 388. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/93597/
                                        p388-consens/p388-consens.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/93597/p388-consens/

.

Agrawal:1990:OUF

[420] R. Agrawal, N. H. Gehani, and J. Srinivasan. OdeView: a user-friendly graphical interface to Ode. In Garcia-Molina and Jagadish [2263], page 389. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/93597/
                                        p389-agrawal/p389-agrawal.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/93597/p389-agrawal/

.

Blum:1990:ISQ

[421] Bruce I. Blum and Ralph D. Semmel. The INA: a simple query language with only attribute names. In Garcia-Molina and Jagadish [2263], page 390. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/93597/
                                        p390-blum/p390-blum.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/93597/p390-blum/.

Kuntz:1990:PGD

[422] Michel Kuntz. Pasta-3: a graphical direct manipulation interface for knowledge base management systems. In Garcia-Molina and Jagadish [2263], page 391. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p391-kuntz/p391-kuntz.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/93597/p391-kuntz/.

Kent:1990:IDS

[423] Bill Kent, Peter Lyngbaek, Samir Mathur, and Kevin Wilkinson. The Iris database system. In Garcia-Molina and Jagadish [2263], page 392. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/93597/
p392-kent/p392-kent.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/93597/p392-kent/.

Kabanza:1990:HIT

[424] F. Kabanza, J.-M. Stevenne, and P. Wolper. Handling infinite temporal data. In ACM [2264], pages 392–403. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26

1990. URL http://www.acm.org/
pubs/articles/proceedings/pods/
298514/p392-kabanza/p392-kabanza.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/298514/

p392-kabanza/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Consens:1990:GVF

[425] Mariano P. Consens and Alberto O. Mendelzon. GraphLog: a visual formalism for real life recursion. In ACM [2264], pages 404–416. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3

                                        A26 1990. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        298514/p404-consens/p404-consens.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/298514/

p404-consens/. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Gyssens:1990:GOOa

[426] M. Gyssens, J. Paredaens, and D. Van Gucht. A graph-oriented object database model. In ACM [2264], pages 417–424. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26 1990. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

Gyssens:1990:GOOb

[427] Marc Gyssens, Jan Paredaens, and Dirk Van Gucht. A graph-oriented object model for database end-user interfaces. In Garcia-Molina and Jagadish [2263], pages 24–33. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/93597/
                                        p24-gyssens/p24-gyssens.pdf;   http:

//www.acm.org/pubs/citations/proceedings/
mod/93597/p24-gyssens/.

Lenat:1991:KAC

[428] Douglas B. Lenat. Keynote address: computers versus common sense. In ACM [2265], page 1. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3

A296 1991. URL http://www.acm.
org/pubs/articles/proceedings/mod/
115790/p1-lenat/p1-lenat.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/115790/p1-lenat/

. ACM order number 475910.

Hillebrand:1991:TDB

[429] Gerd G. Hillebrand, Paris C. Kanellakis, Harry G. Mairson, and Moshe Y. Vardi. Tools for Datalog boundedness. In ACM [2265], pages 1–

12. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p1-hillebrand/p1-hillebrand.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/113413/
p1-hillebrand/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/113413/p1-hillebrand/. ACM

order number 475910.

Stonebraker:1991:MPO

[430] Michael Stonebraker. Managing persistent objects in a multi-level store. In ACM [2265], pages 2–

11. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p2-stonebraker/p2-stonebraker.pdf;

                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/115790/p2-stonebraker/

. ACM order number 475910.

Tsangaris:1991:SAC

[431] Manolis M. Tsangaris and Jeffrey F. Naughton. A stochastic approach for clustering in object bases. In ACM [2265], pages 12–21. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/115790/
                                        p12-tsangaris/p12-tsangaris.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/115790/p12-tsangaris/

. ACM order number 475910.

Afrati:1991:DVP

[432] Foto Afrati, Stavros S. Cosmadakis, and Mihalis Yannakakis. On Datalog vs. polynomial time (extended abstract). In ACM [2265], pages 13–25. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        113413/p13-afrati/p13-afrati.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/113413/p13-afrati/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/113413/

p13-afrati/. ACM order number

475910.

Cheng:1991:ECC

[433] Jia-Bing R. Cheng and A. R. Hurson. Effective clustering of complex objects in object-oriented databases. In ACM [2265], pages 22–31. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/

pubs/articles/proceedings/mod/115790/
p22-cheng/p22-cheng.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/115790/p22-cheng/. ACM order

number 475910.

Vadaparty:1991:PRB

[434] Kumar Vadaparty. On the power of rule-based languages with sets. In ACM [2265], pages 26–36. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296 1991. URL http://

www.acm.org/pubs/articles/proceedings/
pods/113413/p26-vadaparty/p26-vadaparty.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/113413/
p26-vadaparty/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/113413/p26-vadaparty/. ACM

order number 475910.

Perrizo:1991:HHD

[435] William Perrizo, Joseph Rajkumar, and Prabhu Ram. HYDRO: a heterogeneous distributed database system. In ACM [2265], pages 32–

39. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p32-perrizo/p32-perrizo.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/115790/p32-perrizo/. ACM or-

der number 475910.

Immerman:1991:EFF

[436] Neil Immerman, Sushant Patnaik, and David Stemple. The expressiveness of a family of finite set languages. In ACM [2265], pages 37–52. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        113413/p37-immerman/p37-immerman.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/113413/
                                        p37-immerman/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/113413/p37-immerman/. ACM

order number 475910.

Krishnamurthy:1991:LFI

[437] Ravi Krishnamurthy, Witold Litwin, and William Kent. Language features for interoperability of databases with schematic discrepancies. In ACM [2265], pages 40–49. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/115790/
                                        p40-krishnamurthy/p40-krishnamurthy.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/115790/
                                        p40-krishnamurthy/. ACM order

number 475910.

Jajodia:1991:TMS

[438] Sushil Jajodia and Ravi Sandhu. Toward a multilevel secure relational data model. In ACM [2265], pages 50–59. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/115790/
                                        p50-jajodia/p50-jajodia.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/115790/p50-jajodia/. ACM or-

der number 475910.

Garcia-Molina:1991:NDQ

[439] Hector Garcia-Molina and Kenneth Salem. Non-deterministic queue operations. In ACM [2265], pages 53– 62. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p53-garcia-molina/p53-garcia-
molina.pdf;   http://www.acm.org/
pubs/citations/proceedings/pods/
113413/p53-garcia-molina/;   http:
//www.acm.org:80/pubs/citations/
proceedings/pods/113413/p53-garcia-

molina/. ACM order number 475910.

Gordin:1991:SOC

[440] Douglas N. Gordin and Alexander J. Pasik. Set-oriented constructs: from Rete rule bases to database systems. In ACM [2265], pages 60–

67. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p60-gordin/p60-gordin.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/115790/p60-gordin/. ACM or-

der number 475910.

Krishnakumar:1991:BIR

[441] Narayanan Krishnakumar and Arthur J. Bernstein. Bounded ignorance in replicated systems. In ACM [2265], pages 63–74. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p63-krishnakumar/p63-krishnakumar.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/113413/
p63-krishnakumar/;   http://www.
acm.org:80/pubs/citations/proceedings/
pods/113413/p63-krishnakumar/. ACM

order number 475910.

Sudarshan:1991:SOB

[442] S. Sudarshan, Divesh Srivastava, Raghu Ramakrishnan, and Jeffrey F. Naughton. Space optimization in the bottom-up evaluation of logic programs. In ACM [2265], pages 68–

77. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/115790/
                                        p68-sudarshan/p68-sudarshan.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/115790/p68-sudarshan/

. ACM order number 475910.

Johnson:1991:TUB

[443] Donald B. Johnson and Larry Raab. A tight upper bound on the benefits of replication and consistency control protocols. In ACM [2265], pages 75–81. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        113413/p75-johnson/p75-johnson.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/113413/
                                        p75-johnson/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/113413/p75-johnson/. ACM

order number 475910.

Wolfson:1991:IER

[444] Ouri Wolfson, Hasanat M. Dewan, Salvatore J. Stolfo, and Yechiam Yemini. Incremental evaluation of rules and its relationship to parallelism. In ACM [2265], pages 78–87. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/

pubs/articles/proceedings/mod/115790/
p78-wolfson/p78-wolfson.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/115790/p78-wolfson/. ACM or-

der number 475910.

Wang:1991:MHS

[445] Wei hsing Wang, Meichun Hsu, and Eugene Pinsky. Modeling hot spots in database systems (extended abstract). In ACM [2265], pages 82–

91. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p82-wang/p82-wang.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/113413/p82-wang/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/113413/

p82-wang/. ACM order number 475910.

Levy:1991:OCP

[446] Eliezer Levy, Henry F. Korth, and Abraham Silberschatz. An optimistic commit protocol for distributed transaction management. In ACM [2265], pages 88–97. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p88-levy/p88-levy.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/115790/p88-levy/. ACM order

number 475910.

Ross:1991:MAT

[447] Kenneth A. Ross. Modular acyclicity and tail recursion in logic programs. In ACM [2265], pages 92–

101. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p92-ross/p92-ross.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/113413/p92-ross/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/113413/

p92-ross/. ACM order number 475910.

Agrawal:1991:UMD

[448] D. Agrawal and V. Krishnaswamy. Using multiversion data for noninterfering execution of write-only transactions. In ACM [2265], pages 98–107. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p98-agrawal/p98-agrawal.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/115790/p98-agrawal/. ACM or-

der number 475910.

Lakshmanan:1991:SQO

[449] Laks V. S. Lakshmanan and Héctor J. Hernández. Structural query optimization — a uniform framework for semantic query optimization in deductive databases. In ACM [2265], pages 102–114. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p102-lakshmanan/p102-lakshmanan.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/113413/
p102-lakshmanan/;   http://www.
acm.org:80/pubs/citations/proceedings/

pods/113413/p102-lakshmanan/. ACM

order number 475910.

Chrysanthis:1991:ECO

[450] Panos K. Chrysanthis, S. Raghuram, and Krithi Ramamritham. Extracting concurrency from objects: a methodology. In ACM [2265], pages 108–

117. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p108-chrysanthis/p108-chrysanthis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/115790/
p108-chrysanthis/. ACM order num-

ber 475910.

Chaudhuri:1991:DRT

[451] Surajit Chaudhuri. Detecting redundant tuples during query evaluation. In ACM [2265], pages 115–

126. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p115-chaudhuri/p115-chaudhuri.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/113413/
p115-chaudhuri/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/113413/p115-chaudhuri/. ACM

order number 475910.

Neugebauer:1991:OED

[452] Leonore Neugebauer. Optimization and evaluation of database queries including embedded interpolation procedures. In ACM [2265], pages 118–

127. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/115790/
                                        p118-neugebauer/p118-neugebauer.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/115790/
                                        p118-neugebauer/. ACM order num-

ber 475910.

Mumick:1991:ORL

[453] Inderpal Singh Mumick and Hamid Pirahesh. Overbound and right-linear queries. In ACM [2265], pages 127–

141. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        113413/p127-mumick/p127-mumick.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/113413/
                                        p127-mumick/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/113413/p127-mumick/. ACM

order number 475910.

Becker:1991:SPS

[454] Bruno Becker, Hans-Werner Six, and Peter Widmayer. Spatial priority search: an access technique for scaleless maps. In ACM [2265], pages 128–137. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/115790/
                                        p128-becker/p128-becker.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/115790/p128-becker/. ACM or-

der number 475910.

Kolovson:1991:SID

[455] Curtis P. Kolovson and Michael Stonebraker. Segment indexes: dynamic indexing techniques for multidimensional interval data. In ACM [2265], pages 138–147. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3

A296 1991. URL http://www.acm.
org/pubs/articles/proceedings/mod/
115790/p138-kolovson/p138-kolovson.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/115790/

p138-kolovson/. ACM order number 475910.

Tsur:1991:DDA

[456] Shalom Tsur. Deductive databases in action. In ACM [2265], pages 142–153. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p142-tsur/p142-tsur.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/113413/p142-tsur/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/113413/

p142-tsur/. ACM order number 475910.

Keller:1991:EAC

[457] Tom Keller, Goetz Graefe, and David Maier. Efficient assembly for complex objects. In ACM [2265], pages 148–157. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p148-keller/p148-keller.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/115790/p148-keller/. ACM or-

der number 475910.

Ganguly:1991:MMP

[458] Sumit Ganguly, Sergio Greco, and Carlo Zaniolo. Minimum and maximum predicates in logic programming. In ACM [2265], pages 154–

163. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p154-ganguly/p154-ganguly.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/113413/
p154-ganguly/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/113413/p154-ganguly/. ACM

order number 475910.

Vandenberg:1991:ASC

[459] Scott L. Vandenberg and David J. DeWitt. Algebraic support for complex objects with arrays, identity, and inheritance. In ACM [2265], pages 158–167. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p158-vandenberg/p158-vandenberg.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/115790/
p158-vandenberg/. ACM order num-

ber 475910.

Matsliach:1991:PAF

[460] Gabriel Matsliach. Performance analysis of file organizations that use multi-bucket data leaves with partial expansions (extended abstract). In ACM [2265], pages 164–

180. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p164-matsliach/p164-matsliach.

pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/113413/
p164-matsliach/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/113413/p164-matsliach/. ACM

order number 475910.

Ioannidis:1991:LDV

[461] Yannis E. Ioannidis and Younkyung Cha Kang. Left-deep vs. bushy trees: an analysis of strategy spaces and its implications for query optimization. In ACM [2265], pages 168–

177. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p168-ioannidis/p168-ioannidis.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/115790/p168-ioannidis/

. ACM order number 475910.

Guntzer:1991:NDU

[462] U. Güntzer, W. Kießling, and H. Thöne. New direction for uncertainty reasoning in deductive databases. In ACM [2265], pages 178–187. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p178-guntzer/p178-guntzer.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/115790/p178-guntzer/

. ACM order number 475910.

Helm:1991:CBQ

[463] Richard Helm, Kim Marriott, and Martin Odersky. Constraint-based query optimization for spatial databases. In ACM [2265], pages 181–191. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296 1991. URL http://

                                        www.acm.org/pubs/articles/proceedings/
                                        pods/113413/p181-helm/p181-helm.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/113413/
                                        p181-helm/;   http://www.acm.org:
                                        80/pubs/citations/proceedings/pods/
                                        113413/p181-helm/. ACM order num-

ber 475910.

Sheng:1991:NDD

[464] Yeh-Heng Sheng. A non-deterministic deductive database language. In ACM [2265], pages 188–197. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3

                                        A296 1991. URL http://www.acm.
                                        org/pubs/articles/proceedings/mod/
                                        115790/p188-sheng/p188-sheng.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/115790/p188-sheng/

. ACM order number 475910.

Nurmi:1991:UUR

[465] Otto Nurmi and Eljas Soisalon-Soininen. Uncoupling updating and rebalancing in chromatic binary search trees. In ACM [2265], pages 192–

198. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        113413/p192-nurmi/p192-nurmi.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/113413/p192-nurmi/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/113413/

p192-nurmi/. ACM order number

475910.

Lou:1991:LOO

[466] Yanjun Lou and Z. Meral Ozsoyoglu. LLO: an object-oriented deductive language with methods and method inheritance. In ACM [2265], pages 198–207. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p198-lou/p198-lou.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/115790/p198-lou/. ACM order

number 475910.

Jakobsson:1991:MAA

[467] Håkan Jakobsson. Mixed-approach algorithms for transitive closure (extended abstract). In ACM [2265], pages 199–205. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p199-jakobsson/p199-jakobsson.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/113413/
p199-jakobsson/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/113413/p199-jakobsson/. ACM

order number 475910.

Ross:1991:NHE

[468] Kenneth A. Ross. On negation in HiLog (extended abstract). In ACM [2265], pages 206–215. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3

A296 1991. URL http://www.acm.
org/pubs/articles/proceedings/pods/
113413/p206-ross/p206-ross.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/113413/p206-ross/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/113413/

p206-ross/. ACM order number 475910.

Jagadish:1991:RTS

[469] H. V. Jagadish. A retrieval technique for similar shapes. In ACM [2265], pages 208–217. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/115790/
                                        p208-jagadish/p208-jagadish.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/115790/p208-jagadish/

. ACM order number 475910.

Sohn:1991:TDL

[470] Kirack Sohn and Allen Van Gelder. Termination detection in logic programs using argument sizes (extended abstract). In ACM [2265], pages 216–226. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        113413/p216-sohn/p216-sohn.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/113413/p216-sohn/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/113413/

p216-sohn/. ACM order number 475910.

Ahmed:1991:VMC

[471] Rafi Ahmed and Shamkant B. Navathe. Version management of composite objects in CAD databases. In ACM [2265], pages 218–227. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3

                                        A296 1991. URL http://www.acm.
                                        org/pubs/articles/proceedings/mod/
                                        115790/p218-ahmed/p218-ahmed.pdf;

 http://www.acm.org/pubs/citations/
proceedings/mod/115790/p218-ahmed/

. ACM order number 475910.

Brodsky:1991:IIC

[472] Alexander Brodsky and Yehoshua Sagiv. Inference of inequality constraints in logic programs (extended abstracts). In ACM [2265], pages 227–240. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p227-brodsky/p227-brodsky.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/113413/
p227-brodsky/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/113413/p227-brodsky/. ACM

order number 475910.

Chiueh:1991:TAM

[473] Tzi cker Chiueh and Randy Katz. Trait: an attribute management system for VLSI design objects. In ACM [2265], pages 228–237. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3

A296 1991. URL http://www.acm.
org/pubs/articles/proceedings/mod/
115790/p228-chiueh/p228-chiueh.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/115790/

p228-chiueh/. ACM order number 475910.

Abiteboul:1991:OV

[474] Serge Abiteboul and Anthony Bonner. Objects and views. In ACM [2265], pages 238–247. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/

                                        pubs/articles/proceedings/mod/115790/
                                        p238-abiteboul/p238-abiteboul.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/115790/p238-abiteboul/

. ACM order number 475910.

Seib:1991:PDP

[475] Jürgen Seib and Georg Lausen. Parallelizing Datalog programs by generalized pivoting. In ACM [2265], pages 241–251. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        113413/p241-seib/p241-seib.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/113413/p241-seib/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/113413/

p241-seib/. ACM order number

475910.

Barsalou:1991:URD

[476] Thierry Barsalou, Niki Siambela, Arthur M. Keller, and Gio Wiederhold. Updating relational databases through object-based views. In ACM [2265], pages 248–257. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/115790/
                                        p248-barsalou/p248-barsalou.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/115790/p248-barsalou/

. ACM order number 475910.

Malvestuto:1991:SMC

[477] F. M. Malvestuto, M. Moscarini, and M. Rafanelli. Suppressing marginal cells to protect sensitive information in a two-dimensional statistical table (extended abstract). In ACM [2265], pages 252–258. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p252-malvestuto/p252-malvestuto.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/113413/
p252-malvestuto/;   http://www.
acm.org:80/pubs/citations/proceedings/
pods/113413/p252-malvestuto/. ACM

order number 475910.

Kemper:1991:FMO

[478] Alfons Kemper, Christoph Kilger, and Guido Moerkotte. Function materialization in object bases. In ACM [2265], pages 258–267. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p258-kemper/p258-kemper.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/115790/p258-kemper/. ACM or-

der number 475910.

Roy:1991:SCC

[479] Shaibal Roy. Semantic complexity of classes of relational queries and query independent data partitioning. In ACM [2265], pages 259–

267. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p259-roy/p259-roy.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/113413/p259-roy/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/113413/

p259-roy/. ACM order number 475910.

Ioannidis:1991:PES

[480] Yannis E. Ioannidis and Stavros Christodoulakis. On the propagation of errors in the size of join results. In ACM [2265], pages 268–

277. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/115790/
                                        p268-ioannidis/p268-ioannidis.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/115790/p268-ioannidis/

. ACM order number 475910.

Seshadri:1991:ESR

[481] S. Seshadri and Jeffrey F. Naughton. On the expected size of recursive Datalog queries. In ACM [2265], pages 268–279. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        113413/p268-seshadri/p268-seshadri.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/113413/
                                        p268-seshadri/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/113413/p268-seshadri/. ACM

order number 475910.

Hou:1991:ECC

[482] Wen-Chi Hou, Gultekin Ozsoyoglu, and Erdogan Dogdu. Error-constrained COUNT query evaluation in relational databases. In ACM [2265], pages 278–287. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/115790/
                                        p278-hou/p278-hou.pdf;   http://

www.acm.org/pubs/citations/proceedings/
mod/115790/p278-hou/. ACM order

number 475910.

Baudinet:1991:RIT

[483] Marianne Baudinet, Marc Niézette, and Pierre Wolper. On the representation of infinite temporal data and queries (extended abstract). In ACM [2265], pages 280–290. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3

A296 1991. URL http://www.acm.
org/pubs/articles/proceedings/pods/
113413/p280-baudinet/p280-baudinet.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/113413/
p280-baudinet/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/113413/p280-baudinet/. ACM

order number 475910.

Imielinski:1991:IOD

[484] Tomasz Imielinski, Shamim Naqvi, and Kumar Vadaparty. Incomplete object

— a data model for design and planning applications. In ACM [2265], pages 288–297. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p288-imielinski/p288-imielinski.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/115790/
p288-imielinski/. ACM order num-

ber 475910.

VandenBussche:1991:EPS

[485] Jan Van den Bussche and Jan Paredaens. The expressive power structured values in pure OODB’s (extended abstract). In ACM [2265], pages 291–299. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p291-van_den_bussche/p291-
van_den_bussche.pdf;   http://www.
acm.org/pubs/citations/proceedings/
pods/113413/p291-van_den_bussche/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/113413/
p291-van_den_bussche/. ACM order

number 475910.

Richardson:1991:AEO

[486] Joel Richardson and Peter Schwarz. Aspects: extending objects to support multiple, independent roles. In ACM [2265], pages 298–307. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3

A296 1991. URL http://www.acm.
org/pubs/articles/proceedings/mod/
115790/p298-richardson/p298-richardson.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/115790/
p298-richardson/. ACM order num-

ber 475910.

Beeri:1991:SOE

[487] Catriel Beeri and Tova Milo. Subtyping in OODB’s (extended abstract). In ACM [2265], pages 300–

314. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p300-beeri/p300-beeri.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/113413/p300-beeri/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/113413/

p300-beeri/. ACM order number 475910.

Phipps:1991:GND

[488] Geoffrey Phipps, Marcia A. Derr, and Kenneth A. Ross. Glue-nail: a deductive database system. In ACM [2265], pages 308–317. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p308-phipps/p308-phipps.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/115790/p308-phipps/. ACM or-

der number 475910.

Grumbach:1991:TQL

[489] Stéphane Grumbach and Victor Vianu. Tractable query languages for complex object databases. In ACM [2265], pages 315–327. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/pods/
113413/p315-grumbach/p315-grumbach.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/113413/
p315-grumbach/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/113413/p315-grumbach/. ACM

order number 475910.

Annevelink:1991:DPL

[490] Jurgen Annevelink. Database programming languages: a functional approach. In ACM [2265], pages 318–

327. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p318-annevelink/p318-annevelink.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/115790/

                                        p318-annevelink/. ACM order num-

ber 475910.

Chen:1991:NRB

[491] Qiming Chen and Yahiko Kambayashi. Nested relation based database knowledge representation. In ACM [2265], pages 328–337. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/115790/
                                        p328-chen/p328-chen.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/115790/p328-chen/. ACM order

number 475910.

Hull:1991:EDR

[492] Richard Hull and Masatoshi Yoshikawa. On the equivalence of database restructurings involving object identifiers (extended abstract). In ACM [2265], pages 328–340. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        113413/p328-hull/p328-hull.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/113413/p328-hull/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/113413/

p328-hull/. ACM order number

475910.

Shyy:1991:KHL

[493] Yuh-Ming Shyy and Stanley Y. W. Su. K: a high-level knowledge base programming language for advanced database applications. In ACM [2265], pages 338–347. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/

pubs/articles/proceedings/mod/115790/
p338-shyy/p338-shyy.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/115790/p338-shyy/. ACM order

number 475910.

Hansen:1991:EMI

[494] S. C. Hansen and E. A. Unger. An extended memoryless inference control model: accounting for dependence in table-level controls. In ACM [2265], pages 348–356. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p348-hansen/p348-hansen.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/115790/p348-hansen/. ACM or-

der number 475910.

Carey:1991:DCT

[495] Michael J. Carey, Michael J. Franklin, Miron Livny, and Eugene J. Shekita. Data caching tradeoffs in client-server DBMS architectures. In ACM [2265], pages 357–366. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p357-carey/p357-carey.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/115790/p357-carey/. ACM or-

der number 475910.

Wang:1991:CCC

[496] Yongdong Wang and Lawrence A. Rowe. Cache consistency and concurrency control in a client/server DBMS architecture. In ACM [2265], pages 367–376. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/115790/
                                        p367-wang/p367-wang.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/115790/p367-wang/. ACM order

number 475910.

Pu:1991:RCD

[497] Calton Pu and Avraham Leff. Replica control in distributed systems: as asynchronous approach. In ACM [2265], pages 377–386. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/115790/
                                        p377-pu/p377-pu.pdf;   http://www.
                                        acm.org/pubs/citations/proceedings/
                                        mod/115790/p377-pu/. ACM order

number 475910.

Ng:1991:FBA

[498] Raymond Ng, Christos Faloutsos, and Timos Sellis. Flexible buffer allocation based on marginal gains. In ACM [2265], pages 387–396. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3

                                        A296 1991. URL http://www.acm.
                                        org/pubs/articles/proceedings/mod/
                                        115790/p387-ng/p387-ng.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/115790/p387-ng/. ACM order

number 475910.

Gruenwald:1991:MRA

[499] Le Gruenwald and Margaret H. Eich. MMDB reload algorithms. In ACM [2265], pages 397–405. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3

                                        A296 1991. URL http://www.acm.
                                        org/pubs/articles/proceedings/mod/
                                        115790/p397-gruenwald/p397-gruenwald.

pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/115790/
p397-gruenwald/. ACM order number

475910.

Weikum:1991:DFA

[500] Gerhard Weikum, Peter Zabback, and Peter Scheuermann. Dynamic file allocation in disk arrays. In ACM [2265], pages 406–415. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p406-weikum/p406-weikum.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/115790/p406-weikum/. ACM or-

der number 475910.

Srinivasan:1991:PBT

[501] V. Srinivasan and Michael J. Carey. Performance of B-tree concurrency control algorithms. In ACM [2265], pages 416–425. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p416-srinivasan/p416-srinivasan.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/115790/
p416-srinivasan/. ACM order num-

ber 475910.

Lanka:1991:FPB

[502] Sitaram Lanka and Eric Mays. Fully persistent B+-trees. In ACM [2265], pages 426–435. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

1991. URL http://www.acm.org/
pubs/articles/proceedings/mod/115790/
p426-lanka/p426-lanka.pdf;   http:
//www.acm.org/pubs/citations/proceedings/

                                        mod/115790/p426-lanka/. ACM or-

der number 475910.

Seeger:1991:MDB

[503] Bernhard Seeger and Per-Åke Larson. Multi-disk B-trees. In ACM [2265], pages 436–445. ISBN 0-89791-430-

9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/115790/
                                        p436-seeger/p436-seeger.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/115790/p436-seeger/. ACM or-

der number 475910.

Anonymous:1991:SPH

[504] Anonymous. Are standards the panacea for heterogeneous distributed DBMSs? In ACM [2265], page 446. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296 1991. ACM order number 475910.

Lohman:1991:SIE

[505] Guy M. Lohman, George Lapis, Tobin Lehman, Rakesh Agrawal, Roberta Cochrane, John McPherson, C. Mohan, Hamid Pirahesh, and Jennifer Widom. Starburst II: the extender strikes back! In ACM [2265], page

447. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296

                                        1991. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/115790/
                                        p447-lohman/p447-lohman.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/115790/p447-lohman/. ACM or-

der number 475910.

Dozier:1992:KAA

[506] Jeff Dozier. Keynote address: access to data in NASA’s Earth observing system. In Stonebraker [2268], page 1. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p1-dozier/p1-dozier.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p1-dozier/.

Beeri:1992:NDM

[507] Catriel Beeri. New data models and languages — the challenge. In ACM [2269], pages 1–15. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26

1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p1-beeri/p1-beeri.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/137097/p1-beeri/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/137097/

p1-beeri/. ACM order number 475920.

Merz:1992:DQF

[508] Ulla Merz and Roger King. DIRECT: a query facility for multiple databases. In Stonebraker [2268], page 2. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p2-merz/p2-merz.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/130283/p2-merz/.

Chen:1992:PBK

[509] Q. Chen, W. W. Chu, and R.-C. Lee. Pattern-based knowledge induction from databases. In Kim et al. [2267], pages 2–9. ISBN 981-02-1315-

8. LCCN QA76.9.D3 D3589 1992.

Celentano:1992:CDB

[510] A. Celentano, M. G. Fugini, and S. Pozzi. Conceptual document browsing and retrieval in Kabiria. In Stonebraker [2268], page 3. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p3-celentano/p3-celentano.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/130283/p3-celentano/

.

Johnson:1992:UDM

[511] Rowland R. Johnson, Mandy Goldner, Mitch Lee, Keith McKay, Robert Shectman, and John Woodruff. USD

— a database management system for scientific research. In Stonebraker [2268], page 4. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p4-johnson/p4-johnson.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p4-johnson/.

Gesmann:1992:PDS

[512] Michael Gesmann, Andreas Grasnickel, Theo Härder, Christoph Hübel, Wolfgang Käfer, Bernhard Mitschang, and Harald Schöning. PRIMA

— a database system supporting dynamically defined composite objects. In Stonebraker [2268], page 5. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p5-gesmann/p5-gesmann.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p5-gesmann/.

Gray:1992:DTP

[513] Jim Gray. Database and transaction processing benchmarks. In Stonebraker [2268], page 6. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p6-gray/p6-gray.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/130283/p6-gray/.

Bancilhon:1992:OOD

[514] François Bancilhon. The O_{2} objectoriented database system. In Stonebraker [2268], page 7. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p7-bancilhon/p7-bancilhon.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/130283/p7-bancilhon/

.

Orenstein:1992:AOD

[515] Jack Orenstein. Architectures for object data management. In Stonebraker [2268], page 8. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p8-orenstein/p8-orenstein.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/130283/p8-orenstein/

.

Ganguly:1992:QOP

[516] Sumit Ganguly, Waqar Hasan, and Ravi Krishnamurthy. Query optimization for parallel execution. In Stonebraker [2268], pages 9–18. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/
                                        p9-ganguly/p9-ganguly.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/130283/p9-ganguly/.

Murakami:1992:SPK

[517] K. Murakami and T. Aibara. Some properties of knowledge information inferred by two interactive default reasoning systems. In Kim et al. [2267], pages 10–18. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Papadimitriou:1992:TBS

[518] Christos H. Papadimitriou and Mihalis Yannakakis. Tie-breaking semantics and structural totality. In ACM [2269], pages 16–22. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3

                                        A26 1992. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        137097/p16-papadimitriou/p16-papadimitriou.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/137097/
                                        p16-papadimitriou/;   http://www.
                                        acm.org:80/pubs/citations/proceedings/
                                        pods/137097/p16-papadimitriou/.

ACM order number 475920.

Qi:1992:FPE

[519] R. Qi and W. Bibel. A framework for the parallel evaluation of recursive queries in deductive databases. In Kim et al. [2267], pages 19–27. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Hong:1992:EIO

[520] Wei Hong. Exploiting inter-operation parallelism in XPRS. In Stonebraker [2268], pages 19–28. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p19-hong/p19-hong.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/130283/p19-hong/.

Abiteboul:1992:QET

[521] Serge Abiteboul, Kevin Compton, and Victor Vianu. Queries are easier than you thought (probably). In ACM [2269], pages 23–32. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26

1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p23-abiteboul/p23-abiteboul.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/137097/
p23-abiteboul/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/137097/p23-abiteboul/. ACM

order number 475920.

Han:1992:CEA

[522] J. Han and Wen Yu Lu. Compilation and evaluation of asynchronous chain recursions in deductive databases. In Kim et al. [2267], pages 28–41. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Ghandeharizadeh:1992:PAA

[523] Shahram Ghandeharizadeh, David J. DeWitt, and Waheed Qureshi. A performance analysis of alternative multi-attribute declustering strategies. In Stonebraker [2268], pages 29–38. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p29-ghandeharizadeh/p29-ghandeharizadeh.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/130283/
p29-ghandeharizadeh/.

Greiner:1992:LEQ

[524] Russell Greiner. Learning efficient query processing strategies. In ACM [2269], pages 33–46. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26

1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p33-greiner/p33-greiner.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/137097/
p33-greiner/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/137097/p33-greiner/. ACM

order number 475920.

Pirahesh:1992:ERB

[525] Hamid Pirahesh, Joseph M. Hellerstein, and Waqar Hasan. Extensible/rule based query rewrite optimization in Starburst. In Stonebraker [2268], pages 39–48. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p39-pirahesh/p39-pirahesh.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/130283/p39-pirahesh/

.

Kim:1992:MMI

[526] H.-J. Kim, S.-H. Lee, and H. R. Astudillo. Method maintenance issues in dynamic object-oriented databases. In Kim et al. [2267], pages 42–50. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Rotem:1992:ADA

[527] Doron Rotem. Analysis of disk arm movement for large sequential reads. In ACM [2269], pages 47–54. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3

A26 1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p47-rotem/p47-rotem.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/137097/p47-rotem/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/137097/

p47-rotem/. ACM order number 475920.

Hanson:1992:RCT

[528] Eric N. Hanson. Rule condition testing and action execution in Ariel. In Stonebraker [2268], pages 49–58. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p49-hanson/p49-hanson.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p49-hanson/.

Vittori:1992:NOS

[529] E. Vittori, N. Le Thanh, G. Mopolo, Miranda, and S. NICE-OS: An object server for persistent programming in NICE-C++. In Kim et al. [2267], pages 51–59. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Chaudhuri:1992:ERN

[530] Surajit Chaudhuri and Moshe Y. Vardi. On the equivalence of recursive and nonrecursive datalog programs. In ACM [2269], pages 55–66. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3

A26 1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p55-chaudhuri/p55-chaudhuri.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/137097/
p55-chaudhuri/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/137097/p55-chaudhuri/. ACM

order number 475920.

Aiken:1992:BDP

[531] Alexander Aiken, Jennifer Widom, and Joseph M. Hellerstein. Behavior of database production rules: termination, confluence, and observable determinism. In Stonebraker [2268], pages 59–68. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p59-aiken/p59-aiken.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p59-aiken/.

Masai:1992:AFI

[532] K. Masai, S. Yamamoto, H. Ishikawa, Sumiyoshi, and T. Advanced features of integrated DB/DC system “XDM”. In Kim et al. [2267], pages 60–67. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Levy:1992:CRD

[533] Alon Levy and Yehoshua Sagiv. Constraints and redundancy in datalog. In ACM [2269], pages 67–80. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3

A26 1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p67-levy/p67-levy.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/137097/p67-levy/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/137097/

p67-levy/. ACM order number 475920.

Ling:1992:SIC

[534] Tok Wang Ling and Sin Yeung Lee. A survey of integrity constraint checking methods in relational databases. In Kim et al. [2267], pages 68–78. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Wade:1992:FDO

[535] Andrew E. Wade. Full distribution in Objectivity/DB. In Stonebraker [2268], page 69. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p69-wade/p69-wade.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/130283/p69-wade/.

Barry:1992:IDO

[536] Douglas K. Barry. ITASCA Distributed ODBMS. In Stonebraker [2268], page 70. ISBN 0-89791-521-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? URL http://

www.acm.org/pubs/articles/proceedings/
mod/130283/p70-barry/p70-barry.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/130283/
p70-barry/.

Cruz:1992:DVL

[537] Isabel F. Cruz. DOODLE: a visual language for object-oriented databases. In Stonebraker [2268], pages 71–80. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/
                                        p71-cruz/p71-cruz.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/130283/p71-cruz/.

Xu:1992:RCR

[538] H. Xu and Y. Kambayashi. Realization of composite relationship views utilizing regular expressions. In Kim et al. [2267], pages 79–87. ISBN 981-02-1315-

8. LCCN QA76.9.D3 D3589 1992.

Dong:1992:DEC

[539] Guozhu Dong. Datalog expressiveness of chain queries: grammar tools and characterizations. In ACM [2269], pages 81–90. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26

                                        1992. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        137097/p81-dong/p81-dong.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/pods/137097/p81-dong/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/137097/

p81-dong/. ACM order number

475920.

Gehani:1992:ESA

[540] N. H. Gehani, H. V. Jagadish, and O. Shmueli. Event specification in an active object-oriented database. In Stonebraker [2268], pages 81–90. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p81-gehani/p81-gehani.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p81-gehani/.

Unland:1992:NTM

[541] R. Unland. A nested transaction model for engineering applications. In Kim et al. [2267], pages 88–101. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Lieuwen:1992:TBA

[542] Daniel F. Lieuwen and David J. De-Witt. A transformation-based approach to optimizing loops in database programming languages. In Stonebraker [2268], pages 91–100. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p91-lieuwen/p91-lieuwen.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p91-lieuwen/.

Beeri:1992:VMS

[543] Catriel Beeri, Raghu Ramakrishnan, Divesh Srivastava, and S. Sudarshan. The valid model semantics for logic programs. In ACM [2269], pages 91–104. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26

1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p91-beeri/p91-beeri.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/137097/p91-beeri/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/137097/

p91-beeri/. ACM order number 475920.

Rosenthal:1992:WCW

[544] Arnon Rosenthal. What can we do to strengthen the connection between theory and system builders (panel). In Stonebraker [2268], page 101. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/
                                        p101-rosenthal/p101-rosenthal.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/130283/p101-rosenthal/

.

Kasi:1992:HPA

[545] Jay Kasi. High performance and availability through data distribution. In Stonebraker [2268], page 102. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/
                                        p102-kasi/p102-kasi.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/130283/p102-kasi/.

Fang:1992:SIF

[546] D. Fang and D. McLeod. Seamless interconnection in federated database systems. In Kim et al. [2267], pages 102–112. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Davison:1992:PIB

[547] Wayne Davison. Parallel index building in Informix OnLine 6.0. In Stonebraker [2268], page 103. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/
                                        p103-davison/p103-davison.pdf;

http://www.acm.org/pubs/citations/
proceedings/mod/130283/p103-davison/

.

Agrawal:1992:UDC

[548] D. Agrawal, A. El Abbadi, and R. Jeffers. Using delayed commitment in locking protocols for real-time databases. In Stonebraker [2268], pages 104–113. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p104-agrawal/p104-agrawal.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/130283/p104-agrawal/

.

Greco:1992:GC

[549] Sergio Greco, Carlo Zaniolo, and Sumit Ganguly. Greedy by choice. In ACM [2269], pages 105–113. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3

A26 1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p105-greco/p105-greco.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/137097/p105-greco/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/137097/

p105-greco/. ACM order number 475920.

Kim:1992:SOD

[550] J. L. Kim. In support of optimism in a distributed database system. In Kim et al. [2267], pages 113–119. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589

Dan:1992:PAC

[551] Asit Dan and Philip S. Yu. Performance analysis of coherency control policies through lock retention. In Stonebraker [2268], pages 114–123. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/
                                        p114-dan/p114-dan.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/130283/p114-dan/.

Ross:1992:MAD

[552] Kenneth A. Ross and Yehoshua Sagiv. Monotonic aggregation in deductive databases. In ACM [2269], pages 114–126. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26

                                        1992. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        137097/p114-ross/p114-ross.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/137097/p114-ross/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/137097/

p114-ross/. ACM order number

475920.

Son:1992:ADR

[553] Sang Hyuk Son, J. Lee, and H. Kang. Approaches to design of real-time database systems. In Kim et al. [2267], pages 120–131. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Mohan:1992:EFM

[554] C. Mohan, Hamid Pirahesh, and Raymond Lorie. Efficient and flexible methods for transient versioning of records to avoid locking by read-only transactions. In Stonebraker [2268], pages 124–133. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p124-mohan/p124-mohan.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p124-mohan/.

VanGelder:1992:WFS

[555] Allen Van Gelder. The well-founded semantics of aggregation. In ACM [2269], pages 127–138. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26

1992. URL http://www.acm.org/
pubs/citations/proceedings/pods/
137097/p127-van_gelder/. ACM or-

der number 475920.

Son:1992:PER

[556] S. H. Son, C.-H. Chang, and Y.-K. Kim. Performance evaluation of realtime locking protocols. In Kim et al. [2267], pages 132–141. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Low:1992:HTD

[557] Chee Chin Low, Beng Chin Ooi, and Hongjun Lu. H-trees: a dynamic associative search index for OODB. In Stonebraker [2268], pages 134–143. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p134-low/p134-low.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/130283/p134-low/.

Kanasaki:1992:CBE

[558] K. Kanasaki and T. L. Kunii. Casebased evolutionary world model for electronic secretaries. In Kim et al. [2267], pages 142–149. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Rabinovich:1992:FTC

[559] Michael Rabinovich and Edward D. Lazowska. A fault-tolerant commit protocol for replicated databases. In ACM [2269], pages 139–148. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3

A26 1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p139-rabinovich/p139-rabinovich.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/137097/
p139-rabinovich/;   http://www.
acm.org:80/pubs/citations/proceedings/
pods/137097/p144-rabinovich/. ACM

order number 475920.

Tsangaris:1992:POC

[560] Manolis M. Tsangaris and Jeffrey F. Naughton. On the performance of object clustering techniques. In Stonebraker [2268], pages 144–153. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p144-tsangaris/p144-tsangaris.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/130283/p144-tsangaris/

.

Wolfson:1992:DAD

[561] Ouri Wolfson and Sushil Jajodia. Distributed algorithms for dynamic replication of data. In ACM [2269], pages 149–163. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26

1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/

137097/p149-wolfson/p149-wolfson.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/137097/
p149-wolfson/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/137097/p149-wolfson/. ACM

order number 475920.

Shimada:1992:SCS

[562] M. Shimada, H. Nishimoto, T. Ishizaka, and A. Schuetz. Supporting the character sets of Japanese Kanji and Korean Hangul in the ADABAS/NATURAL system. In Kim et al. [2267], pages 150–155. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Ley:1992:TRA

[563] Michael Ley. The term retrieval abstract machine. In Stonebraker [2268], pages 154–163. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p154-ley/p154-ley.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/130283/p154-ley/.

Cicchetti:1992:MMR

[564] R. Cicchetti, L. Lakhal, and S. Miranda. MARELA: a matrix relational model for statistical database management. In Kim et al. [2267], pages 156–169. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Bamford:1992:UMI

[565] Roger Bamford. Using multiversioning to improve performance without loss of consistency. In Stonebraker [2268], page 164. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p164-bamford/p164-bamford.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/130283/p164-bamford/

.

Descollonges:1992:CMT

[566] Marc Descollonges. A concurrency model for transaction management. In Stonebraker [2268], page 164. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p164-descollonges/p164-descollonges.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/130283/
p164-descollonges/.

Mehrotra:1992:ETA

[567] Sharad Mehrotra, Rajeev Rastogi, Yuri Breitbart, Henry F. Korth, and Avi Silberschatz. Ensuring transaction atomicity in multidatabase systems. In ACM [2269], pages 164–175. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3

A26 1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p164-mehrotra/p164-mehrotra.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/137097/
p164-mehrotra/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/137097/p164-mehrotra/. ACM

order number 475920.

Franklin:1992:CRC

[568] Michael J. Franklin, Michael J. Zwilling, C. K. Tan, Michael J. Carey, and David J. DeWitt. Crash recovery in client-server EXODUS. In Stonebraker [2268], pages 165–174. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p165-franklin/p165-franklin.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/130283/p165-franklin/

.

Luan:1992:FAC

[569] Y. Q. Luan, N. Ohbo, H. Kitagawa, and X. Yu. Functional approach to chemical structure databases. In Kim et al. [2267], pages 170–183. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Jhingran:1992:ARD

[570] Anant Jhingran and Pratap Khedkar. Analysis of recovery in a database system using a write-ahead log protocol. In Stonebraker [2268], pages 175–184. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p175-jhingran/p175-jhingran.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/130283/p175-jhingran/

.

Beeri:1992:FPP

[571] Catriel Beeri and Tova Milo. Functional and predictive programming in OODB’s. In ACM [2269], pages 176–190. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26

1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p176-beeri/p176-beeri.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/137097/p176-beeri/
;   http://www.acm.org:80/pubs/

                                        citations/proceedings/pods/137097/

p176-beeri/. ACM order number

475920.

Oomoto:1992:DIV

[572] E. Oomoto and K. Tanaka. Design and implementation of a visual query language for historical databases. In Kim et al. [2267], pages 184–191. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Lomet:1992:MRM

[573] David B. Lomet. MLR: a recovery method for multi-level systems. In Stonebraker [2268], pages 185–194. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/
                                        p185-lomet/p185-lomet.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/130283/p185-lomet/.

VandenBussche:1992:SDE

[574] Jan Van den Bussche and Dirk Van Gucht. Semi-determinism (extended abstract). In ACM [2269], pages 191–201. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26

                                        1992. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        137097/p191-van_den_bussche/p191-
                                        van_den_bussche.pdf;   http://www.
                                        acm.org/pubs/citations/proceedings/
                                        pods/137097/p191-van_den_bussche/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/137097/
                                        p191-van_den_bussche/. ACM order

number 475920.

Kitagawa:1992:MMD

[575] F. Kitagawa and H. Ikeda. Multimedia database language SL/B5 based on screen flow. In Kim et al. [2267], pages 192–200. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Kamel:1992:PRT

[576] Ibrahim Kamel and Christos Faloutsos. Parallel R-trees. In Stonebraker [2268], pages 195–204. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p195-kamel/p195-kamel.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p195-kamel/.

Ohtomo:1992:MOV

[577] M. Ohtomo and H. Ikeda. Maporiented visual language: MOL. In Kim et al. [2267], pages 201–208. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Chan:1992:CMP

[578] Edward P. F. Chan. Containment and minimization of positive conjunctive queries in OODB’s. In ACM [2269], pages 202–211. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26

1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p202-chan/p202-chan.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/137097/p202-chan/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/137097/

p202-chan/. ACM order number 475920.

Hoel:1992:QCS

[579] Erik G. Hoel and Hanan Samet. A qualitative comparison study of data structures for large line segment databases. In Stonebraker [2268], pages 205–214. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p205-hoel/p205-hoel.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p205-hoel/.

Kasahara:1992:NVI

[580] H. Kasahara and T. Kishimoto. Navigation in the visual information database world. In Kim et al. [2267], pages 209–217. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Turek:1992:LBM

[581] John Turek, Dennis Shasha, and Sundeep Prakash. Locking without blocking: making lock based concurrent data structure algorithms nonblocking. In ACM [2269], pages 212–222. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3

A26 1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p212-turek/p212-turek.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/137097/p212-turek/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/137097/

p212-turek/. ACM order number 475920.

Analyti:1992:FSM

[582] Anastasia Analyti and Sakti Pramanik. Fast search in main memory databases. In Stonebraker [2268], pages 215–224.

ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p215-analyti/p215-analyti.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/130283/p215-analyti/

.

Chang:1992:SSS

[583] C. C. Chang and D. J. Buehrer. A survey of some spatial match query algorithms. In Kim et al. [2267], pages 218–223. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Agrawal:1992:AET

[584] D. Agrawal, A. El Abbadi, and R. Jeffers. An approach to eliminate transaction blocking in locking protocols. In ACM [2269], pages 223–235. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3

A26 1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p223-agrawal/p223-agrawal.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/137097/
p223-agrawal/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/137097/p223-agrawal/. ACM

order number 475920.

Embley:1992:IOM

[585] D. W. Embley and G. Nagy. Intersection operations in a multi-layered spatial data model. In Kim et al. [2267], pages 224–230. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Celis:1992:DPA

[586] Pedro Celis. Distribution, parallelism, and availability in nonstop SQL. In Stonebraker [2268], page 225.

ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p225-celis/p225-celis.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p225-celis/.

Rabinovich:1992:IFT

[587] Michael Rabinovich and Edward D. Lazowska. Improving fault tolerance and supporting partial writes in structured coterie protocols for replicated objects. In Stonebraker [2268], pages 226–235. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p226-rabinovich/p226-rabinovich.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/130283/
p226-rabinovich/.

Iwaihara:1992:PNQ

[588] M. Iwaihara and Y. Kambayashi. Processing nondeterministic queries including sets and constraints utilizing data dependencies. In Kim et al. [2267], pages 231–239. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Bhide:1992:ESP

[589] Anupam Bhide, Ambuj Goyal, Hui-I. Hsiao, and Anant Jhingran. An efficient scheme for providing high availability. In Stonebraker [2268], pages 236–245. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p236-bhide/p236-bhide.pdf;   http:

//www.acm.org/pubs/citations/proceedings/
mod/130283/p236-bhide/.

Wong:1992:TBI

[590] M. H. Wong and D. Agrawal. Tolerating bounded inconsistency for increasing concurrency in database systems. In ACM [2269], pages 236–245. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3

A26 1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p236-wong/p236-wong.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/137097/p236-wong/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/137097/

p236-wong/. ACM order number 475920.

Ajitomi:1992:ERA

[591] N. Ajitomi and H. Kurose. An enhanced RETE algorithm for large scale data access. In Kim et al. [2267], pages 240–249. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Polyzois:1992:ERB

[592] Christos A. Polyzois and Hector Garcia-Molina. Evaluation of remote backup algorithms for transaction processing systems. In Stonebraker [2268], pages 246–255. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p246-polyzois/p246-polyzois.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/130283/p246-polyzois/

.

Grahne:1992:KT

[593] Gösta Grahne, Alberto O. Mendelzon, and Peter Z. Revesz. Knowledgebase transformations. In ACM [2269], pages 246–260. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26

1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p246-grahne/p246-grahne.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/137097/
p246-grahne/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/137097/p246-grahne/. ACM

order number 475920.

Ramamohanarao:1992:PMR

[594] K. Ramamohanarao, J. Shepherd, and R. Sacks-Davis. Partial-match retrieval using multiple-key hashing with multiple file copies. In Kim et al. [2267], pages 250–275. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Lanzelotte:1992:OOO

[595] Rosana S. G. Lanzelotte, Patrick Valduriez, and Mohamed Zaït. Optimization of object-oriented recursive queries using cost-controlled strategies. In Stonebraker [2268], pages 256–265. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p256-lanzelotte/p256-lanzelotte.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/130283/
p256-lanzelotte/.

Eiter:1992:CPK

[596] Thomas Eiter and Georg Gottlob. On the complexity of propositional knowledge base revision, updates, and counterfactuals. In ACM [2269], pages 261–273. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26

1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p261-eiter/p261-eiter.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/137097/p261-eiter/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/137097/

p261-eiter/. ACM order number 475920.

Kafer:1992:RTC

[597] Wolfgang Käfer and Harald Schöning. Realizing a temporal complex-object data model. In Stonebraker [2268], pages 266–275. ISBN 0-89791-521-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? URL http://

www.acm.org/pubs/articles/proceedings/
mod/130283/p266-kafer/p266-kafer.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/130283/
p266-kafer/.

Chomicki:1992:RTI

[598] Jan Chomicki. Real-time integrity constraints. In ACM [2269], pages 274–

282. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26 1992. URL http://www.

acm.org/pubs/articles/proceedings/
pods/137097/p274-chomicki/p274-
chomicki.pdf;   http://www.acm.
org/pubs/citations/proceedings/
pods/137097/p274-chomicki/;   http:
//www.acm.org:80/pubs/citations/
proceedings/pods/137097/p274-chomicki/

. ACM order number 475920.

Chang:1992:MAS

[599] J. W. Chang and Y. J. Lee. Multikey access scheme based on term discrimination and signature clustering. In Kim et al. [2267], pages 276–282. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Biliris:1992:PTD

[600] Alexandros Biliris. The performance of three database storage structures for managing large objects. In Stonebraker [2268], pages 276–285. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/
                                        p276-biliris/p276-biliris.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/130283/p276-biliris/

.

Kitsuregawa:1992:OFD

[601] M. Kitsuregawa, M. Nakano, and M. Takagi. Overview of functional disk system. In Kim et al. [2267], pages 283–290. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Bell:1992:IDD

[602] Colin Bell, Anil Nerode, Raymond T. Ng, and V. S. Subrahmanian. Implementing deductive databases by linear programming. In ACM [2269], pages 283–292. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26

                                        1992. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        137097/p283-bell/p283-bell.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/137097/p283-bell/
                                        ;   http://www.acm.org:80/pubs/

citations/proceedings/pods/137097/

p283-bell/. ACM order number 475920.

Wang:1992:ERL

[603] Yun Wang. Experience from a real life query optimizer. In Stonebraker [2268], page 286. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p286-wang/p286-wang.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p286-wang/.

Rengarajan:1992:RVS

[604] T. K. Rengarajan. Rdb/VMS support for multi-media databases. In Stonebraker [2268], page 287. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p287-rengarajan/p287-rengarajan.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/130283/
p287-rengarajan/.

Mehrotra:1992:CCP

[605] Sharad Mehrotra, Rajeev Rastogi, Yuri Breitbart, Henry F. Korth, and Avi Silberschatz. The concurrency control problem in multidatabases: characteristics and solutions. In Stonebraker [2268], pages 288–297. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p288-mehrotra/p288-mehrotra.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/130283/p288-mehrotra/

.

Scheuermann:1992:ATD

[606] P. Scheuermann, G. Weikum, and P. Zabback. Automatic tuning of data placement and load balancing in disk arrays. In Kim et al. [2267], pages 291–301. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Ginsburg:1992:PMR

[607] Seymour Ginsburg and Xiaoyang Wang. Pattern matching by Rsoperations: towards a unified approach to querying sequenced data. In ACM [2269], pages 293–300. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3

                                        A26 1992. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        137097/p293-ginsburg/p293-ginsburg.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/137097/
                                        p293-ginsburg/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/137097/p293-ginsburg/. ACM

order number 475920.

Shasha:1992:SRG

[608] Dennis Shasha, Eric Simon, and Patrick Valduriez. Simple rational guidance for chopping up transactions. In Stonebraker [2268], pages 298–307. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/
                                        p298-shasha/p298-shasha.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/130283/p298-shasha/.

Srivastava:1992:PCS

[609] Divesh Srivastava and Raghu Ramakrishnan. Pushing constraint selections.

In ACM [2269], pages 301–315. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3

A26 1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p301-srivastava/p301-srivastava.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/137097/
p301-srivastava/;   http://www.
acm.org:80/pubs/citations/proceedings/
pods/137097/p301-srivastava/. ACM

order number 475920.

Kiyoki:1992:IPQ

[610] Y. Kiyoki, T. Kurosawa, P. Liu, and K. Kato. Implementation of a parallel query processing system supporting application-specific database operations. In Kim et al. [2267], pages 302–309. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Rahm:1992:PEE

[611] Erhard Rahm. Performance evaluation of extended storage architectures for transaction processing. In Stonebraker [2268], pages 308–317. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p308-rahm/p308-rahm.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p308-rahm/.

Bonner:1992:CRM

[612] Anthony J. Bonner. The complexity of reusing and modifying rulebases. In ACM [2269], pages 316–330. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3

A26 1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/

                                        137097/p316-bonner/p316-bonner.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/137097/
                                        p316-bonner/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/137097/p316-bonner/. ACM

order number 475920.

Cowley:1992:SDM

[613] Paula J. Cowley. Scientific data management: real-world issues and requirements. In Stonebraker [2268], page 318. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/
                                        p318-cowley/p318-cowley.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/130283/p318-cowley/.

Chou:1992:DIP

[614] Hong-Tai Chou. The design and implementation of persistent transactions in an object database system. In Stonebraker [2268], page 319. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/
                                        p319-chou/p319-chou.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/130283/p319-chou/.

Harris:1992:HPM

[615] Craig Harris, Madhu Reddy, and Carl Woolf. A high performance multiversion concurrency control protocol for object databases. In Stonebraker [2268], page 320. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/

p320-harris/p320-harris.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p320-harris/.

Terry:1992:CQA

[616] Douglas Terry, David Goldberg, David Nichols, and Brian Oki. Continuous queries over append-only databases. In Stonebraker [2268], pages 321–330. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p321-terry/p321-terry.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p321-terry/.

Srinivasan:1992:CBL

[617] V. Srinivasan and Michael J. Carey. Compensation-based on-line query processing. In Stonebraker [2268], pages 331–340. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? URL http://

www.acm.org/pubs/articles/proceedings/
mod/130283/p331-srinivasan/p331-
srinivasan.pdf;   http://www.acm.
org/pubs/citations/proceedings/
mod/130283/p331-srinivasan/.

vanderMeyden:1992:CQI

[618] R. van der Meyden. The complexity of querying indefinite data about linearly ordered domains. In ACM [2269], pages 331–345. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3

A26 1992. URL http://www.acm.
org/pubs/citations/proceedings/
pods/137097/p331-van_der_meyden/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/137097/

p331-van_der_meyden/. ACM order

number 475920.

Haas:1992:SSP

[619] Peter J. Haas and Arun N. Swami. Sequential sampling procedures for query size estimation. In Stonebraker [2268], pages 341–350. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p341-haas/p341-haas.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p341-haas/.

Ross:1992:RRN

[620] Kenneth A. Ross. Relations with relation names as arguments: algebra and calculus. In ACM [2269], pages 346–353. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26

1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p346-ross/p346-ross.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/137097/p346-ross/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/137097/

p346-ross/. ACM order number 475920.

Lomet:1992:AMC

[621] David Lomet and Betty Salzberg. Access method concurrency with recovery. In Stonebraker [2268], pages 351–360. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p351-lomet/p351-lomet.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p351-lomet/.

Gupta:1992:MST

[622] Ashish Gupta and Inderpal Singh Mumick. Magic-sets transformation in nonrecursive systems. In ACM [2269], pages 354–367. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26

1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p354-gupta/p354-gupta.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/137097/p354-gupta/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/137097/

p354-gupta/. ACM order number 475920.

Mohan:1992:ACI

[623] C. Mohan and Inderpal Narang. Algorithms for creating indexes for very large tables without quiescing updates. In Stonebraker [2268], pages 361–370. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p361-mohan/p361-mohan.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p361-mohan/.

Morishita:1992:ACP

[624] Shinichi Morishita. Avoiding Cartesian products in programs for multiple joins (extended abstract). In ACM [2269], pages 368–379. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3

A26 1992. URL http://www.acm.org/
pubs/articles/proceedings/pods/
137097/p368-morishita/p368-morishita.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/137097/
p368-morishita/;   http://www.acm.

                                        org:80/pubs/citations/proceedings/
                                        pods/137097/p368-morishita/. ACM

order number 475920.

Mohan:1992:AIE

[625] C. Mohan and Frank Levine. ARIES/ IM: an efficient and high concurrency index management method using write-ahead logging. In Stonebraker [2268], pages 371–380. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/
                                        p371-mohan/p371-mohan.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/130283/p371-mohan/.

Jakobsson:1992:TBT

[626] Håkan Jakobsson. On tree-based techniques for query evaluation. In ACM [2269], pages 380–392. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3

                                        A26 1992. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        137097/p380-jakobsson/p380-jakobsson.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/137097/
                                        p380-jakobsson/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/137097/p380-jakobsson/. ACM

order number 475920.

Carter:1992:EIM

[627] Fred Carter. Extending INGRES with methods and triggers. In Stonebraker [2268], page 381. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/
                                        p381-carter/p381-carter.pdf;   http:

//www.acm.org/pubs/citations/proceedings/
mod/130283/p381-carter/.

Bigelow:1992:IGC

[628] Richard Bigelow. Implementation of general constraints in SIM. In Stonebraker [2268], page 382. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p382-bigelow/p382-bigelow.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/130283/p382-bigelow/

.

Cluet:1992:GFO

[629] Sophie Cluet and Claude Delobel. A general framework for the optimization of object-oriented queries. In Stonebraker [2268], pages 383–392. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p383-cluet/p383-cluet.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p383-cluet/.

Kifer:1992:QOO

[630] Michael Kifer, Won Kim, and Yehoshua Sagiv. Querying objectoriented databases. In Stonebraker [2268], pages 393–402. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

????  URL http://www.acm.org/
pubs/articles/proceedings/mod/130283/
p393-kifer/p393-kifer.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/130283/p393-kifer/.

Orenstein:1992:QPO

[631] Jack Orenstein, Sam Haradhvala, Benson Margulies, and Don Sakahara. Query processing in the ObjectStore database system. In Stonebraker [2268], pages 403–412. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/
                                        p403-orenstein/p403-orenstein.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/130283/p403-orenstein/

.

Melmon:1992:SOS

[632] Paul Melmon. The Sybase Open Server. In Stonebraker [2268], page 413. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/
                                        p413-melmon/p413-melmon.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/130283/p413-melmon/.

Newmann:1992:MVI

[633] Scott Newmann. Multi-vendor interoperability through SQL access. In Stonebraker [2268], page 414. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        ????  URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/130283/
                                        p414-newmann/p414-newmann.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/130283/p414-newmann/

.

Maier:1993:CO

[634] David Maier and Bennet Vance. A call to order. In ACM [2270], pages 1–16. ISBN 0-89791-593-

3. LCCN QA 76.9 D3 A26

1993. URL http://www.acm.org/
pubs/articles/proceedings/pods/
153850/p1-maier/p1-maier.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/153850/p1-maier/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/153850/
p1-maier/.

Stonebraker:1993:SSB

[635] Michael Stonebraker, Jim Frew, Kenn Gardels, and Jeff Meredith. The SE-QUOIA 2000 storage benchmark. In Buneman and Jajodia [2271], pages 2–

11. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p2-stonebraker/p2-stonebraker.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/170035/p2-stonebraker/

.

Carey:1993:B

[636] Michael J. Carey, David J. De-Witt, and Jeffrey F. Naughton. The 007 Benchmark. In Buneman and Jajodia [2271], pages 12–

21. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p12-carey/p12-carey.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p12-carey/.

VandenBussche:1993:RPR

[637] Jan Van den Bussche, Dirk Van Gucht, and Gottfried Vossen. Reflective programming in the relational algebra.

In ACM [2270], pages 17–25. ISBN 0-89791-593-3. LCCN QA 76.9 D3

A26 1993. URL http://www.acm.org/
pubs/articles/proceedings/pods/
153850/p17-van_den_bussche/p17-
van_den_bussche.pdf;   http://www.
acm.org/pubs/citations/proceedings/
pods/153850/p17-van_den_bussche/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/153850/
p17-van_den_bussche/.

Leutenegger:1993:MST

[638] Scott T. Leutenegger and Daniel Dias. A modeling study of the TPC-C benchmark. In Buneman and Jajodia [2271], pages 22–31. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p22-leutenegger/p22-leutenegger.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/170035/
p22-leutenegger/.

Wong:1993:NFC

[639] Limsoon Wong. Normal forms and conservative properties for query languages over collection types. In ACM [2270], pages 26–36. ISBN 0-89791-593-3. LCCN QA 76.9 D3 A26

1993. URL http://www.acm.org/
pubs/articles/proceedings/pods/
153850/p26-wong/p26-wong.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/153850/p26-wong/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/153850/
p26-wong/.

Abiteboul:1993:MR

[640] Serge Abiteboul, Georg Lausen, Heinz Uphoff, and Emmanuel Waller. Methods and rules. In Buneman and Jajodia [2271], pages 32–41. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p32-abiteboul/p32-abiteboul.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/170035/p32-abiteboul/

.

Libkin:1993:SRQ

[641] Leonid Libkin and Limsoon Wong. Semantic representations and query languages for OR-sets. In ACM [2270], pages 37–48. ISBN 0-89791-593-3. LCCN QA 76.9 D3 A26

1993. URL http://www.acm.org/
pubs/articles/proceedings/pods/
153850/p37-libkin/p37-libkin.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/153850/p37-libkin/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/153850/
p37-libkin/.

Brant:1993:ISR

[642] David A. Brant and Daniel P. Miranker. Index support for rule activation. In Buneman and Jajodia [2271], pages 42–48. ISBN 0-89791-592-

5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2 1993. URL http://www.acm.

org/pubs/articles/proceedings/mod/
170035/p42-brant/p42-brant.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/170035/p42-brant/

.

Grumbach:1993:TTA

[643] Stéphane Grumbach and Tova Milo. Towards tractable algebras for bags. In ACM [2270], pages 49–58. ISBN 0-89791-593-3. LCCN QA 76.9 D3

A26 1993. URL http://www.acm.org/
pubs/articles/proceedings/pods/
153850/p49-grumbach/p49-grumbach.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/153850/
p49-grumbach/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/153850/p49-grumbach/.

Gupta:1993:LVG

[644] Ashish Gupta and Jennifer Widom. Local verification of global integrity constraints in distributed databases. In Buneman and Jajodia [2271], pages 49–

58. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p49-gupta/p49-gupta.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p49-gupta/.

Pang:1993:PPH

[645] Hwee Hwa Pang, Michael J. Carey, and Miron Livny. Partially preemptible hash joins. In Buneman and Jajodia [2271], pages 59–68. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p59-pang/p59-pang.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/170035/p59-pang/.

Chaudhuri:1993:ORC

[646] Surajit Chaudhuri and M. Y. Vardi. Optimization of real conjunctive queries. In ACM [2270], pages 59–70. ISBN 0-89791-593-3. LCCN QA 76.9

D3 A26 1993. URL http://www.acm.
org/pubs/articles/proceedings/pods/
153850/p59-chaudhuri/p59-chaudhuri.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/153850/
p59-chaudhuri/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/153850/p59-chaudhuri/.

Lo:1993:OPA

[647] Ming-Ling Lo, Ming-Syan Syan Chen, C. V. Ravishankar, and Philip S. Yu. On optimal processor allocation to support pipelined hash joins. In Buneman and Jajodia [2271], pages 69–

78. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p69-lo/p69-lo.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/170035/p69-lo/.

Revesz:1993:STC

[648] Peter Z. Revesz. On the semantics of theory change: Arbitration between old and new information. In ACM [2270], pages 71–82. ISBN 0-89791-593-3. LCCN QA 76.9 D3 A26

1993. URL http://www.acm.org/
pubs/articles/proceedings/pods/
153850/p71-revesz/p71-revesz.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/153850/p71-revesz/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/153850/
p71-revesz/.

Sun:1993:IAS

[649] Wei Sun, Yibei Ling, Naphtali Rishe, and Yi Deng. An instant and accurate size estimation method for joins and selections in a retrievalintensive environment. In Buneman and Jajodia [2271], pages 79–

88. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p79-sun/p79-sun.pdf;   http://www.
                                        acm.org/pubs/citations/proceedings/
                                        mod/170035/p79-sun/.

Raz:1993:ECO

[650] Yoav Raz. Extended commitment ordering, or guaranteeing global serializability by applying commitment order selectively to global transactions. In ACM [2270], pages 83–96. ISBN 0-89791-593-3. LCCN QA 76.9 D3

                                        A26 1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        153850/p83-raz/p83-raz.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        pods/153850/p83-raz/;   http://
                                        www.acm.org:80/pubs/citations/proceedings/
                                        pods/153850/p83-raz/.

Meseguer:1993:LSO

[651] José Meseguer and Xiaolei Qian. A logical semantics for objectoriented databases. In Buneman and Jajodia [2271], pages 89–98. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p89-meseguer/p89-meseguer.pdf;

http://www.acm.org/pubs/citations/
proceedings/mod/170035/p89-meseguer/

.

Rastogi:1993:CNS

[652] Rajeev Rastogi, Sharad Mehrotra, Yuri Breitbart, Henry F. Korth, and Avi Silberschatz. On correctness of non-serializable executions. In ACM [2270], pages 97–108. ISBN 0-89791-593-3. LCCN QA 76.9 D3 A26

1993. URL http://www.acm.org/
pubs/articles/proceedings/pods/
153850/p97-rastogi/p97-rastogi.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/153850/
p97-rastogi/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/153850/p97-rastogi/.

Anwar:1993:NPR

[653] E. Anwar, L. Maugis, and S. Chakravarthy. A new perspective on rule support for object-oriented databases. In Buneman and Jajodia [2271], pages 99–

108. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p99-anwar/p99-anwar.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p99-anwar/.

Ananthanarayanan:1993:UCE

[654] R. Ananthanarayanan, V. Gottemukkala, W. Kaefer, T. J. Lehman, and H. Pirahesh. Using the coexistence approach to achieve combined functionality of object-oriented and relational systems. In Buneman and Jajodia [2271], pages 109–

118. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p109-ananthanarayanan/p109-ananthanarayanan.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/170035/
                                        p109-ananthanarayanan/.

Levy:1993:EQR

[655] Alon Levy, Inderpal Singh Mumick, Yehoshua Sagiv, and Oded Shmueli. Equivalence, query-reachability and satisfiability in Datalog extensions. In ACM [2270], pages 109–122. ISBN 0-89791-593-3. LCCN QA 76.9 D3

                                        A26 1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        153850/p109-levy/p109-levy.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/153850/p109-levy/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/153850/
                                        p109-levy/.

Shatdal:1993:USV

[656] Ambuj Shatdal and Jeffrey F. Naughton. Using shared virtual memory for parallel join processing. In Buneman and Jajodia [2271], pages 119–128. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p119-shatdal/p119-shatdal.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/170035/p119-shatdal/

.

Morishita:1993:AFT

[657] Shinichi Morishita. An alternating fixpoint tailored to magic programs. In ACM [2270], pages 123–134. ISBN 0-89791-593-3. LCCN QA 76.9 D3

A26 1993. URL http://www.acm.org/
pubs/articles/proceedings/pods/
153850/p123-morishita/p123-morishita.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/153850/
p123-morishita/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/153850/p123-morishita/.

Tomasic:1993:CDS

[658] Anthony Tomasic and Hector Garcia-Molina. Caching and database scaling in distributed shared-nothing information retrieval systems. In Buneman and Jajodia [2271], pages 129–

138. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p129-tomasic/p129-tomasic.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/170035/p129-tomasic/

.

Chaudhuri:1993:FNE

[659] Surajit Chaudhuri. Finding nonrecursive envelopes for Datalog predicates. In ACM [2270], pages 135–146. ISBN 0-89791-593-3. LCCN QA 76.9 D3

A26 1993. URL http://www.acm.org/
pubs/articles/proceedings/pods/
153850/p135-chaudhuri/p135-chaudhuri.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/153850/
p135-chaudhuri/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/153850/p135-chaudhuri/.

Mohan:1993:EFM

[660] C. Mohan and Inderpal Narang. An efficient and flexible method for archiving a data base. In Buneman and Jajodia [2271], pages 139–146. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p139-mohan/p139-mohan.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/170035/p139-mohan/.

Derr:1993:DIG

[661] Marcia A. Derr, Shinichi Morishita, and Geoffrey Phipps. Design and implementation of the gluenail database system. In Buneman and Jajodia [2271], pages 147–156. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p147-derr/p147-derr.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/170035/p147-derr/.

Schaerf:1993:NMN

[662] Marco Schaerf. Negation and minimality in non-Horn databases. In ACM [2270], pages 147–157. ISBN 0-89791-593-3. LCCN QA 76.9 D3

                                        A26 1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        153850/p147-schaerf/p147-schaerf.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/153850/
                                        p147-schaerf/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/153850/p147-schaerf/.

Gupta:1993:MVI

[663] Ashish Gupta, Inderpal Singh Mumick, and V. S. Subrahmanian. Maintaining views incrementally. In Buneman and Jajodia [2271], pages 157–

166. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p157-gupta/p157-gupta.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p157-gupta/.

Eiter:1993:CAV

[664] Thomas Eiter and Georg Gottlob. Complexity aspects of various semantics for disjunctive databases. In ACM [2270], pages 158–167. ISBN 0-89791-593-3. LCCN QA 76.9 D3

A26 1993. URL http://www.acm.org/
pubs/articles/proceedings/pods/
153850/p158-eiter/p158-eiter.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/153850/p158-eiter/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/153850/
p158-eiter/.

Ramakrishnan:1993:ICD

[665] Raghu Ramakrishnan, Divesh Srivastava, S. Sudarshan, and Praveen Seshadri. Implementation of the CORAL deductive database system. In Buneman and Jajodia [2271], pages 167–

176. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p167-ramakrishnan/p167-ramakrishnan.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/170035/
p167-ramakrishnan/.

Chen:1993:QEU

[666] Weidong Chen and David S. Warren. Query evaluation under the well-founded semantics. In ACM [2270], pages 168–179. ISBN 0-89791-593-3. LCCN QA 76.9 D3 A26

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        153850/p168-chen/p168-chen.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/153850/p168-chen/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/153850/
                                        p168-chen/.

Kolodner:1993:AIG

[667] Elliot K. Kolodner and William E. Weihl. Atomic incremental garbage collection and recovery for a large stable heap. In Buneman and Jajodia [2271], pages 177–186. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2 1993. URL http://www.acm.

                                        org/pubs/articles/proceedings/mod/
                                        170035/p177-kolodner/p177-kolodner.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/170035/
                                        p177-kolodner/.

VanGelder:1993:MJS

[668] Allen Van Gelder. Multiple join size estimation by virtual domains (extended abstract). In ACM [2270], pages 180–189. ISBN 0-89791-593-3. LCCN QA 76.9 D3 A26

                                        1993. URL http://www.acm.org/
                                        pubs/citations/proceedings/pods/
                                        153850/p180-van_gelder/.

Keen:1993:PEE

[669] John S. Keen and William J. Dally. Performance evaluation of ephemeral logging. In Buneman and Jajodia [2271], pages 187–196.

ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p187-keen/p187-keen.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p187-keen/.

Haas:1993:FPE

[670] Peter J. Haas, Jeffrey F. Naughton, S. Seshadri, and Arun N. Swami. Fixed-precision estimation of join selectivity. In ACM [2270], pages 190–201. ISBN 0-89791-593-3. LCCN QA 76.9

D3 A26 1993. URL http://www.acm.
org/pubs/articles/proceedings/pods/
153850/p190-haas/p190-haas.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/153850/p190-haas/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/153850/
p190-haas/.

Hong:1993:RTT

[671] D. Hong, T. Johnson, and S. Chakravarthy. Real-time transaction scheduling: a cost conscious approach. In Buneman and Jajodia [2271], pages 197–

206. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p197-hong/p197-hong.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p197-hong/.

Chomicki:1993:FCT

[672] Jan Chomicki and Damian Niwiński. On the feasibility of checking temporal integrity constraints. In ACM [2270], pages 202–213. ISBN 0-89791-593-3. LCCN QA 76.9 D3 A26

1993. URL http://www.acm.org/
pubs/articles/proceedings/pods/
153850/p202-chomicki/p202-chomicki.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/153850/
p202-chomicki/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/153850/p202-chomicki/.

Agrawal:1993:MAR

[673] Rakesh Agrawal, Tomasz Imieliński, and Arun Swami. Mining association rules between sets of items in large databases. In Buneman and Jajodia [2271], pages 207–216. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p207-agrawal/p207-agrawal.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/170035/p207-agrawal/

.

Pagel:1993:TAR

[674] Bernd-Uwe Pagel, Hans-Werner Six, Heinrich Toben, and Peter Widmayer. Towards an analysis of range query performance in spatial data structures. In ACM [2270], pages 214–221. ISBN 0-89791-593-3. LCCN QA 76.9 D3

A26 1993. URL http://www.acm.org/
pubs/articles/proceedings/pods/
153850/p214-pagel/p214-pagel.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/153850/p214-pagel/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/153850/
p214-pagel/.

Borgida:1993:LDD

[675] Alex Borgida and Ronald J. Brachman. Loading data into description reasoners. In Buneman and Jajodia [2271], pages 217–226. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p217-borgida/p217-borgida.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/170035/p217-borgida/

.

Nodine:1993:BEG

[676] Mark H. Nodine, Michael T. Goodrich, and Jeffrey Scott Vitter. Blocking for external graph searching. In ACM [2270], pages 222–232. ISBN 0-89791-593-3. LCCN QA 76.9 D3

A26 1993. URL http://www.acm.org/
pubs/articles/proceedings/pods/
153850/p222-nodine/p222-nodine.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/153850/
p222-nodine/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/153850/p222-nodine/.

Wang:1993:TMA

[677] X. Sean Wang, Sushil Jajodia, and V. S. Subrahmanian. Temporal modules: an approach toward federated temporal databases. In Buneman and Jajodia [2271], pages 227–236. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p227-wang/p227-wang.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p227-wang/.

Kanellakis:1993:IDM

[678] Paris C. Kanellakis, Sridhar Ramaswamy, Darren E. Vengroff, and Jeffrey S. Vitter. Indexing for data models with constraints and classes (extended abstract). In ACM [2270], pages 233–243. ISBN 0-89791-593-3. LCCN QA 76.9 D3 A26

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        153850/p233-kanellakis/p233-kanellakis.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/153850/
                                        p233-kanellakis/;   http://www.
                                        acm.org:80/pubs/citations/proceedings/
                                        pods/153850/p233-kanellakis/.

Brinkhoff:1993:EPS

[679] Thomas Brinkhoff, Hans-Peter Kriegel, and Bernhard Seeger. Efficient processing of spatial joints using R-trees. In Buneman and Jajodia [2271], pages 237–246. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

                                        no.2 1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p237-brinkhoff/p237-brinkhoff.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/170035/p237-brinkhoff/

.

Harel:1993:CRR

[680] D. Harel and T. Hirst. Completeness results for recursive data bases. In ACM [2270], pages 244–252. ISBN 0-89791-593-3. LCCN QA 76.9 D3 A26 1993.

Hirst:1993:CRR

[681] Tirza Hirst and David Harel. Completeness results for recursive data bases. In ACM [2270], pages 244–252.

ISBN 0-89791-593-3. LCCN QA 76.9

D3 A26 1993. URL http://www.acm.
org/pubs/articles/proceedings/pods/
153850/p244-hirst/p244-hirst.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/153850/p244-hirst/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/153850/
p244-hirst/.

Ishikawa:1993:ESF

[682] Yoshiharu Ishikawa, Hiroyuki Kitagawa, and Nobuo Ohbo. Evaluation of signature files as set access facilities in OODBs. In Buneman and Jajodia [2271], pages 247–

256. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p247-ishikawa/p247-ishikawa.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/170035/p247-ishikawa/

.

Escobar-Molano:1993:STC

[683] Martha Escobar-Molano, Richard Hull, and Dean Jacobs. Safety and translation of calculus queries with scalar functions. In ACM [2270], pages 253–264. ISBN 0-89791-593-3. LCCN QA 76.9 D3 A26

1993. URL http://www.acm.org/
pubs/articles/proceedings/pods/
153850/p253-escobar-molano/p253-
escobar-molano.pdf;   http://www.
acm.org/pubs/citations/proceedings/
pods/153850/p253-escobar-molano/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/153850/
p253-escobar-molano/.

Curewitz:1993:PPD

[684] Kenneth M. Curewitz, P. Krishnan, and Jeffrey Scott Vitter. Practical prefetching via data compression. In Buneman and Jajodia [2271], pages 257–266. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

                                        no.2 1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p257-curewitz/p257-curewitz.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/170035/p257-curewitz/

.

Denninghoff:1993:DMS

[685] Karl Denninghoff and Victor Vianu. Database method schemas and object creation. In ACM [2270], pages 265–275. ISBN 0-89791-593-3. LCCN QA 76.9 D3 A26

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        153850/p265-denninghoff/p265-denninghoff.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/153850/
                                        p265-denninghoff/;   http://www.
                                        acm.org:80/pubs/citations/proceedings/
                                        pods/153850/p265-denninghoff/.

Hellerstein:1993:PMO

[686] Joseph M. Hellerstein and Michael Stonebraker. Predicate migration: optimizing queries with expensive predicates. In Buneman and Jajodia [2271], pages 267–276. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p267-hellerstein/p267-hellerstein.
                                        pdf;   http://www.acm.org/pubs/

citations/proceedings/mod/170035/
p267-hellerstein/.

Wong:1993:CBS

[687] Man H. Wong and Divyakant Agrawal. Context-based synchronization: An approach beyond semantics for concurrency control. In ACM [2270], pages 276–287. ISBN 0-89791-593-3. LCCN QA 76.9 D3 A26

1993. URL http://www.acm.org/
pubs/articles/proceedings/pods/
153850/p276-wong/p276-wong.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/153850/p276-wong/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/153850/
p276-wong/.

Guting:1993:SOS

[688] Ralf Hartmut Güting. Secondorder signature: a tool for specifying data models, query processing, and optimization. In Buneman and Jajodia [2271], pages 277–

286. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p277-guting/p277-guting.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p277-guting/.

Blakeley:1993:EBO

[689] José A. Blakeley, William J. McKenna, and Goetz Graefe. Experiences building the open OODB query optimizer. In Buneman and Jajodia [2271], pages 287–296. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

no.2 1993. URL http://www.acm.org/

                                        pubs/articles/proceedings/mod/170035/
                                        p287-blakeley/p287-blakeley.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/170035/p287-blakeley/

.

Rastogi:1993:SHO

[690] Rajeev Rastogi, Henry F. Korth, and Abraham Silberschatz. Strict histories in object-based database systems. In ACM [2270], pages 288–299. ISBN 0-89791-593-3. LCCN QA 76.9 D3

                                        A26 1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        153850/p288-rastogi/p288-rastogi.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/153850/
                                        p288-rastogi/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/153850/p288-rastogi/.

ONeil:1993:LKP

[691] Elizabeth J. O’Neil, Patrick E. O’Neil, and Gerhard Weikum. The LRU-K page replacement algorithm for database disk buffering. In Buneman and Jajodia [2271], pages 297–

306. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p297-o_neil/p297-o_neil.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/170035/p297-o_neil/.

Schek:1993:TUT

[692] Hans-Jörg Schek, Gerhard Weikum, and Haiyan Ye. Towards a unified theory of concurrency control and recovery. In ACM [2270], pages 300–311. ISBN 0-89791-593-3. LCCN QA 76.9 D3 A26

1993. URL http://www.acm.org/
pubs/articles/proceedings/pods/
153850/p300-schek/p300-schek.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/153850/p300-schek/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/153850/
p300-schek/.

Orji:1993:DDM

[693] Cyril U. Orji and Jon A. Solworth. Doubly distorted mirrors. In Buneman and Jajodia [2271], pages 307–

316. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p307-orji/p307-orji.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p307-orji/.

Hou:1993:CRA

[694] Robert Y. Hou and Yale N. Patt. Comparing rebuild algorithms for mirrored and RAID5 disk arrays. In Buneman and Jajodia [2271], pages 317–

326. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p317-hou/p317-hou.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/170035/p317-hou/.

Litwin:1993:LLH

[695] Witold Litwin, Marie-Anne Neimat, and Donovan A. Schneider. LH: Linear Hashing for distributed files. In Buneman and Jajodia [2271], pages 327–336. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

no.2 1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p327-litwin/p327-litwin.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p327-litwin/.

Johnson:1993:LUD

[696] Theodore Johnson and Padmashree Krishna. Lazy updates for distributed search structure. In Buneman and Jajodia [2271], pages 337–

346. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p337-johnson/p337-johnson.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/170035/p337-johnson/

.

Li:1993:ALP

[697] Jianzhong Li, Doron Rotem, and Jaideep Srivastava. Algorithms for loading parallel grid files. In Buneman and Jajodia [2271], pages 347–

356. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p347-li/p347-li.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/170035/p347-li/.

Vadaparty:1993:TUV

[698] K. Vadaparty, Y. A. Aslandogan, and G. Ozsoyoglu. Towards a unified visual database access. In Buneman and Jajodia [2271], pages 357–

366. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p357-vadaparty/p357-vadaparty.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/170035/p357-vadaparty/

.

Watters:1993:IRR

[699] Aaron Watters. Interpreting a reconstructed relational calculus (extended abstract). In Buneman and Jajodia [2271], pages 367–376. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p367-watters/p367-watters.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/170035/p367-watters/

.

Beeri:1993:PAR

[700] Catriel Beeri and Tova Milo. On the power of algebras with recursion. In Buneman and Jajodia [2271], pages 377–387. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

no.2 1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p377-beeri/p377-beeri.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p377-beeri/.

Alonso:1993:DSI

[701] Rafael Alonso and Henry F. Korth. Database system issues in nomadic computing. In Buneman and Jajodia [2271], pages 388–392. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p388-alonso/p388-alonso.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/170035/p388-alonso/.

Dayal:1993:TGT

[702] Umesh Dayal, Hector Garcia-Molina, Mei Hsu, Ben Kao, and Ming-Chien Shan. Third generation TP monitors: a database challenge. In Buneman and Jajodia [2271], pages 393–

397. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p393-dayal/p393-dayal.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/170035/p393-dayal/.

Egenhofer:1993:WSA

[703] Max J. Egenhofer. What’s special about spatial?: database requirements for vehicle navigation in geographic space. In Buneman and Jajodia [2271], pages 398–402. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p398-egenhofer/p398-egenhofer.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/170035/p398-egenhofer/

.

Ordille:1993:DCG

[704] Joann J. Ordille and Barton P. Miller. Database challenges in global information systems. In Buneman and Jajodia [2271], pages 403–407. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p403-ordille/p403-ordille.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/170035/p403-ordille/

.

Zdonik:1993:IDS

[705] Stanley B. Zdonik. Incremental database systems: databases from the ground up. In Buneman and Jajodia [2271], pages 408–412. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p408-zdonik/p408-zdonik.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p408-zdonik/.

Carey:1993:THD

[706] Michael J. Carey, Laura M. Haas, and Miron Livny. Tapes hold data, too: challenges of tuples on tertiary store. In Buneman and Jajodia [2271], pages 413–417. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

no.2 1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p413-carey/p413-carey.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p413-carey/.

Jagadish:1993:IMD

[707] H. V. Jagadish. Issues in multimedia databases. In Buneman and Jajodia [2271], page 419. ISBN 0-89791-592-

5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2 1993. URL http://www.acm.

                                        org/pubs/articles/proceedings/mod/
                                        170035/p419-jagadish/p419-jagadish.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/170035/
                                        p419-jagadish/.

Motro:1993:WTA

[708] Amihai Motro. What to teach about databases. In Buneman and Jajodia [2271], page 420. ISBN 0-89791-592-

5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2 1993. URL http://www.acm.

                                        org/pubs/articles/proceedings/mod/
                                        170035/p420-motro/p420-motro.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/170035/p420-motro/

.

Shan:1993:PAD

[709] Ming-Chien Shan. Pegasus architecture and design principles. In Buneman and Jajodia [2271], pages 422–

425. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p422-shan/p422-shan.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/170035/p422-shan/.

Bukhres:1993:ISB

[710] Omran Bukhres, Jiansan Chen, and Rob Pezzoli. An InterBase system at BNR. In Buneman and Jajodia [2271], pages 426–429. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p426-bukhres/p426-bukhres.pdf;

http://www.acm.org/pubs/citations/
proceedings/mod/170035/p426-bukhres/

.

Klein:1993:ODC

[711] Johannes Klein and Francis Upton. Open DECdtm: constraint based transaction management. In Buneman and Jajodia [2271], pages 430–

433. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p430-klein/p430-klein.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p430-klein/.

Wiederhold:1993:III

[712] Gio Wiederhold. Intelligent integration of information. In Buneman and Jajodia [2271], pages 434–

437. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p434-wiederhold/p434-wiederhold.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/170035/
p434-wiederhold/.

Vieille:1993:DOO

[713] Laurent Vieille. A deductive and object-oriented database system: why and how? In Buneman and Jajodia [2271], page 438. ISBN 0-89791-592-

5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2 1993. URL http://www.acm.

org/pubs/articles/proceedings/mod/
170035/p438-vieille/p438-vieille.
pdf;   http://www.acm.org/pubs/

                                        citations/proceedings/mod/170035/
                                        p438-vieille/.

Stonebraker:1993:MD

[714] Michael Stonebraker. The Miro DBMS. In Buneman and Jajodia [2271], page

439. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p439-stonebraker/p439-stonebraker.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/170035/
                                        p439-stonebraker/.

Velez:1993:MTM

[715] Fernando Vélez. Modularity and tuning mechanisms in the O_{2} System. In Buneman and Jajodia [2271], page

440. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p440-velez/p440-velez.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/170035/p440-velez/.

Wade:1993:SLV

[716] Andrew E. Wade. Single logical view over enterprise-wide distributed databases. In Buneman and Jajodia [2271], page 441. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p441-wade/p441-wade.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/170035/p441-wade/.

Mohan:1993:IRD

[717] C. Mohan. IBM’s relational DBMS products: features and technologies. In Buneman and Jajodia [2271], pages 445–448. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

no.2 1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p445-mohan/p445-mohan.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p445-mohan/.

Fushimi:1993:GCD

[718] Shinya Fushimi and Masaru Kitsuregawa. GREO: a commercial database processor based on a pipelined hardware sorter. In Buneman and Jajodia [2271], pages 449–

452. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p449-fushimi/p449-fushimi.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/170035/p449-fushimi/

.

Tseng:1993:PDP

[719] Emy Tseng and David Reiner. Parallel database processing on the KSR1 computer. In Buneman and Jajodia [2271], pages 453–455. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p453-tseng/p453-tseng.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p453-tseng/.

Jin:1993:CCR

[720] W. Woody Jin, Marek Rusinkiewicz, Linda Ness, and Amit Sheth. Concurrency control and recovery of multidatabase work flows in telecommunication applications. In Buneman and Jajodia [2271], pages 456–459. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p456-jin/p456-jin.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/170035/p456-jin/.

Sherman:1993:AED

[721] Mark Sherman. Architecture of the Encina distributed transaction processing family. In Buneman and Jajodia [2271], pages 460–463. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p460-sherman/p460-sherman.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/170035/p460-sherman/

.

Colton:1993:RDD

[722] Malcolm Colton. Replicated data in a distributed environment. In Buneman and Jajodia [2271], pages 464–

466. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p464-colton/p464-colton.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p464-colton/.

Singhal:1993:DOO

[723] Anoop Singhal, Robert M. Arlein, and Chi-Yuan Lo. DDB: an object oriented design data manager for VLSI CAD. In Buneman and Jajodia [2271], pages 467–470. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

no.2 1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p467-singhal/p467-singhal.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/170035/p467-singhal/

.

Nassif:1993:IAM

[724] Rodolphe Nassif and Don Mitchusson. Issues and approaches for migration/cohabitation between legacy and new systems. In Buneman and Jajodia [2271], pages 471–474. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p471-nassif/p471-nassif.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p471-nassif/.

Thieman:1993:IDN

[725] James R. Thieman. The international directory network and connected data information systems for research in the earth and space sciences. In Buneman and Jajodia [2271], pages 475–

478. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p475-thieman/p475-thieman.pdf;
http://www.acm.org/pubs/citations/

                                        proceedings/mod/170035/p475-thieman/

.

Mukhopadhyay:1993:IUA

[726] Debajyoti Mukhopadhyay. Interoperability using APPC. In Buneman and Jajodia [2271], pages 479–

482. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p479-mukhopadhyay/p479-mukhopadhyay.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/170035/
                                        p479-mukhopadhyay/.

Sheth:1993:MII

[727] Amit P. Sheth and George Karabatis. Multidatabase interdependencies in industry. In Buneman and Jajodia [2271], pages 483–486. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p483-sheth/p483-sheth.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/170035/p483-sheth/.

Cohen:1993:RIB

[728] David Cohen, Gary Larson, and Larry Berke. Role of interoperability in business application development. In Buneman and Jajodia [2271], pages 487–

490. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p487-cohen/p487-cohen.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/170035/p487-cohen/.

Woelk:1993:TSU

[729] Darrell Woelk, Paul Attie, Phil Cannata, Greg Meredith, Amit Sheth, Munindar Singh, and Christine Tomlinson. Task scheduling using intertask dependencies in Carnot. In Buneman and Jajodia [2271], pages 491–

494. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p491-woelk/p491-woelk.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p491-woelk/.

Mannai:1993:EIO

[730] Dhamir N. Mannai and Khaled Bugrara. Enhancing inter-operability and data sharing in medical information systems. In Buneman and Jajodia [2271], pages 495–498. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p495-mannai/p495-mannai.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p495-mannai/.

Woyna:1993:MBS

[731] Mark A. Woyna, John H. Christiansen, Christopher W. Hield, and Kathy Lee Simunich. Modeling battlefield sensor environments with an object database management system. In Buneman and Jajodia [2271], pages 499–501. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

no.2 1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/

                                        p499-woyna/p499-woyna.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/170035/p499-woyna/.

Schaller:1993:ICM

[732] Tony Schaller. The INtersect concept for multidatabase system integration in the pharmaceutical industry. In Buneman and Jajodia [2271], pages 502–

504. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p502-schaller/p502-schaller.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/170035/p502-schaller/

.

Gemis:1993:GGO

[733] Marc Gemis, Jan Paredaens, Inge Thyssens, and Jan Van den Bussche. GOOD: a graph-oriented object database system. In Buneman and Jajodia [2271], pages 505–510. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p505-gemis/p505-gemis.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/170035/p505-gemis/.

Consens:1993:HHB

[734] Mariano Consens and Alberto Mendelzon. Hy+: a Hygraph-based query and visualization system. In Buneman and Jajodia [2271], pages 511–

516. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/

pubs/articles/proceedings/mod/170035/
p511-consens/p511-consens.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/170035/p511-consens/

.

Chu:1993:DIC

[735] Wesley W. Chu, Matthew Merzbacher, and Ladislav Berkovich. The design and implementation of CoBase. In Buneman and Jajodia [2271], pages 517–522. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

no.2 1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p517-chu/p517-chu.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/170035/p517-chu/.

Keller:1993:PSB

[736] Arthur M. Keller, Richard Jensen, and Shailesh Agarwal. Persistence software: bridging object-oriented programming and relational databases. In Buneman and Jajodia [2271], pages 523–

528. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p523-keller/p523-keller.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p523-keller/.

Kupper:1993:NCN

[737] D. Küpper, M. Storbel, and D. Rösner. NAUDA: a cooperative natural language interface to relational databases. In Buneman and Jajodia [2271], pages 529–533. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

no.2 1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p529-kupper/p529-kupper.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p529-kupper/.

Bukhres:1993:IMP

[738] O. Bukhres, J. Chen, A. Elmagarmid, X. Liu, and J. Mullen. InterBase: a multidatabase prototype systems. In Buneman and Jajodia [2271], pages 534–539. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

no.2 1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p534-bukhres/p534-bukhres.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/170035/p534-bukhres/

.

Su:1993:OKO

[739] Stanley Y. W. Su, Herman X. Lam, Srinivasa Eddula, Javier Arroyo, Neeta Prasad, and Ronghao Zhuang. OSAM*.KBMS: an object-oriented knowledge base management system for supporting advanced applications. In Buneman and Jajodia [2271], pages 540–541. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

no.2 1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p540-su/p540-su.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/170035/p540-su/.

Moenkeberg:1993:CPS

[740] Axel Moenkeberg, Peter Zabback, Christof Hasse, and Gerhard Weikum. The COMFORT prototype: a step towards automated database performance tuning. In Buneman and Jajodia [2271], pages 542–543. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p542-moenkeberg/p542-moenkeberg.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/170035/
p542-moenkeberg/.

Ramakrishnan:1993:CDD

[741] Raghu Ramakrishnan, William G. Roth, Praveen Seshadri, Divesh Srivastava, and S. Sudarshan. The CORAL deductive database system. In Buneman and Jajodia [2271], pages 544–

545. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p544-ramakrishnan/p544-ramakrishnan.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/170035/
p544-ramakrishnan/.

Andersson:1993:SP

[742] M. Andersson, A.-M. Auddino, Y. Dupont, E. Fontana, M. Gentile, and S. Spaccapietra. The “SUPER” project. In Buneman and Jajodia [2271], pages 546–547. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

no.2 1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p546-andersson/p546-andersson.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/170035/p546-andersson/

.

Polyachenko:1993:ICP

[743] Boris E. Polyachenko and Filipp I. Andon. Instrumental complex of parallel software system development and operating environment support for distributed processing within multitransputer systems, TRANSSOFT. In Buneman and Jajodia [2271], pages 548–549. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

no.2 1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p548-polyachenko/p548-polyachenko.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/170035/
p548-polyachenko/.

Cacace:1993:LP

[744] F. Cacace, S. Ceri, S. Crespi-Reghizzi, P. Fraternali, S. Paraboschi, and L. Tanca. The LOGRES prototype. In Buneman and Jajodia [2271], pages 550–551. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

no.2 1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p550-cacace/p550-cacace.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p550-cacace/.

Ford:1993:OOM

[745] Steve Ford, José A. Blakeley, and Thomas J. Bannon. Open OODB: a modular object-oriented DBMS. In Buneman and Jajodia [2271], pages 552–553. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

no.2 1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p552-ford/p552-ford.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p552-ford/.

Hasan:1993:PGD

[746] W. Hasan, M. Heytens, C. Kolovson, M.-A. Neimat, S. Potamianos, and D. Schneider. Papyrus GIS demonstration. In Buneman and Jajodia [2271], pages 554–555. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p554-hasan/p554-hasan.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p554-hasan/.

Rakow:1993:VVS

[747] Thomas C. Rakow and Peter Muth. The V^{3} video server–managing analog and digital video clips. In Buneman and Jajodia [2271], pages 556–

557. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p556-rakow/p556-rakow.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p556-rakow/.

Muth:1993:VON

[748] Peter Muth and Thomas C. Rakow. VODAK open nested transactions– visualizing database internals. In Buneman and Jajodia [2271], pages 558–

559. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

1993. URL http://www.acm.org/
pubs/articles/proceedings/mod/170035/
p558-muth/p558-muth.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/170035/p558-muth/.

Luniewski:1993:IOU

[749] Allen Luniewski, Peter Schwarz, Kurt Shoens, Jim Stamos, and John Thomas. Information organization using Rufus. In Buneman and Jajodia [2271], pages 560–561. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2

                                        1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p560-luniewski/p560-luniewski.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/170035/p560-luniewski/

.

Arens:1993:SRI

[750] Yigal Arens and Craig Knoblock. SIMS: Retrieving and integrating information from multiple sources. In Buneman and Jajodia [2271], pages 562–563. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22

                                        no.2 1993. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/170035/
                                        p562-arens/p562-arens.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/170035/p562-arens/.

Harel:1994:WPW

[751] David Harel. Will I be pretty, will I be rich?: some thoughts on theory vs. practice in systems engineering. In ACM [2272], pages 1–3. ISBN 0-89791-642-5. LCCN QA 76.9 D3

                                        A26 1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        182591/p1-harel/p1-harel.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/pods/182591/p1-harel/
                                        ;   http://www.acm.org:80/pubs/

citations/proceedings/pods/182591/
p1-harel/.

Barbara:1994:SWC

[752] Daniel Barbará and Tomasz Imieliński. Sleepers and workaholics: caching strategies in mobile environments. In Snodgrass and Winslett [2273], pages 1–12. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p1-barbara/p1-barbara.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p1-barbara/.

Faloutsos:1994:BUI

[753] Christos Faloutsos and Ibrahim Kamel. Beyond uniformity and independence: Analysis of R-trees using the concept of fractal dimension. In ACM [2272], pages 4–13. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p4-faloutsos/p4-faloutsos.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/182591/
p4-faloutsos/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/182591/p4-faloutsos/.

Huang:1994:DRM

[754] Yixiu Huang, Prasad Sistla, and Ouri Wolfson. Data replication for mobile computers. In Snodgrass and Winslett [2273], pages 13–24. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/

                                        p13-huang/p13-huang.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/191839/p13-huang/.

Haas:1994:RCS

[755] Peter J. Haas, Jeffrey F. Naughton, and Arun N. Swami. On the relative cost of sampling for join selectivity estimation. In ACM [2272], pages 14–24. ISBN 0-89791-642-5. LCCN QA 76.9 D3

                                        A26 1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        182591/p14-haas/p14-haas.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/pods/182591/p14-haas/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/182591/
                                        p14-haas/.

Ramaswamy:1994:PCE

[756] Sridhar Ramaswamy and Sairam Subramanian. Path caching (extended abstract): a technique for optimal external searching. In ACM [2272], pages 25–35. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        182591/p25-ramaswamy/p25-ramaswamy.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/182591/
                                        p25-ramaswamy/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/182591/p25-ramaswamy/.

Ramaswamy:1994:PCT

[757] S. Ramaswamy and S. Subramanian. Path caching: a technique for optimal external searching. In ACM [2272], pages 25–35. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26 1994.

Imielinski:1994:EEI

[758] Tomasz Imielinski, S. Viswanathan, and B. R. Badrinath. Energy efficient indexing on air. In Snodgrass and Winslett [2273], pages 25–

36. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p25-imielinski/p25-imielinski.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/191839/p25-imielinski/

.

Chen:1994:ORT

[759] Ling Tony Chen and Doron Rotem. Optimal response time retrieval of replicated data (extended abstract). In ACM [2272], pages 36–44. ISBN 0-89791-642-5. LCCN QA 76.9 D3

A26 1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p36-chen/p36-chen.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/182591/p36-chen/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/182591/
p36-chen/.

Ahn:1994:SCP

[760] Ilsoo Ahn. SIGMOD challenges paper: database issues in telecommunications network management. In Snodgrass and Winslett [2273], pages 37–

43. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p37-ahn/p37-ahn.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/191839/p37-ahn/.

Biliris:1994:ASS

[761] A. Biliris, S. Dar, N. Gehani, H. V. Jagadish, and K. Ramamritham. AS-SET: a system for supporting extended transactions. In Snodgrass and Winslett [2273], pages 44–54. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p44-biliris/p44-biliris.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/191839/p44-biliris/.

Gupta:1994:CCP

[762] Ashish Gupta, Yehoshua Sagiv, Jeffrey D. Ullman, and Jennifer Widom. Constraint checking with partial information. In ACM [2272], pages 45–55. ISBN 0-89791-642-

5. LCCN QA 76.9 D3 A26

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        182591/p45-gupta/p45-gupta.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/182591/p45-gupta/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/182591/
                                        p45-gupta/.

Mohan:1994:ACM

[763] C. Mohan and Inderpal Narang. ARIES/CSA: a method for database recovery in client-server architectures. In Snodgrass and Winslett [2273], pages 55–66. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2 1994. URL http://www.acm.

                                        org/pubs/articles/proceedings/mod/
                                        191839/p55-mohan/p55-mohan.pdf;

http://www.acm.org/pubs/citations/
proceedings/mod/191839/p55-mohan/

.

Stuckey:1994:CQC

[764] Peter J. Stuckey and S. Sudarshan. Compiling query constraints (extended abstract). In ACM [2272], pages 56–67. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p56-stuckey/p56-stuckey.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/182591/
p56-stuckey/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/182591/p56-stuckey/.

Zhang:1994:ERA

[765] Aidong Zhang, Marian Nodine, Bharat Bhargava, and Omran Bukhres. Ensuring relaxed atomicity for flexible transactions in multidatabase systems. In Snodgrass and Winslett [2273], pages 67–78. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23

no.2 1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p67-zhang/p67-zhang.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p67-zhang/.

Sohn:1994:CAA

[766] Kirack Sohn. Constraints among argument sizes in logic programs (extended abstract). In ACM [2272], pages 68–74. ISBN 0-89791-642-

5. LCCN QA 76.9 D3 A26

1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p68-sohn/p68-sohn.pdf;

http://www.acm.org/pubs/citations/
proceedings/pods/182591/p68-sohn/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/182591/
p68-sohn/.

Agrawal:1994:TDM

[767] Rakesh Agrawal. Tutorial: Data mining. In ACM [2272], pages 75–76. ISBN 0-89791-642-5. LCCN QA 76.9 D3

A26 1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p75-agrawal/p75-agrawal.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/182591/
p75-agrawal/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/182591/p75-agrawal/.

Kivinen:1994:PSK

[768] Jyrki Kivinen and Heikki Mannila. The power of sampling in knowledge discovery. In ACM [2272], pages 77–85. ISBN 0-89791-642-5. LCCN QA 76.9 D3

A26 1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p77-kivinen/p77-kivinen.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/182591/
p77-kivinen/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/182591/p77-kivinen/.

Berson:1994:SSM

[769] Steven Berson, Shahram Ghandeharizadeh, Richard Muntz, and Xiangyu Ju. Staggered striping in multimedia information systems. In Snodgrass and Winslett [2273], pages 79– 90. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/

pubs/articles/proceedings/mod/191839/
p79-berson/p79-berson.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p79-berson/.

Chaudhuri:1994:CDA

[770] Surajit Chaudhuri and Phokion G. Kolaitis. Can Datalog be approximated? In ACM [2272], pages 86–96. ISBN 0-89791-642-5. LCCN QA 76.9 D3

A26 1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p86-chaudhuri/p86-chaudhuri.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/182591/
p86-chaudhuri/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/182591/p86-chaudhuri/.

Gibbs:1994:DMT

[771] Simon Gibbs, Christian Breiteneder, and Dennis Tsichritzis. Data modeling of time-based media. In Snodgrass and Winslett [2273], pages 91–

102. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p91-gibbs/p91-gibbs.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p91-gibbs/.

Afrati:1994:BAD

[772] Foto N. Afrati. Bonded arity Datalog (≠) queries on graphs. In ACM [2272], pages 97–106. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p97-afrati/p97-afrati.pdf;
http://www.acm.org/pubs/citations/

proceedings/pods/182591/p97-afrati/

.

Mumick:1994:IMS

[773] Inderpal Singh Mumick and Hamid Pirahesh. Implementation of magic-sets in a relational database system. In Snodgrass and Winslett [2273], pages 103–114. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23

no.2 1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p103-mumick/p103-mumick.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p103-mumick/.

Chaudhuri:1994:CEB

[774] Surajit Chaudhuri and Moshe Y. Vardi. On the complexity of equivalence between recursive and nonrecursive Datalog programs. In ACM [2272], pages 107–116. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p107-chaudhuri/p107-chaudhuri.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/182591/
p107-chaudhuri/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/182591/p107-chaudhuri/.

Wang:1994:CPD

[775] Jason Tsong-Li Wang, Gung-Wei Chirn, Thomas G. Marr, Bruce Shapiro, Dennis Shasha, and Kaizhong Zhang. Combinatorial pattern discovery for scientific data: some preliminary results. In Snodgrass and Winslett [2273], pages 115–125. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p115-wang/p115-wang.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p115-wang/.

Hua:1994:DBS

[776] Kien A. Hua, S. D. Lang, and Wen K. Lee. A decomposition-based simulated annealing technique for data clustering. In ACM [2272], pages 117–128. ISBN 0-89791-642-5. LCCN QA 76.9 D3

A26 1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p117-hua/p117-hua.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/182591/p117-hua/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/182591/
p117-hua/.

Gravano:1994:EGT

[777] Luis Gravano, Héctor García-Molina, and Anthony Tomasic. The effectiveness of GIOSS for the text database discovery problem. In Snodgrass and Winslett [2273], pages 126–137. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p126-gravano/p126-gravano.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/191839/p126-gravano/

.

Alonso:1994:RRC

[778] G. Alonso, D. Agrawal, and A. El Abbadi. Reducing recovery constraints on locking based protocols. In ACM [2272], pages 129–138. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        182591/p129-alonso/p129-alonso.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/182591/
                                        p129-alonso/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/182591/p129-alonso/.

Ioannidis:1994:IPE

[779] Yannis E. Ioannidis and Yezdi Lashkari. Incomplete path expressions and their disambiguation. In Snodgrass and Winslett [2273], pages 138–

149. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p138-ioannidis/p138-ioannidis.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/191839/p138-ioannidis/

.

Agrawal:1994:RSA

[780] D. Agrawal, J. L. Bruno, A. El Abbadi, and V. Krishnasawamy. Relative serializability: An approach for relaxing the atomicity of transactions. In ACM [2272], pages 139–149. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26 1994.

Agrawal:1994:RSE

[781] D. Agrawal, J. L. Bruno, A. El Abbadi, and V. Krishnaswamy. Relative serializability (extended abstract): an approach for relaxing the atomicity of transactions. In ACM [2272], pages 139–149. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        182591/p139-agrawal/p139-agrawal.

pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/182591/
p139-agrawal/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/182591/p139-agrawal/.

Tannen:1994:TLC

[782] Val Tannen. Tutorial: Languages for collection types. In ACM [2272], pages 150–154. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p150-tannen/p150-tannen.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/182591/
p150-tannen/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/182591/p150-tannen/.

Cole:1994:ODQ

[783] Richard L. Cole and Goetz Graefe. Optimization of dynamic query evaluation plans. In Snodgrass and Winslett [2273], pages 150–160. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p150-cole/p150-cole.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p150-cole/.

Libkin:1994:NTS

[784] Leonid Libkin and Limsoon Wong. New techniques for studying set languages, bag languages, and aggregate functions. In ACM [2272], pages 155–166. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/

                                        182591/p155-libkin/p155-libkin.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/182591/
                                        p155-libkin/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/182591/p155-libkin/.

Chen:1994:ASE

[785] Chungmin Melvin Chen and Nick Roussopoulos. Adaptive selectivity estimation using query feedback. In Snodgrass and Winslett [2273], pages 161–172. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23

                                        no.2 1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p161-chen/p161-chen.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/191839/p161-chen/.

Suciu:1994:QLN

[786] Dan Suciu and Val Breazu-Tannen. A query language for NC. In ACM [2272], pages 167–178. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        182591/p167-suciu/p167-suciu.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/182591/p167-suciu/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/182591/
                                        p167-suciu/.

Swami:1994:EPF

[787] Arun Swami and K. Bernhard Schiefer. Estimating page fetches for index scans with finite LRU buffers. In Snodgrass and Winslett [2273], pages 173–

184. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p173-swami/p173-swami.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p173-swami/.

Colby:1994:QLL

[788] Latha S. Colby, Edward L. Robertson, Lawrence V. Saxton, and Dirk Van Gucht. A query language for list-based complex objects. In ACM [2272], pages 179–189. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p179-colby/p179-colby.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/182591/p179-colby/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/182591/
p179-colby/.

Hsiao:1994:PEM

[789] Hui-I. Hsiao, Ming-Syan Chen, and Philip S. Yu. On parallel execution of multiple pipelined hash joins. In Snodgrass and Winslett [2273], pages 185–196. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23

no.2 1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p185-hsiao/p185-hsiao.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p185-hsiao/.

Mumick:1994:UFS

[790] Inderpal Singh Mumick and Oded Shmueli. Universal finiteness and satisfiability (extended abstract). In ACM [2272], pages 190–200. ISBN 0-89791-642-5. LCCN QA 76.9 D3

A26 1994. URL http://www.acm.org/

                                        pubs/articles/proceedings/pods/
                                        182591/p190-mumick/p190-mumick.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/182591/
                                        p190-mumick/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/182591/p190-mumick/.

Brinkhoff:1994:MSP

[791] Thomas Brinkhoff, Hans-Peter Kriegel, Ralf Schneider, and Bernhard Seeger. Multi-step processing of spatial joins. In Snodgrass and Winslett [2273], pages 197–208. ISBN 0-89791-639-

5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2 1994. URL http://www.acm.

                                        org/pubs/articles/proceedings/mod/
                                        191839/p197-brinkhoff/p197-brinkhoff.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/191839/
                                        p197-brinkhoff/.

Suciu:1994:ACO

[792] Dan Suciu and Jan Paredaens. Any algorithm in the complex object algebra with powerset needs exponential space to compute transitive closure. In ACM [2272], pages 201–209. ISBN 0-89791-642-5. LCCN QA 76.9 D3

                                        A26 1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        182591/p201-suciu/p201-suciu.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/182591/p201-suciu/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/182591/
                                        p201-suciu/.

Lo:1994:SJU

[793] Ming-Ling Lo and Chinya V. Ravishankar. Spatial joins using seeded trees. In Snodgrass and Winslett [2273], pages 209–220. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2 1994. URL http://www.

acm.org/pubs/articles/proceedings/
mod/191839/p209-lo/p209-lo.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/191839/p209-lo/.

Patnaik:1994:DFP

[794] Sushant Patnaik and Neil Immerman. Dyn-FO: a parallel, dynamic complexity class. In ACM [2272], pages 210–221. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p210-patnaik/p210-patnaik.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/182591/
p210-patnaik/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/182591/p210-patnaik/.

Pang:1994:MMR

[795] Hwee Hwa Pang, Michael J. Carey, and Miron Livny. Managing memory for real-time queries. In Snodgrass and Winslett [2273], pages 221–

232. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p221-pang/p221-pang.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p221-pang/.

Hillebrand:1994:FDQ

[796] Gerd G. Hillebrand and Paris C. Kanellakis. Functional database query languages as typed lambda calculi of fixed order (extended abstract). In ACM [2272], pages 222–231. ISBN 0-89791-642-5. LCCN QA 76.9 D3

A26 1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p222-hillebrand/p222-hillebrand.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/182591/
p222-hillebrand/;   http://www.
acm.org:80/pubs/citations/proceedings/
pods/182591/p222-hillebrand/.

Mendelzon:1994:OM

[797] Alberto O. Mendelzon, Tova Milo, and Emmanuel Waller. Object migration. In ACM [2272], pages 232–242. ISBN 0-89791-642-5. LCCN QA 76.9 D3

A26 1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p232-mendelzon/p232-mendelzon.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/182591/
p232-mendelzon/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/182591/p232-mendelzon/.

Nyberg:1994:ARM

[798] Chris Nyberg, Tom Barclay, Zarka Cvetanovic, Jim Gray, and Dave Lomet. AlphaSort: a RISC machine sort. In Snodgrass and Winslett [2273], pages 233–242. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23

no.2 1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p233-nyberg/p233-nyberg.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p233-nyberg/.

Gray:1994:QGB

[799] Jim Gray, Prakash Sundaresan, Susanne Englert, Ken Baclawski, and Peter J. Weinberger. Quickly generating billion-record synthetic databases. In Snodgrass and Winslett [2273], pages 243–252. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23

no.2 1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p243-gray/p243-gray.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p243-gray/.

Calvanese:1994:MOO

[800] Diego Calvanese and Maurizio Lenzerini. Making object-oriented schemas more expressive. In ACM [2272], pages 243–254. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p243-calvanese/p243-calvanese.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/182591/
p243-calvanese/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/182591/p243-calvanese/.

Vingralek:1994:DFO

[801] Radek Vingralek, Yuri Breitbart, and Gerhard Weikum. Distributed file organization with scalable cost/ performance. In Snodgrass and Winslett [2273], pages 253–264. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p253-vingralek/p253-vingralek.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/191839/p253-vingralek/

.

Ohori:1994:PCV

[802] Atsushi Ohori and Keishi Tajima. A polymorphic calculus for views and object sharing (extended abstract). In ACM [2272], pages 255–266. ISBN 0-89791-642-5. LCCN QA 76.9 D3

                                        A26 1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        182591/p255-ohori/p255-ohori.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/182591/p255-ohori/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/182591/
                                        p255-ohori/.

Kroll:1994:DST

[803] Brigitte Kröll and Peter Widmayer. Distributing a search tree among a growing number of processors. In Snodgrass and Winslett [2273], pages 265–

276. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p265-kroll/p265-kroll.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/191839/p265-kroll/.

Eiter:1994:ADD

[804] Thomas Eiter, Georg Gottlob, and Heikki Mannila. Adding disjunction to datalog (extended abstract). In ACM [2272], pages 267–278. ISBN 0-89791-642-5. LCCN QA 76.9 D3

                                        A26 1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        182591/p267-eiter/p267-eiter.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/182591/p267-eiter/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/182591/
                                        p267-eiter/.

Dewan:1994:PDL

[805] Hasanat M. Dewan, Salvatore J. Stolfo, Mauricio Hernández, and Jae-Jun Hwang. Predictive dynamic load balancing of parallel and distributed rule and query processing. In Snodgrass and Winslett [2273], pages 277–

288. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p277-dewan/p277-dewan.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p277-dewan/.

Paredaens:1994:TTS

[806] Jan Paredaens, Jan Van den Bussche, and Dirk Van Gucht. Towards a theory of spatial database queries (extended abstract). In ACM [2272], pages 279–288. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p279-paredaens/p279-paredaens.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/182591/
p279-paredaens/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/182591/p279-paredaens/.

Grumbach:1994:FRD

[807] Stéphane Grumbach and Jianwen Su. Finitely representable databases (extended abstract). In ACM [2272], pages 289–300. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p289-grumbach/p289-grumbach.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/182591/

                                        p289-grumbach/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/182591/p289-grumbach/.

Tomasic:1994:IUI

[808] Anthony Tomasic, Héctor García-Molina, and Kurt Shoens. Incremental updates of inverted lists for text document retrieval. In Snodgrass and Winslett [2273], pages 289–

300. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p289-tomasic/p289-tomasic.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/191839/p289-tomasic/

.

Gonnet:1994:TDD

[809] Gaston H. Gonnet. Text dominated databases, theory practice and experience (abstract). In ACM [2272], pages 301–302. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        182591/p301-gonnet/p301-gonnet.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/182591/
                                        p301-gonnet/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/182591/p301-gonnet/.

Gonnet:1994:TTD

[810] G. H. Gonnet. Tutorial: Text dominated databases, theory practice and experience. In ACM [2272], pages 301–

302. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26 1994.

Consens:1994:OQF

[811] Mariano P. Consens and Tova Milo. Optimizing queries on files. In Snodgrass and Winslett [2273], pages 301–

312. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p301-consens/p301-consens.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/191839/p301-consens/

.

Grahne:1994:RAS

[812] Gösta Grahne, Matti Nykänen, and Esko Ukkonen. Reasoning about strings in databases. In ACM [2272], pages 303–312. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26

1994. URL http://www.acm.org/
pubs/articles/proceedings/pods/
182591/p303-grahne/p303-grahne.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/182591/
p303-grahne/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/182591/p303-grahne/.

Christophides:1994:SDN

[813] V. Christophides, S. Abiteboul, S. Cluet, and M. Scholl. From structured documents to novel query facilities. In Snodgrass and Winslett [2273], pages 313–324. ISBN 0-89791-639-

5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2 1994. URL http://www.acm.

org/pubs/articles/proceedings/mod/
191839/p313-christophides/p313-
christophides.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/191839/p313-christophides/.

Hellerstein:1994:PPP

[814] Joseph M. Hellerstein. Practical predicate placement. In Snodgrass and Winslett [2273], pages 325–335. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p325-hellerstein/p325-hellerstein.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/191839/
                                        p325-hellerstein/.

Kemper:1994:ODQ

[815] A. Kemper, G. Moerkotte, K. Peithner, and M. Steinbrunn. Optimizing disjunctive queries with expensive predicates. In Snodgrass and Winslett [2273], pages 336–347. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p336-kemper/p336-kemper.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/191839/p336-kemper/.

Galindo-Legaria:1994:OD

[816] César A. Galindo-Legaria. Outerjoins as disjunctions. In Snodgrass and Winslett [2273], pages 348–358. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p348-galindo-legaria/p348-galindo-
                                        legaria.pdf;   http://www.acm.org/
                                        pubs/citations/proceedings/mod/
                                        191839/p348-galindo-legaria/.

Carey:1994:FGS

[817] Michael J. Carey, Michael J. Franklin, and Markos Zaharioudakis. Finegrained sharing in a page server OODBMS. In Snodgrass and Winslett [2273], pages 359–370. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p359-carey/p359-carey.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p359-carey/.

Cook:1994:PSP

[818] Jonathan E. Cook, Alexander L. Wolf, and Benjamin G. Zorn. Partition selection policies in object database garbage collection. In Snodgrass and Winslett [2273], pages 371–382. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p371-cook/p371-cook.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p371-cook/.

Carey:1994:SPA

[819] Michael J. Carey, David J. De-Witt, Michael J. Franklin, Nancy E. Hall, Mark L. McAuliffe, Jeffrey F. Naughton, Daniel T. Schuh, Marvin H. Solomon, C. K. Tan, Odysseas G. Tsatalos, Seth J. White, and Michael J. Zwilling. Shoring up persistent applications. In Snodgrass and Winslett [2273], pages 383–394. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p383-carey/p383-carey.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/191839/p383-carey/.

White:1994:QHP

[820] Seth J. White and David J. De-Witt. QuickStore: a high performance mapped object store. In Snodgrass and Winslett [2273], pages 395–

406. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p395-white/p395-white.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/191839/p395-white/.

McIver:1994:SAL

[821] William J. McIver and Roger King. Self-adaptive, on-line reclustering of complex object data. In Snodgrass and Winslett [2273], pages 407–418. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p407-mciver/p407-mciver.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/191839/p407-mciver/.

Faloutsos:1994:FSM

[822] Christos Faloutsos, M. Ranganathan, and Yannis Manolopoulos. Fast subsequence matching in time-series databases. In Snodgrass and Winslett [2273], pages 419–429. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p419-faloutsos/p419-faloutsos.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/191839/p419-faloutsos/

.

Seshadri:1994:SQP

[823] Praveen Seshadri, Miron Livny, and Raghu Ramakrishnan. Sequence query processing. In Snodgrass and Winslett [2273], pages 430–441. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p430-seshadri/p430-seshadri.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/191839/p430-seshadri/

.

Sagonas:1994:XED

[824] Konstantinos Sagonas, Terrance Swift, and David S. Warren. XSB as an efficient deductive database engine. In Snodgrass and Winslett [2273], pages 442–453. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23

no.2 1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p442-sagonas/p442-sagonas.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/191839/p442-sagonas/

.

Dar:1994:PST

[825] Shaul Dar and Raghu Ramakrishnan. A performance study of transitive closure algorithms. In Snodgrass and Winslett [2273], pages 454–

465. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p454-dar/p454-dar.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/191839/p454-dar/.

Anonymous:1994:PDS

[826] Anonymous. Parallel database systems in the 1990’s. In Snodgrass and Winslett [2273], page 466. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2 1994.

Daniels:1994:OSR

[827] Dean Daniels, Lip Boon Doo, Alan Downing, Curtis Elsbernd, Gary Hallmark, Sandeep Jain, Bob Jenkins, Peter Lim, Gordon Smith, Benny Souder, and Jim Stamos. Oracle’s symmetric replication technology and implications for application design. In Snodgrass and Winslett [2273], page

467. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p467-daniels/p467-daniels.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/191839/p467-daniels/

.

Dietterich:1994:DDD

[828] Daniel J. Dietterich. DEC data distributor: for data replication and data warehousing. In Snodgrass and Winslett [2273], page 468. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/

p468-dietterich/p468-dietterich.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/191839/
p468-dietterich/.

Gorelik:1994:SRS

[829] Alex Gorelik, Yongdong Wang, and Mark Deppe. Sybase replication server. In Snodgrass and Winslett [2273], page

469. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p469-gorelik/p469-gorelik.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/191839/p469-gorelik/

.

Laursen:1994:OMS

[830] Andrew Laursen, Jeffrey Olkin, and Mark Porter. Oracle media server: providing consumer based interactive access to multimedia data. In Snodgrass and Winslett [2273], pages 470–

477. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p470-laursen/p470-laursen.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/191839/p470-laursen/

.

Kulkarni:1994:OOE

[831] Krishna G. Kulkarni. Object-oriented extensions in SQL3: a status report. In Snodgrass and Winslett [2273], page

478. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/

                                        p478-kulkarni/p478-kulkarni.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/191839/p478-kulkarni/

.

Martin:1994:CCO

[832] Bruce E. Martin. COSS: the common object services specifications. In Snodgrass and Winslett [2273], page

479. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p479-martin/p479-martin.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/191839/p479-martin/.

Cattell:1994:OSO

[833] R. G. G. Cattell. ODMG-93: a standard for object-oriented DBMSs. In Snodgrass and Winslett [2273], page

480. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p480-cattell/p480-cattell.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/191839/p480-cattell/

.

Kim:1994:UXU

[834] Won Kim. UniSQL/X unified relational and object-oriented database system. In Snodgrass and Winslett [2273], page 481. ISBN 0-89791-639-

5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2 1994. URL http://www.acm.

                                        org/pubs/articles/proceedings/mod/
                                        191839/p481-kim/p481-kim.pdf;
                                        http://www.acm.org/pubs/citations/

proceedings/mod/191839/p481-kim/

.

Ubell:1994:MED

[835] Michael Ubell. The Montage extensible DataBlade architecture. In Snodgrass and Winslett [2273], page

482. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p482-ubell/p482-ubell.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p482-ubell/.

Pirahesh:1994:OOF

[836] Hamid Pirahesh. Object-oriented features of DB2 client/server. In Snodgrass and Winslett [2273], page 483. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p483-pirahesh/p483-pirahesh.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/191839/p483-pirahesh/

.

Vaskevitch:1994:DCT

[837] David Vaskevitch. Database in crisis and transition: a technical agenda for the year 2001. In Snodgrass and Winslett [2273], pages 484–489. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p484-vaskevitch/p484-vaskevitch.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/191839/
p484-vaskevitch/.

Ballinger:1994:ETD

[838] Carrie Ballinger. Evolving teradata decision support for massively parallel processing with UNIX. In Snodgrass and Winslett [2273], page 490. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p490-ballinger/p490-ballinger.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/191839/p490-ballinger/

.

Englert:1994:NSS

[839] Susanne Englert. Nonstop SQL: scalability and availability for decision support. In Snodgrass and Winslett [2273], page 491. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23

                                        no.2 1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p491-englert/p491-englert.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/191839/p491-englert/

.

Fernandez:1994:RBW

[840] Phillip M. Fernandez. Red brick warehouse: a read-mostly RDBMS for open SMP platforms. In Snodgrass and Winslett [2273], page 492. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p492-fernandez/p492-fernandez.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/191839/p492-fernandez/

.

Melling:1994:EIA

[841] Wesley P. Melling. Enterprise information architectures–they’re finally changing. In Snodgrass and Winslett [2273], pages 493–504. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p493-melling/p493-melling.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/191839/p493-melling/

.

Singh:1994:RTP

[842] Munindar P. Singh, Christine Tomlinson, and Darrell Woelk. Relaxed transaction processing. In Snodgrass and Winslett [2273], page 505. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p505-singh/p505-singh.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p505-singh/.

Flokstra:1994:IDD

[843] Jan Flokstra, Maurice van Keulen, and Jacek Skowronek. The IMPRESS DDT: a database design toolbox based on a formal specification language. In Snodgrass and Winslett [2273], page

506. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p506-flokstra/p506-flokstra.pdf;
 http://www.acm.org/pubs/citations/

                                        proceedings/mod/191839/p506-flokstra/

.

Bryan:1994:MPA

[844] G. M. Bryan, W. E. Moore, B. J. Curry, K. W. Lodge, and J. Geyer. The MEDUSA project: autonomous data management in a shared-nothing parallel database machine. In Snodgrass and Winslett [2273], page 507. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p507-bryan/p507-bryan.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/191839/p507-bryan/.

Missikoff:1994:MSC

[845] M. Missikoff and M. Toiati. MOSAICO— a system for conceptual modeling and rapid prototyping of objectoriented database application. In Snodgrass and Winslett [2273], page

508. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p508-missikoff/p508-missikoff.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/191839/p508-missikoff/

.

Kuhn:1994:LBM

[846] Eva Kühn, Thomas Tschernko, and Konrad Schwarz. A language based multidatabase system. In Snodgrass and Winslett [2273], page 509. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p509-kuhn/p509-kuhn.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p509-kuhn/.

Grossman:1994:PSP

[847] R. L. Grossman and X. Qin. Ptool: a scalable persistent object manager. In Snodgrass and Winslett [2273], page

510. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p510-grossman/p510-grossman.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/191839/p510-grossman/

.

Theodoulidis:1994:OTD

[848] Babis Theodoulidis, Aziz Ait-Braham, George Andrianopoulos, Jayant Chaudhary, George Karvelis, and Simon Sou. The ORES temporal database management system. In Snodgrass and Winslett [2273], page 511. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p511-theodoulidis/p511-theodoulidis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/191839/
p511-theodoulidis/.

Sagonas:1994:XDD

[849] Konstantinos Sagonas, Terrance Swift, and David S. Warren. XSB as a deductive database. In Snodgrass and Winslett [2273], page 512. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p512-sagonas/p512-sagonas.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/191839/p512-sagonas/

.

Dogac:1994:MOO

[850] Asuman Dogac, Budak Arpinar, Cem Evrendilek, Cetin Ozkan, Ilker Altintas, Ilker Durusoy, Mehmet Altinel, Tansel Okay, and Yuksel Saygin. METU object-oriented DBMS. In Snodgrass and Winslett [2273], page

513. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p513-dogac/p513-dogac.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p513-dogac/.

Agrawal:1994:QPD

[851] R. Agrawal, M. Carey, C. Faloutsos, S. Ghosh, M. Houtsma, T. Imieliński, B. Iyer, A. Mahboob, H. Miranda, R. Srikant, and A. Swami. Quest: a project on database mining. In Snodgrass and Winslett [2273], page

514. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p514-agrawal/p514-agrawal.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/191839/p514-agrawal/

.

Catarci:1994:QDG

[852] Tiziana Catarci and Giuseppe Santucci. Query by diagram: a graphical environment for querying databases. In Snodgrass and Winslett [2273], page

515. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p515-catarci/p515-catarci.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/191839/p515-catarci/

.

Han:1994:DSP

[853] Jiawei Han, Yongjian Fu, Yue Huang, Yandong Cai, and Nick Cercone. DBLearn: a system prototype for knowledge discovery in relational databases. In Snodgrass and Winslett [2273], page 516. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p516-han/p516-han.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/191839/p516-han/.

Biliris:1994:EEO

[854] Alexandros Biliris and Euthimios Panagos. EOS: an extensible object store. In Snodgrass and Winslett [2273], page 517. ISBN 0-89791-639-

5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2 1994. URL http://www.acm.

org/pubs/articles/proceedings/mod/
191839/p517-biliris/p517-biliris.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/191839/
p517-biliris/.

Hwang:1994:MFD

[855] S.-Y. Hwang, E.-P. Lim, H.-R.

Yang, S. Musukula, K. Mediratta, M. Ganesh, D. Clements, J. Stenoien, and J. Srivastava. The MYRIAD federated database prototype. In Snodgrass and Winslett [2273], page 518. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p518-hwang/p518-hwang.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/191839/p518-hwang/.

Brinkhoff:1994:GSE

[856] Thomas Brinkhoff, Hans-Peter Kriegel, Ralf Schneider, and Bernhard Seeger. GENESYS: a system for efficient spatial query processing. In Snodgrass and Winslett [2273], page 519. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p519-brinkhoff/p519-brinkhoff.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/191839/p519-brinkhoff/

.

Blakeley:1994:OOD

[857] José A. Blakeley. Open object database management systems. In Snodgrass and Winslett [2273], page

520. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

                                        1994. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/191839/
                                        p520-blakeley/p520-blakeley.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/191839/p520-blakeley/

.

Mohan:1994:SCA

[858] C. Mohan. A survey and critique of advanced transaction models. In Snodgrass and Winslett [2273], page

521. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2

1994. URL http://www.acm.org/
pubs/articles/proceedings/mod/191839/
p521-mohan/p521-mohan.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/191839/p521-mohan/.

Jagadish:1994:DN

[859] H. V. Jagadish. Databases for networks. In Snodgrass and Winslett [2273], page 522. ISBN 0-89791-639-

5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2 1994. URL http://www.acm.

org/pubs/articles/proceedings/mod/
191839/p522-jagadish/p522-jagadish.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/191839/
p522-jagadish/.

Berenson:1995:CAS

[860] Hal Berenson, Phil Bernstein, Jim Gray, Jim Melton, Elizabeth O’Neil, and Patrick O’Neil. A critique of ANSI SQL isolation levels. In ACM [2275], pages 1–10. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p1-berenson/p1-berenson.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p1-berenson/. ACM or-

der number 472950.

Papadimitriou:1995:DMA

[861] Christos H. Papadimitriou. Database metatheory: Asking the big queries. In ACM [2274], pages 1–10. ISBN 0-89791-730-8. LCCN QA 76.9 D3

                                        A26 1995. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        212433/p1-papadimitriou/p1-papadimitriou.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/212433/
                                        p1-papadimitriou/;   http://www.
                                        acm.org:80/pubs/citations/proceedings/
                                        pods/212433/p1-papadimitriou/.

Kabanza:1995:HIT

[862] F. Kabanza, J.-M. Stevenne, and P. Wolper. Handling infinite temporal data. In ACM [2274], pages 3–17. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Consens:1995:AQT

[863] Mariano P. Consens and Tova Milo. Algebras for querying text regions (extended abstract). In ACM [2274], pages 11–22. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26

                                        1995. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        212433/p11-consens/p11-consens.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/212433/
                                        p11-consens/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/212433/p11-consens/.

Molesky:1995:RPS

[864] Lory D. Molesky and Krithi Ramamritham. Recovery protocols for shared memory database systems. In ACM [2275], pages 11–22. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p11-molesky/p11-molesky.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p11-molesky/. ACM or-

der number 472950.

Lipton:1995:QSE

[865] R. J. Lipton and J. F. Naughton. Query size estimation by adaptive sampling. In ACM [2274], pages 18–25. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Adya:1995:EOC

[866] Atul Adya, Robert Gruber, Barbara Liskov, and Umesh Maheshwari. Efficient optimistic concurrency control using loosely synchronized clocks. In ACM [2275], pages 23–34. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p23-adya/p23-adya.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/223784/p23-adya/. ACM order

number 472950.

Mecca:1995:SDT

[867] Giansalvatore Mecca and Anthony J. Bonner. Sequences, Datalog and transducers. In ACM [2274], pages 23–35. ISBN 0-89791-730-8. LCCN QA 76.9

D3 A26 1995. URL http://www.acm.
org/pubs/articles/proceedings/pods/
212433/p23-mecca/p23-mecca.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/212433/p23-mecca/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/212433/
p23-mecca/.

Kanellakis:1995:CQL

[868] P. C. Kanellakis, G. M. Kuper, and P. Z. Revesz. Constraint query languages. In ACM [2274], pages 26–52. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Brodsky:1995:LLQ

[869] Alexander Brodsky and Yoram Kornatzky. The LyriC language: querying constraint objects. In ACM [2275], pages 35–46. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p35-brodsky/p35-brodsky.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/223784/p35-brodsky/. ACM or-

der number 472950.

Jagadish:1995:SBQ

[870] H. V. Jagadish, Alberto O. Mendelzon, and Tova Milo. Similaritybased queries. In ACM [2274], pages 36–45. ISBN 0-89791-730-

8. LCCN QA 76.9 D3 A26

                                        1995. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        212433/p36-jagadish/p36-jagadish.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/212433/
                                        p36-jagadish/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/212433/p36-jagadish/.

Kanellakis:1995:CPD

[871] Paris Kanellakis. Constraint programming and database languages: a tutorial. In ACM [2274], pages 46–53. ISBN 0-89791-730-8. LCCN QA 76.9 D3

                                        A26 1995. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/

212433/p46-kanellakis/p46-kanellakis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/212433/
p46-kanellakis/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/212433/p46-kanellakis/.

Kanellakis:1995:TCP

[872] P. Kanellakis. Tutorial: Constraint programming and database languages. In ACM [2274], pages 46–53. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Fegaras:1995:TEC

[873] Leonidas Fegaras and David Maier. Towards an effective calculus for object query languages. In ACM [2275], pages 47–58. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p47-fegaras/p47-fegaras.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p47-fegaras/. ACM or-

der number 472950.

Agrawal:1995:CSL

[874] D. Agrawal and A. El Abbadi. Constrained shared locks for increasing concurrency in databases. In ACM [2274], pages 53–63. ISBN 0-89791-730-

8. LCCN QA 76.9 D3 A26 1995.

Brodsky:1995:SPO

[875] Alexander Brodsky, Catherine Lassez, Jean-Louis Lassez, and Michael J. Maher. Separability of polyhedra for optimal filtering of spatial and constraint data. In ACM [2274], pages 54–65. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26

1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p54-brodsky/p54-brodsky.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/212433/
p54-brodsky/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/212433/p54-brodsky/.

Gardarin:1995:OFE

[876] Georges Gardarin, Fernando Machuca, and Philippe Pucheral. OFL: a functional execution model for object query languages. In ACM [2275], pages 59–70. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p59-gardarin/p59-gardarin.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/223784/p59-gardarin/

. ACM order number 472950.

Schlipf:1995:EPL

[877] J. S. Schlipf. The expressive powers of the logic programming semantics. In ACM [2274], pages 64–86. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Grumbach:1995:DOC

[878] Stéphane Grumbach and Jianwen Su. Dense-order constraint databases (extended abstract). In ACM [2274], pages 66–77. ISBN 0-89791-730-

8. LCCN QA 76.9 D3 A26

1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p66-grumbach/p66-grumbach.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/212433/
p66-grumbach/;   http://www.acm.

org:80/pubs/citations/proceedings/
pods/212433/p66-grumbach/.

Roussopoulos:1995:NNQ

[879] Nick Roussopoulos, Stephen Kelley, and Frédéric Vincent. Nearest neighbor queries. In ACM [2275], pages 71–79. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p71-roussopoulos/p71-roussopoulos.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/223784/
p71-roussopoulos/. ACM order num-

ber 472950.

Chomicki:1995:MIR

[880] Jan Chomicki and Gabriel Kuper. Measuring infinite relations. In ACM [2274], pages 78–85. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26

1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p78-chomicki/p78-chomicki.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/212433/
p78-chomicki/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/212433/p78-chomicki/.

Freeston:1995:GSD

[881] Michael Freeston. A general solution of the n-dimensional B-tree problem. In ACM [2275], pages 80–91. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p80-freeston/p80-freeston.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/223784/p80-freeston/

. ACM order number 472950.

Pagel:1995:WQO

[882] Bernd-Uwe Pagel, Hans-Werner Six, and Mario Winter. Window queryoptimal clustering of spatial objects. In ACM [2274], pages 86–94. ISBN 0-89791-730-8. LCCN QA 76.9 D3

                                        A26 1995. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        212433/p86-pagel/p86-pagel.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/212433/p86-pagel/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/212433/
                                        p86-pagel/.

Saraiya:1995:ETD

[883] Y. P. Saraiya. On the efficiency of transforming database logic programs. In ACM [2274], pages 87–109. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Papadias:1995:TRW

[884] Dimitris Papadias, Timos Sellis, Yannis Theodoridis, and Max J. Egenhofer. Topological relations in the world of minimum bounding rectangles: a study with R-trees. In ACM [2275], pages 92–103. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p92-papadias/p92-papadias.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/223784/p92-papadias/

. ACM order number 472950.

Levy:1995:AQU

[885] Alon Y. Levy, Alberto O. Mendelzon, Yehoshua Sagiv, and D. Srivastava. Answering queries using views (extended abstract). In ACM [2274], pages 95–104. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26

1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p95-levy/p95-levy.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/212433/p95-levy/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/212433/
p95-levy/.

Shatdal:1995:APA

[886] Ambuj Shatdal and Jeffrey F. Naughton. Adaptive parallel aggregation algorithms. In ACM [2275], pages 104–114. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

QA1 .A87. URL http://www.acm.
org/pubs/articles/proceedings/mod/
223784/p104-shatdal/p104-shatdal.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/223784/

p104-shatdal/. ACM order number 472950.

Rajaraman:1995:AQU

[887] Anand Rajaraman, Yehoshua Sagiv, and Jeffrey D. Ullman. Answering queries using templates with binding patterns (extended abstract). In ACM [2274], pages 105–112. ISBN 0-89791-730-8. LCCN QA 76.9 D3

A26 1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p105-rajaraman/p105-rajaraman.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/212433/
p105-rajaraman/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/212433/p105-rajaraman/.

Kolaitis:1995:EPD

[888] P. G. Kolaitis and M. Y. Vardi. On the expressive power of datalog: Tools and a case study. In ACM [2274], pages 110–134. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Jagadish:1995:VMI

[889] H. V. Jagadish, Inderpal Singh Mumick, and Abraham Silberschatz. View maintenance issues for the chronicle data model (extended abstract). In ACM [2274], pages 113–124. ISBN 0-89791-730-8. LCCN QA 76.9 D3

                                        A26 1995. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        212433/p113-jagadish/p113-jagadish.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/212433/
                                        p113-jagadish/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/212433/p113-jagadish/.

Wilschut:1995:PEM

[890] Annita N. Wilschut, Jan Flokstra, and Peter M. G. Apers. Parallel evaluation of multi-join queries. In ACM [2275], pages 115–126. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p115-wilschut/p115-wilschut.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/223784/p115-wilschut/

. ACM order number 472950.

Goodman:1995:RPG

[891] Nathan Goodman. Research problems in genome databases. In ACM [2274], page 125. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26

                                        1995. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        212433/p125-goodman/p125-goodman.
                                        pdf;   http://www.acm.org/pubs/

citations/proceedings/pods/212433/
p125-goodman/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/212433/p125-goodman/.

Picouet:1995:SEI

[892] Phillippe Picouet and Victor Vianu. Semantics and expressiveness issues in active databases (extended abstract). In ACM [2274], pages 126–138. ISBN 0-89791-730-8. LCCN QA 76.9 D3

A26 1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p126-picouet/p126-picouet.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/212433/
p126-picouet/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/212433/p126-picouet/.

Hernandez:1995:MPP

[893] Mauricio A. Hernández and Salvatore J. Stolfo. The merge/purge problem for large databases. In ACM [2275], pages 127–138. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p127-hernandez/p127-hernandez.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/223784/p127-hernandez/

. ACM order number 472950.

Seshadri:1995:ESR

[894] S. Seshadri and J. F. Naughton. On the expected size of recursive datalog queries. In ACM [2274], pages 137–148. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Dong:1995:SBF

[895] Guozhu Dong and Jianwen Su. Spacebounded FOIES (extended abstract).

In ACM [2274], pages 139–150. ISBN 0-89791-730-8. LCCN QA 76.9 D3

A26 1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p139-dong/p139-dong.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/212433/p139-dong/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/212433/
p139-dong/.

Ramaswamy:1995:OIC

[896] Sridhar Ramaswamy and Paris C. Kanellakis. OODB indexing by class-division. In ACM [2275], pages 139–150. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p139-ramaswamy/p139-ramaswamy.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/223784/p139-ramaswamy/

. ACM order number 472950.

Grumbach:1995:TQL

[897] S. Grumbach and V. Vianu. Tractable query languages for complex object databases. In ACM [2274], pages 149–

167. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Aref:1995:HTI

[898] Walid Aref, Daniel Barbará, and Padmavathi Vallabhaneni. The handwritten trie: indexing electronic ink. In ACM [2275], pages 151–162. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p151-aref/p151-aref.pdf;   http:
//www.acm.org/pubs/citations/proceedings/

mod/223784/p151-aref/. ACM order

number 472950.

Cadoli:1995:SRK

[899] Marco Cadoli, Francesco M. Donini, Paolo Liberatore, and Marco Schaerf. The size of a revised knowledge base. In ACM [2274], pages 151–162. ISBN 0-89791-730-8. LCCN QA 76.9 D3

A26 1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p151-cadoli/p151-cadoli.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/212433/
p151-cadoli/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/212433/p151-cadoli/.

Levy:1995:SQO

[900] Alon Y. Levy and Yehoshua Sagiv. Semantic query optimization in Datalog programs (extended abstract). In ACM [2274], pages 163–173. ISBN 0-89791-730-8. LCCN QA 76.9 D3

A26 1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p163-levy/p163-levy.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/212433/p163-levy/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/212433/
p163-levy/.

Faloutsos:1995:FFA

[901] Christos Faloutsos and King-Ip Lin. FastMap: a fast algorithm for indexing, data-mining and visualization of traditional and multimedia datasets. In ACM [2275], pages 163–174. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/

                                        p163-faloutsos/p163-faloutsos.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/223784/p163-faloutsos/

. ACM order number 472950.

Johnson:1995:TUB

[902] D. B. Johnson and L. Raab. A tight upper bound on the benefits of replica control protocols. In ACM [2274], pages 168–176. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Wood:1995:MFC

[903] Peter T. Wood. Magic factoring of closure programs (extended abstract). In ACM [2274], pages 174–183. ISBN 0-89791-730-8. LCCN QA 76.9 D3

                                        A26 1995. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        212433/p174-wood/p174-wood.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/212433/p174-wood/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/212433/
                                        p174-wood/.

Park:1995:EHB

[904] Jong Soo Park, Ming-Syan Chen, and Philip S. Yu. An effective hash-based algorithm for mining association rules. In ACM [2275], pages 175–186. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p175-park/p175-park.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/223784/p175-park/. ACM order

number 472950.

Afrati:1995:DVP

[905] F. Afrati, S. S. Cosmadakis, and M. Yannakakis. On datalog vs. polynomial time. In ACM [2274], pages 177– 196. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Ross:1995:STC

[906] Kenneth A. Ross. Structural totality and constraint stratification. In ACM [2274], pages 184–195. ISBN 0-89791-730-8. LCCN QA 76.9 D3

A26 1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p184-ross/p184-ross.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/212433/p184-ross/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/212433/
p184-ross/.

White:1995:ICR

[907] Seth J. White and David J. De-Witt. Implementing crash recovery in QuickStore: a performance study. In ACM [2275], pages 187–198. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p187-white/p187-white.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p187-white/. ACM or-

der number 472950.

Chang:1995:UWG

[908] Ti-Pin Chang and Richard Hull. Using witness generators to support bidirectional update between objectbased databases (extended abstract). In ACM [2274], pages 196–207. ISBN 0-89791-730-8. LCCN QA 76.9 D3

A26 1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p196-chang/p196-chang.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/212433/p196-chang/
;   http://www.acm.org:80/pubs/

                                        citations/proceedings/pods/212433/
                                        p196-chang/.

Helm:1995:SQO

[909] R. Helm, K. Marriott, and M. Odersky. Spatial query optimization: From Boolean constraints to range queries. In ACM [2274], pages 197–210. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Acharya:1995:BDD

[910] Swarup Acharya, Rafael Alonso, Michael Franklin, and Stanley Zdonik. Broadcast disks: data management for asymmetric communication environments. In ACM [2275], pages 199–210. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p199-acharya/p199-acharya.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/223784/p199-acharya/

. ACM order number 472950.

Andries:1995:AUM

[911] Marc Andries, Luca Cabibbo, Jan Paredaens, and Jan Van den Bussche. Applying an update method to a set of receivers (extended abstract). In ACM [2274], pages 208–218. ISBN 0-89791-730-8. LCCN QA 76.9 D3

                                        A26 1995. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        212433/p208-andries/p208-andries.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/212433/
                                        p208-andries/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/212433/p208-andries/.

Garcia-Molina:1995:NDQ

[912] H. García-Molina and K. Salem. Nondeterministic queue operations. In ACM [2274], pages 211–222. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Gupta:1995:AMV

[913] Ashish Gupta, Inderpal S. Mumick, and Kenneth A. Ross. Adapting materialized views after redefinitions. In ACM [2275], pages 211–222. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p211-gupta/p211-gupta.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p211-gupta/. ACM or-

der number 472950.

Libkin:1995:NID

[914] Leonid Libkin. Normalizing incomplete databases. In ACM [2274], pages 219–230. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26

1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p219-libkin/p219-libkin.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/212433/
p219-libkin/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/212433/p219-libkin/.

Hou:1995:EDC

[915] Wen-Chi Hou and Zhongyang Zhang. Enhancing database correctness: a statistical approach. In ACM [2275], pages 223–232. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p223-hou/p223-hou.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/223784/p223-hou/. ACM order

number 472950.

Beeri:1995:SO

[916] C. Beeri and T. Milo. Subtyping in OODBs. In ACM [2274], pages 223–

243. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Kolaitis:1995:CGD

[917] Phokion G. Kolaitis. Combinatorial games in database theory. In ACM [2274], pages 231–232. ISBN 0-89791-730-8. LCCN QA 76.9 D3

A26 1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p231-kolaitis/p231-kolaitis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/212433/
p231-kolaitis/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/212433/p231-kolaitis/.

Kolaitis:1995:TCG

[918] P. Kolaitis. Tutorial:combinatorial games in database theory. In ACM [2274], pages 231–232. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Chaudhri:1995:SLP

[919] Vinay K. Chaudhri and Vassos Hadzilacos. Safe locking policies for dynamic databases. In ACM [2274], pages 233–244. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26

1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p233-chaudhri/p233-chaudhri.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/212433/

p233-chaudhri/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/212433/p233-chaudhri/.

Ioannidis:1995:BHO

[920] Yannis E. Ioannidis and Viswanath Poosala. Balancing histogram optimality and practicality for query result size estimation. In ACM [2275], pages 233–244. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p233-ioannidis/p233-ioannidis.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/223784/p233-ioannidis/

. ACM order number 472950.

Ganguly:1995:EPD

[921] S. Ganguly, S. Greco, and C. Zaniolo. Extrema predicates in deductive databases. In ACM [2274], pages 244–

259. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Keidar:1995:IRA

[922] Idit Keidar and Danny Dolev. Increasing the resilience of atomic commit, at no additional cost. In ACM [2274], pages 245–254. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26

1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p245-keidar/p245-keidar.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/212433/
p245-keidar/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/212433/p245-keidar/.

Adelberg:1995:AUS

[923] B. Adelberg, H. Garcia-Molina, and B. Kao. Applying update streams in a soft real-time database system. In ACM [2275], pages 245–256. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p245-adelberg/p245-adelberg.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/223784/p245-adelberg/

. ACM order number 472950.

Chekuri:1995:SPP

[924] Chandra Chekuri, Waqar Hasan, and Rajeev Motwani. Scheduling problems in parallel query optimization. In ACM [2274], pages 255–265. ISBN 0-89791-730-8. LCCN QA 76.9 D3

A26 1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p255-chekuri/p255-chekuri.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/212433/
p255-chekuri/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/212433/p255-chekuri/.

Bettini:1995:SAQ

[925] Claudio Bettini, X. Sean Wang, Elisa Bertino, and Sushil Jajodia. Semantic assumptions and query evaluation in temporal databases. In ACM [2275], pages 257–268. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p257-bettini/p257-bettini.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/223784/p257-bettini/

. ACM order number 472950.

Vardi:1995:CBV

[926] Moshe Y. Vardi. On the complexity of bounded-variable queries (extended abstract). In ACM [2274], pages 266–276. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26

1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p266-vardi/p266-vardi.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/212433/p266-vardi/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/212433/
p266-vardi/.

Sistla:1995:TCI

[927] A. Prasad Sistla and Ouri Wolfson. Temporal conditions and integrity constraints in active database systems. In ACM [2275], pages 269–280. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p269-sistla/p269-sistla.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p269-sistla/. ACM or-

der number 472950.

Stolboushkin:1995:FQD

[928] Alexei P. Stolboushkin and Michael A. Taitslin. Finite queries do not have effective syntax. In ACM [2274], pages 277–285. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26

1995. URL http://www.acm.org/
pubs/articles/proceedings/pods/
212433/p277-stolboushkin/p277-stolboushkin.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/212433/
p277-stolboushkin/;   http://www.
acm.org:80/pubs/citations/proceedings/
pods/212433/p277-stolboushkin/.

Davison:1995:DRB

[929] Diane L. Davison and Goetz Graefe. Dynamic resource brokering for multiuser query execution. In ACM [2275], pages 281–292. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p281-davison/p281-davison.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/223784/p281-davison/

. ACM order number 472950.

DeRougemont:1995:RQ

[930] M. De Rougemont. The reliability of queries. In ACM [2274], pages 286–291. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

deRougemont:1995:RQE

[931] Michel de Rougemont. The reliability of queries (extended abstract). In ACM [2274], pages 286–291. ISBN 0-89791-730-8. LCCN QA 76.9 D3

                                        A26 1995. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        pods/212433/p286-de_rougemont/;
                                         http://www.acm.org:80/pubs/citations/
                                        proceedings/pods/212433/p286-de_
                                        rougemont/.

Du:1995:RMQ

[932] Weimin Du, Ming-Chien Shan, and Umeshwar Dayal. Reducing multidatabase query response time by tree balancing. In ACM [2275], pages 293–303. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p293-du/p293-du.pdf;   http://www.
                                        acm.org/pubs/citations/proceedings/
                                        mod/223784/p293-du/. ACM order

number 472950.

Bhargava:1995:HBR

[933] Gautam Bhargava, Piyush Goel, and Bala Iyer. Hypergraph based reorderings of outer join queries with complex predicates. In ACM [2275], pages 304–315. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p304-bhargava/p304-bhargava.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/223784/p304-bhargava/

. ACM order number 472950.

Zhuge:1995:VMW

[934] Yue Zhuge, Héctor García-Molina, Joachim Hammer, and Jennifer Widom. View maintenance in a warehousing environment. In ACM [2275], pages 316–327. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p316-zhuge/p316-zhuge.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p316-zhuge/. ACM or-

der number 472950.

Griffin:1995:IMV

[935] Timothy Griffin and Leonid Libkin. Incremental maintenance of views with duplicates. In ACM [2275], pages 328–339. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p328-griffin/p328-griffin.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/223784/p328-griffin/

. ACM order number 472950.

Lu:1995:EMM

[936] James J. Lu, Guido Moerkotte, Joachim Schue, and V. S. Subrahmanian. Efficient maintenance of materialized mediated views. In ACM [2275], pages 340–351. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p340-lu/p340-lu.pdf;   http://www.
                                        acm.org/pubs/citations/proceedings/
                                        mod/223784/p340-lu/. ACM order

number 472950.

Freedman:1995:SSV

[937] Craig S. Freedman and David J. De-Witt. The SPIFFI scalable videoon-demand system. In ACM [2275], pages 352–363. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p352-freedman/p352-freedman.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/223784/p352-freedman/

. ACM order number 472950.

Berson:1995:FTD

[938] Steven Berson, Leana Golubchik, and Richard R. Muntz. Fault tolerant design of multimedia servers. In ACM [2275], pages 364–375. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p364-berson/p364-berson.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/223784/p364-berson/. ACM or-

der number 472950.

Dan:1995:OVP

[939] Asit Dan and Dinkar Sitaram. An online video placement policy based on bandwidth to space ratio (BSR). In ACM [2275], pages 376–385. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p376-dan/p376-dan.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/223784/p376-dan/. ACM order

number 472950.

Abiteboul:1995:DIF

[940] Serge Abiteboul, Sophie Cluet, and Tova Milo. A database interface for file update. In ACM [2275], pages 386–397. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p386-abiteboul/p386-abiteboul.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/223784/p386-abiteboul/

. ACM order number 472950.

Brin:1995:CDM

[941] Sergey Brin, James Davis, and Héctor García-Molina. Copy detection mechanisms for digital documents. In ACM [2275], pages 398–409. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p398-brin/p398-brin.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p398-brin/. ACM order

number 472950.

Chaudhuri:1995:JQE

[942] Surajit Chaudhuri, Umeshwar Dayal, and Tak W. Yan. Join queries with external text sources: execution and optimization techniques. In ACM [2275], pages 410–422. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p410-chaudhuri/p410-chaudhuri.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/223784/p410-chaudhuri/

. ACM order number 472950.

Microsoft:1995:OOR

[943] Fox Development Team Microsoft. Object-oriented, rapid application development in a PC database environment. In ACM [2275], pages 423–424. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        QA1 .A87. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        mod/223784/p423-fox_development_
                                        team_microsoft/. ACM order number

472950.

Microsoft:1995:UFF

[944] The Access Team Microsoft. Upsizing form file server to client server architectures. In ACM [2275], pages 425–426. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        QA1 .A87. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        mod/223784/p425-the_access_team_

microsoft/. ACM order number

472950.

Moore:1995:LNS

[945] Kenneth Moore. The Lotus notes storage system. In ACM [2275], pages 427–428. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p427-moore/p427-moore.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p427-moore/. ACM or-

der number 472950.

Gettys:1995:DOR

[946] William L. Gettys. DIRECTV and Oracle Rdb: the challenge of VLDB transaction processing. In ACM [2275], pages 429–430. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p429-gettys/p429-gettys.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p429-gettys/. ACM or-

der number 472950.

Hope:1995:ETP

[947] Greg Hope. Enterprise transaction processing on Windows NT. In ACM [2275], pages 431–432. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p431-hope/p431-hope.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p431-hope/. ACM order

number 472950.

Ivinskis:1995:HAC

[948] Kestutis Ivinskis. High availability of commercial applications. In ACM [2275], pages 433–434. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/

                                        pubs/articles/proceedings/mod/223784/
                                        p433-ivinskis/p433-ivinskis.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/223784/p433-ivinskis/

. ACM order number 472950.

Buneman:1995:DYW

[949] Peter Buneman and David Maier. The data that you won’t find in databases: tutorial panel on data exchange formats. In ACM [2275], page 435. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p435-buneman/p435-buneman.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/223784/p435-buneman/

. ACM order number 472950.

Gray:1995:PDS

[950] Jim Gray. Parallel database systems 101. In ACM [2275], page 436. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p436-gray/p436-gray.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/223784/p436-gray/. ACM order

number 472950.

Ellison:1995:KA

[951] Larry J. Ellison. Keynote address. In ACM [2275], page 437. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p437-ellison/p437-ellison.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/223784/p437-ellison/

. ACM order number 472950.

Epstein:1995:KA

[952] Robert S. Epstein. Keynote Address. In ACM [2275], page 438. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p438-epstein/p438-epstein.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/223784/p438-epstein/

. ACM order number 472950.

Goldring:1995:TEU

[953] Rob Goldring. Things every update replication customer should know (abstract). In ACM [2275], pages 439–440. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p439-goldring/p439-goldring.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/223784/p439-goldring/

. ACM order number 472950.

Shyy:1995:VRS

[954] Yuh-Ming Shyy, H. Stephen Au-Yeung, and C. P. Chou. VER-SANT replication: supporting faulttolerant object databases. In ACM [2275], pages 441–442. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p441-shyy/p441-shyy.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p441-shyy/. ACM order

number 472950.

Woelk:1995:CID

[955] D. Woelk, B. Bohrer, N. Jacobs, K. Ong, C. Tomlinson, and C. Unnikrishnan. Carnot and InfoSleuth: database technology and the World Wide Web. In ACM [2275], pages 443–444. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p443-woelk/p443-woelk.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p443-woelk/. ACM or-

der number 472950.

Edelstein:1995:RPT

[956] Herb Edelstein. Research and products–are they relevant to each other? (panel session). In ACM [2275], page 445. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p445-edelstein/p445-edelstein.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/223784/p445-edelstein/

. ACM order number 472950.

Squire:1995:DET

[957] Case Squire. Data extraction and transformation for the data warehouse. In ACM [2275], pages 446–447. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p446-squire/p446-squire.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p446-squire/. ACM or-

der number 472950.

Bansal:1995:RWR

[958] Sanju K. Bansal. Real world requirements for decision support– implications for RDBMS. In ACM [2275], page 448. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p448-bansal/p448-bansal.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p448-bansal/. ACM or-

der number 472950.

French:1995:OSF

[959] Clark D. French. “one size fits all” database architectures do not work for DSS. In ACM [2275], pages 449–450. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p449-french/p449-french.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p449-french/. ACM or-

der number 472950.

Perna:1995:LIA

[960] Janet Perna. Leveraging the information asset. In ACM [2275], pages 451–452. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p451-perna/p451-perna.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p451-perna/. ACM or-

der number 472950.

Olson:1995:CYA

[961] Michael A. Olson. Cover your assets. In ACM [2275], page 453. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/

                                        pubs/articles/proceedings/mod/223784/
                                        p453-olson/p453-olson.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/223784/p453-olson/. ACM or-

der number 472950.

Atkinson:1995:UCA

[962] Robert Atkinson. Use of a component architecture in integrating relational and non-relational. In ACM [2275], page 454. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p454-atkinson/p454-atkinson.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/223784/p454-atkinson/

. ACM order number 472950.

Kleissner:1995:EOF

[963] Charly Kleissner. Enterprise Objects Framework: a second generation object-relational enabler. In ACM [2275], pages 455–459. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p455-kleissner/p455-kleissner.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/223784/p455-kleissner/

. ACM order number 472950.

Baru:1995:ODP

[964] Chaitanya Baru and Gilles Fecteau. An overview of DB2 parallel edition. In ACM [2275], pages 460–462. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p460-baru/p460-baru.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/

mod/223784/p460-baru/. ACM order

number 472950.

Gerber:1995:IOX

[965] Bob Gerber. Informix online XPS. In ACM [2275], page 463. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p463-gerber/p463-gerber.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p463-gerber/. ACM or-

der number 472950.

Levine:1995:OMA

[966] Charles Levine. Order-of-magnitude advantage on TPC-C through massive parallelism. In ACM [2275], pages 464–465. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p464-levine/p464-levine.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p464-levine/. ACM or-

der number 472950.

Deutsch:1995:OSP

[967] Donald R. Deutsch. Objects and SQL (panel session): strange relations? In ACM [2275], page 466. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p466-deutsch/p466-deutsch.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/223784/p466-deutsch/

. ACM order number 472950.

Faloutsos:1995:IMD

[968] Christos Faloutsos. Indexing multimedia databases. In ACM [2275], page 467. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p467-faloutsos/p467-faloutsos.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/223784/p467-faloutsos/

. ACM order number 472950.

Melton:1995:OET

[969] Jim Melton and Nelson Mendoca Mattos. An overview of the emerging third-generation SQL standard. In ACM [2275], page 468. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p468-melton/p468-melton.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/223784/p468-melton/. ACM or-

der number 472950.

Sheth:1995:WAA

[970] Amit Sheth. Workflow automation: applications, technology and research. In ACM [2275], page 469. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p469-sheth/p469-sheth.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/223784/p469-sheth/. ACM or-

der number 472950.

Norrie:1995:IAC

[971] Moira C. Norrie. Integration approaches for CIM. In ACM [2275], page 470. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p470-norrie/p470-norrie.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p470-norrie/. ACM or-

der number 472950.

Gibson:1995:TST

[972] Garth A. Gibson. Tutorial on storage technology: RAID and beyond. In ACM [2275], page 471. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p471-gibson/p471-gibson.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p471-gibson/. ACM or-

der number 472950.

Imielinski:1995:DIR

[973] T. Imielinski and A. Virmani. DataMine– interactive rule discovery system. In ACM [2275], page 472. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p472-imielinski/p472-imielinski.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/223784/
p472-imielinski/. ACM order num-

ber 472950.

Ceri:1995:ATC

[974] Stefano Ceri, Piero Fraternali, Stefano Paraboschi, and Giuseppe Psaila. The Algres testbed of CHIMERA: an active object-oriented database system. In ACM [2275], page 473. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p473-ceri/p473-ceri.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/223784/p473-ceri/. ACM order

number 472950.

Collet:1995:NS

[975] C. Collet and T. Coupaye. The NAOS system. In ACM [2275], page 474. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p474-collet/p474-collet.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/223784/p474-collet/. ACM or-

der number 472950.

Ashley:1995:QIC

[976] Jonathan Ashley, Myron Flickner, James Hafner, Denis Lee, Wayne Niblack, and Dragutin Petkovic. The query by image content (QBIC) system. In ACM [2275], page 475. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p475-ashley/p475-ashley.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/223784/p475-ashley/. ACM or-

der number 472950.

Buchmann:1995:RAO

[977] A. P. Buchmann, A. Deutsch, J. Zimmermann, and M. Higa. The REACH active OODBMS. In ACM [2275], page 476. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/

pubs/articles/proceedings/mod/223784/
p476-buchmann/p476-buchmann.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/223784/p476-buchmann/

. ACM order number 472950.

Massari:1995:QQI

[978] Antonio Massari, Stefano Pavani, Lorenzo Saladini, and Panos K. Chrysanthis. QBI: query by icons. In ACM [2275], page 477. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p477-massari/p477-massari.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/223784/p477-massari/

. ACM order number 472950.

Shklar:1995:ISS

[979] Leon Shklar, Amit Sheth, Vipul Kashyap, and Satish Thatte. Info-Harness: a system for search and retrieval of heterogeneous information. In ACM [2275], page 478. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p478-shklar/p478-shklar.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p478-shklar/. ACM or-

der number 472950.

Thomas:1995:DIA

[980] J. Thomas, S. Deßloch, and N. Mattos. Design and implementation of advanced knowledge processing in the KBMS KRISYS. In ACM [2275], page 479. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/

pubs/articles/proceedings/mod/223784/
p479-thomas/p479-thomas.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p479-thomas/. ACM or-

der number 472950.

Gatziu:1995:SAD

[981] Stella Gatziu, Andreas Geppert, and Klaus R. Dittrich. The SAMOS active DBMS prototype. In ACM [2275], page 480. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p480-gatziu/p480-gatziu.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p480-gatziu/. ACM or-

der number 472950.

Radeke:1995:EFD

[982] E. Radeke, R. Böttger, B. Burkert, Y. Engel, G. Kachel, S. Kolmschlag, and D. Nolte. Efendi: federated database system of Cadlab. In ACM [2275], page 481. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p481-radeke/p481-radeke.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p481-radeke/. ACM or-

der number 472950.

Keim:1995:VSV

[983] Daniel A. Keim and Hans-Peter Kriegel. VisDB: a system for visualizing large databases. In ACM [2275], page 482. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/

p482-keim/p482-keim.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p482-keim/. ACM order

number 472950.

Hammer:1995:ITM

[984] Joachim Hammer, Héctor García-Molina, Kelly Ireland, Yannis Papakonstantinou, Jeffrey Ullman, and Jennifer Widom. Information translation, mediation, and mosaic-based browsing in the TSIMMIS system. In ACM [2275], page 483. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p483-hammer/p483-hammer.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p483-hammer/. ACM or-

der number 472950.

Li:1995:SSP

[985] Wen-Syan Li and Chris Clifton. Semint: a system prototype for semantic integration in heterogeneous databases. In ACM [2275], page 484. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p484-li/p484-li.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/223784/p484-li/. ACM order

number 472950.

Team:1995:PDS

[986] The Paradise Team. Paradise: a database system for GIS applications. In ACM [2275], page 485. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/

                                        pubs/articles/proceedings/mod/223784/
                                        p485-the_paradise_team/p485-the_
                                        paradise_team.pdf;   http://www.
                                        acm.org/pubs/citations/proceedings/
                                        mod/223784/p485-the_paradise_team/

. ACM order number 472950.

Team:1995:SCB

[987] The SHORE Team. Shore: combining the best features of OODBMS and file systems. In ACM [2275], page 486. ISBN 0-89791-731-

6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p486-the_shore_team/p486-the_shore_
                                        team.pdf;   http://www.acm.org/
                                        pubs/citations/proceedings/mod/
                                        223784/p486-the_shore_team/. ACM

order number 472950.

Wang:1995:PMP

[988] Jason T. L. Wang, Kaizhong Zhang, and Dennis Shasha. Pattern matching and pattern discovery in scientific, program, and document databases. In ACM [2275], page 487. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/223784/
                                        p487-wang/p487-wang.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/223784/p487-wang/. ACM order

number 472950.

Grossman:1995:PLW

[989] R. L. Grossman, D. Hanley, and X. Qin. PTool: a light weight persistent object manager. In ACM [2275], page 488. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p488-grossman/p488-grossman.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/223784/p488-grossman/

. ACM order number 472950.

Dreyer:1995:UCT

[990] Werner Dreyer, Angelika Kotz Dittrich, and Duri Schmidt. Using the CA-LANDA time series management system. In ACM [2275], page 489. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p489-dreyer/p489-dreyer.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p489-dreyer/. ACM or-

der number 472950.

Jonker:1995:EMD

[991] Willem Jonker and Heribert Schütz. The ECRC multi database system. In ACM [2275], page 490. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/223784/
p490-jonker/p490-jonker.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/223784/p490-jonker/. ACM or-

der number 472950.

Goodman:1995:TRP

[992] N. Goodman. Tutorial: Research problems in genome databases p. 125. In ACM [2274], page ?? ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

Aho:1996:AIG

[993] Alfred V. Aho. Accessing information from globally distributed knowledge repositories (extended abstract).

In ACM [2276], pages 1–4. ISBN 0-89791-781-2. LCCN QA 76.9 D3

A26 1996. URL http://www.acm.org/
pubs/articles/proceedings/pods/
237661/p1-aho/p1-aho.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
pods/237661/p1-aho/;   http://www.
acm.org:80/pubs/citations/proceedings/
pods/237661/p1-aho/. ACM order

number 475960.

Srikant:1996:MQA

[994] Ramakrishnan Srikant and Rakesh Agrawal. Mining quantitative association rules in large relational tables. In Jagadish and Mumick [2277], pages 1–12. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p1-srikant/p1-srikant.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p1-srikant/.

Benedikt:1996:REP

[995] Michael Benedikt, Guozhu Dong, Leonid Libkin, and Limsoon Wong. Relational expressive power of constraint query languages. In ACM [2276], pages 5–16. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26

1996. URL http://www.acm.org/
pubs/articles/proceedings/pods/
237661/p5-benedikt/p5-benedikt.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/237661/
p5-benedikt/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/237661/p5-benedikt/. ACM

order number 475960.

Fukuda:1996:DMU

[996] Takeshi Fukuda, Yasukiko Morimoto, Shinichi Morishita, and Takeshi Tokuyama. Data mining using two-dimensional optimized association rules: scheme, algorithms, and visualization. In Jagadish and Mumick [2277], pages 13–23. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p13-fukuda/p13-fukuda.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p13-fukuda/.

Stolboushkin:1996:LVO

[997] Alexei P. Stolboushkin and Michael A. Taitslin. Linear vs. order constraint queries over rational databases (extended abstract). In ACM [2276], pages 17–27. ISBN 0-89791-781-

2. LCCN QA 76.9 D3 A26

1996. URL http://www.acm.org/
pubs/articles/proceedings/pods/
237661/p17-stolboushkin/p17-stolboushkin.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/237661/
p17-stolboushkin/;   http://www.
acm.org:80/pubs/citations/proceedings/
pods/237661/p17-stolboushkin/. ACM

order number 475960.

Selfridge:1996:IID

[998] Peter G. Selfridge, Divesh Srivastava, and Lynn O. Wilson. IDEA: interactive data exploration and analysis. In Jagadish and Mumick [2277], pages 24–34. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p24-selfridge/p24-selfridge.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/233269/p24-selfridge/

.

Grumbach:1996:TPC

[999] Stéphane Grumbach and Jianwen Su. Towards practical constraint databases (extended abstract). In ACM [2276], pages 28–39. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26

                                        1996. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        237661/p28-grumbach/p28-grumbach.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/237661/
                                        p28-grumbach/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/237661/p28-grumbach/. ACM

order number 475960.

Vance:1996:RBJ

[1000] Bennet Vance and David Maier. Rapid bushy join-order optimization with Cartesian products. In Jagadish and Mumick [2277], pages 35–46. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p35-vance/p35-vance.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/233269/p35-vance/.

Chomicki:1996:VIA

[1001] J. Chomicki, D. Q. Goldin, and G. M. Kuper. Variable independence and aggregation closure. In ACM [2276], pages 40–48. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26

                                        1996. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        237661/p40-chomicki/p40-chomicki.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/237661/
                                        p40-chomicki/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/

pods/237661/p40-chomicki/. ACM

order number 475960.

Goel:1996:SQO

[1002] Piyush Goel and Bala Iyer. SQL query optimization: reordering for a general class of queries. In Jagadish and Mumick [2277], pages 47–

56. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p47-goel/p47-goel.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/233269/p47-goel/.

Abiteboul:1996:TVF

[1003] Serge Abiteboul, Laurent Herr, and Jan Van den Bussche. Temporal versus first-order logic to query temporal databases. In ACM [2276], pages 49–57. ISBN 0-89791-781-

2. LCCN QA 76.9 D3 A26

1996. URL http://www.acm.org/
pubs/articles/proceedings/pods/
237661/p49-abiteboul/p49-abiteboul.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/237661/
p49-abiteboul/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/237661/p49-abiteboul/. ACM

order number 475960.

Simmen:1996:FTO

[1004] David Simmen, Eugene Shekita, and Timothy Malkemus. Fundamental techniques for order optimization. In Jagadish and Mumick [2277], pages 57–67. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p57-simmen/p57-simmen.pdf;   http:

                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/233269/p57-simmen/.

Toman:1996:PVI

[1005] David Toman. Point vs. intervalbased query languages for temporal databases (extended abstract). In ACM [2276], pages 58–67. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26

                                        1996. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        237661/p58-toman/p58-toman.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/237661/p58-toman/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/237661/

p58-toman/. ACM order number

475960.

Bettini:1996:TCT

[1006] Claudio Bettini, X. Sean Wang, and Sushil Jajodia. Testing complex temporal relationships involving multiple granularities and its application to data mining (extended abstract). In ACM [2276], pages 68–78. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26

                                        1996. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        237661/p68-bettini/p68-bettini.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/237661/
                                        p68-bettini/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/237661/p68-bettini/. ACM

order number 475960.

OConnell:1996:TCB

[1007] W. O’Connell, I. T. Ieong, D. Schrader, C. Watson, G. Au, A. Biliris, S. Choo, P. Colin, G. Linderman, E. Panagos, J. Wang, and T. Walter. A Teradata content-based multimedia object manager for massively parallel architectures. In Jagadish and Mumick [2277], pages 68–78. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996;

QA1 .A87. URL http://www.acm.
org/pubs/citations/proceedings/
mod/233269/p68-o_connell/.

Anonymous:1996:MPC

[1008] Anonymous. In memoriam: Paris C.

Kanellakis. In ACM [2276], page 79. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26 1996. URL http:

//www.acm.org:80/pubs/citations/
proceedings/pods/237661/p79-????

/. ACM order number 475960.

Hull:1996:MPC

[1009] Richard Hull. In memoriam: Paris C.

Kanellakis. In ACM [2276], page 79. ISBN 0-89791-781-2. LCCN QA 76.9

D3 A26 1996. URL http://www.acm.
org/pubs/articles/proceedings/pods/
237661/p79-hull/p79-hull.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/237661/p79-hull/

. ACM order number 475960.

Ozden:1996:FTA

[1010] Banu Özden, Rajeev Rastogi, Prashant Shenoy, and Avi Silberschatz. Faulttolerant architectures for continuous media servers. In Jagadish and Mumick [2277], pages 79–90. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p79-ozden/p79-ozden.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p79-ozden/.

Egenhofer:1996:GDS

[1011] Max J. Egenhofer. Geographic database systems: issues and research needs. In ACM [2276], page 80. ISBN 0-89791-781-2. LCCN QA 76.9 D3

A26 1996. URL http://www.acm.org/
pubs/articles/proceedings/pods/
237661/p80-egenhofer/p80-egenhofer.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/237661/
p80-egenhofer/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/237661/p80-egenhofer/. ACM

order number 475960.

Papadimitriou:1996:TQS

[1012] C. H. Papadimitriou, D. Suciu, and V. Vianu. Topological queries in spatial databases. In ACM [2276], pages 81–92. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26

1996. URL http://www.acm.org/
pubs/articles/proceedings/pods/
237661/p81-papadimitriou/p81-papadimitriou.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/237661/
p81-papadimitriou/;   http://www.
acm.org:80/pubs/citations/proceedings/
pods/237661/p81-papadimitriou/.

ACM order number 475960.

Chaudhuri:1996:OQM

[1013] Surajit Chaudhuri and Luis Gravano. Optimizing queries over multimedia repositories. In Jagadish and Mumick [2277], pages 91–102. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p91-chaudhuri/p91-chaudhuri.pdf;
 http://www.acm.org/pubs/citations/

proceedings/mod/233269/p91-chaudhuri/

.

Gyssens:1996:TPQ

[1014] Marc Gyssens, Laks V. S. Lakshmanan, and Iyer N. Subramanian. Tables as a paradigm for querying and restructuring (extended abstract). In ACM [2276], pages 93–103. ISBN 0-89791-781-2. LCCN QA 76.9 D3

A26 1996. URL http://www.acm.org/
pubs/articles/proceedings/pods/
237661/p93-gyssens/p93-gyssens.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/237661/
p93-gyssens/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/237661/p93-gyssens/. ACM

order number 475960.

Zhang:1996:BED

[1015] Tian Zhang, Raghu Ramakrishnan, and Miron Livny. BIRCH: an efficient data clustering method for very large databases. In Jagadish and Mumick [2277], pages 103–114. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p103-zhang/p103-zhang.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p103-zhang/.

Beeri:1996:GP

[1016] C. Beeri, T. Milo, and P. Ta-Shma. On genericity and parametricity. In ACM [2276], pages 104–116. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26 1996. ACM order number 475960.

Beeri:1996:GPE

[1017] Catriel Beeri, Tova Milo, and Paula Ta-Shma. On genericity and parametricity (extended abstract). In ACM [2276], pages 104–116. ISBN 0-89791-781-2. LCCN QA 76.9 D3

A26 1996. URL http://www.acm.org/
pubs/articles/proceedings/pods/
237661/p104-beeri/p104-beeri.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/237661/p104-beeri/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/237661/

p104-beeri/. ACM order number 475960.

Zou:1996:LRS

[1018] Chendong Zou and Betty Salzberg. On-line reorganization of sparselypopulated B^{+}-trees. In Jagadish and Mumick [2277], pages 115–124. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p115-zou/p115-zou.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/233269/p115-zou/.

Benedikt:1996:VPD

[1019] Michael Benedikt, Timothy Griffin, and Leonid Libkin. Verifiable properties of database transactions. In ACM [2276], pages 117–127. ISBN 0-89791-781-2. LCCN QA 76.9 D3

A26 1996. URL http://www.acm.org/
pubs/articles/proceedings/pods/
237661/p117-benedikt/p117-benedikt.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/237661/
p117-benedikt/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/237661/p117-benedikt/. ACM

order number 475960.

Achyutuni:1996:TTL

[1020] Kiran J. Achyutuni, Edward Omiecinski, and Shamkant B. Navathe. Two techniques for on-line index modification in shared nothing parallel databases. In Jagadish and Mumick [2277], pages 125–136. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p125-achyutuni/p125-achyutuni.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/233269/p125-achyutuni/

.

Maher:1996:CCT

[1021] Michael J. Maher and Divesh Srivastava. Chasing constrained tuplegenerating dependencies. In ACM [2276], pages 128–138 (or 127–138??). ISBN 0-89791-781-2. LCCN QA 76.9

D3 A26 1996. URL http://www.acm.
org/pubs/articles/proceedings/pods/
237661/p128-maher/p128-maher.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/237661/p128-maher/

. ACM order number 475960.

Adali:1996:QCO

[1022] S. Adali, K. S. Candan, Y. Papakonstantinou, and V. S. Subrahmanian. Query caching and optimization in distributed mediator systems. In Jagadish and Mumick [2277], pages 137–

146. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/citations/proceedings/mod/
233269/p137-adali/.

Wong:1996:RTF

[1023] Man Hon Wong. Recovery for transaction failures in object-based databases. In ACM [2276], pages 139–149. ISBN 0-89791-781-2. LCCN QA 76.9 D3

                                        A26 1996. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        237661/p139-wong/p139-wong.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/237661/p139-wong/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/237661/

p139-wong/. ACM order number 475960.

Franklin:1996:PTC

[1024] Michael J. Franklin, Björn Thór Jónsson, and Donald Kossmann. Performance tradeoffs for client-server query processing. In Jagadish and Mumick [2277], pages 149–160. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p149-franklin/p149-franklin.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/233269/p149-franklin/

.

Pagel:1996:WQR

[1025] Bernd-Uwe Pagel and Hans-Werner Six. Are window queries representative for arbitrary range queries? In ACM [2276], pages 150–160. ISBN 0-89791-781-2. LCCN QA 76.9 D3

                                        A26 1996. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        237661/p150-pagel/p150-pagel.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/237661/p150-pagel/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/237661/

p150-pagel/. ACM order number 475960.

Theodoridis:1996:MPR

[1026] Yannis Theodoridis and Timos Sellis. A model for the prediction of R-tree performance. In ACM [2276], pages 161–171. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26

1996. URL http://www.acm.org/
pubs/articles/proceedings/pods/
237661/p161-theodoridis/p161-theodoridis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/237661/
p161-theodoridis/;   http://www.
acm.org:80/pubs/citations/proceedings/
pods/237661/p161-theodoridis/. ACM

order number 475960.

Blakeley:1996:DAM

[1027] José A. Blakeley. Data access for the masses through OLE DB. In Jagadish and Mumick [2277], pages 161–

172. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p161-blakeley/p161-blakeley.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/233269/p161-blakeley/

.

Ganguly:1996:EAC

[1028] Sumit Ganguly, Akshay Goel, and Avi Silberschatz. Efficient and accurate cost models for parallel query optimization (extended abstract). In ACM [2276], pages 172–181. ISBN 0-89791-781-2. LCCN QA 76.9 D3

A26 1996. URL http://www.acm.org/
pubs/articles/proceedings/pods/
237661/p172-ganguly/p172-ganguly.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/237661/
p172-ganguly/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/237661/p172-ganguly/. ACM

order number 475960.

Gray:1996:DRS

[1029] Jim Gray, Pat Helland, Patrick O’Neil, and Dennis Shasha. The dangers of replication and a solution. In Jagadish and Mumick [2277], pages 173–

182. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p173-gray/p173-gray.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/233269/p173-gray/.

Fukuda:1996:MOA

[1030] Takeshi Fukuda, Yasuhido Morimoto, Shinichi Morishita, and Takeshi Tokuyama. Mining optimized association rules for numeric attributes. In ACM [2276], pages 182–191. ISBN 0-89791-781-2. LCCN QA 76.9 D3

                                        A26 1996. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        237661/p182-fukuda/p182-fukuda.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/237661/
                                        p182-fukuda/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/237661/p182-fukuda/. ACM

order number 475960.

Mogi:1996:HMM

[1031] Kazuhiko Mogi and Masaru Kitsuregawa. Hot mirroring: a method of hiding parity update penalty and degradation during rebuilds for RAID5. In Jagadish and Mumick [2277], pages 183–194. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p183-mogi/p183-mogi.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/233269/p183-mogi/.

Jagadish:1996:MCB

[1032] H. V. Jagadish, Alberto O. Mendelzon, and Inderpal Singh Mumick. Managing conflicts between rules (extended abstract). In ACM [2276], pages 192–201. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26

1996. URL http://www.acm.org/
pubs/articles/proceedings/pods/
237661/p192-jagadish/p192-jagadish.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/237661/
p192-jagadish/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/237661/p192-jagadish/. ACM

order number 475960.

Jagadish:1996:MRC

[1033] H. V. Jagadish, A. O. Mendelzon, and I. S. Mumick. Managing rule conflicts in an active database. In ACM [2276], pages 192–201. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26 1996. ACM order number 475960.

Hillyer:1996:RSO

[1034] Bruce K. Hillyer and Avi Silberschatz. Random I/O scheduling in online tertiary storage systems. In Jagadish and Mumick [2277], pages 195–204. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p195-hillyer/p195-hillyer.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/233269/p195-hillyer/

.

Bertino:1996:SAI

[1035] Elisa Bertino and Barbara Catania.

Static analysis of intensional databases in U-Datalog (extended abstract). In ACM [2276], pages 202–212. ISBN 0-89791-781-2. LCCN QA 76.9 D3

A26 1996. URL http://www.acm.org/
pubs/articles/proceedings/pods/
237661/p202-bertino/p202-bertino.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/237661/
p202-bertino/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/237661/p202-bertino/. ACM

order number 475960.

Harinarayan:1996:IDC

[1036] Venky Harinarayan, Anand Rajaraman, and Jeffrey D. Ullman. Implementing data cubes efficiently. In Jagadish and Mumick [2277], pages 205–

216. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p205-harinarayan/p205-harinarayan.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/233269/
p205-harinarayan/.

Manber:1996:FDR

[1037] Udi Manber. Future directions and research problems in the World Wide Web. In ACM [2276], pages 213–215. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26

1996. URL http://www.acm.org/
pubs/articles/proceedings/pods/
237661/p213-manber/p213-manber.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/237661/
p213-manber/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/237661/p213-manber/. ACM

order number 475960.

Fagin:1996:CFI

[1038] Ronald Fagin. Combining fuzzy information from multiple systems (extended abstract). In ACM [2276], pages 216–226. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26

1996. URL http://www.acm.org/
pubs/articles/proceedings/pods/
237661/p216-fagin/p216-fagin.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/237661/p216-fagin/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/237661/

p216-fagin/. ACM order number 475960.

Rao:1996:PBS

[1039] Sudhir G. Rao, Antonio Badia, and Dirk van Gucht. Providing better support for a class of decision support queries. In Jagadish and Mumick [2277], pages 217–227. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p217-rao/p217-rao.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/233269/p217-rao/.

Levy:1996:AQU

[1040] Alon Y. Levy, Anand Rajaraman, and Jeffrey D. Ullman. Answering queries using limited external query processors (extended abstract). In ACM [2276], pages 227–237. ISBN 0-89791-781-2. LCCN QA 76.9 D3

A26 1996. URL http://www.acm.org/
pubs/articles/proceedings/pods/
237661/p227-levy/p227-levy.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/237661/p227-levy/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/237661/

p227-levy/. ACM order number 475960.

Libkin:1996:QLM

[1041] Leonid Libkin, Rona Machlin, and Limsoon Wong. A query language for multidimensional arrays: design, implementation, and optimization techniques. In Jagadish and Mumick [2277], pages 228–239. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p228-libkin/p228-libkin.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/233269/p228-libkin/.

Rajaraman:1996:IIO

[1042] Anand Rajaraman and Jeffrey D. Ullman. Integrating information by outerjoins and full disjunctions (extended abstract). In ACM [2276], pages 238–248. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26

                                        1996. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        237661/p238-rajaraman/p238-rajaraman.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/237661/
                                        p238-rajaraman/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/237661/p238-rajaraman/. ACM

order number 475960.

Agarwal:1996:SSS

[1043] Ramesh C. Agarwal. A super scalar sort algorithm for RISC processors. In Jagadish and Mumick [2277], pages 240–246. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p240-agarwal/p240-agarwal.pdf;
                                        http://www.acm.org/pubs/citations/

proceedings/mod/233269/p240-agarwal/

.

Lo:1996:SHJ

[1044] Ming-Ling Lo and Chinya V. Ravishankar. Spatial hash-joins. In Jagadish and Mumick [2277], pages 247–

258. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p247-lo/p247-lo.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/233269/p247-lo/.

Patel:1996:PBS

[1045] Jignesh M. Patel and David J. DeWitt.

Partition based spatial-merge join. In Jagadish and Mumick [2277], pages 259–270. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p259-patel/p259-patel.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p259-patel/.

Ganguly:1996:BSS

[1046] Sumit Ganguly, Phillip B. Gibbons, Yossi Matias, and Avi Silberschatz. Bifocal sampling for skew-resistant join size estimation. In Jagadish and Mumick [2277], pages 271–281. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p271-ganguly/p271-ganguly.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/233269/p271-ganguly/

.

Krishnan:1996:EAS

[1047] P. Krishnan, Jeffrey Scott Vitter, and Bala Iyer. Estimating alphanumeric selectivity in the presence of wildcards. In Jagadish and Mumick [2277], pages 282–293. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p282-krishnan/p282-krishnan.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/233269/p282-krishnan/

.

Poosala:1996:IHS

[1048] Viswanath Poosala, Peter J. Haas, Yannis E. Ioannidis, and Eugene J. Shekita. Improved histograms for selectivity estimation of range predicates. In Jagadish and Mumick [2277], pages 294–305. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p294-poosala/p294-poosala.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/233269/p294-poosala/

.

Doherty:1996:SMP

[1049] Michael Doherty, Richard Hull, and Mohammed Rupawalla. Structures for manipulating proposed updates in object-oriented databases. In Jagadish and Mumick [2277], pages 306–

317. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p306-doherty/p306-doherty.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/233269/p306-doherty/

.

Liskov:1996:SES

[1050] B. Liskov, A. Adya, M. Castro, S. Ghemawat, R. Gruber, U. Maheshwari, A. C. Myers, M. Day, and L. Shrira. Safe and efficient sharing of persistent objects in Thor. In Jagadish and Mumick [2277], pages 318–329. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p318-liskov/p318-liskov.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p318-liskov/.

Blott:1996:OAO

[1051] Stephen Blott, Lukas Relly, and Hans-Jörg Schek. An open abstractobject storage system. In Jagadish and Mumick [2277], pages 330–340. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p330-blott/p330-blott.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p330-blott/.

Tajima:1996:SDS

[1052] Keishi Tajima. Static detection of security flaws in object-oriented databases. In Jagadish and Mumick [2277], pages 341–352. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p341-tajima/p341-tajima.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p341-tajima/.

Brown:1996:GOB

[1053] Kurt P. Brown, Michael J. Carey, and Miron Livny. Goal-oriented buffer management revisited. In Jagadish and Mumick [2277], pages 353–364. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p353-brown/p353-brown.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p353-brown/.

Garofalakis:1996:MDR

[1054] Minos N. Garofalakis and Yannis E. Ioannidis. Multi-dimensional resource scheduling for parallel queries. In Jagadish and Mumick [2277], pages 365–

376. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p365-garofalakis/p365-garofalakis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/233269/
p365-garofalakis/.

Cook:1996:SAS

[1055] Jonathan E. Cook, Artur W. Klauser, Alexander L. Wolf, and Benjamin G. Zorn. Semi-automatic, self-adaptive control of garbage collection rates in object databases. In Jagadish and Mumick [2277], pages 377–388. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p377-cook/p377-cook.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p377-cook/.

McAuliffe:1996:TEE

[1056] Mark L. McAuliffe, Michael J. Carey, and Marvin H. Solomon. Towards effective and efficient free space management. In Jagadish and Mumick [2277], pages 389–400. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p389-mcauliffe/p389-mcauliffe.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/233269/p389-mcauliffe/

.

Cherniack:1996:RLI

[1057] Mitch Cherniack and Stanley B.

Zdonik. Rule languages and internal algebras for rule-based optimizers. In Jagadish and Mumick [2277], pages 401–412. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p401-cherniack/p401-cherniack.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/233269/p401-cherniack/

.

Christophides:1996:EQG

[1058] Vassilis Christophides, Sophie Cluet, and Guido Moerkotte. Evaluating queries with generalized path expressions. In Jagadish and Mumick [2277], pages 413–422. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p413-christophides/p413-christophides.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/233269/
p413-christophides/.

Hellerstein:1996:QET

[1059] Joseph M. Hellerstein and Jeffrey F.

Naughton. Query execution techniques for caching expensive methods. In Jagadish and Mumick [2277], pages 423–434. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p423-hellerstein/p423-hellerstein.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/233269/
p423-hellerstein/.

Seshadri:1996:CBO

[1060] Praveen Seshadri, Joseph M. Hellerstein, Hamid Pirahesh, T. Y. Cliff Leung, Raghu Ramakrishnan, Divesh Srivastava, Peter J. Stuckey, and S. Sudarshan. Cost-based optimization for magic: algebra and implementation. In Jagadish and Mumick [2277], pages 435–446. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p435-seshadri/p435-seshadri.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/233269/p435-seshadri/

.

Ross:1996:MVM

[1061] Kenneth A. Ross, Divesh Srivastava, and S. Sudarshan. Materialized view maintenance and integrity constraint checking: trading space for time. In Jagadish and Mumick [2277], pages 447–458. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p447-ross/p447-ross.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p447-ross/.

Morpain:1996:MDC

[1062] Claire Morpain, Michéle Cart, Jean Ferrié, and Jean-François Pons. Maintaining database consistency in presence of value dependencies in multidatabase systems. In Jagadish and Mumick [2277], pages 459–468. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p459-morpain/p459-morpain.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/233269/p459-morpain/

.

Colby:1996:ADV

[1063] Latha S. Colby, Timothy Griffin, Leonid Libkin, Inderpal Singh Mumick, and Howard Trickey. Algorithms for deferred view maintenance. In Jagadish and Mumick [2277], pages 469–

480. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p469-colby/p469-colby.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p469-colby/.

Hull:1996:FSD

[1064] Richard Hull and Gang Zhou. A framework for supporting data integration using the materialized and virtual approaches. In Jagadish and Mumick [2277], pages 481–492. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p481-hull/p481-hull.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p481-hull/.

Chawathe:1996:CDH

[1065] Sudarshan S. Chawathe, Anand Rajaraman, Hector Garcia-Molina, and Jennifer Widom. Change detection in hierarchically structured information. In Jagadish and Mumick [2277], pages 493–504. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p493-chawathe/p493-chawathe.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/233269/p493-chawathe/

.

Buneman:1996:QLO

[1066] Peter Buneman, Susan Davidson, Gerd Hillebrand, and Dan Suciu. A query language and optimization techniques for unstructured data. In Jagadish and Mumick [2277], pages 505–516. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p505-buneman/p505-buneman.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/233269/p505-buneman/

.

Goyal:1996:GPD

[1067] Nita Goyal, Charles Hoch, Ravi Krishnamurthy, Brian Meckler, and Michael Suckow. Is GUI programming a database research problem? In Jagadish and Mumick [2277], pages 517–

528. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p517-goyal/p517-goyal.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/233269/p517-goyal/.

Nguyen:1996:ARD

[1068] Tam Nguyen and V. Srinivasan. Accessing relational databases from the World Wide Web. In Jagadish and Mumick [2277], pages 529–540. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p529-nguyen/p529-nguyen.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p529-nguyen/.

Fernandez:1996:IOE

[1069] Phil Fernandez and Donovan Schneider. The ins and outs (and everything in between) of data warehousing. In Jagadish and Mumick [2277], page 541. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p541-fernandez/p541-fernandez.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/233269/p541-fernandez/

.

Bernstein:1996:RSE

[1070] Pillip A. Bernstein. Repository system engineering. In Jagadish and Mumick [2277], page 542. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p542-bernstein/p542-bernstein.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/233269/p542-bernstein/

.

Keim:1996:DV

[1071] Daniel A. Keim. databases and visualization. In Jagadish and Mumick [2277], page 543. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p543-keim/p543-keim.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/233269/p543-keim/.

Mohan:1996:SAW

[1072] C. Mohan. State of the art in workflow management research and products. In Jagadish and Mumick [2277], page 544. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p544-mohan/p544-mohan.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/233269/p544-mohan/.

Han:1996:DMT

[1073] Jiawei Han. Data mining techniques. In Jagadish and Mumick [2277], page

545. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p545-han/p545-han.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/233269/p545-han/.

Piatko:1996:TTT

[1074] Peter Piatko, Roman Yangarber, Daoi Lin, and Dennis Shasha. Thinksheet: a tool for tailoring complex documents. In Jagadish and Mumick [2277], page

546. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p546-piatko/p546-piatko.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/233269/p546-piatko/.

Bohm:1996:HAS

[1075] Klemens Böhm and Karl Aberer.

HyperStorM–administering structured documents using object-oriented database technology. In Jagadish and Mumick [2277], page 547. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p547-bohm/p547-bohm.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p547-bohm/.

Lefler:1996:DST

[1076] Mark Lefler, Mark Stokrp, and Craig Wong. DBSim: a simulation tool for predicting database performance. In Jagadish and Mumick [2277], page

548. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p548-lefler/p548-lefler.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p548-lefler/.

Quass:1996:LLO

[1077] Dallan Quass, Jennifer Widom, Roy Goldman, Kevin Haas, Qingshan Luo, Jason McHugh, Svetlozar Nestorov, Anand Rajaraman, Hugo Rivero, Serge Abiteboul, Jeff Ullman, and Janet Wiener. LORE: a Lightweight Object REpository for semistructured data. In Jagadish and Mumick [2277], page

549. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p549-quass/p549-quass.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p549-quass/.

Han:1996:DIM

[1078] Jaiwei Han, Youngjian Fu, Wei Wang, Jenny Chiang, Osmar R. Zaïane, and Krzysztof Koperski. DBMiner: interactive mining of multiple-level knowledge in relational databases. In Jagadish and Mumick [2277], page 550. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p550-han/p550-han.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/233269/p550-han/.

Choo:1996:PCB

[1079] S. Choo, W. O’Connell, G. Linerman, H. Chen, K. Ganapathy, A. Biliris, E. Panagos, and D. Schrader. prospector: a content-based multimedia server for massively parallel architectures. In Jagadish and Mumick [2277], page

551. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p551-choo/p551-choo.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/233269/p551-choo/.

Dogac:1996:MID

[1080] Asuman Dogac, Ugur Halici, Ebru Kilic, Gokhan Ozhan, Fatma Ozcan, Sena Nural, Cevdet Dengi, Sema Mancuhan, Budak Arpinar, Pinar Koksal, and Cem Evrendilek. METU interoperable database system. In Jagadish and Mumick [2277], page

552. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p552-dogac/p552-dogac.pdf;   http:

//www.acm.org/pubs/citations/proceedings/
mod/233269/p552-dogac/.

Fukuda:1996:SSO

[1081] Takeshi Fukuda, Yasuhiko Morimoto, Shinichi Morishita, and Takeshi Tokuyama. SONAR: system for optimized numeric association rules. In Jagadish and Mumick [2277], page

553. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p553-fukuda/p553-fukuda.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p553-fukuda/.

Hung:1996:CAC

[1082] Patrick C. K. Hung, Helen P. Yeung, and Kamalakar Karlapalem. CapBasED-AMS: a capability-based and event-driven activity management system. In Jagadish and Mumick [2277], page 554. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/
p554-hung/p554-hung.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/233269/p554-hung/.

Rundensteiner:1996:MPO

[1083] E. A. Rundensteiner, H. A. Kuno, Y.-G. Ra, V. Crestana-Taube, M. C. Jones, and P. J. Marron. The MultiView project: object-oriented view technology and applications. In Jagadish and Mumick [2277], page 555. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/233269/

                                        p555-rundensteiner/p555-rundensteiner.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/233269/
                                        p555-rundensteiner/.

Biliris:1996:BSS

[1084] A. Biliris, T. A. Funkhouser, W. O’Connell, and E. Panagos. BeSS: storage support for interactive visualization systems. In Jagadish and Mumick [2277], page 556. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p556-biliris/p556-biliris.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/233269/p556-biliris/

.

Roth:1996:GP

[1085] M. Tork Roth, M. Arya, L. Haas, M. Carey, W. Cody, R. Fagin, P. Schwarz, J. Thomas, and E. Wimmers. The Garlic project. In Jagadish and Mumick [2277], page 557. ISBN 0-89791-794-4. LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/233269/
                                        p557-roth/p557-roth.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/233269/p557-roth/.

Ozden:1997:MSD

[1086] Banu Özden, Rajeev Rastogi, and Avi Silberschatz. Multimedia support for databases. In ACM [2279], pages 1–11. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26

                                        1997. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        263661/p1-ozden/p1-ozden.pdf;
                                        http://www.acm.org/pubs/citations/

proceedings/pods/263661/p1-ozden/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/263661/
p1-ozden/.

Berchtold:1997:FPS

[1087] Stefan Berchtold, Christian Böhm, Bernhard Braunmüller, Daniel A. Keim, and Hans-Peter Kriegel. Fast parallel similarity search in multimedia databases. In Peckman [2278], pages 1–12. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p1-berchtold/p1-berchtold.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/253260/p1-berchtold/

.

Papadimitriou:1997:CDQ

[1088] Christos H. Papadimitriou and Mihalis Yannakakis. On the complexity of database queries (extended abstract). In ACM [2279], pages 12–19. ISBN 0-89791-910-6. LCCN QA 76.9 D3

A26 1997. URL http://www.acm.org/
pubs/articles/proceedings/pods/
263661/p12-papadimitriou/p12-papadimitriou.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/263661/
p12-papadimitriou/;   http://www.
acm.org:80/pubs/citations/proceedings/
pods/263661/p12-papadimitriou/.

Rafiei:1997:SBQ

[1089] Davood Rafiei and Alberto Mendelzon. Similarity-based queries for time series data. In Peckman [2278], pages 13–25. ISBN 0-89791-911-

4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p13-rafiei/p13-rafiei.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/253260/p13-rafiei/.

Levy:1997:DCQ

[1090] Alon Y. Levy and Dan Suciu. Deciding containment for queries with complex objects (extended abstract). In ACM [2279], pages 20–31. ISBN 0-89791-910-6. LCCN QA 76.9 D3

                                        A26 1997. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        263661/p20-levy/p20-levy.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/pods/263661/p20-levy/
                                        ;   http://www.acm.org:80/pubs/
                                        citations/proceedings/pods/263661/
                                        p20-levy/.

Chawathe:1997:MCD

[1091] Sudarshan S. Chawathe and Hector Garcia-Molina. Meaningful change detection in structured data. In Peckman [2278], pages 26–37. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p26-chawathe/p26-chawathe.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/253260/p26-chawathe/

.

Ibarra:1997:CED

[1092] Oscar H. Ibarra and Jianwen Su. On the containment and equivalence of database queries with linear constraints (extended abstract). In ACM [2279], pages 32–43. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26

                                        1997. URL http://www.acm.org/

pubs/articles/proceedings/pods/
263661/p32-ibarra/p32-ibarra.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/263661/p32-ibarra/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/263661/
p32-ibarra/.

ONeil:1997:IQP

[1093] Patrick O’Neil and Dallan Quass. Improved query performance with variant indexes. In Peckman [2278], pages 38–49. ISBN 0-89791-911-

4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p38-o_neil/p38-o_neil.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p38-o_neil/.

Albert:1997:CQE

[1094] Joseph Albert, Yanis Ioannidis, and Raghu Ramakrishnan. Conjunctive query equivalence of keyed relational schemas (extended abstract). In ACM [2279], pages 44–50. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26

1997. URL http://www.acm.org/
pubs/articles/proceedings/pods/
263661/p44-albert/p44-albert.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/263661/p44-albert/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/263661/
p44-albert/.

Zaharioudakis:1997:HCC

[1095] Markos Zaharioudakis and Michael J.

Carey. Highly concurrent cache consistency for indices in clientserver database systems. In Peckman [2278], pages 50–61. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p50-zaharioudakis/p50-zaharioudakis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/253260/
p50-zaharioudakis/.

Hull:1997:MSH

[1096] Richard Hull. Managing semantic heterogeneity in databases: a theoretical prospective. In ACM [2279], pages 51–61. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26

1997. URL http://www.acm.org/
pubs/articles/proceedings/pods/
263661/p51-hull/p51-hull.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/263661/p51-hull/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/263661/
p51-hull/.

Gyssens:1997:CGQ

[1097] Marc Gyssens, Jan Van den Bussche, and Dirk Van Gucht. Complete geometrical query languages (extended abstract). In ACM [2279], pages 62–67. ISBN 0-89791-910-

6. LCCN QA 76.9 D3 A26

1997. URL http://www.acm.org/
pubs/articles/proceedings/pods/
263661/p62-gyssens/p62-gyssens.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/263661/
p62-gyssens/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/263661/p62-gyssens/.

Kornacker:1997:CRG

[1098] Marcel Kornacker, C. Mohan, and Joseph M. Hellerstein. Concurrency and recovery in generalized search trees. In Peckman [2278], pages 62–72. ISBN 0-89791-911-

4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p62-kornacker/p62-kornacker.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/253260/p62-kornacker/

.

Dumortier:1997:DSL

[1099] Freddy Dumortier, Marc Gyssens, Luc Vandeurzen, and Dirk Van Gucht. On the decidability of semi-linearity for semi-algebraic sets and its implications for spatial databases (extended abstract). In ACM [2279], pages 68–77. ISBN 0-89791-910-

6. LCCN QA 76.9 D3 A26

1997. URL http://www.acm.org/
pubs/articles/proceedings/pods/
263661/p68-dumortier/p68-dumortier.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/263661/
p68-dumortier/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/263661/p68-dumortier/.

Ho:1997:RQO

[1100] Ching-Tien Ho, Rakesh Agrawal, Nimrod Megiddo, and Ramakrishnan Srikant. Range queries in OLAP data cubes. In Peckman [2278], pages 73–88. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p73-ho/p73-ho.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/253260/p73-ho/.

Berchtold:1997:CMN

[1101] Stefan Berchtold, Christian Böhm, Daniel A. Keim, and Hans-Peter Kriegel. A cost model for nearest neighbor search in high-dimensional data space. In ACM [2279], pages 78–86. ISBN 0-89791-910-

6. LCCN QA 76.9 D3 A26

                                        1997. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        263661/p78-berchtold/p78-berchtold.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/263661/
                                        p78-berchtold/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/263661/p78-berchtold/.

Benedikt:1997:LRD

[1102] Michael Benedikt and Leonid Libkin. Languages for relational databases over interpreted structures. In ACM [2279], pages 87–98. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26

                                        1997. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        263661/p87-benedikt/p87-benedikt.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/263661/
                                        p87-benedikt/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/263661/p87-benedikt/.

Roussopoulos:1997:COB

[1103] Nick Roussopoulos, Yannis Kotidis, and Mema Roussopoulos. Cubetree: organization of and bulk incremental updates on the data cube. In Peckman [2278], pages 89–99. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p89-roussopoulos/p89-roussopoulos.

pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/253260/
p89-roussopoulos/.

Beeri:1997:RQU

[1104] Catriel Beeri, Alon Y. Levy, and Marie-Christine Rousset. Rewriting queries using views in description logics. In ACM [2279], pages 99–108. ISBN 0-89791-910-6. LCCN QA 76.9 D3

A26 1997. URL http://www.acm.org/
pubs/articles/proceedings/pods/
263661/p99-beeri/p99-beeri.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/263661/p99-beeri/
;   http://www.acm.org:80/pubs/
citations/proceedings/pods/263661/
p99-beeri/.

Mumick:1997:MDC

[1105] Inderpal Singh Mumick, Dallan Quass, and Barinderpal Singh Mumick. Maintenance of data cubes and summary tables in a warehouse. In Peckman [2278], pages 100–111. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p100-mumick/p100-mumick.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p100-mumick/.

Duschka:1997:ARQ

[1106] Oliver M. Duschka and Michael R.

Genesereth. Answering recursive queries using views. In ACM [2279], pages 109–116. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26

1997. URL http://www.acm.org/
pubs/articles/proceedings/pods/
263661/p109-duschka/p109-duschka.
pdf;   http://www.acm.org/pubs/

                                        citations/proceedings/pods/263661/
                                        p109-duschka/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/263661/p109-duschka/.

Tsuei:1997:DBS

[1107] Thin-Fong Tsuei, Allan N. Packer, and Keng-Tai Ko. Database buffer size investigation for OLTP workloads. In Peckman [2278], pages 112–122. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p112-tsuei/p112-tsuei.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/253260/p112-tsuei/.

Buneman:1997:SD

[1108] Peter Buneman. Semistructured data. In ACM [2279], pages 117–121. ISBN 0-89791-910-6. LCCN QA 76.9 D3

                                        A26 1997. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        263661/p117-buneman/p117-buneman.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/263661/
                                        p117-buneman/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/263661/p117-buneman/.

Abiteboul:1997:RPQ

[1109] Serge Abiteboul and Victor Vianu. Regular path queries with constraints. In ACM [2279], pages 122–133. ISBN 0-89791-910-6. LCCN QA 76.9 D3

                                        A26 1997. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        263661/p122-abiteboul/p122-abiteboul.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/263661/
                                        p122-abiteboul/;   http://www.acm.

org:80/pubs/citations/proceedings/
pods/263661/p122-abiteboul/.

Doppelhammer:1997:DPR

[1110] Joachen Doppelhammer, Thomas Höppler, Alfons Kemper, and Donald Kossmann. Database performance in the real world: TPC-D and SAP R/3. In Peckman [2278], pages 123–134. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p123-doppelhammer/p123-doppelhammer.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/253260/
p123-doppelhammer/.

Mendelzon:1997:FMW

[1111] Alberto O. Mendelzon and Tova Milo.

Formal models of Web queries. In ACM [2279], pages 134–143. ISBN 0-89791-910-6. LCCN QA 76.9 D3

A26 1997. URL http://www.acm.org/
pubs/articles/proceedings/pods/
263661/p134-mendelzon/p134-mendelzon.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/263661/
p134-mendelzon/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/263661/p134-mendelzon/.

Carey:1997:BOR

[1112] Michael J. Carey, David J. DeWitt, Jeffrey F. Naughton, Mohammad Asgarian, Paul Brown, Johannes E. Gehrke, and Dhaval N. Shah. The BUCKY object-relational benchmark. In Peckman [2278], pages 135–146. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/

                                        pubs/articles/proceedings/mod/253260/
                                        p135-carey/p135-carey.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/253260/p135-carey/.

Atzeni:1997:CP

[1113] Paolo Atzeni and Giansalvatore Mecca. Cut and paste. In ACM [2279], pages 144–153. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26

                                        1997. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        263661/p144-atzeni/p144-atzeni.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/263661/
                                        p144-atzeni/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/263661/p144-atzeni/.

Adelberg:1997:SRS

[1114] Brad Adelberg, Hector Garcia-Molina, and Jennifer Widom. The STRIP rule system for efficiently maintaining derived data. In Peckman [2278], pages 147–158. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p147-adelberg/p147-adelberg.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/253260/p147-adelberg/

.

Nerjes:1997:SSG

[1115] Guido Nerjes, Peter Muth, and Gerhard Weikum. Stochastic service guarantees for continuous data on multi-zone disks. In ACM [2279], pages 154–160. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26

                                        1997. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/

263661/p154-nerjes/p154-nerjes.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/263661/
p154-nerjes/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/263661/p154-nerjes/.

Zhao:1997:ABA

[1116] Yihong Zhao, Prasad M. Deshpande, and Jeffrey F. Naughton. An arraybased algorithm for simultaneous multidimensional aggregates. In Peckman [2278], pages 159–170. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p159-zhao/p159-zhao.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p159-zhao/.

Agrawal:1997:EAR

[1117] D. Agrawal, A. El Abbadi, and R. C.

Steinke. Epidemic algorithms in replicated databases (extended abstract). In ACM [2279], pages 161–172. ISBN 0-89791-910-6. LCCN QA 76.9 D3

A26 1997. URL http://www.acm.org/
pubs/articles/proceedings/pods/
263661/p161-agrawal/p161-agrawal.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/263661/
p161-agrawal/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/263661/p161-agrawal/.

Hellerstein:1997:OA

[1118] Joseph M. Hellerstein, Peter J. Haas, and Helen J. Wang. Online aggregation. In Peckman [2278], pages 171–182. ISBN 0-89791-911-

4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p171-hellerstein/p171-hellerstein.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/253260/
                                        p171-hellerstein/.

Breitbart:1997:RCB

[1119] Yuri Breitbart and Henry F. Korth. Replication and consistency: being lazy helps sometimes. In ACM [2279], pages 173–184. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26

                                        1997. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        263661/p173-breitbart/p173-breitbart.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/263661/
                                        p173-breitbart/;   http://www.acm.
                                        org:80/pubs/citations/proceedings/
                                        pods/263661/p173-breitbart/.

Acharya:1997:BPP

[1120] Swarup Acharya, Michael Franklin, and Stanley Zdonik. Balancing push and pull for data broadcast. In Peckman [2278], pages 183–194. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p183-acharya/p183-acharya.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/253260/p183-acharya/

.

Shoshani:1997:OSD

[1121] Arie Shoshani. OLAP and statistical databases: similarities and differences. In ACM [2279], pages 185–196. ISBN 0-89791-910-6. LCCN QA 76.9 D3

                                        A26 1997. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/

263661/p185-shoshani/p185-shoshani.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/263661/
p185-shoshani/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/263661/p185-shoshani/.

Bayardo:1997:IAB

[1122] R. J. Bayardo, W. Bohrer, R. Brice, A. Cichocki, J. Fowler, A. Helal, V. Kashyap, T. Ksiezyk, G. Martin, M. Nodine, M. Rashid, M. Rusinkiewicz, R. Shea, C. Unnikrishnan, A. Unruh, and D. Woelk. InfoSleuth: agent-based semantic integration of information in open and dynamic environments. In Peckman [2278], pages 195–206. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p195-bayardo/p195-bayardo.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/253260/p195-bayardo/

.

Alonso:1997:CPC

[1123] Gustavo Alonso, Stephen Blott, Armin Fessler, and Hans-Jörg Schek. Correctness and parallelism in composite systems. In ACM [2279], pages 197–208. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26

1997. URL http://www.acm.org/
pubs/articles/proceedings/pods/
263661/p197-alonso/p197-alonso.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/263661/
p197-alonso/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/263661/p197-alonso/.

Gravano:1997:SSP

[1124] Luis Gravano, Chen-Chuan K. Chang, Héctor García-Molina, and Andreas Paepcke. STARTS: Stanford proposal for Internet meta-searching. In Peckman [2278], pages 207–218. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p207-gravano/p207-gravano.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/253260/p207-gravano/

.

Gunopulos:1997:DMH

[1125] Dimitrios Gunopulos, Heikki Mannila, Roni Khardon, and Hannu Toivonen. Data mining, hypergraph transversals, and machine learning (extended abstract). In ACM [2279], pages 209–216. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26

1997. URL http://www.acm.org/
pubs/articles/proceedings/pods/
263661/p209-gunopulos/p209-gunopulos.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/263661/
p209-gunopulos/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/263661/p209-gunopulos/.

Ludascher:1997:RAL

[1126] Bertram Ludäscher, Wolfgang May, and Georg Lausen. Referential actions as logical rules. In ACM [2279], pages 217–227. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26

1997. URL http://www.acm.org/
pubs/articles/proceedings/pods/
263661/p217-ludascher/p217-ludascher.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/263661/
p217-ludascher/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/263661/p217-ludascher/.

Carey:1997:SEA

[1127] Michael J. Carey and Donald Kossmann. On saying “Enough already!” in SQL. In Peckman [2278], pages 219–230. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p219-carey/p219-carey.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p219-carey/.

Ho:1997:PSQ

[1128] Ching-Tien Ho, Jehoshua Bruck, and Rakesh Agrawal. Partial-sum queries in OLAP data cubes using covering codes. In ACM [2279], pages 228–237. ISBN 0-89791-910-6. LCCN QA 76.9 D3

A26 1997. URL http://www.acm.org/
pubs/articles/proceedings/pods/
263661/p228-ho/p228-ho.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
pods/263661/p228-ho/;   http://
www.acm.org:80/pubs/citations/proceedings/
pods/263661/p228-ho/.

Griffin:1997:FIH

[1129] Timothy Griffin and Richard Hull.

A framework for implementing hypothetical queries. In Peckman [2278], pages 231–242. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p231-griffin/p231-griffin.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/253260/p231-griffin/

.

Scheufele:1997:CGO

[1130] Wolfgang Scheufele and Guido Moerkotte. On the complexity of generating optimal plans with cross products (extended abstract). In ACM [2279], pages 238–248. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26

1997. URL http://www.acm.org/
pubs/articles/proceedings/pods/
263661/p238-scheufele/p238-scheufele.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/263661/
p238-scheufele/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/263661/p238-scheufele/.

Arpaci-Dusseau:1997:HPS

[1131] Andrea C. Arpaci-Dusseau, Remzi H. Arpaci-Dusseau, David E. Culler, Joseph M. Hellerstein, and David A. Patterson. High-performance sorting on networks of workstations. In Peckman [2278], pages 243–254. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p243-arpaci-dusseau/p243-arpaci-
dusseau.pdf;   http://www.acm.org/
pubs/citations/proceedings/mod/
253260/p243-arpaci-dusseau/.

Hellerstein:1997:AIS

[1132] Joseph M. Hellerstein, Elias Koutsoupias, and Christos H. Papadimitriou. On the analysis of indexing schemes. In ACM [2279], pages 249–256. ISBN 0-89791-910-6. LCCN QA 76.9 D3

A26 1997. URL http://www.acm.org/
pubs/articles/proceedings/pods/
263661/p249-hellerstein/p249-hellerstein.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/263661/

p249-hellerstein/;   http://www.
acm.org:80/pubs/citations/proceedings/
pods/263661/p249-hellerstein/.

Brin:1997:DIC

[1133] Sergey Brin, Rajeev Motwani, Jeffrey D. Ullman, and Shalom Tsur. Dynamic itemset counting and implication rules for market basket data. In Peckman [2278], pages 255–264. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p255-brin/p255-brin.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p255-brin/.

Goldstein:1997:PQL

[1134] Jonathan Goldstein, Raghu Ramakrishnan, Uri Shaft, and Jie-Bing Yu. Processing queries by linear constraints. In ACM [2279], pages 257–267. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26

1997. URL http://www.acm.org/
pubs/articles/proceedings/pods/
263661/p257-goldstein/p257-goldstein.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/263661/
p257-goldstein/;   http://www.acm.
org:80/pubs/citations/proceedings/
pods/263661/p257-goldstein/.

Brin:1997:BMB

[1135] Sergey Brin, Rajeev Motwani, and Craig Silverstein. Beyond market baskets: generalizing association rules to correlations. In Peckman [2278], pages 265–276. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/

                                        pubs/articles/proceedings/mod/253260/
                                        p265-brin/p265-brin.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/253260/p265-brin/.

Han:1997:SPD

[1136] Eui-Hong Han, George Karypis, and Vipin Kumar. Scalable parallel data mining for association rules. In Peckman [2278], pages 277–288. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p277-han/p277-han.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/253260/p277-han/.

Korn:1997:ESA

[1137] Flip Korn, H. V. Jagadish, and Christos Faloutsos. Efficiently supporting ad hoc queries in large datasets of time sequences. In Peckman [2278], pages 289–300. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p289-korn/p289-korn.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/253260/p289-korn/.

Livny:1997:DIQ

[1138] M. Livny, R. Ramakrishnan, K. Beyer, G. Chen, D. Donjerkovic, S. Lawande, J. Myllymaki, and K. Wenger. DE-Vise: integrated querying and visual exploration of large datasets. In Peckman [2278], pages 301–312. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/

p301-livny/p301-livny.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p301-livny/.

Maheshwari:1997:PGC

[1139] Umesh Maheshwari and Barbara Liskov. Partitioned garbage collection of a large object store. In Peckman [2278], pages 313–323. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p313-maheshwari/p313-maheshwari.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/253260/
p313-maheshwari/.

Koudas:1997:SSS

[1140] Nick Koudas and Kenneth C. Sevcik.

Size separation spatial join. In Peckman [2278], pages 324–335. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p324-koudas/p324-koudas.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p324-koudas/.

Patel:1997:BSG

[1141] Jignesh Patel, JieBing Yu, Navin Kabra, Kristin Tufte, Biswadeep Nag, Josef Burger, Nancy Hall, Karthikeyan Ramasamy, Roger Lueder, Curt Ellmann, Jim Kupsch, Shelly Guo, Johan Larson, David De Witt, and Jeffrey Naughton. Building a scaleable geo-spatial DBMS: technology, implementation, and evaluation. In Peckman [2278], pages 336–347. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p336-patel/p336-patel.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/253260/p336-patel/.

Gebhardt:1997:TNS

[1142] Michael Gebhardt, Matthias Jarke, and Stephan Jacobs. A toolkit for negotiation support interfaces to multi-dimensional data. In Peckman [2278], pages 348–356. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p348-gebhardt/p348-gebhardt.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/253260/p348-gebhardt/

.

Bozkaya:1997:DBI

[1143] Tolga Bozkaya and Meral Ozsoyoglu. Distance-based indexing for highdimensional metric spaces. In Peckman [2278], pages 357–368. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p357-bozkaya/p357-bozkaya.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/253260/p357-bozkaya/

.

Katayama:1997:STI

[1144] Norio Katayama and Shin’ichi Satoh. The SR-tree: an index structure for high-dimensional nearest neighbor queries. In Peckman [2278], pages 369–380. ISBN 0-89791-911-

4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p369-katayama/p369-katayama.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/253260/p369-katayama/

.

Shivakumar:1997:WII

[1145] Narayanan Shivakumar and Héctor García-Molina. Wave-indices: indexing evolving databases. In Peckman [2278], pages 381–392. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p381-shivakumar/p381-shivakumar.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/253260/
p381-shivakumar/.

Quass:1997:LWV

[1146] Dallan Quass and Jennifer Widom. Online warehouse view maintenance. In Peckman [2278], pages 393–404. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p393-quass/p393-quass.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p393-quass/.

Colby:1997:SMV

[1147] Latha S. Colby, Akira Kawaguchi, Daniel F. Lieuwen, Inderpal Singh Mumick, and Kenneth A. Ross. Supporting multiple view maintenance policies. In Peckman [2278], pages 405–416. ISBN 0-89791-911-

4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/

pubs/articles/proceedings/mod/253260/
p405-colby/p405-colby.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p405-colby/.

Agrawal:1997:EVM

[1148] D. Agrawal, A. El Abbadi, A. Singh, and T. Yurek. Efficient view maintenance at data warehouses. In Peckman [2278], pages 417–427. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p417-agrawal/p417-agrawal.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/253260/p417-agrawal/

.

Llirbat:1997:ECR

[1149] François Llirbat, Françoise Fabret, and Eric Simon. Eliminating costly redundant computations from SQL trigger executions. In Peckman [2278], pages 428–439. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p428-llirbat/p428-llirbat.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/253260/p428-llirbat/

.

Motakis:1997:TAA

[1150] Iakovos Motakis and Carlo Zaniolo. Temporal aggregation in active database rules. In Peckman [2278], pages 440–451. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p440-motakis/p440-motakis.pdf;

http://www.acm.org/pubs/citations/
proceedings/mod/253260/p440-motakis/

.

Miller:1997:ARI

[1151] R. J. Miller and Y. Yang. Association rules over interval data. In Peckman [2278], pages 452–461. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p452-miller/p452-miller.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p452-miller/.

George:1997:STP

[1152] Binto George and Jayant Haritsa. Secure transaction processing in firm real-time database systems. In Peckman [2278], pages 462–473. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p462-george/p462-george.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p462-george/.

Jajodia:1997:UFE

[1153] Sushil Jajodia, Pierangela Samarati, V. S. Subrahmanian, and Eliza Bertino. A unified framework for enforcing multiple access control policies. In Peckman [2278], pages 474–485. ISBN 0-89791-911-

4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p474-jajodia/p474-jajodia.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/253260/p474-jajodia/

.

Gupta:1997:RCP

[1154] Ramesh Gupta, Jayant Haritsa, and Krithi Ramamritham. Revisiting commit processing in distributed database systems. In Peckman [2278], pages 486–497. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p486-gupta/p486-gupta.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/253260/p486-gupta/.

Shasha:1997:LWS

[1155] Dennis Shasha. Lessons from Wall Street: case studies in configuration, tuning, and distribution. In Peckman [2278], pages 498–501. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p498-shasha/p498-shasha.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/253260/p498-shasha/.

Carey:1997:ORD

[1156] Michael J. Carey, Nelson M. Mattos, and Anil K. Nori. Object-relational database systems (tutorial): principles, products and challenges. In Peckman [2278], page 502. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p502-carey/p502-carey.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/253260/p502-carey/.

Kramer:1997:DWT

[1157] Ralf Kramer. Databases on the Web:

technologies for federation architectures and case studies. In Peckman [2278], pages 503–506. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p503-kramer/p503-kramer.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p503-kramer/.

Chaudhuri:1997:DWO

[1158] Surajit Chaudhuri and Umeshwar Dayal. Data warehousing and OLAP for decision support. In Peckman [2278], pages 507–508. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p507-chaudhuri/p507-chaudhuri.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/253260/p507-chaudhuri/

.

Chaudhuri:1997:QOC

[1159] Surajit Chaudhuri. Query optimization at the crossroads. In Peckman [2278], page 509. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p509-chaudhuri/p509-chaudhuri.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/253260/p509-chaudhuri/

.

Cruz:1997:DDV

[1160] Isabel F. Cruz, M. Averbuch, Wendy T.

Lucas, Melissa Radzyminski, and Kirby Zhang. Delaunay: a database visualization system. In Peckman [2278], pages 510–513. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p510-cruz/p510-cruz.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p510-cruz/.

Goyal:1997:PPP

[1161] Nita Goyal, Charles Hoch, Ravi Krishnamurthy, Brian Meckler, Michael Suchow, and Moshe Zloof. Picture programming project. In Peckman [2278], pages 514–516. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p514-goyal/p514-goyal.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p514-goyal/.

Livny:1997:DDA

[1162] M. Livny, R. Ramakrishnan, K. Beyer, G. Chen, D. Donjerkovic, S. Lawande, J. Myllymaki, and K. Wenger. DE-Vise (demo abstract): integrated querying and visual exploration of large datasets. In Peckman [2278], pages 517–520. ISBN 0-89791-911-

4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p517-livny/p517-livny.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p517-livny/.

Li:1997:SOB

[1163] Wen-Syan Li, K. Selçuk Candan, Kyoji Hirata, and Yoshinori Hara. SEMCOG: an object-based image retrieval system and its visual query interface. In Peckman [2278], pages 521–524. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p521-li/p521-li.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/253260/p521-li/.

Bressan:1997:CIM

[1164] S. Bressan, C. H. Goh, K. Fynn, M. Jakobisiak, K. Hussein, H. Kon, T. Lee, S. Madnick, T. Pena, J. Qu, A. Shum, and M. Siegel. The Context Interchange mediator prototype. In Peckman [2278], pages 525–527. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p525-bressan/p525-bressan.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/253260/p525-bressan/

.

Atzeni:1997:MMD

[1165] Paolo Atzeni and Riccardo Torlone.

MDM: a multiple-data model tool for the management of heterogeneous database schemes. In Peckman [2278], pages 528–531. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p528-atzeni/p528-atzeni.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p528-atzeni/.

Hammer:1997:TBW

[1166] Joachim Hammer, Héctor García-Molina, Svetlozar Nestorov, Ramana Yerneni, Marcus Breunig, and Vasilis Vassalos. Template-based wrappers in the TSIMMIS system. In Peckman [2278], pages 532–535. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p532-hammer/p532-hammer.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p532-hammer/.

Gingras:1997:LMD

[1167] Frédéric Gingras, Laks V. S. Lakshmanan, Iyer N. Subramanian, Despina Papoulis, and Nematollaah Shiri. Languages for multi-database interoperability. In Peckman [2278], pages 536–538. ISBN 0-89791-911-

4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p536-gingras/p536-gingras.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/253260/p536-gingras/

.

Genesereth:1997:III

[1168] Michael R. Genesereth, Arthur M. Keller, and Oliver M. Duschka. Infomaster: an information integration system. In Peckman [2278], pages 539–542. ISBN 0-89791-911-

4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p539-genesereth/p539-genesereth.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/253260/
p539-genesereth/.

Bayardo:1997:IP

[1169] R. J. Bayardo, W. Bohrer, R. Brice, A. Cichocki, J. Fowler, A. Halal, V. Kashyap, T. Ksiezyk, G. Martin, M. Nodine, M. Rashid, M. Rusinkiewicz, R. Shea, C. Unnikrishnan, A. Unruh, and D. Woelk. The InfoSleuth Project. In Peckman [2278], pages 543–545. ISBN 0-89791-911-

4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p543-bayardo/p543-bayardo.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/253260/p543-bayardo/

.

Tomasic:1997:DIS

[1170] Anthony Tomasic, Rémy Amouroux, Philippe Bonnet, Olga Kapitskaia, Hubert Naacke, and Louiqa Raschid. The distributed information search component (Disco) and the World Wide Web. In Peckman [2278], pages 546–548. ISBN 0-89791-911-

4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p546-tomasic/p546-tomasic.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/253260/p546-tomasic/

.

Fernandez:1997:SWS

[1171] Mary Fernandez, Daniela Florescu, Jaewoo Kang, Alon Levy, and Dan Suciu. STRUDEL: a Web site management system. In Peckman [2278], pages 549–552. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p549-fernandez/p549-fernandez.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/253260/p549-fernandez/

.

Han:1997:GSP

[1172] Jaiwei Han, Krzysztof Koperski, and Nebojsa Stefanovic. GeoMiner: a system prototype for spatial data mining. In Peckman [2278], pages 553–556. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p553-han/p553-han.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/253260/p553-han/.

Labio:1997:WPD

[1173] Wilburt J. Labio, Yue Zhuge, Janet L. Wiener, Himanshu Gupta, Héctor García-Molina, and Jennifer Widom. The WHIPS prototype for data warehouse creation and maintenance. In Peckman [2278], pages 557–559. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/253260/
                                        p557-labio/p557-labio.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/253260/p557-labio/.

Wang:1997:SMD

[1174] Jason Tsong-Li Wang, Dennis Shasha, George J. S. Chang, Liam Relihan, Kaizhong Zhang, and Girish Patel. Structural matching and discovery in document databases. In Peckman [2278], pages 560–563. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p560-wang/p560-wang.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p560-wang/.

Berchtold:1997:SSS

[1175] Stefan Berchtold and Hans-Peter Kriegel. S3: similarity search in CAD database systems. In Peckman [2278], pages 564–567. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p564-berchtold/p564-berchtold.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/253260/p564-berchtold/

.

Seshadri:1997:PDE

[1176] Praveen Seshadri and Mark Paskin.

PREDATOR: an OR-DBMS with enhanced data types. In Peckman [2278], pages 568–571. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p568-seshadri/p568-seshadri.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/253260/p568-seshadri/

.

Chakravarthy:1997:SOO

[1177] S. Chakravarthy. Sentinel: an object-oriented DBMS with eventbased rules. In Peckman [2278], pages 572–575. ISBN 0-89791-911-

4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/

p572-chakravarthy/p572-chakravarthy.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/253260/
p572-chakravarthy/.

Wodtke:1997:MWE

[1178] Dirk Wodtke, Jeanine Weissenfels, Gerhard Weikum, Angelika Kotz Dittrich, and Peter Muth. The MEN-TOR workbench for enterprise-wide workflow management. In Peckman [2278], pages 576–579. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p576-wodtke/p576-wodtke.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/253260/p576-wodtke/.

Ioannidis:1997:ZDE

[1179] Yannis E. Ioannidis, Miron Livny, Anastassia Ailamaki, Anand Narayanan, and Andrew Therber. Zoo: a desktop experiment management environment. In Peckman [2278], pages 580–583. ISBN 0-89791-911-

4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/253260/
p580-ioannidis/p580-ioannidis.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/253260/p580-ioannidis/

.

Fagin:1998:FQM

[1180] Ronald Fagin. Fuzzy queries in multimedia database systems. In ACM [2281], pages 1–10. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

1998. URL http://www.acm.org/
pubs/articles/proceedings/pods/
275487/p1-fagin/p1-fagin.pdf;

http://www.acm.org/pubs/citations/
proceedings/pods/275487/p1-fagin/

. ACM order number 475980.

Tsur:1998:QFG

[1181] Dick Tsur, Jeffrey D. Ullman, Serge Abiteboul, Chris Clifton, Rajeev Motwani, Svetlozar Nestorov, and Arnon Rosenthal. Query flocks: a generalization of association-rule mining. In Haas and Tiwary [2280], pages 1–12. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p1-tsur/p1-tsur.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/276304/p1-tsur/.

Neven:1998:ESD

[1182] Frank Neven and Jan Van den Bussche. Expressiveness of structured document query languages based on attribute grammars. In ACM [2281], pages 11–17. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

1998. URL http://www.acm.org/
pubs/articles/proceedings/pods/
275487/p11-neven/p11-neven.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/275487/p11-neven/

. ACM order number 475980.

Ng:1998:EMP

[1183] Raymond T. Ng, Laks V. S. Lakshmanan, Jiawei Han, and Alex Pang. Exploratory mining and pruning optimizations of constrained associations rules. In Haas and Tiwary [2280], pages 13–24. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/

                                        p13-ng/p13-ng.pdf;   http://www.
                                        acm.org/pubs/citations/proceedings/
                                        mod/276304/p13-ng/.

Aggarwal:1998:NFI

[1184] Charu C. Aggarwal and Philip S. Yu. A new framework for itemset generation. In ACM [2281], pages 18–24. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

                                        1998. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        275487/p18-aggarwal/p18-aggarwal.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/275487/

p18-aggarwal/. ACM order number 475980.

Davulcu:1998:LBM

[1185] Hasan Davulcu, Michael Kifer, C. R. Ramakrishnan, and I. V. Ramakrishnan. Logic based modeling and analysis of workflows. In ACM [2281], pages 25–33. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

                                        1998. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        275487/p25-davulcu/p25-davulcu.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/275487/

p25-davulcu/. ACM order number 475980.

Shintani:1998:PMA

[1186] Takahiko Shintani and Masaru Kitsuregawa. Parallel mining algorithms for generalized association rules with classification hierarchy. In Haas and Tiwary [2280], pages 25–36. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p25-shintani/p25-shintani.pdf;

http://www.acm.org/pubs/citations/
proceedings/mod/276304/p25-shintani/

.

Chaudhuri:1998:OQO

[1187] Surajit Chaudhuri. An overview of query optimization in relational systems. In ACM [2281], pages 34–

43. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296 1998. URL http://

www.acm.org/pubs/articles/proceedings/
pods/275487/p34-chaudhuri/p34-chaudhuri.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/275487/

p34-chaudhuri/. ACM order number 475980.

Rao:1998:RIN

[1188] Jun Rao and Kenneth A. Ross. Reusing invariants: a new strategy for correlated queries. In Haas and Tiwary [2280], pages 37–48. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p37-rao/p37-rao.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/276304/p37-rao/.

Samoladas:1998:LBT

[1189] Vasilis Samoladas and Daniel P. Miranker. A lower bound theorem for indexing schemes and its application to multidimensional range queries. In ACM [2281], pages 44–51. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

1998. URL http://www.acm.org/
pubs/articles/proceedings/pods/
275487/p44-samoladas/p44-samoladas.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/275487/

p44-samoladas/. ACM order number 475980.

Fegaras:1998:QUO

[1190] Leonidas Fegaras. Query unnesting in object-oriented databases. In Haas and Tiwary [2280], pages 49–60. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p49-fegaras/p49-fegaras.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/276304/p49-fegaras/.

Koutsoupias:1998:TBD

[1191] Elias Koutsoupias and D. Scot Taylor. Tight bounds for 2dimensional indexing schemes. In ACM [2281], pages 52–58. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

                                        1998. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        275487/p52-koutsoupias/p52-koutsoupias.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/275487/
                                        p52-koutsoupias/. ACM order num-

ber 475980.

Ciaccia:1998:CMS

[1192] Paolo Ciaccia, Marco Patella, and Pavel Zezula. A cost model for similarity queries in metric spaces. In ACM [2281], pages 59–68. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

                                        1998. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        275487/p59-ciaccia/p59-ciaccia.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/275487/

p59-ciaccia/. ACM order number

475980.

Cherniack:1998:CRT

[1193] Mitch Cherniack and Stan Zdonik. Changing the rules: transformations for rule-based optimizers. In Haas and Tiwary [2280], pages 61–72. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p61-cherniack/p61-cherniack.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/276304/p61-cherniack/

.

Christodoulakis:1998:DBD

[1194] Stavros Christodoulakis and Fenia A.

Zioga. Data base design principles for striping and placement of delay-sensitive data on disks. In ACM [2281], pages 69–78. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

1998. URL http://www.acm.org/
pubs/articles/proceedings/pods/
275487/p69-christodoulakis/p69-
christodoulakis.pdf;   http://www.
acm.org/pubs/citations/proceedings/
pods/275487/p69-christodoulakis/

. ACM order number 475980.

Guha:1998:CEC

[1195] Sudipto Guha, Rajeev Rastogi, and Kyuseok Shim. CURE: an efficient clustering algorithm for large databases. In Haas and Tiwary [2280], pages 73–84. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p73-guha/p73-guha.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/276304/p73-guha/.

Garofalakis:1998:TCA

[1196] Minos N. Garofalakis, Yannis E. Ioannidis, Banu Özden, and Avi Silberschatz. Throughput-competitive admission control for continuous media databases. In ACM [2281], pages 79–88. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296 1998. URL http://

www.acm.org/pubs/articles/proceedings/
pods/275487/p79-garofalakis/p79-
garofalakis.pdf;   http://www.acm.
org/pubs/citations/proceedings/
pods/275487/p79-garofalakis/. ACM

order number 475980.

Bayardo:1998:EML

[1197] Roberto J. Bayardo. Efficiently mining long patterns from databases. In Haas and Tiwary [2280], pages 85–93. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p85-bayardo/p85-bayardo.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p85-bayardo/.

Segoufin:1998:QSD

[1198] Luc Segoufin and Victor Vianu. Querying spatial databases via topological invariants. In ACM [2281], pages 89–98. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296 1998. URL http://

www.acm.org/pubs/articles/proceedings/
pods/275487/p89-segoufin/p89-segoufin.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/275487/

p89-segoufin/. ACM order number 475980.

Agrawal:1998:ASC

[1199] Rakesh Agrawal, Johannes Gehrke, Dimitrios Gunopulos, and Prabhakar Raghavan. Automatic subspace clustering of high dimensional data for data mining applications. In Haas and Tiwary [2280], pages 94–105. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p94-agrawal/p94-agrawal.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p94-agrawal/.

Benedikt:1998:SCQ

[1200] Michael Benedikt and Leonid Libkin.

Safe constraint queries. In ACM [2281], pages 99–108. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

1998. URL http://www.acm.org/
pubs/articles/proceedings/pods/
275487/p99-benedikt/p99-benedikt.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/275487/

p99-benedikt/. ACM order number 475980.

Kabra:1998:EMQ

[1201] Navin Kabra and David J. De-Witt. Efficient mid-query reoptimization of sub-optimal query execution plans. In Haas and Tiwary [2280], pages 106–117. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p106-kabra/p106-kabra.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p106-kabra/.

Vandeurzen:1998:ELL

[1202] Luc Vandeurzen, Marc Gyssens, and Dirk Van Gucht. An expressive language for linear spatial database queries. In ACM [2281], pages 109–118. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

1998. URL http://www.acm.org/
pubs/articles/proceedings/pods/

275487/p109-vandeurzen/p109-vandeurzen.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/275487/
p109-vandeurzen/. ACM order num-

ber 475980.

Jonsson:1998:IQE

[1203] Björn T. Jónsson, Michael J. Franklin, and Divesh Srivastava. Interaction of query evaluation and buffer management for information retrieval. In Haas and Tiwary [2280], pages 118–129. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p118-jonsson/p118-jonsson.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/276304/p118-jonsson/

.

Vitter:1998:EMA

[1204] Jeffrey Scott Vitter. External memory algorithms. In ACM [2281], pages 119–128. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

1998. URL http://www.acm.org/
pubs/articles/proceedings/pods/
275487/p119-vitter/p119-vitter.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/275487/

p119-vitter/. ACM order number 475980.

Buneman:1998:PCS

[1205] Peter Buneman, Wenfei Fan, and Scott Weinstein. Path constraints on semistructured and structured data. In ACM [2281], pages 129–138. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

1998. URL http://www.acm.org/
pubs/articles/proceedings/pods/
275487/p129-buneman/p129-buneman.
pdf;   http://www.acm.org/pubs/

citations/proceedings/pods/275487/

p129-buneman/. ACM order number 475980.

Urhan:1998:CBQ

[1206] Tolga Urhan, Michael J. Franklin, and Laurent Amsaleg. Cost-based query scrambling for initial delays. In Haas and Tiwary [2280], pages 130–141. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p130-urhan/p130-urhan.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p130-urhan/.

Florescu:1998:QCC

[1207] Daniela Florescu, Alon Levy, and Dan Suciu. Query containment for conjunctive queries with regular expressions. In ACM [2281], pages 139–148. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

1998. URL http://www.acm.org/
pubs/articles/proceedings/pods/
275487/p139-florescu/p139-florescu.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/275487/

p139-florescu/. ACM order number 475980.

Berchtold:1998:PTT

[1208] Stefan Berchtold, Christian Böhm, and Hans-Peter Kriegal. The pyramidtechnique: towards breaking the curse of dimensionality. In Haas and Tiwary [2280], pages 142–153. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p142-berchtold/p142-berchtold.pdf;
http://www.acm.org/pubs/citations/

proceedings/mod/276304/p142-berchtold/

.

Calvanese:1998:DQC

[1209] Diego Calvanese, Giuseppe De Giacomo, and Maurizio Lenzerini. On the decidability of query containment under constraints. In ACM [2281], pages 149–158. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

1998. URL http://www.acm.org/
pubs/articles/proceedings/pods/
275487/p149-calvanese/p149-calvanese.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/275487/
p149-calvanese/. ACM order number

475980.

Seidl:1998:OMS

[1210] Thomas Seidl and Hans-Peter Kriegel. Optimal multi-step k-nearest neighbor search. In Haas and Tiwary [2280], pages 154–165. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p154-seidl/p154-seidl.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p154-seidl/.

Papadimitriou:1998:LSI

[1211] Christos H. Papadimitriou, Hisao Tamaki, Prabhakar Raghavan, and Santosh Vempala. Latent semantic indexing: a probabilistic analysis. In ACM [2281], pages 159–168. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

1998. URL http://www.acm.org/
pubs/articles/proceedings/pods/
275487/p159-papadimitriou/p159-
papadimitriou.pdf;   http://www.
acm.org/pubs/citations/proceedings/

pods/275487/p159-papadimitriou/.

ACM order number 475980.

Kanth:1998:DRS

[1212] K. V. Ravi Kanth, Divyakant Agrawal, and Ambuj Singh. Dimensionality reduction for similarity searching in dynamic databases. In Haas and Tiwary [2280], pages 166–176. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p166-ravi_kanth/p166-ravi_kanth.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/276304/
p166-ravi_kanth/.

Agarwal:1998:ESL

[1213] Pankaj K. Agarwal, Lars Arge, Jeff Erickson, Paolo G. Franciosa, and Jeffry Scott Vitter. Efficient searching with linear constraints. In ACM [2281], pages 169–178. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

1998. URL http://www.acm.org/
pubs/articles/proceedings/pods/
275487/p169-agarwal/p169-agarwal.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/275487/

p169-agarwal/. ACM order number 475980.

Cluet:1998:YMN

[1214] Sophie Cluet, Claude Delobel, Jérôme Siméon, and Katarzyna Smaga. Your mediators need data conversion! In Haas and Tiwary [2280], pages 177–188. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p177-cluet/p177-cluet.pdf;   http:

                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/276304/p177-cluet/.

Abiteboul:1998:RTE

[1215] Serge Abiteboul, Victor Vianu, Brad Fordham, and Yelena Yesha. Relational transducers for electronic commerce. In ACM [2281], pages 179–187. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

                                        1998. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        275487/p179-abiteboul/p179-abiteboul.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/275487/
                                        p179-abiteboul/. ACM order number

475980.

Samarati:1998:GDP

[1216] Pierangela Samarati and Latanya Sweeney. Generalizing data to provide anonymity when disclosing information (abstract). In ACM [2281], page 188. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

                                        1998. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        275487/p188-samarati/p188-samarati.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/275487/

p188-samarati/. ACM order number

475980.

Neven:1998:TQL

[1217] Frank Neven, Dirk Van Gucht, Jan Van den Bussche, and Gottfried Vossen. Typed query languages for databases containing queries. In ACM [2281], pages 189–196. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

                                        1998. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        275487/p189-neven/p189-neven.pdf;
                                         http://www.acm.org/pubs/citations/

proceedings/pods/275487/p189-neven/

. ACM order number 475980.

Miller:1998:USH

[1218] Reée J. Miller. Using schematically heterogeneous structures. In Haas and Tiwary [2280], pages 189–200. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p189-miller/p189-miller.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p189-miller/.

Kolaitis:1998:CCP

[1219] Phokion G. Kolaitis, David L. Martin, and Madhukar N. Thakur. On the complexity of the containment problem for conjunctive queries with built-in predicates. In ACM [2281], pages 197–204. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

1998. URL http://www.acm.org/
pubs/articles/proceedings/pods/
275487/p197-kolaitis/p197-kolaitis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/275487/

p197-kolaitis/. ACM order number 475980.

Cohen:1998:IHD

[1220] William W. Cohen. Integration of heterogeneous databases without common domains using queries based on textual similarity. In Haas and Tiwary [2280], pages 201–212. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p201-cohen/p201-cohen.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p201-cohen/.

Kolaitis:1998:CQC

[1221] Phokion G. Kolaitis and Moshe Y. Vardi. Conjunctive-query containment and constraint satisfaction. In ACM [2281], pages 205–213. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

                                        1998. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        275487/p205-kolaitis/p205-kolaitis.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/275487/

p205-kolaitis/. ACM order number 475980.

Grumbach:1998:DSC

[1222] Stéphane Grumbach, Philippe Rigaux, and Luc Segoufin. The DEDALE system for complex spatial queries. In Haas and Tiwary [2280], pages 213– 224. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        QA1 .A87. URL http://www.acm.
                                        org/pubs/articles/proceedings/mod/
                                        276304/p213-grumbach/p213-grumbach.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/276304/
                                        p213-grumbach/.

Nutt:1998:DEA

[1223] Werner Nutt, Yehoshus Sagiv, and Sara Shurin. Deciding equivalences among aggregate queries. In ACM [2281], pages 214–223. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

                                        1998. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        275487/p214-nutt/p214-nutt.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/275487/p214-nutt/

. ACM order number 475980.

Klavans:1998:DBD

[1224] Judith Klavans. Data bases in digital libraries: Where computer science and information management meet. In ACM [2281], pages 224–226. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

1998. URL http://www.acm.org/
pubs/articles/proceedings/pods/
275487/p224-klavans/p224-klavans.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/275487/

p224-klavans/. ACM order number 475980.

Papadopoulos:1998:SQP

[1225] Apostolos N. Papadopoulos and Yannis Manolopoulos. Similarity query processing using disk arrays. In Haas and Tiwary [2280], pages 225–236. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p225-papadopoulos/p225-papadopoulos.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/276304/
p225-papadopoulos/.

Gradel:1998:CQR

[1226] Erich Grädel, Yuri Gurevich, and Colin Hirsch. The complexity of query reliability. In ACM [2281], pages 227–234. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

1998. URL http://www.acm.org/
pubs/articles/proceedings/pods/
275487/p227-gradel/p227-gradel.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/275487/

p227-gradel/. ACM order number 475980.

Graedel:1998:CQR

[1227] E. Graedel, Y. Gurevich, and C. Hirsch. The complexity of query reliability. In ACM [2281], pages 227–

234. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296 1998. ACM order number 475980.

Etessami:1998:DTI

[1228] Kousha Etessami. Dynamic tree isomorphism via first-order updates to a relational database. In ACM [2281], pages 235–243. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

                                        1998. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        275487/p235-etessami/p235-etessami.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/275487/

p235-etessami/. ACM order number

475980.

Hjaltason:1998:IDJ

[1229] Gísli R. Hjaltason and Hanan Samet. Incremental distance join algorithms for spatial databases. In Haas and Tiwary [2280], pages 237–248. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p237-hjaltason/p237-hjaltason.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/276304/p237-hjaltason/

.

Vorobyov:1998:CNL

[1230] Sergei Vorobyov and Andrie Voronkov. Complexity of nonrecursive logic programs with complex values. In ACM [2281], pages 244–253. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

                                        1998. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        275487/p244-vorobyov/p244-vorobyov.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/275487/

p244-vorobyov/. ACM order number 475980.

Kotidis:1998:ASO

[1231] Yannis Kotidis and Nick Roussopoulos. An alternative storage organization for ROLAP aggregate views based on cubetrees. In Haas and Tiwary [2280], pages 249–258. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p249-kotidis/p249-kotidis.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/276304/p249-kotidis/

.

Abiteboul:1998:CAQ

[1232] Serge Abiteboul and Oliver M.

Duschka. Complexity of answering queries using materialized views. In ACM [2281], pages 254–263. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

1998. URL http://www.acm.org/
pubs/articles/proceedings/pods/
275487/p254-abiteboul/p254-abiteboul.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/275487/
p254-abiteboul/. ACM order number

475980.

Deshpande:1998:CMQ

[1233] Prasad M. Deshpande, Karthikeyan Ramasamy, Amit Shukla, and Jeffrey F. Naughton. Caching multidimensional queries using chunks. In Haas and Tiwary [2280], pages 259–270. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p259-deshpande/p259-deshpande.pdf;
 http://www.acm.org/pubs/citations/

                                        proceedings/mod/276304/p259-deshpande/

.

Bailey:1998:DUR

[1234] James Bailey, Guozhu Dong, and Kotagiri Ramamohanarao. Decidability and undecidability results for the termination problem of active database rules. In ACM [2281], pages 264–273. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

                                        1998. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        275487/p264-bailey/p264-bailey.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/275487/

p264-bailey/. ACM order number

475980.

Zhao:1998:SOE

[1235] Yihong Zhao, Prasad M. Deshpande, Jeffrey F. Naughton, and Amit Shukla. Simultaneous optimization and evaluation of multiple dimensional queries. In Haas and Tiwary [2280], pages 271–282. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p271-zhao/p271-zhao.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/276304/p271-zhao/.

Smith:1998:DAV

[1236] John R. Smith, Vittorio Castelli, Anant Jhingran, and Chung-Sheng Li. Dynamic assembly of views in data cubes. In ACM [2281], pages 274–283. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296

                                        1998. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        275487/p274-smith/p274-smith.pdf;
                                         http://www.acm.org/pubs/citations/

proceedings/pods/275487/p274-smith/

. ACM order number 475980.

Adelberg:1998:NTS

[1237] Brad Adelberg. NoDoSE–a tool for semi-automatically extracting structured and semistructured data from text documents. In Haas and Tiwary [2280], pages 283–294. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p283-adelberg/p283-adelberg.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/276304/p283-adelberg/

.

Nestorov:1998:ESS

[1238] Svetlozar Nestorov, Serge Abiteboul, and Rajeev Motwani. Extracting schema from semistructured data. In Haas and Tiwary [2280], pages 295–306. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p295-nestorov/p295-nestorov.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/276304/p295-nestorov/

.

Chakrabarti:1998:EHC

[1239] Soumen Chakrabarti, Byron Dom, and Piotr Indyk. Enhanced hypertext categorization using hyperlinks. In Haas and Tiwary [2280], pages 307–318. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p307-chakrabarti/p307-chakrabarti.
pdf;   http://www.acm.org/pubs/

citations/proceedings/mod/276304/
p307-chakrabarti/.

Subramanian:1998:CBO

[1240] Subbu N. Subramanian and Shivakumar Venkataraman. Cost-based optimization of decision support queries using transient-views. In Haas and Tiwary [2280], pages 319–330. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p319-subramanian/p319-subramanian.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/276304/
p319-subramanian/.

Gibbons:1998:NSB

[1241] Phillip B. Gibbons and Yossi Matias. New sampling-based summary statistics for improving approximate query answers. In Haas and Tiwary [2280], pages 331–342. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p331-gibbons/p331-gibbons.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/276304/p331-gibbons/

.

Sarawagi:1998:IAR

[1242] Sunita Sarawagi, Shiby Thomas, and Rakesh Agrawal. Integrating association rule mining with relational database systems: alternatives and implications. In Haas and Tiwary [2280], pages 343–354. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/

pubs/articles/proceedings/mod/276304/
p343-sarawagi/p343-sarawagi.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/276304/p343-sarawagi/

.

Chan:1998:BID

[1243] Chee-Yong Chan and Yannis E.

Ioannidis. Bitmap index design and evaluation. In Haas and Tiwary [2280], pages 355–366. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p355-chan/p355-chan.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p355-chan/.

Chaudhuri:1998:AWI

[1244] Surajit Chaudhuri and Vivek Narasayya.

AutoAdmin “what-if” index analysis utility. In Haas and Tiwary [2280], pages 367–378. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p367-chaudhuri/p367-chaudhuri.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/276304/p367-chaudhuri/

.

Jaedicke:1998:PPA

[1245] Michael Jaedicke and Bernhard Mitschang.

On parallel processing of aggregate and scalar functions in objectrelational DBMS. In Haas and Tiwary [2280], pages 379–389. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p379-jaedicke/p379-jaedicke.pdf;

 http://www.acm.org/pubs/citations/
proceedings/mod/276304/p379-jaedicke/

.

Godfrey:1998:SPD

[1246] Michael Godfrey, Tobias Mayr, Praveen Seshadri, and Thorsten von Eicken. Secure and portable database extensibility. In Haas and Tiwary [2280], pages 390–401. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p390-godfrey/p390-godfrey.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/276304/p390-godfrey/

.

Adali:1998:MSA

[1247] S. Adali, P. Bonatti, M. L. Sapino, and V. S. Subrahmanian. A multisimilarity algebra. In Haas and Tiwary [2280], pages 402–413. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p402-adali/p402-adali.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p402-adali/.

Fernandez:1998:CBS

[1248] Mary Fernández, Daniela Florescu, Jaewoo Kang, Alon Levy, and Dan Suciu. Catching the boat with Strudel: experiences with a Website management system. In Haas and Tiwary [2280], pages 414–425. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p414-fernandez/p414-fernandez.pdf;

 http://www.acm.org/pubs/citations/
proceedings/mod/276304/p414-fernandez/

.

Manku:1998:AMO

[1249] Gurmeet Singh Manku, Sridhar Rajagopalan, and Bruce G. Lindsay. Approximate medians and other quantiles in one pass and with limited memory. In Haas and Tiwary [2280], pages 426–435. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p426-manku/p426-manku.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p426-manku/.

Chaudhuri:1998:RSH

[1250] Surajit Chaudhuri, Rajeev Motwani, and Vivek Narasayya. Random sampling for histogram construction: how much is enough? In Haas and Tiwary [2280], pages 436–447. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p436-chaudhuri/p436-chaudhuri.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/276304/p436-chaudhuri/

.

Matias:1998:WBH

[1251] Yossi Matias, Jeffrey Scott Vitter, and Min Wang. Wavelet-based histograms for selectivity estimation. In Haas and Tiwary [2280], pages 448–459. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p448-matias/p448-matias.pdf;   http:

                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/276304/p448-matias/.

Lomet:1998:ETA

[1252] David Lomet and Gerhard Weikum. Efficient transparent application recovery in client-server information systems. In Haas and Tiwary [2280], pages 460–471. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p460-lomet/p460-lomet.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/276304/p460-lomet/.

Larson:1998:MMD

[1253] Per-Åke Larson and Goetz Graefe. Memory management during run generation in external sorting. In Haas and Tiwary [2280], pages 472–483. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p472-larson/p472-larson.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/276304/p472-larson/.

Anderson:1998:RCP

[1254] Todd Anderson, Yuri Breitbart, Henry F. Korth, and Avishai Wool. Replication, consistency, and practicality: are these mutually exclusive? In Haas and Tiwary [2280], pages 484–495. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p484-anderson/p484-anderson.pdf;
                                         http://www.acm.org/pubs/citations/

proceedings/mod/276304/p484-anderson/

.

Stonebraker:1998:WWR

[1255] Michael Stonebraker. Are we working on the right problems? (panel). In Haas and Tiwary [2280], page 496. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p496-stonebraker/p496-stonebraker.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/276304/
p496-stonebraker/.

Mylopoulos:1998:NGD

[1256] John Mylopoulos. Next generation database systems won’t work without semantics! (panel). In Haas and Tiwary [2280], page 497. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p497-mylopoulos/p497-mylopoulos.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/276304/
p497-mylopoulos/.

Adam:1998:ECT

[1257] Nabil R. Adam and Yelena Yesha.

Electronic commerce: tutorial. In Haas and Tiwary [2280], page 498. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p498-adam/p498-adam.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p498-adam/.

Kemper:1998:SRT

[1258] Alfons Kemper, Donald Kossmann, and Florian Matthes. SAP R/3 (tutorial): a database application system. In Haas and Tiwary [2280], page 499. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p499-kemper/p499-kemper.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/276304/p499-kemper/.

Clossman:1998:JRD

[1259] Gray Clossman, Phil Shaw, Mark Hapner, Johannes Klein, Richard Pledereder, and Brian Becker. Java and relational databases (tutorial): SQLJ. In Haas and Tiwary [2280], page 500. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p500-clossman/p500-clossman.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/276304/p500-clossman/

.

Berchtold:1998:HDI

[1260] Stefan Berchtold and Daniel A. Keim. High-dimensional index structures database support for next decade’s applications (tutorial). In Haas and Tiwary [2280], page 501. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p501-berchtold/p501-berchtold.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/276304/p501-berchtold/

.

Blakeley:1998:MUD

[1261] José A. Blakeley and Michael J.

Pizzo. Microsoft universal data access platform. In Haas and Tiwary [2280], pages 502–503. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p502-blakeley/p502-blakeley.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/276304/p502-blakeley/

.

White:1998:EJP

[1262] Seth White, Rick Cattell, and Shel Finkelstein. Enterprise Java platform data access. In Haas and Tiwary [2280], pages 504–505. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p504-white/p504-white.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p504-white/.

Reinwald:1998:SOH

[1263] Berthold Reinwald and Hamid Pirahesh. SQL open heterogeneous data access. In Haas and Tiwary [2280], pages 506–507. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p506-reinwald/p506-reinwald.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/276304/p506-reinwald/

.

Eaton:1998:MLS

[1264] Chris Eaton. Managing large systems with DB2 UDB. In Haas and Tiwary [2280], pages 508–509. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p508-eaton/p508-eaton.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p508-eaton/.

Doherty:1998:DSM

[1265] C. Gregory Doherty. Database systems management and Oracle8. In Haas and Tiwary [2280], pages 510–511. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p510-doherty/p510-doherty.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/276304/p510-doherty/

.

Spiro:1998:UST

[1266] Peter Spiro. Ubiquitous, self-tuning, scalable servers. In Haas and Tiwary [2280], pages 512–515. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p512-spiro/p512-spiro.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p512-spiro/.

Franklin:1998:DYF

[1267] Michael Franklin and Stan Zdonik. “data in your face”: push technology in perspective. In Haas and Tiwary [2280], pages 516–519. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/

pubs/articles/proceedings/mod/276304/
p516-franklin/p516-franklin.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/276304/p516-franklin/

.

Ramakrishnan:1998:PNA

[1268] Satish Ramakrishnan and Vibha Dayal.

The PointCast network (abstract). In Haas and Tiwary [2280], page 520. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p520-ramakrishnan/p520-ramakrishnan.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/276304/
p520-ramakrishnan/.

Chan:1998:TPS

[1269] Arvola Chan. Transactional publish/subscribe: the proactive multicast of database changes (abstract). In Haas and Tiwary [2280], page 521. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p521-chan/p521-chan.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p521-chan/.

Ng:1998:DMA

[1270] KianSing Ng, Huan Liu, and Hwee-Bong Kwah. A data mining application: customer retention at the Port of Singapore Authority (PSA). In Haas and Tiwary [2280], pages 522–525. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p522-ng/p522-ng.pdf;   http://www.

                                        acm.org/pubs/citations/proceedings/
                                        mod/276304/p522-ng/.

Anderson:1998:ORR

[1271] Richard Anderson, Gopalan Arun, and Richard Frank. Oracle Rdb’s record caching model. In Haas and Tiwary [2280], pages 526–527. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p526-anderson/p526-anderson.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/276304/p526-anderson/

.

Lahiri:1998:UOU

[1272] Tirthankar Lahiri, Ashok Joshi, Amit Jasuja, and Sumanta Chatterjee. 50,000 users on an Oracle8 universal server database. In Haas and Tiwary [2280], pages 528–530. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p528-lahiri/p528-lahiri.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/276304/p528-lahiri/.

Aulakh:1998:AQD

[1273] Kamar Aulakh. About Quark Digital Media System. In Haas and Tiwary [2280], pages 531–532. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p531-aulakh/p531-aulakh.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/276304/p531-aulakh/.

Whelan:1998:FID

[1274] Daniel S. Whelan. FileNet integrated document management database usage and issues. In Haas and Tiwary [2280], page 533. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p533-whelan/p533-whelan.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p533-whelan/.

Nauman:1998:DHT

[1275] John Nauman and Ray Suorsa.

Developing a high traffic, readonly Web site. In Haas and Tiwary [2280], pages 534–535. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p534-nauman/p534-nauman.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p534-nauman/.

Chong:1998:RBP

[1276] James Chong. Real business processing meets the Web. In Haas and Tiwary [2280], page 536. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p536-chong/p536-chong.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p536-chong/.

Lassettre:1998:ORD

[1277] Edwin R. Lassettre. Olympic records for data at the 1998 Nagano games. In Haas and Tiwary [2280], page 537.

ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p537-lassettre/p537-lassettre.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/276304/p537-lassettre/

.

Brewer:1998:DHA

[1278] Eric A. Brewer. Delivering high availability for Inktomi search engines. In Haas and Tiwary [2280], page 538. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p538-brewer/p538-brewer.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p538-brewer/.

Kennamer:1998:MCH

[1279] Sherri Kennamer. Microsoft.com: a high-scale data management and transaction processing solution. In Haas and Tiwary [2280], pages 539–540. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p539-kennamer/p539-kennamer.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/276304/p539-kennamer/

.

Li:1998:FPD

[1280] Bin Li and Dennis Shasha. Free parallel data mining. In Haas and Tiwary [2280], pages 541–543. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p541-li/p541-li.pdf;   http://www.

acm.org/pubs/citations/proceedings/
mod/276304/p541-li/.

Mecca:1998:AWB

[1281] G. Mecca, P. Atzeni, A. Masci, G. Sindoni, and P. Merialdo. The Araneus Web-based management system. In Haas and Tiwary [2280], pages 544–546. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p544-mecca/p544-mecca.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p544-mecca/.

Liu:1998:CPU

[1282] Ling Liu, Calton Pu, Wei Tang, David Buttler, John Biggs, Tong Zhou, Paul Benninghoff, Wei Han, and Fenghua Yu. CQ: a personalized update monitoring toolkit. In Haas and Tiwary [2280], pages 547–549. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p547-liu/p547-liu.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/276304/p547-liu/.

Olston:1998:D

[1283] Chris Olston, Allison Woodruff, Alexander Aiken, Michael Chu, Vuk Ercegovac, Mark Lin, Mybrid Spalding, and Michael Stonebraker. DataSplash. In Haas and Tiwary [2280], pages 550–552. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/

                                        p550-olston/p550-olston.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/276304/p550-olston/.

Chaudhuri:1998:MIT

[1284] Surajit Chaudhuri and Vivek Narasayya. Microsoft index turning wizard for SQL Server 7.0. In Haas and Tiwary [2280], pages 553–554. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p553-chaudhuri/p553-chaudhuri.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/276304/p553-chaudhuri/

.

Kiepuszewski:1998:FPB

[1285] Bartek Kiepuszewski, Ralf Muhlberger, and Maria E. Orlowska. FlowBack: providing backward recovery for workflow management systems. In Haas and Tiwary [2280], pages 555–557. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p555-kiepuszewski/p555-kiepuszewski.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/276304/
                                        p555-kiepuszewski/.

Cohen:1998:PDL

[1286] William W. Cohen. Providing database-like access to the Web using queries based on textual similarity. In Haas and Tiwary [2280], pages 558–560. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/

p558-cohen/p558-cohen.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p558-cohen/.

Ambite:1998:ASC

[1287] José Luis Ambite, Naveen Ashish, Greg Barish, Craig A. Knoblock, Steven Minton, Pragnesh J. Modi, Ion Muslea, Andrew Philpot, and Sheila Tejada. Ariadne: a system for constructing mediators for Internet sources. In Haas and Tiwary [2280], pages 561–563. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p561-ambite/p561-ambite.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p561-ambite/.

Li:1998:CBM

[1288] Chen Li, Ramana Yerneni, Vasilis Vassalos, Hector Garcia-Molina, Yannis Papakonstantinou, Jeffrey Ullman, and Murty Valiveti. Capability based mediation in TSIMMIS. In Haas and Tiwary [2280], pages 564–566. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p564-li/p564-li.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/276304/p564-li/.

Avnur:1998:CCO

[1289] Ron Avnur, Joseph M. Hellerstein, Bruce Lo, Chris Olston, Bhaskaran Raman, Vijayshankar Raman, Tali Roth, and Kirk Wylie. CONTROL: continuous output and navigation technology with refinement on-line. In Haas and Tiwary [2280], pages 567–569.

ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p567-avnur/p567-avnur.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/276304/p567-avnur/.

Kornacker:1998:AAM

[1290] Marcel Kornacker, Mehul Shah, and Joseph M. Hellerstein. AMDB: an access method debugging tool. In Haas and Tiwary [2280], pages 570–571. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p570-kornacker/p570-kornacker.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/276304/p570-kornacker/

.

Lacroix:1998:UOS

[1291] Zoé Lacroix, Arnaud Sahuguet, and Raman Chandrasekar. User-oriented smart-cache for the Web: what you seek is what you get! In Haas and Tiwary [2280], pages 572–574. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p572-lacroix/p572-lacroix.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/276304/p572-lacroix/

.

Baumann:1998:MDS

[1292] P. Baumann, A. Dehmel, P. Furtado, R. Ritsch, and N. Widmann. The multidimensional database system RasDaMan. In Haas and Tiwary [2280], pages 575–577. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p575-baumann/p575-baumann.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/276304/p575-baumann/

.

Park:1998:XEM

[1293] Jang Ho Park, Yong Sik Kwon, Ki Hong Kim, Sang Ho Lee, Byoung Dae Park, and Sang Kyun Cha. Xmas: an extensible mainmemory storage system for highperformance applications. In Haas and Tiwary [2280], pages 578–580. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p578-park/p578-park.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p578-park/.

Zaiane:1998:MSP

[1294] Osmar R. Zaïane, Jiawei Han, Ze-Nian Li, Sonny H. Chee, and Jenny Y. Chiang. MultiMediaMiner: a system prototype for multimedia data mining. In Haas and Tiwary [2280], pages 581–583. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p581-zaiane/p581-zaiane.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p581-zaiane/.

Toyama:1998:SES

[1295] Motomichi Toyama. SuperSQL: an extended SQL for database publishing and presentation. In Haas and Tiwary [2280], pages 584–586. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p584-toyama/p584-toyama.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/276304/p584-toyama/.

Ceri:1998:IWL

[1296] Stefano Ceri, Piero Fraternali, and Stefano Paraboschi. The IDEA Web lab. In Haas and Tiwary [2280], pages 587–589. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p587-ceri/p587-ceri.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/276304/p587-ceri/.

Dar:1998:DDD

[1297] Shaul Dar, Gadi Entin, Shai Geva, and Eran Palmon. DTL’s DataSpot: database exploration as easy as browsing the Web …. In Haas and Tiwary [2280], pages 590–592. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/276304/
                                        p590-dar/p590-dar.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/276304/p590-dar/.

Chen:1998:PPQ

[1298] Jing Chen, Limsoon Wong, and Louxin Zhang. A protein patent query system powered by Kleisli. In Haas and Tiwary [2280], pages 593–595. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/276304/
p593-chen/p593-chen.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/276304/p593-chen/.

Abiteboul:1999:VX

[1299] Serge Abiteboul. On views and XML.

In ACM [2283], pages 1–9. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p1-abiteboul/p1-abiteboul.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/303976/

p1-abiteboul/. ACM order number 475990.

Mamoulis:1999:ISJ

[1300] Nikos Mamoulis and Dimitris Papadias. Integration of spatial join algorithms for processing multiple inputs. In Delis et al. [2282], pages 1–12. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p1-mamoulis/p1-mamoulis.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p1-mamoulis/.

Alon:1999:TJS

[1301] Noga Alon, Phillip B. Gibbons, Yossi Matias, and Mario Szegedy. Tracking join and self-join sizes in limited storage. In ACM [2283], pages 10–20. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p10-alon/p10-alon.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/303976/p10-alon/

. ACM order number 475990.

Acharya:1999:SES

[1302] Swarup Acharya, Viswanath Poosala, and Sridhar Ramaswamy. Selectivity estimation in spatial databases. In Delis et al. [2282], pages 13–24. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p13-acharya/p13-acharya.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/304182/p13-acharya/.

Gottlob:1999:HDT

[1303] Georg Gottlob, Nicola Leone, and Francesco Scarcello. Hypertree decompositions and tractable queries. In ACM [2283], pages 21–32. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

                                        1999. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        303976/p21-gottlob/p21-gottlob.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/303976/

p21-gottlob/. ACM order number

475990.

Chakrabarti:1999:ECC

[1304] Kaushik Chakrabarti and Sharad Mehrotra. Efficient concurrency control in multidimensional access methods. In Delis et al. [2282], pages 25–36. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p25-chakrabarti/p25-chakrabarti.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/304182/
                                        p25-chakrabarti/.

More:1999:EST

[1305] Sachin More, S. Muthukrishnan, and Elizabeth Shriver. Efficiently sequencing tape-resident jobs. In ACM [2283], pages 33–43. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p33-more/p33-more.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/303976/p33-more/

. ACM order number 475990.

Jagadish:1999:SSO

[1306] H. V. Jagadish, Laks V. S. Lakshmanan, and Divesh Srivastava. Snakes and sandwiches: optimal clustering strategies for a data warehouse. In Delis et al. [2282], pages 37–48. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p37-jagadish/p37-jagadish.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/304182/p37-jagadish/

.

Papadias:1999:POM

[1307] Dimitris Papadias, Nikos Mamoulis, and Yannis Theodoridis. Processing and optimization of multiway spatial joins using R-trees. In ACM [2283], pages 44–55. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p44-papadias/p44-papadias.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/303976/

p44-papadias/. ACM order number 475990.

Ankerst:1999:OOP

[1308] Mihael Ankerst, Markus M. Breunig, Hans-Peter Kriegel, and Jörg Sander. OPTICS: ordering points to identify the clustering structure. In Delis et al. [2282], pages 49–60. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p49-ankerst/p49-ankerst.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/304182/p49-ankerst/.

Buneman:1999:IBP

[1309] Peter Buneman, Wenfie Fan, and Scott Weinstein. Interaction between path and type constraints. In ACM [2283], pages 56–67. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

                                        1999. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        303976/p56-buneman/p56-buneman.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/303976/

p56-buneman/. ACM order number

475990.

Aggarwal:1999:FAP

[1310] Charu C. Aggarwal, Joel L. Wolf, Philip S. Yu, Cecilia Procopiuc, and Jong Soo Park. Fast algorithms for projected clustering. In Delis et al. [2282], pages 61–72. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p61-aggarwal/p61-aggarwal.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/304182/p61-aggarwal/

.

Arenas:1999:CQA

[1311] Marcelo Arenas, Leopoldo Bertossi, and Jan Chomicki. Consistent query answers in inconsistent databases. In ACM [2283], pages 68–79. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p68-arenas/p68-arenas.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/303976/p68-arenas/

. ACM order number 475990.

Lomet:1999:LLE

[1312] David Lomet and Mark Tuttle. Logical logging to extend recovery to new domains. In Delis et al. [2282], pages 73–

84. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p73-lomet/p73-lomet.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p73-lomet/.

VandenBussche:1999:TIP

[1313] Jan Van den Bussche and Emmanuel Waller. Type inference in the polymorphic relational algebra. In ACM [2283], pages 80–90. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p80-van_den_bussche/p80-
van_den_bussche.pdf;   http://www.
acm.org/pubs/citations/proceedings/
pods/303976/p80-van_den_bussche/

. ACM order number 475990.

Shanmugasundaram:1999:ECC

[1314] Jayavel Shanmugasundaram, Arvind Nithrakashyap, Rajendran Sivasankaran, and Krithi Ramamritham. Efficient concurrency control for broadcast environments. In Delis et al. [2282], pages 85–96. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

QA1 .A87. URL http://www.acm.
org/pubs/articles/proceedings/mod/
304182/p85-shanmugasundaram/p85-
shanmugasundaram.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/304182/p85-shanmugasundaram/

.

Hara:1999:RAN

[1315] Carmem S. Hara and Susan B. Davidson. Reasoning about nested functional dependencies. In ACM [2283], pages 91–100. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p91-hara/p91-hara.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/303976/p91-hara/

. ACM order number 475990.

Breitbart:1999:UPP

[1316] Yuri Breitbart, Raghavan Komondoor, Rajeev Rastogi, S. Seshadri, and Avi Silberschatz. Update propagation protocols for replicated databases. In Delis et al. [2282], pages 97–108. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p97-breitbart/p97-breitbart.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/304182/p97-breitbart/

.

Tygar:1999:OPE

[1317] J. D. Tygar. Open problems in electronic commerce. In ACM [2283], page 101. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p101-tygar/p101-tygar.pdf;

 http://www.acm.org/pubs/citations/
proceedings/pods/303976/p101-tygar/

. ACM order number 475990.

Benedikt:1999:EAA

[1318] Michael Benedikt and Leonid Libkin.

Exact and approximate aggregation in constraint query languages. In ACM [2283], pages 102–113. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p102-benedikt/p102-benedikt.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/303976/

p102-benedikt/. ACM order number 475990.

Jamil:1999:BRM

[1319] Hasan M. Jamil. Belief reasoning in MLS deductive databases. In Delis et al. [2282], pages 109–120. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p109-jamil/p109-jamil.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p109-jamil/.

Dawson:1999:MDU

[1320] Steven Dawson, Sabrina De Capitani di Vimercati, Patrick Lincoln, and Pierangela Samarati. Minimal data upgrading to prevent inference and association attacks. In ACM [2283], pages 114–125. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p114-dawson/p114-dawson.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/303976/

p114-dawson/. ACM order number

475990.

Adali:1999:MPA

[1321] S. Adali, M. L. Sapino, and V. S. Subrahmanian. A multimedia presentation algebra. In Delis et al. [2282], pages 121–132. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p121-adali/p121-adali.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/304182/p121-adali/.

Ganti:1999:FMC

[1322] Venkatesh Ganti, Johannes Gehrke, and Raghu Ramakrishnan. A framework for measuring changes in data characteristics. In ACM [2283], pages 126–137. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

                                        1999. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        303976/p126-ganti/p126-ganti.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/303976/p126-ganti/

. ACM order number 475990.

Jagadish:1999:QND

[1323] H. V. Jagadish, Laks V. S. Lakshmanan, Tova Milo, Divesh Srivastava, and Dimitra Vista. Querying network directories. In Delis et al. [2282], pages 133–144. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p133-jagadish/p133-jagadish.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/304182/p133-jagadish/

.

Chu:1999:LEC

[1324] Francis Chu, Joseph Y. Halpern, and Praveen Seshadri. Least expected cost query optimization: An exercise in utility. In ACM [2283], pages 138–147. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p138-chu/p138-chu.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/303976/p138-chu/

. ACM order number 475990.

Hidber:1999:OAR

[1325] Christian Hidber. Online association rule mining. In Delis et al. [2282], pages 145–156. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p145-hidber/p145-hidber.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p145-hidber/.

Cosmadakis:1999:ICR

[1326] Stavros Cosmadakis. Inherent complexity of recursive queries. In ACM [2283], pages 148–154. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p148-cosmadakis/p148-cosmadakis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/303976/
p148-cosmadakis/. ACM order num-

ber 475990.

Cohen:1999:RAQ

[1327] Sara Cohen, Werner Nutt, and Alexander Serebrenik. Rewriting aggregate queries using views. In ACM [2283], pages 155–166. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p155-cohen/p155-cohen.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/303976/p155-cohen/

. ACM order number 475990.

Lakshmanan:1999:OCF

[1328] Laks V. S. Lakshmanan, Raymond Ng, Jiawei Han, and Alex Pang. Optimization of constrained frequent set queries with 2-variable constraints. In Delis et al. [2282], pages 157–168. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p157-lakshmanan/p157-lakshmanan.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/304182/
p157-lakshmanan/.

Karloff:1999:CVS

[1329] Howard Karloff and Milena Mihail. On the complexity of the view-selection problem. In ACM [2283], pages 167–173. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p167-karloff/p167-karloff.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/303976/

p167-karloff/. ACM order number 475990.

Gehrke:1999:BOD

[1330] Johannes Gehrke, Venkatesh Ganti, Raghu Ramakrishnan, and Wei-Yin Loh. BOAT — optimistic decision tree construction. In Delis et al. [2282], pages 169–180. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p169-gehrke/p169-gehrke.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p169-gehrke/.

Grumbach:1999:QAD

[1331] Stéphane Grumbach, Maurizio Rafanelli, and Leonardo Tininini. Querying aggregate data. In ACM [2283], pages 174–184. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p174-grumbach/p174-grumbach.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/303976/

p174-grumbach/. ACM order number 475990.

Aboulnaga:1999:STH

[1332] Ashraf Aboulnaga and Surajit Chaudhuri. Self-tuning histograms: building histograms without looking at data. In Delis et al. [2282], pages 181–192. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p181-aboulnaga/p181-aboulnaga.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/304182/p181-aboulnaga/

.

Kleinberg:1999:ALA

[1333] Jon Kleinberg and Andrew Tomkins.

Applications of linear algebra in information retrieval and hypertext analysis. In ACM [2283], pages 185–193. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p185-kleinberg/p185-kleinberg.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/303976/
p185-kleinberg/. ACM order number

475990.

Vitter:1999:ACM

[1334] Jeffrey Scott Vitter and Min Wang. Approximate computation of multidimensional aggregates of sparse data using wavelets. In Delis et al. [2282], pages 193–204. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p193-vitter/p193-vitter.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p193-vitter/.

Calvanese:1999:RRE

[1335] Diego Calvanese, Giuseppe De Giacomo, Maurizio Lenzerini, and Moshe Y. Vardi. Rewriting of regular expressions and regular path queries. In ACM [2283], pages 194–204. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p194-calvanese/p194-calvanese.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/303976/
p194-calvanese/. ACM order number

475990.

Lee:1999:MDS

[1336] Ju-Hong Lee, Deok-Hwan Kim, and Chin-Wan Chung. Multi-dimensional selectivity estimation using compressed histogram information. In Delis et al. [2282], pages 205–214. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p205-lee/p205-lee.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/304182/p205-lee/.

Neven:1999:QA

[1337] Frank Neven and Thomas Schwentick.

Query automata. In ACM [2283], pages 205–214. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p205-neven/p205-neven.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/303976/p205-neven/

. ACM order number 475990.

Chan:1999:EBE

[1338] Chee-Yong Chan and Yannis E. Ioannidis. An efficient bitmap encoding scheme for selection queries. In Delis et al. [2282], pages 215–226. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p215-chan/p215-chan.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p215-chan/.

Milo:1999:TIQ

[1339] Tova Milo and Dan Suciu. Type inference for queries on semistructured data. In ACM [2283], pages 215–226. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p215-milo/p215-milo.pdf;
 http://www.acm.org/pubs/citations/
proceedings/pods/303976/p215-milo/

. ACM order number 475990.

Kanza:1999:QIA

[1340] Yaron Kanza, Werner Nutt, and Yehoshua Sagiv. Queries with incomplete answers over semistructured data. In ACM [2283], pages 227–236. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

                                        1999. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        303976/p227-kanza/p227-kanza.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/303976/p227-kanza/

. ACM order number 475990.

Wu:1999:QOS

[1341] Ming-Chuan Wu. Query optimization for selections using bitmaps. In Delis et al. [2282], pages 227–238. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p227-wu/p227-wu.pdf;   http://www.
                                        acm.org/pubs/citations/proceedings/
                                        mod/304182/p227-wu/.

Chu:1999:FTS

[1342] Kelvin Kam Wing Chu and Man Hon Wong. Fast time-series searching with scaling and shifting. In ACM [2283], pages 237–248. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

                                        1999. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        303976/p237-chu/p237-chu.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/pods/303976/p237-chu/

. ACM order number 475990.

Blohsfeld:1999:CSE

[1343] Björn Blohsfeld, Dieter Korus, and Bernhard Seeger. A comparison of selectivity estimators for range queries on metric attributes. In Delis et al. [2282], pages 239–250. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p239-blohsfeld/p239-blohsfeld.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/304182/p239-blohsfeld/

.

Jagadish:1999:SSE

[1344] H. V. Jagadish, Raymond T. Ng, and Divesh Srivastava. Substring selectivity estimation. In ACM [2283], pages 249–260. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p249-jagadish/p249-jagadish.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/303976/

p249-jagadish/. ACM order number 475990.

Manku:1999:RST

[1345] Gurmeet Singh Manku, Sridhar Rajagopalan, and Bruce G. Lindsay. Random sampling techniques for space efficient online computation of order statistics of large datasets. In Delis et al. [2282], pages 251–262. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p251-manku/p251-manku.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p251-manku/.

Kollios:1999:IMO

[1346] George Kollios, Dimitrios Gunopulos, and Vassilis J. Tsotras. On indexing mobile objects. In ACM [2283], pages 261–272. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

                                        1999. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        303976/p261-kollios/p261-kollios.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/303976/

p261-kollios/. ACM order number

475990.

Chaudhuri:1999:RSJ

[1347] Surajit Chaudhuri, Rajeev Motwani, and Vivek Narasayya. On randsom sampling over joins. In Delis et al. [2282], pages 263–274. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p263-chaudhuri/p263-chaudhuri.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/304182/p263-chaudhuri/

.

Cluet:1999:ULD

[1348] Sophie Cluet, Olga Kapitskaia, and Divesh Srivastava. Using LDAP directory caches. In ACM [2283], pages 273–284. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

                                        1999. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        303976/p273-cluet/p273-cluet.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/303976/p273-cluet/

. ACM order number 475990.

Acharya:1999:JSA

[1349] Swarup Acharya, Phillip B. Gibbons, Viswanath Poosala, and Sridhar Ramaswamy. Join synopses for approximate query answering. In Delis et al. [2282], pages 275–286. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p275-acharya/p275-acharya.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/304182/p275-acharya/

.

Alonso:1999:CGC

[1350] Gustavo Alonso, Armin Feßler, Guy Pardon, and Hans-Jörg Schek. Correctness in general configurations of transactional components. In ACM [2283], pages 285–293. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p285-alonso/p285-alonso.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/303976/

p285-alonso/. ACM order number 475990.

Haas:1999:RJO

[1351] Peter J. Haas and Joseph M.

Hellerstein. Ripple joins for online aggregation. In Delis et al. [2282], pages 287–298. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p287-haas/p287-haas.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p287-haas/.

Bonner:1999:WTD

[1352] Anthony J. Bonner. Workflow, transactions, and datalog. In ACM [2283], pages 294–305. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p294-bonner/p294-bonner.

                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/pods/303976/

p294-bonner/. ACM order number

475990.

Ives:1999:AQE

[1353] Zachary G. Ives, Daniela Florescu, Marc Friedman, Alon Levy, and Daniel S. Weld. An adaptive query execution system for data integration. In Delis et al. [2282], pages 299–310. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p299-ives/p299-ives.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/304182/p299-ives/.

Al-Houmaily:1999:AIP

[1354] Yousef J. Al-Houmaily and Panos K. Chrysanthis. Atomicity with incompatible presumptions. In ACM [2283], pages 306–315. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

                                        1999. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        303976/p306-al-houmaily/p306-al-
                                        houmaily.pdf;   http://www.acm.
                                        org/pubs/citations/proceedings/
                                        pods/303976/p306-al-houmaily/. ACM

order number 475990.

Florescu:1999:QOP

[1355] Daniela Florescu, Alon Levy, Ioana Manolescu, and Dan Suciu. Query optimization in the presence of limited access patterns. In Delis et al. [2282], pages 311–322. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p311-florescu/p311-florescu.pdf;

 http://www.acm.org/pubs/citations/
proceedings/mod/304182/p311-florescu/

.

Schuldt:1999:CCR

[1356] Heiko Schuldt, Gustavo Alonso, and Hans-Jörg Schek. Concurrency control and recovery in transactional process management. In ACM [2283], pages 316–326. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p316-schuldt/p316-schuldt.
pdf;   http://www.acm.org/pubs/
citations/proceedings/pods/303976/

p316-schuldt/. ACM order number 475990.

Marathe:1999:QPT

[1357] Arunprasad P. Marathe and Kenneth Salem. Query processing techniques for arrays. In Delis et al. [2282], pages 323–334. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p323-marathe/p323-marathe.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/304182/p323-marathe/

.

Boasson:1999:WAS

[1358] Luc Boasson, Patrick Cegielski, Irène Guessarian, and Yuri Matiyasevich. Window-accumulated subsequence matching problem is linear. In ACM [2283], pages 327–336. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p327-boasson/p327-boasson.
pdf;   http://www.acm.org/pubs/

                                        citations/proceedings/pods/303976/

p327-boasson/. ACM order number

475990.

Chang:1999:MYV

[1359] Chen-Chuan K. Chang and Héctor García-Molina. Mind your vocabulary: query mapping across heterogeneous information sources. In Delis et al. [2282], pages 335–346. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p335-chang/p335-chang.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/304182/p335-chang/.

Raman:1999:LPD

[1360] Vijayshankar Raman. Localitypreserving dictionaries: theory and application to clustering in databases. In ACM [2283], pages 337–345. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

                                        1999. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        303976/p337-raman/p337-raman.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/303976/p337-raman/

. ACM order number 475990.

Arge:1999:TDI

[1361] Lars Arge, Vasilis Samoladas, and Jeffrey Scott Vitter. On twodimensional indexability and optimal range search indexing. In ACM [2283], pages 346–357. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

                                        1999. URL http://www.acm.org/
                                        pubs/articles/proceedings/pods/
                                        303976/p346-arge/p346-arge.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/pods/303976/p346-arge/

. ACM order number 475990.

Mayr:1999:CSQ

[1362] Tobias Mayr and Praveen Seshadri.

Client-site query extensions. In Delis et al. [2282], pages 347–358. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p347-mayr/p347-mayr.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p347-mayr/.

Malmi:1999:GUR

[1363] Lauri Malmi and Eljas Soisalon-Soininen. Group updates for relaxed height-balanced trees. In ACM [2283], pages 358–367. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296

1999. URL http://www.acm.org/
pubs/articles/proceedings/pods/
303976/p358-malmi/p358-malmi.pdf;
http://www.acm.org/pubs/citations/
proceedings/pods/303976/p358-malmi/

. ACM order number 475990.

Beyer:1999:BCS

[1364] Kevin Beyer and Raghu Ramakrishnan. Bottom-up computation of sparse and Iceberg CUBE. In Delis et al. [2282], pages 359–370. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p359-beyer/p359-beyer.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p359-beyer/.

Kotidis:1999:DDV

[1365] Yannis Kotidis and Nick Roussopoulos. DynaMat: a dynamic view management system for data warehouses. In Delis et al. [2282], pages 371–382.

ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p371-kotidis/p371-kotidis.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/304182/p371-kotidis/

.

Labio:1999:SWU

[1366] Wilburt Juan Labio, Ramana Yerneni, and Hector Garcia-Molina. Shrinking the warehouse update Window. In Delis et al. [2282], pages 383–394. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p383-labio/p383-labio.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p383-labio/.

Natsev:1999:WSR

[1367] Apostol Natsev, Rajeev Rastogi, and Kyuseok Shim. WALRUS: a similarity retrieval algorithm for image databases. In Delis et al. [2282], pages 395–406. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p395-natsev/p395-natsev.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p395-natsev/.

Aggarwal:1999:NMS

[1368] Charu C. Aggarwal, Joel L. Wolf, and Philip S. Yu. A new method for similarity indexing of market basket data. In Delis et al. [2282], pages 407–418. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p407-aggarwal/p407-aggarwal.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/304182/p407-aggarwal/

.

Keim:1999:EGB

[1369] Daniel A. Keim. Efficient geometrybased similarity search of 3D spatial databases. In Delis et al. [2282], pages 419–430. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p419-keim/p419-keim.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p419-keim/.

Deutsch:1999:SSD

[1370] Alin Deutsch, Mary Fernandez, and Dan Suciu. Storing semistructured data with STORED. In Delis et al. [2282], pages 431–442. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p431-deutsch/p431-deutsch.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/304182/p431-deutsch/

.

Yerneni:1999:CCM

[1371] Ramana Yerneni, Chen Li, Hector Garcia-Molina, and Jeffrey Ullman. Computing capabilities of mediators. In Delis et al. [2282], pages 443–454. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p443-yerneni/p443-yerneni.pdf;

                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/304182/p443-yerneni/

.

Papakonstantinou:1999:QRS

[1372] Yannis Papakonstantinou and Vasilis Vassalos. Query rewriting for semistructured data. In Delis et al. [2282], pages 455–466. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p455-papakonstantinou/p455-papakonstantinou.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/304182/
                                        p455-papakonstantinou/.

Embley:1999:RBD

[1373] D. W. Embley, Y. Jiang, and Y.- K. Ng. Record-boundary discovery in Web documents. In Delis et al. [2282], pages 467–478. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p467-embley/p467-embley.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/304182/p467-embley/.

Callan:1999:ADL

[1374] Jamie Callan, Margaret Connell, and Aiqun Du. Automatic discovery of language models for text databases. In Delis et al. [2282], pages 479–490. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p479-callan/p479-callan.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/304182/p479-callan/.

Davulcu:1999:LAQ

[1375] Hasan Davulcu, Juliana Freire, Michael Kifer, and I. V. Ramakrishnan. A layered architecture for querying dynamic Web content. In Delis et al. [2282], pages 491–502. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p491-davulcu/p491-davulcu.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/304182/p491-davulcu/

.

Seshadri:1999:HSD

[1376] Praveen Seshadri. “Honey, I shrunk the database”: footprint, mobility, and beyond. In Delis et al. [2282], page 503. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p503-seshadri/p503-seshadri.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/304182/p503-seshadri/

.

Garofalakis:1999:CPM

[1377] Minos N. Garofalakis, Sridhar Ramaswamy, Rajeev Rastogi, and Kyuseok Shim. Of crawlers, portals, mice, and men: is there more to mining the Web? In Delis et al. [2282], page 504. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p504-garofalakis/p504-garofalakis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/304182/
p504-garofalakis/.

Ozsu:1999:DMI

[1378] M. Tamer Özsu. Data management issues in electronic commerce. In Delis et al. [2282], page 505. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p505-ozsu/p505-ozsu.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/304182/p505-ozsu/.

Dullmann:1999:PD

[1379] Dirk Düllmann. Petabyte databases. In Delis et al. [2282], page 506. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p506-dullmann/p506-dullmann.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/304182/p506-dullmann/

.

Mohan:1999:DPL

[1380] C. Mohan. A database perspective on Lotus Domino/Notes. In Delis et al. [2282], page 507. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p507-mohan/p507-mohan.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/304182/p507-mohan/.

Chakrabarti:1999:HDD

[1381] Soumen Chakrabarti. Hypertext databases and data mining. In Delis et al. [2282], page 508. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p508-chakrabarti/p508-chakrabarti.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/304182/
p508-chakrabarti/.

Hinneburg:1999:CML

[1382] Alexander Hinneburg and Daniel A.

Keim. Clustering methods for large databases: from the past to the future. In Delis et al. [2282], page 509. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p509-hinneburg/p509-hinneburg.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/304182/p509-hinneburg/

.

Suciu:1999:MWD

[1383] Dan Suciu. Managing Web data.

In Delis et al. [2282], page 510. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p510-suciu/p510-suciu.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p510-suciu/.

Carey:1999:WHD

[1384] M. Carey, D. Chamberlin, D. Doole, S. Rielau, N. Mattos, S. Narayanan, B. Vance, and R. Swagerman. O-O, what’s happening to DB2? In Delis et al. [2282], pages 511–512. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p511-carey/p511-carey.pdf;   http:

//www.acm.org/pubs/citations/proceedings/
mod/304182/p511-carey/.

Krishnamurthy:1999:BOR

[1385] Vishu Krishnamurthy, Sandeepan Banerjee, and Anil Nori. Bringing object-relational technology to the mainstream. In Delis et al. [2282], pages 513–514. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p513-krishnamurthy/p513-krishnamurthy.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/304182/
p513-krishnamurthy/.

Brown:1999:ISS

[1386] Paul Brown. Implementing the spirit of SQL-99. In Delis et al. [2282], pages 515–518. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p515-brown/p515-brown.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p515-brown/.

Baulier:1999:DSM

[1387] J. Baulier, P. Bohannon, S. Gogate, C. Gupta, and S. Haldar. DataBlitz storage manager: mainmemory database performance for critical applications. In Delis et al. [2282], pages 519–520. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p519-baulier/p519-baulier.pdf;
http://www.acm.org/pubs/citations/

proceedings/mod/304182/p519-baulier/

.

Kanth:1999:IMD

[1388] K. V. Ravi Kanth, Siva Ravada, Jayant Sharma, and Jay Banerjee. Indexing medium-dimensionality data in Oracle. In Delis et al. [2282], pages 521–522. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p521-ravi_kanth/p521-ravi_kanth.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/304182/
p521-ravi_kanth/.

Kohler:1999:EIS

[1389] Walt Kohler. EMC information sharing: direct access to MVS data from UNIX and NT. In Delis et al. [2282], pages 523–524. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p523-kohler/p523-kohler.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p523-kohler/.

Greer:1999:DFG

[1390] Rick Greer. Daytona and the foruthgeneration language cymbal. In Delis et al. [2282], pages 525–526. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p525-greer/p525-greer.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p525-greer/.

Hammond:1999:MRM

[1391] Brad Hammond. Merge replication in Microsoft’s SQL server 7.0. In Delis et al. [2282], page 527. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p527-hammond/p527-hammond.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/304182/p527-hammond/

.

Team:1999:MDM

[1392] TimesTen Team. In-memory data management for consumer transactions: the TimesTen approach. In Delis et al. [2282], pages 528–529. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p528-timesten_team/p528-timesten_
                                        team.pdf;   http://www.acm.org/
                                        pubs/citations/proceedings/mod/
                                        304182/p528-timesten_team/.

Ojjeh:1999:MSS

[1393] Bassel Ojjeh. Microsoft site server (commerce ed.): Talk-slides available at the conference. In Delis et al. [2282], page 530. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p530-ojjeh/p530-ojjeh.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/304182/p530-ojjeh/.

Rajaraman:1999:CDI

[1394] Anand Rajaraman. E-commerce database issues and experience: (talkslides available at the conference). In Delis et al. [2282], page 531. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p531-rajaraman/p531-rajaraman.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/304182/p531-rajaraman/

.

Bergstraesser:1999:VWM

[1395] Thomas Bergstraesser, Philip A.

Bernstein, Shankar Pal, and David Shutt. Versions and workspaces in Microsoft repository. In Delis et al. [2282], pages 532–533. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p532-bergstraesser/p532-bergstraesser.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/304182/
p532-bergstraesser/.

Do:1999:NDA

[1396] Lyman Do, Prabhu Ram, and Pamela Drew. The need for distributed asynchronous transactions. In Delis et al. [2282], pages 534–535. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p534-do/p534-do.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/304182/p534-do/.

Jarke:1999:IOD

[1397] Matthias Jarke, Christoph Quix, Guido Blees, Dirk Lehmann, Gunter Michalk, and Stefan Stierl. Improving OLTP data quality using data warehouse mechanisms. In Delis et al. [2282], pages 536–537. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p536-jarke/p536-jarke.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/304182/p536-jarke/.

Trisolini:1999:DIW

[1398] Stefano M. Trisolini, Maurizio Lenzerini, and Daniele Nardi. Data integration and warehousing in Telecom Italia. In Delis et al. [2282], pages 538–539. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p538-trisolini/p538-trisolini.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/304182/p538-trisolini/

.

Liu:1999:XBW

[1399] Ling Liu, Wei Han, David Buttler, Calton Pu, and Wei Tang. An XJML-based wrapper generator for Web information extraction. In Delis et al. [2282], pages 540–543. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p540-liu/p540-liu.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/304182/p540-liu/.

Altinel:1999:DTA

[1400] Mehmet Altinel, Demet Aksoy, Thomas Baby, Michael Franklin, William Shapiro, and Stan Zdonik. DBIStoolkit: adaptable middleware for large scale data delivery. In Delis et al. [2282], pages 544–546. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p544-altinel/p544-altinel.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/304182/p544-altinel/

.

Wolfson:1999:DDM

[1401] Ouri Wolfson, Prasad Sistla, Bo Xu, Jutai Zhou, and Sam Chamberlain. DOMINO: Databases fOr Mov-INg Objects tracking. In Delis et al. [2282], pages 547–549. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p547-wolfson/p547-wolfson.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/304182/p547-wolfson/

.

Braumandl:1999:DPI

[1402] Reinhard Braumandl, Alfons Kemper, and Donald Kossmann. Database patchwork on the Internet. In Delis et al. [2282], pages 550–552. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p550-braumandl/p550-braumandl.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/304182/p550-braumandl/

.

Rundensteiner:1999:EVE

[1403] E. A. Rundensteiner, A. Koeller, X. Zhang, A. J. Lee, A. Nica, A. Van Wyk, and Y. Lee. Evolvable view environment (EVE): non-equivalent view maintenance under schema changes. In Delis et al. [2282], pages 553–555. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p553-rundensteiner/p553-rundensteiner.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/proceedings/mod/304182/
                                        p553-rundensteiner/.

Ng:1999:EMC

[1404] Raymond Ng, Laks V. S. Lakshmanan, Jiawei Han, and Teresa Mah. Exploratory mining via constrained frequent set queries. In Delis et al. [2282], pages 556–558. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p556-ng/p556-ng.pdf;   http://www.
                                        acm.org/pubs/citations/proceedings/
                                        mod/304182/p556-ng/.

Adelberg:1999:NV

[1405] Brad Adelberg and Matthew Denny. Nodose version 2.0. In Delis et al. [2282], pages 559–561. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p559-adelberg/p559-adelberg.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/304182/p559-adelberg/

.

Barga:1999:PMA

[1406] Roger Barga and David B. Lomet. Phoenix: making applications robust. In Delis et al. [2282], pages 562–564.

ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p562-barga/p562-barga.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p562-barga/.

Li:1999:PSP

[1407] Wen-Syan Li, Quoc Vu, Edward Chang, Divyakant Agrawal, Kyoji Hirata, Sougata Mukherjea, Yi-Leh Wu, Corey Bufi, Chen-Chuan Kevin Chang, Yoshinori Hara, Reiko Ito, Yutaka Kimura, Kezuyuki Shimazu, and Yukiyoshi Saito. PowerBookmarks: a system for personalizable Web information organization, sharing, and management. In Delis et al. [2282], pages 565–567. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p565-li/p565-li.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/304182/p565-li/.

Rundensteiner:1999:SOB

[1408] E. A. Rundensteiner, K. Claypool, M. Li, L. Chen, Z. Zhang, C. Natarajan, J. Jin, S. De Lima, and S. Weiner. SERF: ODMG-based generic re-structuring facility. In Delis et al. [2282], pages 568–570. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p568-rundensteiner/p568-rundensteiner.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/304182/
p568-rundensteiner/.

Zhou:1999:TSD

[1409] Tong Zhou, Ling Liu, and Calton Pu. TAM: a system for dynamic transactional activity management. In Delis et al. [2282], pages 571–573. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p571-zhou/p571-zhou.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/304182/p571-zhou/.

Acharya:1999:AAQ

[1410] Swarup Acharya, Phillip B. Gibbons, Viswanath Poosala, and Sridhar Ramaswamy. The Aqua approximate query answering system. In Delis et al. [2282], pages 574–576. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p574-acharya/p574-acharya.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/304182/p574-acharya/

.

Brodsky:1999:CCO

[1411] Alexander Brodsky, Victor E. Segal, Jia Chen, and Paval A. Exarkhopoulo. The CCUBE constraint object-oriented database system. In Delis et al. [2282], pages 577–579. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p577-brodsky/p577-brodsky.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/304182/p577-brodsky/

.

Bonnet:1999:CJP

[1412] Phillippe Bonnet, Kyle Buza, Zhiyuan Chan, Victor Cheng, Randolph Chung, Takako Hickey, Ryan Kennedy, Daniel Mahashin, Tobias Mayr, Ivan Oprencak, Praveen Seshadri, and Hubert Siu. The Cornell Jaguar project: adding mobility to PREDATOR. In Delis et al. [2282], pages 580–581. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p580-bonnet/p580-bonnet.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p580-bonnet/.

Roussopoulos:1999:AMC

[1413] Nick Roussopoulos, Yannis Kotidis, and Yannis Sismanis. The active MultiSync controller of the cubetree storage organization. In Delis et al. [2282], pages 582–583. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p582-roussopoulos/p582-roussopoulos.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/304182/
p582-roussopoulos/.

Bohlen:1999:JDS

[1414] Michael Böhlen, Linas Bukauskas, and Curtis Dyreson. The Jungle database search engine. In Delis et al. [2282], pages 584–586. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p584-bohlen/p584-bohlen.pdf;   http:

                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/304182/p584-bohlen/.

Vossen:1999:WOO

[1415] Gottfried Vossen and Mathias Weske. The WASA2 object-oriented workflow management system. In Delis et al. [2282], pages 587–589. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p587-vossen/p587-vossen.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/304182/p587-vossen/.

Cruz:1999:UCI

[1416] Isabel F. Cruz and Kimberly M. James. A user-centered interface for querying distributed multimedia databases. In Delis et al. [2282], pages 590–592. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p590-cruz/p590-cruz.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/304182/p590-cruz/.

Bouguettaya:1999:WWD

[1417] Athman Bouguettaya, Boualem Benatallah, Lily Hendra, James Beard, Kevin Smith, and Mourad Quzzani. World Wide Database — integrating the Web, CORBA and databases. In Delis et al. [2282], pages 594–596. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/304182/
                                        p594-bouguettaya/p594-bouguettaya.
                                        pdf;   http://www.acm.org/pubs/

citations/proceedings/mod/304182/
p594-bouguettaya/.

Baru:1999:XBI

[1418] Chaitan Baru, Amarnath Gupta, Bertram Ludäscher, Richard Marciano, Yannis Papakonstantinou, Pavel Velikhov, and Vincent Chu. XML-based information mediation with MIX. In Delis et al. [2282], pages 597–599. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/304182/
p597-baru/p597-baru.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/304182/p597-baru/.

Han:2000:MFP

[1419] Jiawei Han, Jian Pei, and Yiwen Yin. Mining frequent patterns without candidate generation0. In Chen et al. [2284], pages 1–12. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p1-han/p1-han.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/342009/p1-han/.

Riedel:2000:DMO

[1420] Erik Riedel, Christos Faloutsos, Gregory R. Ganger, and David F. Nagle. Data mining on an OLTP system (nearly) for free. In Chen et al. [2284], pages 13–21. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p13-riedel/p13-riedel.pdf;   http:

//www.acm.org/pubs/citations/proceedings/
mod/342009/p13-riedel/.

Shenoy:2000:TCV

[1421] Pradeep Shenoy, Jayant R. Haritsa, S. Sundarshan, Gaurav Bhalotia, Mayank Bawa, and Devavrat Shah. Turbo-charging vertical mining of large databases. In Chen et al. [2284], pages 22–33. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p22-shenoy/p22-shenoy.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p22-shenoy/.

Lomet:2000:HSL

[1422] David B. Lomet. High speed on-line backup when using logical log operations. In Chen et al. [2284], pages 34–

45. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p34-lomet/p34-lomet.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p34-lomet/.

Labio:2000:ERI

[1423] Wilburt Juan Labio, Janet L. Wiener, Hector Garcia-Molina, and Vlad Gorelik. Efficient resumption of interrupted warehouse loads. In Chen et al. [2284], pages 46–57. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p46-labio/p46-labio.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p46-labio/.

Lakhamraju:2000:LRO

[1424] Mohana K. Lakhamraju, Rajeev Rastogi, S. Seshadri, and S. Sudarshan. On-line reorganization in object databases. In Chen et al. [2284], pages 58–69. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p58-lakhamraju/p58-lakhamraju.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/342009/p58-lakhamraju/

.

Aggarwal:2000:FGP

[1425] Charu C. Aggarwal and Philip S.

Yu. Finding generalized projected clusters in high dimensional spaces. In Chen et al. [2284], pages 70–81. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p70-aggarwal/p70-aggarwal.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/342009/p70-aggarwal/

.

Palmer:2000:DBS

[1426] Christopher R. Palmer and Christos Faloutsos. Density biased sampling: an improved method for data mining and clustering. In Chen et al. [2284], pages 82–92. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p82-palmer/p82-palmer.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p82-palmer/.

Breunig:2000:LID

[1427] Markus M. Breunig, Hans-Peter Kriegel, Raymond T. Ng, and Jörg Sander. LOF: identifying densitybased local outliers. In Chen et al. [2284], pages 93–104. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p93-breunig/p93-breunig.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p93-breunig/.

Zaharioudakis:2000:ACS

[1428] Markos Zaharioudakis, Roberta Cochrane, George Lapis, Hamid Pirahesh, and Monica Urata. Answering complex SQL queries using automatic summary tables. In Chen et al. [2284], pages 105–116. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p105-zaharioudakis/p105-zaharioudakis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/342009/
p105-zaharioudakis/.

Cho:2000:SDI

[1429] Junghoo Cho and Hector Garcia-Molina. Synchronizing a database to improve freshness. In Chen et al. [2284], pages 117–128. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p117-cho/p117-cho.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/342009/p117-cho/.

Salem:2000:HRJ

[1430] Kenneth Salem, Kevin Beyer, Bruce Lindsay, and Roberta Cochrane. How to roll a join: asynchronous incremental view maintenance. In Chen et al. [2284], pages 129–140. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p129-salem/p129-salem.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p129-salem/.

Christophides:2000:WQL

[1431] Vassilis Christophides, Sophie Cluet, and Jérôme Simèon. On wrapping query languages and efficient XML integration. In Chen et al. [2284], pages 141–152. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p141-christophides/p141-christophides.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/342009/
p141-christophides/.

Liefke:2000:XEC

[1432] Hartmut Liefke and Dan Suciu. XMill:

an efficient compressor for XML data. In Chen et al. [2284], pages 153–164. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p153-liefke/p153-liefke.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p153-liefke/.

Garofalakis:2000:XSE

[1433] Minos Garofalakis, Aristides Gionis, Rajeev Rastogi, S. Seshadri, and Kyuseok Shim. XTRACT: a system for extracting document type descriptors from XML documents. In Chen et al. [2284], pages 165–176. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p165-garofalakis/p165-garofalakis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/342009/
p165-garofalakis/.

Faloutsos:2000:SJS

[1434] Christos Faloutsos, Bernhard Seeger, Agma Traina, and Caetano Traina. Spatial join selectivity using power laws. In Chen et al. [2284], pages 177–188. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

QA1 .A87. URL http://www.acm.
org/pubs/articles/proceedings/mod/
342009/p177-faloutsos/p177-faloutsos.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/342009/
p177-faloutsos/.

Corral:2000:CPQ

[1435] Antonio Corral, Yannis Manolopoulos, Yannis Theodoridis, and Michael Vassilakopoulos. Closest pair queries in spatial databases. In Chen et al. [2284], pages 189–200. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p189-corral/p189-corral.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p189-corral/.

Korn:2000:ISB

[1436] Flip Korn and S. Muthukrishnan. Influence sets based on reverse nearest neighbor queries. In Chen et al. [2284], pages 201–212. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p201-korn/p201-korn.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p201-korn/.

Rodriguez-Martinez:2000:MSE

[1437] Manuel Rodríguez-Martínez and Nick Roussopoulos. MOCHA: a selfextensible database middleware system for distributed data sources. In Chen et al. [2284], pages 213–224. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p213-rodriguez-martinez/p213-rodriguez-
martinez.pdf;   http://www.acm.
org/pubs/citations/proceedings/
mod/342009/p213-rodriguez-martinez/

.

Lee:2000:TST

[1438] Mong Li Lee, Masaru Kitsuregawa, Beng Chin Ooi, Kian-Lee Tan, and Anirban Mondal. Towards selftuning data placement in parallel database systems. In Chen et al. [2284], pages 225–236. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p225-lee/p225-lee.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/342009/p225-lee/.

Litwin:2000:LHA

[1439] Witold Litwin and Thomas Schwarz. LH_{RS}^{∗}: a high-availability scalable distributed data structure using Reed Solomon Codes. In Chen et al. [2284], pages 237–248. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/342009/
                                        p237-litwin/p237-litwin.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/342009/p237-litwin/.

Roy:2000:EEA

[1440] Prasan Roy, S. Seshadri, S. Sudarshan, and Siddhesh Bhobe. Efficient and extensible algorithms for multi query optimization. In Chen et al. [2284], pages 249–260. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/342009/
                                        p249-roy/p249-roy.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/342009/p249-roy/.

Avnur:2000:ECA

[1441] Ron Avnur and Joseph M. Hellerstein. Eddies: continuously adaptive query processing. In Chen et al. [2284], pages 261–272. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/342009/
                                        p261-avnur/p261-avnur.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/342009/p261-avnur/.

Popa:2000:CTF

[1442] Lucian Popa, Alin Deutsch, Arnaud Sahuguet, and Val Tannen. A chase too far? In Chen et al. [2284], pages 273–284. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p273-popa/p273-popa.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p273-popa/.

Goldman:2000:WDP

[1443] Roy Goldman and Jennifer Widom.

WSQ/DSQ: a practical approach for combined querying of databases and the Web. In Chen et al. [2284], pages 285–296. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p285-goldman/p285-goldman.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/342009/p285-goldman/

.

Agrawal:2000:FEC

[1444] Rakesh Agrawal and Edward L. Wimmers. A framework for expressing and combining preferences. In Chen et al. [2284], pages 297–306. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p297-agrawal/p297-agrawal.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/342009/p297-agrawal/

.

Barclay:2000:MTS

[1445] Tom Barclay, Jim Gray, and Don Slutz. Microsoft TerraServer: a spatial data warehouse. In Chen et al. [2284], pages 307–318. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/342009/
                                        p307-barclay/p307-barclay.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/342009/p307-barclay/

.

Forlizzi:2000:DMD

[1446] Luca Forlizzi, Ralf Hartmut Güting, Enrico Nardelli, and Markus Schneider. A data model and data structures for moving objects databases. In Chen et al. [2284], pages 319–330. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/342009/
                                        p319-forlizzi/p319-forlizzi.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/342009/p319-forlizzi/

.

Saltenis:2000:IPC

[1447] Simonas Šaltenis, Christian S. Jensen, Scott T. Leutenegger, and Mario A. Lopez. Indexing the positions of continuously moving objects. In Chen et al. [2284], pages 331–342. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/342009/
                                        p331-saltenis/p331-saltenis.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/342009/p331-saltenis/

.

Shin:2000:AMS

[1448] Hyoseop Shin, Bongki Moon, and Sukho Lee. Adaptive multi-stage distance join processing. In Chen et al. [2284], pages 343–354. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p343-shin/p343-shin.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p343-shin/.

Cho:2000:FRW

[1449] Junghoo Cho, Narayanan Shivakumar, and Hector Garcia-Molina. Finding replicated Web collections. In Chen et al. [2284], pages 355–366. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p355-cho/p355-cho.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/342009/p355-cho/.

Labrinidis:2000:WM

[1450] Alexandros Labrinidis and Nick Roussopoulos. WebView materialization. In Chen et al. [2284], pages 367–378. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p367-labrinidis/p367-labrinidis.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/342009/
p367-labrinidis/.

Chen:2000:NSC

[1451] Jianjun Chen, David J. DeWitt, Feng Tian, and Yuan Wang. NiagaraCQ: a scalable continuous query system for Internet databases. In Chen et al. [2284], pages 379–390. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p379-chen/p379-chen.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p379-chen/.

Chang:2000:OTI

[1452] Yuan-Chi Chang, Lawrence Bergman, Vittorio Castelli, Chung-Sheng Li, Ming-Ling Lo, and John R. Smith. The onion technique: indexing for linear optimization queries. In Chen et al. [2284], pages 391–402. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p391-chang/p391-chang.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p391-chang/.

Jagadish:2000:EMD

[1453] H. V. Jagadish, Nick Koudas, and Divesh Srivastava. On effective multidimensional indexing for strings. In Chen et al. [2284], pages 403–414. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p403-jagadish/p403-jagadish.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/342009/p403-jagadish/

.

Oh:2000:ECE

[1454] JungHwan Oh and Kien A. Hua. Efficient and cost-effective techniques for browsing and indexing large video databases. In Chen et al. [2284], pages 415–426. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p415-oh/p415-oh.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/342009/p415-oh/.

Ramaswamy:2000:EAM

[1455] Sridhar Ramaswamy, Rajeev Rastogi, and Kyuseok Shim. Efficient algorithms for mining outliers from large data sets. In Chen et al. [2284], pages 427–438. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p427-ramaswamy/p427-ramaswamy.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/342009/p427-ramaswamy/

.

Agrawal:2000:PPD

[1456] Rakesh Agrawal and Ramakrishnan Srikant. Privacy-preserving data mining. In Chen et al. [2284], pages 439–450. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p439-agrawal/p439-agrawal.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/342009/p439-agrawal/

.

Szalay:2000:DMM

[1457] Alexander S. Szalay, Peter Z. Kunszt, Ani Thakar, Jim Gray, Don Slutz, and Robert J. Brunner. Designing and mining multi-terabyte astronomy archives: the Sloan Digital Sky Survey. In Chen et al. [2284], pages 451–462. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/342009/
                                        p451-szalay/p451-szalay.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/342009/p451-szalay/.

Gunopulos:2000:AMD

[1458] Dimitrios Gunopulos, George Kollios, Vassilis J. Tsotras, and Carlotta Domeniconi. Approximating multi-dimensional aggregate range queries over real attributes. In Chen et al. [2284], pages 463–474. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/342009/
                                        p463-gunopulos/p463-gunopulos.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/342009/p463-gunopulos/

.

Rao:2000:MBT

[1459] Jun Rao and Kenneth A. Ross. Making B^{+}-trees cache conscious in main memory. In Chen et al. [2284], pages 475–486. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/342009/
                                        p475-rao/p475-rao.pdf;   http://
                                        www.acm.org/pubs/citations/proceedings/
                                        mod/342009/p475-rao/.

Acharya:2000:CSA

[1460] Swarup Acharya, Phillip B. Gibbons, and Viswanath Poosala. Congressional samples for approximate answering of group-by queries. In Chen et al. [2284], pages 487–498. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/citations/proceedings/mod/
342009/p487-acharya/.

Waas:2000:CES

[1461] Florian Waas and César Galindo-Legaria. Counting, enumerating, and sampling of execution plans in a cost-based query optimizer. In Chen et al. [2284], pages 499–509. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p499-waas/p499-waas.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p499-waas/.

Wattez:2000:BQT

[1462] Fanny Wattez, Sophie Cluet, Véronique Benzaken, Guy Ferran, and Christian Fiegel. Benchmarking queries over trees: learning the hard truth the hard way. In Chen et al. [2284], pages 510–

511. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

QA1 .A87. URL http://www.acm.
org/pubs/citations/proceedings/
mod/342009/p510-wattez/.

Lehner:2000:MCA

[1463] Wolfgang Lehner, Richard Sidle, Hamid Pirahesh, and Roberta Wolfgang Cochrane. Maintenance of cube automatic summary tables. In Chen et al. [2284], pages 512–513. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/

                                        p512-lehner/p512-lehner.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/342009/p512-lehner/.

Huber:2000:CAD

[1464] Val Huber. Challenges in automating declarative business rules to enable rapid business response. In Chen et al. [2284], page 514. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/342009/
                                        p514-huber/p514-huber.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/342009/p514-huber/.

Ross:2000:EBR

[1465] Ronald G. Ross. Expressing business rules. In Chen et al. [2284], pages 515–516. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

                                        QA1 .A87. URL http://www.acm.
                                        org/pubs/articles/proceedings/mod/
                                        342009/p515-ross/p515-ross.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/342009/p515-ross/

.

Kintzer:2000:GBP

[1466] Eric Kintzer. Going beyond personalization: rule engines at work. In Chen et al. [2284], page 517. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/342009/
                                        p517-kintzer/p517-kintzer.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/mod/342009/p517-kintzer/

.

Hsiao:2000:DTR

[1467] Hui-I. Hsiao and Inderpal Narang.

DLFM: a transactional resource manager. In Chen et al. [2284], pages 518–

528. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

QA1 .A87. URL http://www.acm.
org/pubs/articles/proceedings/mod/
342009/p518-hsiao/p518-hsiao.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/342009/p518-hsiao/

.

Ponnekanti:2000:OIR

[1468] Nagavamsi Ponnekanti and Hanuma Kodavalla. Online index rebuild. In Chen et al. [2284], pages 529–538. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p529-ponnekanti/p529-ponnekanti.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/342009/
p529-ponnekanti/.

Annamalai:2000:IIO

[1469] Melliyal Annamalai, Rajiv Chopra, Samuel DeFazio, and Susan Mavris. Indexing images in Oracle8i. In Chen et al. [2284], pages 539–547. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p539-annamalai/p539-annamalai.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/342009/p539-annamalai/

.

Weininger:2000:HVL

[1470] Andreas Weininger. Handling very large databases with Informix extended parallel server. In Chen et al. [2284], pages 548–549. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p548-weininger/p548-weininger.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/342009/p548-weininger/

.

Chen:2000:ITW

[1471] Chung-Min Chen, Munir Cochinwala, Claudio Petrone, Marc Pucci, Sunil Samtani, and Patrizia Santa. Internet traffic warehouse. In Chen et al. [2284], pages 550–558. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p550-chen/p550-chen.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p550-chen/.

Ordonez:2000:SFC

[1472] Carlos Ordonez and Paul Cereghini. SQLEM: fast clustering in SQL using the EM algorithm. In Chen et al. [2284], pages 559–570. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p559-ordonez/p559-ordonez.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/342009/p559-ordonez/

.

Jhingran:2000:ARC

[1473] Anant Jhingran. Anatomy of a real E-commerce system. In Chen et al. [2284], pages 571–572. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/citations/proceedings/mod/
342009/p571-jhingran/.

Ramakrishnan:2000:BID

[1474] Raghu Ramakrishnan. From browsing to interacting: DBMS support for responsive websites. In Chen et al. [2284], page 573. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p573-ramakrishnan/p573-ramakrishnan.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/342009/
p573-ramakrishnan/.

Hellerstein:2000:IRP

[1475] Joseph M. Hellerstein. Index research (panel session): forest or trees? In Chen et al. [2284], page 574. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p574-hellerstein/p574-hellerstein.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/342009/
p574-hellerstein/.

Nori:2000:AAP

[1476] Anil K. Nori. Application architecture (panel session): 2Tier or 3Tier? What is DBMS’s role? In Chen et al. [2284], page 575. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p575-nori/p575-nori.pdf;   http:

//www.acm.org/pubs/citations/proceedings/
mod/342009/p575-nori/.

Carey:2000:XDP

[1477] Michael J. Carey, Jennifer Widom, Adam Bosworth, Bruce Lindsay, Michael Stonebraker, and Dan Suciu. Of XML and databases (panel session): where’s the beef? In Chen et al. [2284], page 576. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p576-carey/p576-carey.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p576-carey/.

Bratsberg:2000:DUH

[1478] Svein Erik Bratsberg and Øystein Torbjørnsen. Designing an ultra highly available DBMS (tutorial session). In Chen et al. [2284], page 577. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p577-bratsberg/p577-bratsberg.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/342009/p577-bratsberg/

.

Gal:2000:DME

[1479] Avigdor Gal. Data management in eCommerce (tutorial session): the good, the bad, and the ugly. In Chen et al. [2284], page 578. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p578-gal/p578-gal.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/342009/p578-gal/.

Blakeley:2000:DAT

[1480] José A. Blakeley and Anand Deshpande. Data access (tutorial session). In Chen et al. [2284], page 579. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p579-blakeley/p579-blakeley.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/342009/p579-blakeley/

.

Shukla:2000:LDS

[1481] Shridhar Shukla and Anand Deshpande. LDAP directory services- just another database application? (tutorial session). In Chen et al. [2284], page 580. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

QA1 .A87. URL http://www.acm.
org/pubs/articles/proceedings/mod/
342009/p580-shukla/p580-shukla.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/342009/
p580-shukla/.

Wolfson:2000:RIM

[1482] Ouri Wolfson. Research issues in moving objects databases (tutorial session). In Chen et al. [2284], page 581. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p581-wolfson/p581-wolfson.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/342009/p581-wolfson/

.

Ives:2000:SOD

[1483] Zachary Ives, Alon Levy, Jayant Madhavan, Rachel Pottinger, Stefan Saroiu, Igor Tatarinov, Shiori Betzler, Qiong Chen, Ewa Jaslikowska, Jing Su, and Wai Tak Theodora Yeung. Self-organizing data sharing communities with SAGRES. In Chen et al. [2284], page 582. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p582-ives/p582-ives.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p582-ives/.

Fegaras:2000:DOB

[1484] Leonidas Fegaras, Chandrasekhar Srinivasan, Arvind Rajendran, and David Maier. -DB: an ODMGbased object-oriented DBMS. In Chen et al. [2284], page 583. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p583-fegaras/p583-fegaras.pdf;
http://www.acm.org/pubs/citations/
proceedings/mod/342009/p583-fegaras/

.

Wang:2000:ASE

[1485] Jason T. L. Wang, Xiong Wang, Dennis Shasha, Bruce A. Shapiro, Kaizhong Zhang, Qicheng Ma, and Zasha Weinberg. An approximate search engine for structural databases. In Chen et al. [2284], page 584. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p584-wang/p584-wang.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p584-wang/.

Rundensteiner:2000:SWW

[1486] Elke A. Rundensteiner, Kajal T. Claypool, Li Chen, Hong Su, and Keiji Oenoki. SERFing the Web: Web site management made easy. In Chen et al. [2284], page 585. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p585-rundensteiner/p585-rundensteiner.
pdf;   http://www.acm.org/pubs/
citations/proceedings/mod/342009/
p585-rundensteiner/.

Merialdo:2000:HMB

[1487] Paolo Merialdo, Paolo Atzeni, Marco Magnante, Giansalvatore Mecca, and Marco Pecorone. HOMER: a modelbased CASE tool for data-intensive Web sites. In Chen et al. [2284], page

586. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN

QA1 .A87. URL http://www.acm.
org/pubs/citations/proceedings/
mod/342009/p586-merialdo/.

Chen:2000:FLB

[1488] Songting Chen, Yanlei Diao, Hongjun Lu, and Zengping Tian. FACT: a learning based Web query processing system. In Chen et al. [2284], page 587. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p587-chen/p587-chen.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p587-chen/.

vandenBercken:2000:JXP

[1489] Jochen van den Bercken, Jens-Peter Dittrich, and Bernhard Seeger. javax.XXL: a prototype for a library of query processing algorithms. In Chen et al. [2284], page 588. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/citations/proceedings/mod/
342009/p588-van_den_bercken/.

Sarawagi:2000:III

[1490] Sunita Sarawagi and Gayatri Sathe. i^{3}: intelligent, interactive investigation of OLAP data cubes. In Chen et al. [2284], page 589. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p589-sarawagi/p589-sarawagi.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/342009/p589-sarawagi/

.

Galhardas:2000:AED

[1491] Helena Galhardas, Daniela Florescu, Dennis Shasha, and Eric Simon. AJAX: an extensible data cleaning tool. In Chen et al. [2284], page 590. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p590-galhardas/p590-galhardas.pdf;
 http://www.acm.org/pubs/citations/
proceedings/mod/342009/p590-galhardas/

.

Jarke:2000:CBD

[1492] M. Jarke, C. Quix, D. Calvanese, M. Lenzerini, E. Franconi, S. Ligoudistianos, P. Vassiliadis, and Y. Vassiliou. Concept based design of data warehouses: the DWQ demonstrators. In Chen et al. [2284], page 591.

ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p591-jarke/p591-jarke.pdf;   http:
//www.acm.org/pubs/citations/proceedings/
mod/342009/p591-jarke/.

Pei:2000:TDM

[1493] Jian Pei, Runying Mao, Kan Hu, and Hua Zhu. Towards data mining benchmarking: a test bed for performance study of frequent pattern mining. In Chen et al. [2284], page 592. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p592-pei/p592-pei.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/342009/p592-pei/.

Hsu:2000:IMI

[1494] Wynne Hsu, Mong Li Lee, and Kheng Guan Goh. Image mining in IRIS: integrated retinal information system. In Chen et al. [2284], page 593. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p593-hsu/p593-hsu.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/342009/p593-hsu/.

Rodriguez-Martinez:2000:MDM

[1495] Manuel Rodríguez-Martínez, Nick Roussopoulos, John M. McGann, Stephen Kelley, Vadim Katz, Zhexuan Song, and Joseph JáJá. MOCHA: a database middleware system featuring automatic deployment of applicationspecific functionality. In Chen et al. [2284], page 594. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/citations/proceedings/mod/
                                        342009/p594-rodriguez-martinez/.

Gillmann:2000:GDA

[1496] Michael Gillmann, Jeanine Weissenfels, German Shegalov, Wolfgang Wonner, and Gerhard Weikum. A goal-driven auto-configuration tool for the distributed workflow management system mentorlite. In Chen et al. [2284], page 595. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/342009/
                                        p595-gillmann/p595-gillmann.pdf;
                                         http://www.acm.org/pubs/citations/
                                        proceedings/mod/342009/p595-gillmann/

.

Yang:2000:TTE

[1497] Jun Yang, Huacheng C. Ying, and Jennifer Widom. TIP: a temporal extension to Informix. In Chen et al. [2284], page 596. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/342009/
                                        p596-yang/p596-yang.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/342009/p596-yang/.

Liu:2000:ATA

[1498] Ling Liu, Calton Pu, David Buttler, Wei Han, Henrique Paques, and Wei Tang. AQR-toolkit: an adaptive query routing middleware for distributed data intensive systems.

In Chen et al. [2284], page 597. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p597-liu/p597-liu.pdf;   http://
www.acm.org/pubs/citations/proceedings/
mod/342009/p597-liu/.

Li:2000:SPC

[1499] Chung-Sheng Li, Lawrence D. Bergman, Yuan-Chi Chang, Vittorio Castelli, and John R. Smith. SPIRE: a progressive content-based spatial image retrieval engine. In Chen et al. [2284], page 598. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/articles/proceedings/mod/342009/
p598-li/p598-li.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/342009/p598-li/.

Goh:2000:IRP

[1500] Chong Leng Goh, Beng Chin Ooi, Stephane Bressan, and Kian-Lee Tan. Integrating replacement policies in StorM: an extensible approach. In Chen et al. [2284], page 599. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.acm.org/
pubs/citations/proceedings/mod/
342009/p599-goh/.

Oria:2000:DDI

[1501] Vincent Oria, M. Tamer Özsu, Paul J.

Iglinski, Shu Lin, and Bin Yao. DIS-IMA: a distributed and interoperable image database system. In Chen et al. [2284], page 600. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/342009/
                                        p600-oria/p600-oria.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/342009/p600-oria/.

Revesz:2000:MGC

[1502] Peter Revesz, Rui Chen, Pradip Kanjamala, Yiming Li, Yuguo Liu, and Yonghui Wang. The MLPQ/ GIS constraint database system. In Chen et al. [2284], page 601. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

                                        .A87. URL http://www.acm.org/
                                        pubs/articles/proceedings/mod/342009/
                                        p601-revesz/p601-revesz.pdf;   http:
                                        //www.acm.org/pubs/citations/proceedings/
                                        mod/342009/p601-revesz/.

Han:2001:ECI

[1503] Jiawei Han, Jian Pei, Guozhu Dong, and Ke Wang. Efficient computation of Iceberg cubes with complex measures. In Sellis and Mehrotra [2286], pages 1–12. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Vianu:2001:WOC

[1504] Victor Vianu. A Web odyssey: from Codd to XML. In ACM [2285], pages 1–15. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Gehrke:2001:CCA

[1505] Johannes Gehrke, Flip Korn, and Divesh Srivastava. On computing correlated aggregates over continual data streams. In Sellis and Mehrotra [2286], pages 13–24. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic).

LCCN ???? ACM order number 472010.

Rajaraman:2001:QWU

[1506] Anand Rajaraman and Jeffrey D. Ullmann. Querying Websites using compact skeletons. In ACM [2285], pages 16–27. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Ng:2001:ICC

[1507] Raymond T. Ng, Alan Wagner, and Yu Yin. Iceberg-cube computation with PC clusters. In Sellis and Mehrotra [2286], pages 25–36. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Jim:2001:DDQ

[1508] Trevor Jim and Dan Suciu. Dynamically distributed query evaluation. In ACM [2285], pages 28–39. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Aggarwal:2001:ODH

[1509] Charu C. Aggarwal and Philip S. Yu.

Outlier detection for high dimensional data. In Sellis and Mehrotra [2286], pages 37–46. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Kanza:2001:FQS

[1510] Yaron Kanza and Yehoshua Sagiv.

Flexible queries over semistructured data. In ACM [2285], pages 40–51. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Rinfret:2001:BSI

[1511] Denis Rinfret, Patrick O’Neil, and Elizabeth O’Neil. Bit-sliced index arithmetic. In Sellis and Mehrotra [2286], pages 47–57. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Papadimitriou:2001:MQO

[1512] Christos H. Papadimitriou and Mihalis Yannakakis. Multiobjective query optimization. In ACM [2285], pages 52–59. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Greenwald:2001:SEO

[1513] Michael Greenwald and Sanjeev Khanna. Space-efficient online computation of quantile summaries. In Sellis and Mehrotra [2286], pages 58–66. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Dalvi:2001:PMQ

[1514] Nilesh N. Dalvi, Sumit K. Sanghai, Prasan Roy, and S. Sudarshan. Pipelining in multi-query optimization. In ACM [2285], pages 59–70. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Ipeirotis:2001:PCC

[1515] Panagiotis G. Ipeirotis, Luis Gravano, and Mehran Sahami. Probe, count, and classify: categorizing hidden web databases. In Sellis and Mehrotra [2286], pages 67–78. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Sadri:2001:OSQ

[1516] Reza Sadri, Carlo Zaniolo, Amir Zarkesh, and Jafar Adibi. Optimization of sequence queries in database systems. In ACM [2285], pages 71–81. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Breunig:2001:DBQ

[1517] Markus M. Breunig, Hans-Peter Kriegel, Peer Kröger, and Jörg Sander. Data bubbles: quality preserving performance boosting for hierarchical clustering. In Sellis and Mehrotra [2286], pages 79–90. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Grohe:2001:PCD

[1518] Martin Grohe. The parameterized complexity of database queries. In ACM [2285], pages 82–92. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Joshi:2001:MNH

[1519] Mahesh V. Joshi, Ramesh C. Agarwal, and Vipin Kumar. Mining needle in a haystack: classifying rare classes via two-phase rule induction. In Sellis and Mehrotra [2286], pages 91–102. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Larsen:2001:RMW

[1520] Kim S. Larsen. Relaxed multi-way trees with group updates. In ACM [2285], pages 93–101. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Fagin:2001:OAA

[1521] Ronald Fagin, Amnon Lotem, and Moni Naor. Optimal aggregation algorithms for middleware. In ACM [2285], pages 102–113. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Fernandez:2001:EEX

[1522] Mary Fernandez, Atsuyuki Morishima, and Dan Suciu. Efficient evaluation of XML middle-ware queries. In Sellis and Mehrotra [2286], pages 103–114. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Fan:2001:XIC

[1523] Wenfei Fan and Leonid Libkin. On XML integrity constraints in the presence of DTDs. In ACM [2285], pages 114–125. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Fabret:2001:FAI

[1524] Françoise Fabret, H. Arno Jacobsen, François Llirbat, João Pereira, Kenneth A. Ross, and Dennis Shasha. Filtering algorithms and implementation for very fast publish/subscribe systems.

In Sellis and Mehrotra [2286], pages 115–126. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Murata:2001:EPE

[1525] Makoto Murata. Extended path expressions of XML. In ACM [2285], pages 126–137. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Slivinskas:2001:AQO

[1526] Giedrius Slivinskas, Christian S.

Jensen, and Richard T. Snodgrass. Adaptable query optimization and evaluation in temporal middleware. In Sellis and Mehrotra [2286], pages 127– 138. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Alon:2001:XDV

[1527] Noga Alon, Tova Milo, Frank Neven, Dan Suciu, and Victor Vianu. XML with data values: typechecking revisited. In ACM [2285], pages 138–149. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Kim:2001:OMI

[1528] Kihong Kim, Sang K. Cha, and Keunjoo Kwon. Optimizing multidimensional index trees for main memory access. In Sellis and Mehrotra [2286], pages 139–150. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Abiteboul:2001:RQX

[1529] Serge Abiteboul, Luc Segoufin, and Victor Vianu. Representing and querying XML with incomplete information. In ACM [2285], pages 150–161. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Keogh:2001:LAD

[1530] Eamonn Keogh, Kaushik Chakrabarti, Michael Pazzani, and Sharad Mehrotra. Locally adaptive dimensionality reduction for indexing large time series databases. In Sellis and Mehrotra [2286], pages 151–162. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Mendelzon:2001:QPS

[1531] Alberto O. Mendelzon and George A. Mihaila. Querying partially sound and complete data sources. In ACM [2285], pages 162–170. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Bohannon:2001:MMI

[1532] Philip Bohannon, Peter Mcllroy, and Rajeev Rastogi. Main-memory index structures with fixed-size partial keys. In Sellis and Mehrotra [2286], pages 163–174. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Khanna:2001:CFU

[1533] Sanjeev Khanna and Wang-Chiew Tan. On computing functions with uncertainty. In ACM [2285], pages 171–182.

ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Borkar:2001:AST

[1534] Vinayak Borkar, Kaustubh Deshmukh, and Sunita Sarawagi. Automatic segmentation of text into structured records. In Sellis and Mehrotra [2286], pages 175–186. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Benedikt:2001:SOQ

[1535] Michael Benedikt, Leonid Libkin, Thomas Schwentick, and Luc Segoufin. String operations in query languages. In ACM [2285], pages 183–194. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Yu:2001:EEM

[1536] Clement Yu, Weiyi Meng, Wensheng Wu, and King-Lup Liu. Efficient and effective metasearch for text databases incorporating linkages among documents. In Sellis and Mehrotra [2286], pages 187–198. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Gottlob:2001:RMG

[1537] Georg Gottlob, Nicola Leone, and Francesco Scarcello. Robbers, marshals, and guards: game theoretic and logical characterizations of hypertree width. In ACM [2285], pages 195–206. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Deshpande:2001:IGD

[1538] Amol Deshpande, Minos Garofalakis, and Rajeev Rastogi. Independence is good: dependency-based histogram synopses for high-dimensional data. In Sellis and Mehrotra [2286], pages 199– 210. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Cai:2001:CJP

[1539] Jin-Yi Cai, Venkatesan T. Chakaravarthy, Raghav Kaushik, and Jeffrey F. Naughton. On the complexity of join predicates. In ACM [2285], pages 207–214. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Bruno:2001:SMW

[1540] Nicolas Bruno, Surajit Chaudhuri, and Luis Gravano. STHoles: a multidimensional workload-aware histogram. In Sellis and Mehrotra [2286], pages 211– 222. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Cohen:2001:EAA

[1541] Sara Cohen, Werner Nutt, and Yehoshua Sagiv. Equivalences among aggregate queries with negation. In ACM [2285], pages 215–226. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Jagadish:2001:GOH

[1542] H. V. Jagadish, Hui Jin, Beng Chin Ooi, and Kian-Lee Tan. Global optimization of histograms. In Sellis and Mehrotra [2286], pages 223–234. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Gilbert:2001:OAC

[1543] Anna C. Gilbert, Yannis Kotidis, S. Muthukrishnan, and Marin J. Strauss. Optimal and approximate computation of summary statistics for range aggregates. In ACM [2285], pages 227–236. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Chen:2001:IIP

[1544] Shimin Chen, Phillip B. Gibbons, and Todd C. Mowry. Improving index performance through prefetching. In Sellis and Mehrotra [2286], pages 235– 246. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Zhang:2001:ECT

[1545] Donhui Zhang, Alexander Markowetz, Vassilis Tsotras, Dimitrios Gunopulos, and Bernhard Seeger. Efficient computation of temporal aggregates with range predicates. In ACM [2285], pages 237–245. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Leighton:2001:CDC

[1546] Tom Leighton. The challenges of delivering content on the Internet. In ACM [2285], page 246. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Agrawal:2001:DQP

[1547] Dakshi Agrawal and Charu C. Aggarwal. On the design and quantification of privacy preserving data mining algorithms. In ACM [2285], pages 247–255. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Gionis:2001:ETS

[1548] Aristides Gionis, Dimitrios Gunopulos, and Nick Koudas. Efficient and tumble similar set retrieval. In Sellis and Mehrotra [2286], pages 247–258. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Aggarwal:2001:EDR

[1549] Charu C. Aggarwal. On the effects of dimensionality reduction on high dimensional similarity search. In ACM [2285], pages 256–266. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Hristidis:2001:PSE

[1550] Vagelis Hristidis, Nick Koudas, and Yannis Papakonstantinou. PREFER: a system for the efficient execution of multi-parametric ranked queries. In Sellis and Mehrotra [2286], pages 259– 270. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Bykowski:2001:CRF

[1551] Artur Bykowski and Christophe Rigotti. A condensed representation to find frequent patterns. In ACM [2285], pages 267–273. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Chen:2001:QOC

[1552] Zhiyuan Chen, Johannes Gehrke, and Flip Korn. Query optimization in compressed database systems. In Sellis and Mehrotra [2286], pages 271–282. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Achlioptas:2001:DFR

[1553] Dimitris Achlioptas. Database-friendly random projections. In ACM [2285], pages 274–281. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Ferragina:2001:TDS

[1554] Paolo Ferragina, Nick Koudas, Divesh Srivastava, and S. Muthukrishnan. Two-dimensional substring indexing. In ACM [2285], pages 282–288. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Babu:2001:SMB

[1555] Shivnath Babu, Minos Garofalakis, and Rajeev Rastogi. SPARTAN: a modelbased semantic compression system for massive data tables. In Sellis and Mehrotra [2286], pages 283–294. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Schuldt:2001:PLP

[1556] Heiko Schuldt. Process locking: a protocol based on ordered shared locks for the execution of transactional processes. In ACM [2285], pages 289–300. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Chaudhuri:2001:ROB

[1557] Surajit Chaudhuri, Gautam Das, and Vivek Narasayya. A robust, optimization-based approach for approximate answering of aggregate queries. In Sellis and Mehrotra [2286], pages 295–306. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Mistry:2001:MVS

[1558] Hoshi Mistry, Prasan Roy, S. Sudarshan, and Krithi Ramamritham. Materialized view selection and maintenance using multi-query optimization. In Sellis and Mehrotra [2286], pages 307–318. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Afrati:2001:GEP

[1559] Foto N. Afrati, Chen Li, and Jeffrey D. Ullman. Generating efficient plans for queries using views. In Sellis and Mehrotra [2286], pages 319–330. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Goldstein:2001:OQU

[1560] Jonathan Goldstein and Per-Åke Larson. Optimizing queries using materialized views: a practical, scalable solution. In Sellis and Mehrotra [2286], pages 331–342. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic).

LCCN ???? ACM order number 472010.

Lee:2001:DBA

[1561] Sang-Ho Lee, Kyu-Young Whang, Yang-Sae Moon, and Il-Yeol Song. Dynamic buffer allocation in video-on-demand systems. In Sellis and Mehrotra [2286], pages 343–354. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Olston:2001:APS

[1562] Chris Olston, Boon Thau Loo, and Jennifer Widom. Adaptive precision setting for cached approximate values. In Sellis and Mehrotra [2286], pages 355– 366. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Kalnis:2001:PSA

[1563] Panos Kalnis and Dimitris Papadias.

Proxy-server architectures for OLAP. In Sellis and Mehrotra [2286], pages 367–378. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Bohm:2001:EGO

[1564] Christian Böhm, Bernhard Braunmüller, Florian Krebs, and Hans-Peter Kriegel. Epsilon grid order: an algorithm for the similarity join on massive highdimensional data. In Sellis and Mehrotra [2286], pages 379–388. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Lang:2001:MHD

[1565] Christian A. Lang and Ambuj K.

Singh. Modeling high-dimensional index structures using sampling. In Sellis and Mehrotra [2286], pages 389– 400. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Lazaridis:2001:PAA

[1566] Iosif Lazaridis and Sharad Mehrotra. Progressive approximate aggregate queries with a multi-resolution tree structure. In Sellis and Mehrotra [2286], pages 401–412. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Tatarinov:2001:UX

[1567] Igor Tatarinov, Zachary G. Ives, Alon Y. Halevy, and Daniel S. Weld. Updating XML. In Sellis and Mehrotra [2286], pages 413–424. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Zhang:2001:SCQ

[1568] Chun Zhang, Jeffrey Naughton, David DeWitt, Qiong Luo, and Guy Lohman. On supporting containment queries in relational database management systems. In Sellis and Mehrotra [2286], pages 425–436. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Nguyen:2001:MXD

[1569] Benjamin Nguyen, Serge Abiteboul, Grégory Cobena, and Mihaí Preda. Monitoring XML data on the Web. In Sellis and Mehrotra [2286], pages 437– 448. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Wu:2001:AGR

[1570] Yi-Leh Wu, Divyakant Agrawal, and Amr El Abbadi. Applying the golden rule of sampling for query estimation. In Sellis and Mehrotra [2286], pages 449–460. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Getoor:2001:SEU

[1571] Lise Getoor, Benjamin Taskar, and Daphne Koller. Selectivity estimation using probabilistic models. In Sellis and Mehrotra [2286], pages 461–472. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Schuster:2001:CED

[1572] Assaf Schuster and Ran Wolff.

Communication-efficient distributed mining of association rules. In Sellis and Mehrotra [2286], pages 473–484. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Yan:2001:DDU

[1573] Ling Ling Yan, Renée J. Miller, Laura M. Haas, and Ronald Fagin. Data-driven understanding and refinement of schema mappings. In Sellis and Mehrotra [2286], pages 485–496. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Amer-Yahia:2001:MTP

[1574] Sihem Amer-Yahia, SungRan Cho, Laks V. S. Lakshmanan, and Divesh Srivastava. Minimization of tree pattern queries. In Sellis and Mehrotra [2286], pages 497–508. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Doan:2001:RSD

[1575] AnHai Doan, Pedro Domingos, and Alon Y. Halevy. Reconciling schemas of disparate data sources: a machinelearning approach. In Sellis and Mehrotra [2286], pages 509–520. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Carino:2001:SMN

[1576] Felipe Cariño, Jr., Pekka Kostamaa, Art Kaufmann, and John Burgess. StorHouse metanoia — new applications for database, storage & data warehousing. In Sellis and Mehrotra [2286], pages 521–531. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Candan:2001:EDC

[1577] K. Selçuk Candan, Wen-Syan Li, Qiong Luo, Wang-Pin Hsiung, and Divyakant Agrawal. Enabling dynamic content caching for database-driven web sites. In Sellis and Mehrotra [2286], pages 532–543. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Navas:2001:NDD

[1578] Julio C. Navas and Michael Wynblatt. The network is the database: data management for highly distributed systems. In Sellis and Mehrotra [2286], pages 544–551. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Stonebraker:2001:CIB

[1579] Michael Stonebraker and Joseph M.

Hellerstein. Content integration for ebusiness. In Sellis and Mehrotra [2286], pages 552–560. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Maguire:2001:CMW

[1580] Thomas Maguire. Catalog management in websphere commerce suite. In Sellis and Mehrotra [2286], page 561. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Nazeri:2001:EMA

[1581] Zohreh Nazeri, Eric Bloedorn, and Paul Ostwald. Experiences in mining aviation safety data. In Sellis and Mehrotra [2286], pages 562–566. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Draper:2001:NIE

[1582] Denise Draper, Alon Y. Halevy, and Daniel S. Weld. The nimble integration engine. In Sellis and Mehrotra [2286], pages 567–568. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Meseck:2001:DML

[1583] Reed M. Meseck. Data management:

lasting impact on wild, wild, web. In Sellis and Mehrotra [2286], pages 569– 570. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Galindo-Legaria:2001:OOS

[1584] César Galindo-Legaria and Milind Joshi. Orthogonal optimization of subqueries and aggregation. In Sellis and Mehrotra [2286], pages 571–581. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Godfrey:2001:ECL

[1585] Parke Godfrey, Jarek Gryz, and Calisto Zuzarte. Exploiting constraint-like data characterizations in query optimization. In Sellis and Mehrotra [2286], pages 582–592. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Lahiri:2001:FSQ

[1586] Tirthankar Lahiri, Amit Ganesh, Ron Weiss, and Ashok Joshi. Fast-start: quick fault recovery in oracle. In Sellis and Mehrotra [2286], pages 593– 598. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Deolasee:2001:DDD

[1587] Pavan Deolasee, Amol Katkar, Ankur Panchbudhe, Krithi Ramamritham, and Prashant Shenoy. Dissemination of dynamic data. In Sellis and Mehrotra [2286], page 599. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Kiessling:2001:CYS

[1588] Werner Kießling, Stefan Holland, Stefan Fischer, and Thorsten Ehm. COSIMA—your smart, speaking Esalesperson. In Sellis and Mehrotra [2286], page 600. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Shou:2001:RRT

[1589] L. Shou, C. H. Chionh, Z. Huang, Y. Ruan, and Kian-Lee Tan. RE-VIEW: a real-time virtual walkthrough system. In Sellis and Mehrotra [2286], page 601. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Sahuguet:2001:KMT

[1590] Arnaud Sahuguet. Kweelt: more than just “yet another framework to query XML!”. In Sellis and Mehrotra [2286], page 602. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Chawathe:2001:VES

[1591] Sudarshan S. Chawathe, Thomas Baby, and Jihwang Yoo. VQBD: exploring semistructured data. In Sellis and Mehrotra [2286], page 603. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Buttler:2001:OMS

[1592] David Buttler, Ling Liu, Calton Pu, Henrique Paques, Wei Han, and Wei Tang. OminiSearch: a method for searching dynamic content on the Web. In Sellis and Mehrotra [2286], page 604. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Bertino:2001:SXD

[1593] Elisa Bertino, Silvana Castano, and Elena Ferrari. Securing XML documents: the author-X project demonstration. In Sellis and Mehrotra [2286], page 605. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Claypool:2001:SSS

[1594] Kajal T. Claypool, Elke A. Rundensteiner, Xin Zhang, Su Hong, Harumi Kuno, Wang chien Lee, and Gail Mitchell. Sangam — a solution to support multiple data models, their mappings and maintenance. In Sellis and Mehrotra [2286], page 606. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Hernandez:2001:CSA

[1595] Mauricio A. Hernández, Renée J. Miller, and Laura M. Haas. Clio: a semi-automatic tool for schema mapping. In Sellis and Mehrotra [2286], page 607. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Agrawal:2001:MVI

[1596] Sanjay Agrawal, Surajit Chaudhuri, and Vivek Narasayya. Materialized view and index selection tool for Microsoft SQL Server 2000. In Sellis and Mehrotra [2286], page 608. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Catarci:2001:PDS

[1597] Tiziana Catarci and Giuseppe Santucci. The prototype of the DARE system. In Sellis and Mehrotra [2286], page 609. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Adii:2001:MBP

[1598] Asaf Adii, David Botzer, Opher Etzion, and Tali Yatzkar-Haham. Monitoring business processes through event correlation based on dependency model. In Sellis and Mehrotra [2286], page 610. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Shah:2001:FTL

[1599] Mehul A. Shah and Sirish Chandrasekaran. Fault-tolerant, loadbalancing queries in telegraph. In Sellis and Mehrotra [2286], page 611. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Agichtein:2001:SPS

[1600] Eugene Agichtein, Luis Gravano, Jeff Pavel, Viktoriya Sokolova, and Aleksandr Voskoboynik. Snowball: a prototype system for extracting relations from large text collections. In Sellis and Mehrotra [2286], page 612. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Chang:2001:PPB

[1601] Edward Chang, Kwang-Ting Cheng, and Lihyuarn L. Chang. PBIR — perception-based image retrieval. In Sellis and Mehrotra [2286], page 613. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Kriegel:2001:SDM

[1602] Hans-Peter Kriegel, Andreas Müller, Marco Pötke, and Thomas Seidl. Spatial data management for computer aided design. In Sellis and Mehrotra [2286], page 614. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Lam:2001:RRT

[1603] Kam-Yiu Lam, Edward Chan, Tei-Wei Kuo, S. W. Ng, and Dick Hung. RETINA: a real-time traffic navigation system. In Sellis and Mehrotra [2286], page 615. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Datta:2001:DCA

[1604] Anindya Datta, Kaushik Dutta, Krithi Ramamritham, Helen Thomas, and Debra VanderMeer. Dynamic content acceleration: a caching solution to enable scalable dynamic Web page generation. In Sellis and Mehrotra [2286], page 616. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Whitney:2001:LOR

[1605] Arthur Whitney and Dennis Shasha. Lots o’Ticks: real time high performance time series queries on billions of trades and quotes. In Sellis and Mehrotra [2286], page 617. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Han:2001:DMS

[1606] Jiawei Han, Hasan Jamil, Ying Lu, Liangyou Chen, Yaqin Liao, and Jian Pei. DNA-miner: a system prototype for mining DNA sequences. In Sellis and Mehrotra [2286], page 618. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Chen:2001:DDW

[1607] Jun Chen, Xin Zhang, Songting Chen, Andreas Koeller, and Elke A. Rundensteiner. Dyda: data warehouse maintenance in fully concurrent environments. In Sellis and Mehrotra [2286], page 619. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Larson:2001:XDM

[1608] Per-Åke Larson, Dana Florescu, Goetz Graefe, Guido Moerkotte, Hamid Pirahesh, and Harald Schöning. XML data management (panel session): go native or spruce up relational systems? In Sellis and Mehrotra [2286], page 620. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Rosenthal:2001:WDR

[1609] Arnon Rosenthal, Klaus Dittrich, Jim Donahue, and Bill Maimone. Will database researchers have any role in data security? (panel session). In Sellis and Mehrotra [2286], page 621. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Mohan:2001:ASP

[1610] C. Mohan, Larry Cable, Matthieu Devin, Scott Dietzen, Pat Helland, and Dan Wolfson. Application servers (panel session): born-again TP monitors for the Web. In Sellis and Mehrotra [2286], page 622. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Haas:2001:OQP

[1611] Peter J. Haas and Joseph M. Hellerstein. Online query processing: a tutorial. In Sellis and Mehrotra [2286], page 623. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Gunopulos:2001:TSS

[1612] Dimitrios Gunopulos and Gautam Das. Time series similarity measures and time series indexing. In Sellis and Mehrotra [2286], page 624. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Bussler:2001:SBI

[1613] Christoph Bussler. Semantic B2B integration. In Sellis and Mehrotra [2286], page 625. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Casati:2001:MLD

[1614] Fabio Casati and Ming-Chien Shan. Models and languages for describing and discovering E-services. In Sellis and Mehrotra [2286], page 626. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Smith:2001:SMD

[1615] John R. Smith. Standard for multimedia databases. In Sellis and Mehrotra [2286], page 627. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Buneman:2002:ASD

[1616] Peter Buneman, Sanjeev Khanna, Keishi Tajima, and Wang-Chiew Tan.

Archiving scientific data. In Franklin et al. [2287], pages 1–12. ISBN ???? LCCN ???? ACM order number 475020.

Babcock:2002:MID

[1617] Brian Babcock, Shivnath Babu, Mayur Datar, Rajeev Motwani, and Jennifer Widom. Models and issues in data stream systems. In ACM [2288], pages 1–16. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Riedewald:2002:EIA

[1618] Mirek Riedewald, Divyakant Agrawal, and Amr El Abbadi. Efficient integration and aggregation of historical information. In Franklin et al. [2287], pages 13–24. ISBN ???? LCCN ???? ACM order number 475020.

Gottlob:2002:MDE

[1619] Georg Gottlob and Christoph Koch.

Monadic datalog and the expressive power of languages for web information extraction. In ACM [2288], pages 17–28. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Kalnis:2002:APP

[1620] Panos Kalnis, Wee Siong Ng, Beng Chin Ooi, Dimitris Papadias, and Kian-Lee Tan. An adaptive peer-to-peer network for distributed caching of OLAP results. In Franklin et al. [2287], pages 25–36. ISBN ???? LCCN ???? ACM order number 475020.

Chen:2002:DDN

[1621] Chung-Min Chen and Christine T.

Cheng. From discrepancy to declustering: near-optimal multidimensional declustering strategies for range queries. In ACM [2288], pages 29– 38. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Viglas:2002:RBQ

[1622] Stratis D. Viglas and Jeffrey F. Naughton. Rate-based query optimization for streaming information sources. In Franklin et al. [2287], pages 37–48. ISBN ???? LCCN ???? ACM order number 475020.

Shasha:2002:AAT

[1623] Dennis Shasha, Jason T. L. Wang, and Rosalba Giugno. Algorithmics and applications of tree and graph searching. In ACM [2288], pages 39–52. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Madden:2002:CAC

[1624] Samuel Madden, Mehul Shah, Joseph M. Hellerstein, and Vijayshankar Raman. Continuously adaptive continuous queries over streams. In Franklin et al. [2287], pages 49–60. ISBN ???? LCCN ???? ACM order number 475020.

Segoufin:2002:VSX

[1625] Luc Segoufin and Victor Vianu. Validating streaming XML documents. In ACM [2288], pages 53–64. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Dobra:2002:PCA

[1626] Alin Dobra, Minos Garofalakis, Johannes Gehrke, and Rajeev Rastogi. Processing complex aggregate queries over data streams. In Franklin et al. [2287], pages 61–72. ISBN ???? LCCN ???? ACM order number 475020.

Miklau:2002:CEX

[1627] Gerome Miklau and Dan Suciu. Containment and equivalence for an XPath fragment. In ACM [2288], pages 65– 76. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Olston:2002:BEC

[1628] Chris Olston and Jennifer Widom.

Best-effort cache synchronization with source cooperation. In Franklin et al. [2287], pages 73–84. ISBN ???? LCCN ???? ACM order number 475020.

Neven:2002:PWQ

[1629] Frank Neven. On the power of walking for querying tree-structured data. In ACM [2288], pages 77–84. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Arenas:2002:NFX

[1630] Marcelo Arenas and Leonid Libkin. A normal form for XML documents. In ACM [2288], pages 85–96. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Zadorozhny:2002:EEQ

[1631] Vladimir Zadorozhny, Louiqa Raschid, Maria Esther Vidal, Tolga Urhan, and Laura Bright. Efficient evaluation of queries in a mediator for WebSources. In Franklin et al. [2287], pages 85–96. ISBN ???? LCCN ???? ACM order number 475020.

Datta:2002:PBA

[1632] Anindya Datta, Kaushik Dutta, Helen Thomas, Debra VanderMeer, Suresha, and Krithi Ramamritham. Proxybased acceleration of dynamically generated content on the world wide web: an approach and implementation. In Franklin et al. [2287], pages 97–108. ISBN ???? LCCN ???? ACM order number 475020.

Spielmann:2002:DCW

[1633] Marc Spielmann, Jerzy Tyszkiewicz, and Jan Van den Bussche. Distributed computation of web queries using automata. In ACM [2288], pages 97– 108. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Grust:2002:AXL

[1634] Torsten Grust. Accelerating XPath location steps. In Franklin et al. [2287], pages 109–120. ISBN ???? LCCN ???? ACM order number 475020.

Ross:2002:CSC

[1635] Kenneth A. Ross. Conjunctive selection conditions in main memory. In ACM [2288], pages 109–120. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Chung:2002:AAP

[1636] Chin-Wan Chung, Jun-Ki Min, and Kyuseok Shim. APEX: an adaptive path index for XML data. In Franklin et al. [2287], pages 121–132. ISBN ???? LCCN ???? ACM order number 475020.

Zhang:2002:EAO

[1637] Donghui Zhang, Vassilis J. Tsotras, and Dimitrios Gunopulos. Efficient aggregation over objects with extent. In ACM [2288], pages 121–132. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Schmidt:2002:HEM

[1638] Rolfe R. Schmidt and Cyrus Shahabi.

How to evaluate multiple range-sum queries progressively. In ACM [2288], pages 133–141. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Kaushik:2002:CIB

[1639] Raghav Kaushik, Philip Bohannon, Jeffrey F. Naughton, and Henry F. Korth. Covering indexes for branching path queries. In Franklin et al. [2287], pages 133–144. ISBN ???? LCCN ???? ACM order number 475020.

Lechtenborger:2002:CRV

[1640] Jens Lechtenbörger and Gottfried Vossen. On the computation of relational view complements. In ACM [2288], pages 142–149. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Zhou:2002:IDO

[1641] Jingren Zhou and Kenneth A. Ross.

Implementing database operations using SIMD instructions. In Franklin et al. [2287], pages 145–156. ISBN ???? LCCN ???? ACM order number 475020.

Buneman:2002:PDA

[1642] Peter Buneman, Sanjeev Khanna, and Wang-Chiew Tan. On propagation of deletions and annotations through views. In ACM [2288], pages 150– 158. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Chen:2002:FPB

[1643] Shimin Chen, Phillip B. Gibbons, Todd C. Mowry, and Gary Valentin.

Fractal prefetching B^{+}-trees: optimizing both cache and disk performance. In Franklin et al. [2287], pages 157–168. ISBN ???? LCCN ???? ACM order number 475020.

Chirkova:2002:VSP

[1644] Rada Chirkova. The view-selection problem has an exponential-time lower bound for conjunctive queries and views. In ACM [2288], pages 159– 168. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Hurtado:2002:ODC

[1645] Carlos A. Hurtado and Alberto O. Mendelzon. OLAP dimension constraints. In ACM [2288], pages 169– 179. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Li:2002:SHT

[1646] Wei Li, Dengfeng Gao, and Richard T. Snodgrass. Skew handling techniques in sort-merge join. In Franklin et al. [2287], pages 169–180. ISBN ???? LCCN ???? ACM order number 475020.

Guha:2002:FAH

[1647] Sudipto Guha, Nick Koudas, and Divesh Srivastava. Fast algorithms for hierarchical range histogram construction. In ACM [2288], pages 180– 187. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Freire:2002:SMX

[1648] Juliana Freire, Jayant R. Haritsa, Maya Ramanath, Prasan Roy, and Jérôme Siméon. StatiX: making XML count.

In Franklin et al. [2287], pages 181–191. ISBN ???? LCCN ???? ACM order number 475020.

Mokhtar:2002:MOQ

[1649] Hoda Mokhtar, Jianwen Su, and Oscar Ibarra. On moving object queries: (extended abstract). In ACM [2288], pages 188–198. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Papakonstantinou:2002:QQR

[1650] Yannis Papakonstantinou, Michalis Petropoulos, and Vasilis Vassalos. QURSED: querying and reporting semistructured data. In Franklin et al. [2287], pages 192–203. ISBN ???? LCCN ???? ACM order number 475020.

Cadoli:2002:KCQ

[1651] Marco Cadoli and Toni Mancini.

Knowledge compilation = query rewriting + view synthesis. In ACM [2288], pages 199–208. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Tatarinov:2002:SQO

[1652] Igor Tatarinov, Stratis D. Viglas, Kevin Beyer, Jayavel Shanmugasundaram, Eugene Shekita, and Chun Zhang. Storing and querying ordered XML using a relational database system. In Franklin et al. [2287], pages 204–215. ISBN ???? LCCN ???? ACM order number 475020.

Afrati:2002:AQU

[1653] Foto Afrati, Chen Li, and Prasenjit Mitra. Answering queries using views with arithmetic comparisons. In ACM [2288], pages 209–220. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Hacigumus:2002:ESE

[1654] Hakan Hacigümüş, Bala Iyer, Chen Li, and Sharad Mehrotra. Executing SQL over encrypted data in the databaseservice-provider model. In Franklin et al. [2287], pages 216–227. ISBN ???? LCCN ???? ACM order number 475020.

Arasu:2002:CMR

[1655] Arvind Arasu, Brian Babcock, Shivnath Babu, Jon McAlister, and Jennifer Widom. Characterizing memory requirements for queries over continuous data streams. In ACM [2288], pages 221–232. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Gillmann:2002:WMS

[1656] Michael Gillmann, Gerhard Weikum, and Wolfgang Wonner. Workflow management with service quality guarantees. In Franklin et al. [2287], pages 228–239. ISBN ???? LCCN ???? ACM order number 475020.

Lenzerini:2002:DIT

[1657] Maurizio Lenzerini. Data integration: a theoretical perspective. In ACM [2288], pages 233–246. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Dasu:2002:MDS

[1658] Tamraparni Dasu, Theodore Johnson, S. Muthukrishnan, and Vladislav Shkapenyuk. Mining database structure; or, how to build a data quality browser. In Franklin et al. [2287], pages 240–251. ISBN ???? LCCN ???? ACM order number 475020.

Calvanese:2002:LRV

[1659] Diego Calvanese, Giuseppe De Giacomo, Maurizio Lenzerini, and Moshe Y. Vardi. Lossless regular views. In ACM [2288], pages 247–258. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Luo:2002:SHR

[1660] Gang Luo, Curt J. Ellmann, Peter J.

Haas, and Jeffrey F. Naughton. A scalable hash ripple join algorithm. In Franklin et al. [2287], pages 252–262. ISBN ???? LCCN ???? ACM order number 475020.

Arenas:2002:VCX

[1661] Marcelo Arenas, Wenfei Fan, and Leonid Libkin. On verifying consistency of XML specifications. In ACM [2288], pages 259–270. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Bruno:2002:ESQ

[1662] Nicolas Bruno and Surajit Chaudhuri.

Exploiting statistics on query expressions for optimization. In Franklin et al. [2287], pages 263–274. ISBN ???? LCCN ???? ACM order number 475020.

Cohen:2002:LDX

[1663] Edith Cohen, Haim Kaplan, and Tova Milo. Labeling dynamic XML trees. In ACM [2288], pages 271–281. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Raman:2002:PRO

[1664] Vijayshankar Raman and Joseph M.

Hellerstein. Partial results for online query processing. In Franklin et al. [2287], pages 275–286. ISBN ???? LCCN ???? ACM order number 475020.

Chatterji:2002:CAQ

[1665] S. Chatterji, S. S. K. Evani, S. Ganguly, and M. D. Yemmanuru. On the complexity of approximate query optimization. In ACM [2288], pages 282–292. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Guha:2002:AXJ

[1666] Sudipto Guha, H. V. Jagadish, Nick Koudas, Divesh Srivastava, and Ting Yu. Approximate XML joins. In Franklin et al. [2287], pages 287–298. ISBN ???? LCCN ???? ACM order number 475020.

Chu:2002:LEC

[1667] Francis Chu, Joseph Halpern, and Johannes Gehrke. Least expected cost query optimization: what can we expect? In ACM [2288], pages 293– 302. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

Ramanan:2002:EAM

[1668] Prakash Ramanan. Efficient algorithms for minimizing tree pattern queries. In Franklin et al. [2287], pages 299–309. ISBN ???? LCCN ???? ACM order number 475020.

Bruno:2002:HTJ

[1669] Nicolas Bruno, Nick Koudas, and Divesh Srivastava. Holistic twig joins: optimal XML pattern matching. In Franklin et al. [2287], pages 310–321. ISBN ???? LCCN ???? ACM order number 475020.

deVries:2002:ENS

[1670] Arjen P. de Vries, Nikos Mamoulis, Niels Nes, and Martin Kersten. Efficient k-NN search on vertically decomposed data. In Franklin et al. [2287], pages 322–333. ISBN ???? LCCN ???? ACM order number 475020.

Tao:2002:TPQ

[1671] Yufei Tao and Dimitris Papadias.

Time-parameterized queries in spatiotemporal databases. In Franklin et al. [2287], pages 334–345. ISBN ???? LCCN ???? ACM order number 475020.

Chang:2002:MPS

[1672] Kevin Chen-Chuan Chang and Seung won Hwang. Minimal probing: supporting expensive predicates for top-k queries. In Franklin et al. [2287], pages 346–357. ISBN ???? LCCN ???? ACM order number 475020.

Polyzotis:2002:SSG

[1673] Neoklis Polyzotis and Minos Garofalakis. Statistical synopses for graphstructured XML databases. In Franklin et al. [2287], pages 358–369. ISBN ???? LCCN ???? ACM order number 475020.

Gao:2002:CES

[1674] Like Gao and X. Sean Wang. Continually evaluating similarity-based pattern queries on a streaming time series. In Franklin et al. [2287], pages 370–381. ISBN ???? LCCN ???? ACM order number 475020.

Moon:2002:GMS

[1675] Yang-Sae Moon, Kyu-Young Whang, and Wook-Shin Han. General match: a subsequence matching method in timeseries databases based on generalized windows. In Franklin et al. [2287], pages 382–393. ISBN ???? LCCN ???? ACM order number 475020.

Wang:2002:CPS

[1676] Haixun Wang, Wei Wang, Jiong Yang, and Philip S. Yu. Clustering by pattern similarity in large data sets. In Franklin et al. [2287], pages 394–405. ISBN ???? LCCN ???? ACM order number 475020.

Yang:2002:MLS

[1677] Jiong Yang, Wei Wang, Philip S. Yu, and Jiawei Han. Mining long sequential patterns in a noisy environment. In Franklin et al. [2287], pages 406–417. ISBN ???? LCCN ???? ACM order number 475020.

Procopiuc:2002:MCA

[1678] Cecilia M. Procopiuc, Michael Jones, Pankaj K. Agarwal, and T. M. Murali. A Monte Carlo algorithm for fast projective clustering. In Franklin et al. [2287], pages 418–427. ISBN ???? LCCN ???? ACM order number 475020.

Thaper:2002:DMH

[1679] Nitin Thaper, Sudipto Guha, Piotr Indyk, and Nick Koudas. Dynamic multidimensional histograms. In Franklin et al. [2287], pages 428–439. ISBN ???? LCCN ???? ACM order number 475020.

Choi:2002:SES

[1680] Yong-Jin Choi and Chin-Wan Chung. Selectivity estimation for spatiotemporal queries to moving objects. In Franklin et al. [2287], pages 440–451.

ISBN ???? LCCN ???? ACM order number 475020.

Aggarwal:2002:HSS

[1681] Charu C. Aggarwal. Hierarchical subspace sampling: a unified framework for high dimensional data reduction, selectivity estimation and nearest neighbor search. In Franklin et al. [2287], pages 452–463. ISBN ???? LCCN ???? ACM order number 475020.

Sismanis:2002:DSP

[1682] Yannis Sismanis, Antonios Deligiannakis, Nick Roussopoulos, and Yannis Kotidis. Dwarf: shrinking the PetaCube. In Franklin et al. [2287], pages 464–475. ISBN ???? LCCN ???? ACM order number 475020.

Garofalakis:2002:WSE

[1683] Minos Garofalakis and Phillip B. Gibbons. Wavelet synopses with error guarantees. In Franklin et al. [2287], pages 476–487. ISBN ???? LCCN ???? ACM order number 475020.

Chaudhuri:2002:CSW

[1684] Surajit Chaudhuri, Ashish Kumar Gupta, and Vivek Narasayya. Compressing SQL workloads. In Franklin et al. [2287], pages 488–499. ISBN ???? LCCN ???? ACM order number 475020.

Bhattacharya:2002:CBR

[1685] Suparna Bhattacharya, C. Mohan, Karen W. Brannon, Inderpal Narang, Hui-I Hsiao, and Mahadevan Subramanian. Coordinating backup/recovery and data consistency between database and file systems. In Franklin et al. [2287], pages 500–511. ISBN ???? LCCN ???? ACM order number 475020.

Dogac:2002:EII

[1686] Asuman Dogac, Yusuf Tambag, Pinar Pembecioglu, Sait Pektas, Gokce Laleci, Gokhan Kurt, Serkan Toprak, and Yildiray Kabak. An ebXML infrastructure implementation through UDDI registries and RosettaNet PIPs. In Franklin et al. [2287], pages 512–523. ISBN ???? LCCN ???? ACM order number 475020.

Josifovski:2002:GNF

[1687] Vanja Josifovski, Peter Schwarz, Laura Haas, and Eileen Lin. Garlic: a new flavor of federated query processing for DB2. In Franklin et al. [2287], pages 524–532. ISBN ???? LCCN ???? ACM order number 475020.

Bumbulis:2002:CBT

[1688] Peter Bumbulis and Ivan T. Bowman. A compact B-tree. In Franklin et al. [2287], pages 533–541. ISBN ???? LCCN ???? ACM order number 475020.

Weininger:2002:EEJ

[1689] Andreas Weininger. Efficient execution of joins in a star schema. In Franklin et al. [2287], pages 542–545. ISBN ???? LCCN ???? ACM order number 475020.

Kothuri:2002:QRT

[1690] Ravi Kanth V. Kothuri, Siva Ravada, and Daniel Abugov. Quadtree and Rtree indexes in Oracle Spatial: a comparison using GIS data. In Franklin et al. [2287], pages 546–557. ISBN ???? LCCN ???? ACM order number 475020.

Rao:2002:APD

[1691] Jun Rao, Chun Zhang, Nimrod Megiddo, and Guy Lohman. Automating physical database design in a parallel database. In Franklin et al. [2287], pages 558–569. ISBN ???? LCCN ???? ACM order number 475020.

Szalay:2002:SSP

[1692] Alexander S. Szalay, Jim Gray, Ani R.

Thakar, Peter Z. Kunszt, Tanu Malik, Jordan Raddick, Christopher Stoughton, and Jan vandenBerg. The SDSS Skyserver: public access to the Sloan Digital Sky Server data. In Franklin et al. [2287], pages 570–581. ISBN ???? LCCN ???? ACM order number 475020.

Poess:2002:TDT

[1693] Meikel Poess, Bryan Smith, Lubor Kollar, and Paul Larson. TPC-DS, taking decision support benchmarking to the next level. In Franklin et al. [2287], pages 582–587. ISBN ???? LCCN ???? ACM order number 475020.

Team:2002:MTC

[1694] The TimesTen Team. Mid-tier caching:

the TimesTen approach. In Franklin et al. [2287], pages 588–593. ISBN ???? LCCN ???? ACM order number 475020.

Anton:2002:WCD

[1695] Jesse Anton, Lawrence Jacobs, Xiang Liu, Jordan Parker, Zheng Zeng, and Tie Zhong. Web caching for database applications with Oracle Web Cache. In Franklin et al. [2287], pages 594–599. ISBN ???? LCCN ???? ACM order number 475020.

Luo:2002:MTD

[1696] Qiong Luo, Sailesh Krishnamurthy, C. Mohan, Hamid Pirahesh, Honguk Woo, Bruce G. Lindsay, and Jeffrey F. Naughton. Middle-tier database caching for e-business. In Franklin et al. [2287], pages 600–611. ISBN ???? LCCN ???? ACM order number 475020.

Altinel:2002:DDC

[1697] Mehmet Altinel, Qiong Luo, Sailesh Krishnamurthy, C. Mohan, Hamid Pirahesh, Bruce G. Lindsay, Honguk Woo, and Larry Brown. DBCache: database caching for web application servers. In Franklin et al. [2287], page 612. ISBN ???? LCCN ???? ACM order number 475020.

Markl:2002:LTA

[1698] Volker Markl and Guy Lohman. Learning table access cardinalities with LEO. In Franklin et al. [2287], page 613. ISBN ???? LCCN ???? ACM order number 475020.

Zhang:2002:RMD

[1699] Xin Zhang, Mukesh Mulchandani, Steffen Christ, Brian Murphy, and Elke A. Rundensteiner. Rainbow: mappingdriven XQuery processing system. In Franklin et al. [2287], page 614. ISBN ???? LCCN ???? ACM order number 475020.

Theobald:2002:XSE

[1700] Anja Theobald and Gerhard Weikum. The XXL search engine: ranked retrieval of XML data using indexes and ontologies. In Franklin et al. [2287], page 615. ISBN ???? LCCN ???? ACM order number 475020.

Barbosa:2002:TTB

[1701] Denilson Barbosa, Alberto Mendelzon, John Keenleyside, and Kelly Lyons. Toxgene: a template-based data generator for XML. In Franklin et al. [2287], page 616. ISBN ???? LCCN ???? ACM order number 475020.

Abadi:2002:VCD

[1702] Daniel J. Abadi and Mitch Cherniack.

Visual COKO: a debugger for query optimizer development. In Franklin et al. [2287], page 617. ISBN ???? LCCN ???? ACM order number 475020.

Chen:2002:XSC

[1703] Li Chen, Elke A. Rundensteiner, and Song Wang. XCache: a semantic caching system for XML queries. In Franklin et al. [2287], page 618. ISBN ???? LCCN ???? ACM order number 475020.

Karbhari:2002:ACD

[1704] Pradnya Karbhari, Michael Rabinovich, Zhen Xiao, and Fred Douglis. ACDN: a content delivery network for applications. In Franklin et al. [2287], page 619. ISBN ???? LCCN ???? ACM order number 475020.

Wang:2002:CTE

[1705] Tengjiao Wang, Shiwei Tang, Dongqing Yang, Jun Gao, Yuqing Wu, and Jian Pei. COMMIX: towards effective web information extraction, integration and query answering. In Franklin et al. [2287], page 620. ISBN ???? LCCN ???? ACM order number 475020.

Fung:2002:CND

[1706] Wai Fu Fung, David Sun, and Johannes Gehrke. COUGAR: the network is the database. In Franklin et al. [2287], page 621. ISBN ???? LCCN ???? ACM order number 475020.

Madden:2002:DQL

[1707] Samuel Madden and Joseph M. Hellerstein. Distributing queries over lowpower wireless sensor networks. In Franklin et al. [2287], page 622. ISBN ???? LCCN ???? ACM order number 475020.

Cranor:2002:GHP

[1708] Chuck Cranor, Yuan Gao, Theodore Johnson, Vlaidslav Shkapenyuk, and Oliver Spatscheck. Gigascope: high performance network monitoring with an SQL interface. In Franklin et al. [2287], page 623. ISBN ???? LCCN ???? ACM order number 475020.

Crescenzi:2002:RAD

[1709] Valter Crescenzi, Giansalvatore Mecca, and Paolo Merialdo. Roadrunner: automatic data extraction from dataintensive web sites. In Franklin et al. [2287], page 624. ISBN ???? LCCN ???? ACM order number 475020.

Florescu:2002:XPW

[1710] Daniela Florescu, Andreas Grünhagen, Donald Kossmann, and Steffen Rost. XL: a platform for web services. In Franklin et al. [2287], page 625. ISBN ???? LCCN ???? ACM order number 475020.

Han:2002:COE

[1711] Jiawei Han, Jianyong Wang, Guozhu Dong, Jian Pei, and Ke Wang. Cube-Explorer: online exploration of data cubes. In Franklin et al. [2287], page 626. ISBN ???? LCCN ???? ACM order number 475020.

Agrawal:2002:DEK

[1712] Sanjay Agrawal, Surajit Chaudhuri, and Gautam Das. DBXplorer: enabling keyword search over relational databases. In Franklin et al. [2287], page 627. ISBN ???? LCCN ???? ACM order number 475020.

Phan:2002:GTG

[1713] Jessica M. Phan and Raymond Ng.

GEA: a toolkit for gene expression analysis. In Franklin et al. [2287], page 628. ISBN ???? LCCN ???? ACM order number 475020.

Hinneburg:2002:HEV

[1714] Alexander Hinneburg, Daniel A. Keim, and Markus Wawryniuk. HD-Eye: visual clustering of high dimensional data. In Franklin et al. [2287], page 629. ISBN ???? LCCN ???? ACM order number 475020.

Lu:2002:XMY

[1715] Hongjun Lu, Guoren Wang, Ge Yu, Yubin Bao, Jianhua Lv, and Yaxin Yu. XBase: making your gigabyte disk queriable. In Franklin et al. [2287], page 630. ISBN ???? LCCN ???? ACM order number 475020.

Rundensteiner:2002:XVI

[1716] Elke A. Rundensteiner, Matthew O.

Ward, Jing Yang, and Punit R. Doshi. Xmdvtool: visual interactive data exploration and trend discovery of highdimensional data sets. In Franklin et al. [2287], page 631. ISBN ???? LCCN ???? ACM order number 475020.

Hellerstein:2002:DSF

[1717] Joseph M. Hellerstein and Jennifer Widom. Data streams: fresh current or stagnant backwater? (panel). In Franklin et al. [2287], page 632. ISBN ???? LCCN ???? ACM order number 475020.

Bonnet:2002:GPO

[1718] Philippe Bonnet. Going public: opensource databases and database research. In Franklin et al. [2287], page 633. ISBN ???? LCCN ???? ACM order number 475020.

Cotton:2002:IX

[1719] Paul Cotton. Implementing XQuery. In Franklin et al. [2287], page 634. ISBN ???? LCCN ???? ACM order number 475020.

Garofalakis:2002:QMD

[1720] Minos Garofalakis, Johannes Gehrke, and Rajeev Rastogi. Querying and mining data streams: you only get one look a tutorial. In Franklin et al. [2287], page 635. ISBN ???? LCCN ???? ACM order number 475020.

Mohan:2002:TAS

[1721] C. Mohan. Tutorial: application servers and associated technologies. In Franklin et al. [2287], page 636. ISBN ???? LCCN ???? ACM order number 475020.

Shasha:2002:DTP

[1722] Dennis Shasha and Philippe Bonnet. Database tuning: principles, experiments, and troubleshooting techniques (part I). In Franklin et al. [2287], page 637. ISBN ???? LCCN ???? ACM order number 475020.

Bussler:2002:SSA

[1723] Christoph Bussler. Software as a service: ASP and ASP aggregation. In Franklin et al. [2287], page 638. ISBN ???? LCCN ???? ACM order number 475020.

Ullman:2003:IED

[1724] Jeffrey D. Ullman. Improving the efficiency of database-system teaching. In ACM [2290], pages 1–3. ISBN 1-58113-634-X. LCCN ????

Hull:2003:SLB

[1725] Richard Hull, Michael Benedikt, Vassilis Christophides, and Jianwen Su. Eservices: a look behind the curtain. In ACM [2289], pages 1–14. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Al-Khalifa:2003:QST

[1726] Shurug Al-Khalifa, Cong Yu, and H. V.

Jagadish. Querying structured text in an XML database. In ACM [2290], pages 4–15. ISBN 1-58113-634-X. LCCN ????

Arenas:2003:ITA

[1727] Marcelo Arenas and Leonid Libkin. An information-theoretic approach to normal forms for relational and XML data. In ACM [2289], pages 15–26. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Guo:2003:XRK

[1728] Lin Guo, Feng Shao, Chavdar Botev, and Jayavel Shanmugasundaram. XRANK: ranked keyword search over XML documents. In ACM [2290], pages 16–27. ISBN 1-58113-634-X. LCCN ????

Khuller:2003:ADM

[1729] Samir Khuller, Yoo-Ah Kim, and Yung-Chun (Justin) Wan. Algorithms for data migration with cloning. In ACM [2289], pages 27–36. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Babcock:2003:DTK

[1730] Brian Babcock and Chris Olston. Distributed top-k monitoring. In ACM [2290], pages 28–39. ISBN 1-58113-634-X. LCCN ????

Agrawal:2003:PDS

[1731] Rakesh Agrawal. Privacy in data systems. In ACM [2289], page 37. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Chirkova:2003:MVM

[1732] Rada Chirkova and Chen Li. Materializing views with minimal size to answer queries. In ACM [2289], pages 38–48. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Das:2003:AJP

[1733] Abhinandan Das, Johannes Gehrke, and Mirek Riedewald. Approximate join processing over data streams. In ACM [2290], pages 40–51. ISBN 1-58113-634-X. LCCN ????

Lechtenborger:2003:ICC

[1734] Jens Lechtenbörger. The impact of the constant complement approach towards view updating. In ACM [2289], pages 49–55. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Witkowski:2003:SRO

[1735] Andrew Witkowski, Srikanth Bellamkonda, Tolga Bozkaya, Gregory Dorman, Nathan Folkert, Abhinav Gupta, Lei Shen, and Sankar Subramanian. Spreadsheets in RDBMS for OLAP. In ACM [2290], pages 52–63. ISBN 1-58113-634-X. LCCN ????

Calvanese:2003:VBQ

[1736] Diego Calvanese, Giuseppe De Giacomo, Maurizio Lenzerini, and Moshe Y. Vardi. View-based query containment. In ACM [2289], pages 56–67. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Lakshmanan:2003:QTE

[1737] Laks V. S. Lakshmanan, Jian Pei, and Yan Zhao. QC-trees: an efficient summary structure for semantic OLAP. In ACM [2290], pages 64–75. ISBN 1-58113-634-X. LCCN ????

Gupta:2003:VSP

[1738] Ashish Kumar Gupta, Dan Suciu, and Alon Y. Halevy. The view selection problem for XML content based routing. In ACM [2289], pages 68–77. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Schleimer:2003:WLA

[1739] Saul Schleimer, Daniel S. Wilkerson, and Alex Aiken. Winnowing: local algorithms for document fingerprinting. In ACM [2290], pages 76–85. ISBN 1-58113-634-X. LCCN ????

Kanza:2003:CFD

[1740] Yaron Kanza and Yehoshua Sagiv.

Computing full disjunctions. In ACM [2289], pages 78–89. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Agrawal:2003:ISA

[1741] Rakesh Agrawal, Alexandre Evfimievski, and Ramakrishnan Srikant. Information sharing across private databases. In ACM [2290], pages 86– 97. ISBN 1-58113-634-X. LCCN ????

Fagin:2003:DEG

[1742] Ronald Fagin, Phokion G. Kolaitis, and Lucian Popa. Data exchange: getting to the core. In ACM [2289], pages 90–101. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Sion:2003:RPR

[1743] Radu Sion, Mikhail Atallah, and Sunil Prabhakar. Rights protection for relational data. In ACM [2290], pages 98– 109. ISBN 1-58113-634-X. LCCN ????

Behrend:2003:SSM

[1744] Andreas Behrend. Soft stratification for magic set based query evaluation in deductive databases. In ACM [2289], pages 102–110. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Wang:2003:VDI

[1745] Haixun Wang, Sanghyun Park, Wei Fan, and Philip S. Yu. ViST: a dynamic index method for querying XML data by tree structures. In ACM [2290], pages 110–121. ISBN 1-58113-634-X. LCCN ????

Grahne:2003:QCR

[1746] Gösta Grahne and Alex Thomo. Query containment and rewriting using views for regular path queries under constraints. In ACM [2289], pages 111– 122. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Min:2003:XQC

[1747] Jun-Ki Min, Myung-Jae Park, and Chin-Wan Chung. XPRESS: a queriable compression for XML data. In ACM [2290], pages 122–133. ISBN 1-58113-634-X. LCCN ????

Mendelzon:2003:CDS

[1748] Alberto O. Mendelzon and Ken Q. Pu.

Concise descriptions of subsets of structured sets. In ACM [2289], pages 123–133. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Dittrich:2003:PJR

[1749] Jens-Peter Dittrich, Bernhard Seeger, David Scot Taylor, and Peter Widmayer. On producing join results early. In ACM [2289], pages 134–142. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Chen:2003:DKI

[1750] Qun Chen, Andrew Lim, and Kian Win Ong. D(k)-index: an adaptive structural summary for graph-structured data. In ACM [2290], pages 134–144. ISBN 1-58113-634-X. LCCN ????

Garofalakis:2003:CXD

[1751] Minos Garofalakis and Amit Kumar.

Correlating XML data streams using tree-edit distance embeddings. In ACM [2289], pages 143–154. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Wang:2003:CJS

[1752] Wei Wang, Haifeng Jiang, Hongjun Lu, and Jeffrey Xu Yu. Containment join size estimation: models and methods. In ACM [2290], pages 145–156. ISBN 1-58113-634-X. LCCN ????

Seidl:2003:NDQ

[1753] Helmut Seidl, Thomas Schwentick, and Anca Muscholl. Numerical document queries. In ACM [2289], pages 155– 166. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Mamoulis:2003:EPJ

[1754] Nikos Mamoulis. Efficient processing of joins on set-valued attributes. In ACM [2290], pages 157–168. ISBN 1-58113-634-X. LCCN ????

Segoufin:2003:TQX

[1755] Luc Segoufin. Typing and querying XML documents: some complexity bounds. In ACM [2289], pages 167–178. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Dyreson:2003:TCN

[1756] Curtis E. Dyreson. Temporal coalescing with now granularity, and incomplete information. In ACM [2290], pages 169–180. ISBN 1-58113-634-X. LCCN ????

Gottlob:2003:CXQ

[1757] Georg Gottlob, Christoph Koch, and Reinhard Pichler. The complexity of XPath query evaluation. In ACM [2289], pages 179–190. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Zhu:2003:WIE

[1758] Yunyue Zhu and Dennis Shasha. Warping indexes with envelope transforms for query by humming. In ACM [2290], pages 181–192. ISBN 1-58113-634-X. LCCN ????

Gross-Amblard:2003:QPW

[1759] David Gross-Amblard. Querypreserving watermarking of relational databases and XML documents. In ACM [2289], pages 191–201. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Melnik:2003:RPP

[1760] Sergey Melnik, Erhard Rahm, and Philip A. Bernstein. Rondo: a programming platform for generic model management. In ACM [2290], pages 193–204. ISBN 1-58113-634-X. LCCN ????

Dinur:2003:RIW

[1761] Irit Dinur and Kobbi Nissim. Revealing information while preserving privacy. In ACM [2289], pages 202–210. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Kang:2003:SMO

[1762] Jaewoo Kang and Jeffrey F. Naughton.

On schema matching with opaque column names and data values. In ACM [2290], pages 205–216. ISBN 1-58113-634-X. LCCN ????

Evfimievski:2003:LPB

[1763] Alexandre Evfimievski, Johannes Gehrke, and Ramakrishnan Srikant. Limiting privacy breaches in privacy preserving data mining. In ACM [2289], pages 211–222. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

He:2003:SSM

[1764] Bin He and Kevin Chen-Chuan Chang. Statistical schema matching across web query interfaces. In ACM [2290], pages 217–228. ISBN 1-58113-634-X. LCCN ????

Cohen:2003:MTD

[1765] Edith Cohen and Martin Strauss. Maintaining time-decaying stream aggregates. In ACM [2289], pages 223– 233. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Deligiannakis:2003:EWM

[1766] Antonios Deligiannakis and Nick Roussopoulos. Extended wavelets for multiple measures. In ACM [2290], pages 229–240. ISBN 1-58113-634-X. LCCN ????

Babcock:2003:MVM

[1767] Brain Babcock, Mayur Datar, Rajeev Motwani, and Liadan O’Callaghan. Maintaining variance and k-medians over data stream windows. In ACM [2289], pages 234–243. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Cohen:2003:SBF

[1768] Saar Cohen and Yossi Matias. Spectral bloom filters. In ACM [2290], pages 241–252. ISBN 1-58113-634-X. LCCN ????

Heeren:2003:OIU

[1769] C. Heeren, H. V. Jagadish, and L. Pitt. Optimal indexing using near-minimal space. In ACM [2289], pages 244– 251. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Aggarwal:2003:NNI

[1770] Charu C. Aggarwal and Dakshi Agrawal. On nearest neighbor indexing of nonlinear trajectories. In ACM [2289], pages 252–259. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Babcock:2003:COS

[1771] Brian Babcock, Shivnath Babu, Rajeev Motwani, and Mayur Datar. Chain: operator scheduling for memory minimization in data stream systems. In ACM [2290], pages 253–264. ISBN 1-58113-634-X. LCCN ????

Cali:2003:DCQ

[1772] Andrea Calì, Domenico Lembo, and Riccardo Rosati. On the decidability and complexity of query answering over inconsistent and incomplete databases. In ACM [2289], pages 260–271. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Ganguly:2003:PSE

[1773] Sumit Ganguly, Minos Garofalakis, and Rajeev Rastogi. Processing set expressions over continuous update streams. In ACM [2290], pages 265–276. ISBN 1-58113-634-X. LCCN ????

Kifer:2003:HQF

[1774] Daniel Kifer, Johannes Gehrke, Cristian Bucila, and Walker White. How to quickly find a witness. In ACM [2289], pages 272–283. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Benedikt:2003:CBT

[1775] Michael Benedikt, Chee-Yong Chan, Wenfei Fan, Juliana Freire, and Rajeev Rastogi. Capturing both types and constraints in data integration. In ACM [2290], pages 277–288. ISBN 1-58113-634-X. LCCN ????

Ramesh:2003:FID

[1776] Ganesh Ramesh, William A. Maniatty, and Mohammed J. Zaki. Feasible itemset distributions in data mining: theory and application. In ACM [2289], pages 284–295. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Milo:2003:EIX

[1777] Tova Milo, Serge Abiteboul, Bernd Amann, Omar Benjelloun, and Fred Dang Ngoc. Exchanging intensional XML data. In ACM [2290], pages 289–300. ISBN 1-58113-634-X. LCCN ????

Cormode:2003:WHW

[1778] Graham Cormode and S. Muthukrishnan. What’s hot and what’s not: tracking most frequent items dynamically. In ACM [2289], pages 296–306. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

Fagin:2003:ESS

[1779] Ronald Fagin, Ravi Kumar, and D. Sivakumar. Efficient similarity search and classification via rank aggregation. In ACM [2290], pages 301–312. ISBN 1-58113-634-X. LCCN ????

Chaudhuri:2003:REF

[1780] Surajit Chaudhuri, Kris Ganjam, Venkatesh Ganti, and Rajeev Motwani. Robust and efficient fuzzy match for online data cleaning. In ACM [2290], pages 313–324. ISBN 1-58113-634-X. LCCN ????

Kementsietsidis:2003:MDP

[1781] Anastasios Kementsietsidis, Marcelo Arenas, and Renée J. Miller. Mapping data in peer-to-peer systems: semantics and algorithmic issues. In ACM [2290], pages 325–336. ISBN 1-58113-634-X. LCCN ????

Arasu:2003:ESD

[1782] Arvind Arasu, Hector Garcia-Molina, and Stanford University. Extracting structured data from Web pages. In ACM [2290], pages 337–348. ISBN 1-58113-634-X. LCCN ????

Stolte:2003:SDR

[1783] Etzard Stolte, Christoph von Praun, Gustavo Alonso, and Thomas Gross. Scientific data repositories: designing for a moving target. In ACM [2290], pages 349–360. ISBN 1-58113-634-X. LCCN ????

Chaudhuri:2003:FCP

[1784] Surajit Chaudhuri, Prasanna Ganesan, and Sunita Sarawagi. Factorizing complex predicates in queries to exploit indexes. In ACM [2290], pages 361–372. ISBN 1-58113-634-X. LCCN ????

Ilyas:2003:ECT

[1785] Ihab F. Ilyas, Jun Rao, Guy Lohman, Dengfeng Gao, and Eileen Lin. Estimating compilation time of a query optimizer. In ACM [2290], pages 373–384. ISBN 1-58113-634-X. LCCN ????

Reiss:2003:CSQ

[1786] Frederick R. Reiss and Tapas Kanungo.

A characterization of the sensitivity of query optimization to storage access cost parameters. In ACM [2290], pages 385–396. ISBN 1-58113-634-X. LCCN ????

Lomet:2003:TRR

[1787] David Lomet and Mark Tuttle. A theory of redo recovery. In ACM [2290], pages 397–406. ISBN 1-58113-634-X. LCCN ????

Bierman:2003:FSA

[1788] G. M. Bierman. Formal semantics and analysis of object queries. In ACM [2290], pages 407–418. ISBN 1-58113-634-X. LCCN ????

Gupta:2003:SPX

[1789] Ashish Kumar Gupta and Dan Suciu. Stream processing of XPath queries with predicates. In ACM [2290], pages 419–430. ISBN 1-58113-634-X. LCCN ????

Peng:2003:XQS

[1790] Feng Peng and Sudarshan S. Chawathe. XPath queries on streaming data. In ACM [2290], pages 431–442. ISBN 1-58113-634-X. LCCN ????

Zhang:2003:LBS

[1791] Jun Zhang, Manli Zhu, Dimitris Papadias, Yufei Tao, and Dik Lun Lee. Location-based spatial queries. In ACM [2290], pages 443–454. ISBN 1-58113-634-X. LCCN ????

Sun:2003:HAS

[1792] Chengyu Sun, Divyakant Agrawal, and Amr El Abbadi. Hardware acceleration for spatial selections and joins. In ACM [2290], pages 455–466. ISBN 1-58113-634-X. LCCN ????

Papadias:2003:OPA

[1793] Dimitris Papadias, Yufei Tao, Greg Fu, and Bernhard Seeger. An optimal and progressive algorithm for skyline queries. In ACM [2290], pages 467–478. ISBN 1-58113-634-X. LCCN ????

Cui:2003:CHD

[1794] Bin Cui, Beng Chin Ooi, Jianwen Su, and Kian-Lee Tan. Contorting high dimensional data for efficient main memory KNN processing. In ACM [2290], pages 479–490. ISBN 1-58113-634-X. LCCN ????

Madden:2003:DAQ

[1795] Samuel Madden, Michael J. Franklin, Joseph M. Hellerstein, and Wei Hong. The design of an acquisitional query processor for sensor networks. In ACM [2290], pages 491–502. ISBN 1-58113-634-X. LCCN ????

Deshpande:2003:CQW

[1796] Amol Deshpande, Suman Nath, Phillip B. Gibbons, and Srinivasan Seshan. Cache-and-query for wide area sensor databases. In ACM [2290], pages 503–514. ISBN 1-58113-634-X. LCCN ????

Li:2003:CXT

[1797] Chengkai Li, Philip Bohannon, and P. P. S. Narayan. Composing XSL transformations with XML publishing views. In ACM [2290], pages 515–526. ISBN 1-58113-634-X. LCCN ????

Abiteboul:2003:DXD

[1798] Serge Abiteboul, Angela Bonifati, Grégory Cobéna, Ioana Manolescu, and Tova Milo. Dynamic XML documents with distribution and replication. In ACM [2290], pages 527–538. ISBN 1-58113-634-X. LCCN ????

Babcock:2003:DSS

[1799] Brian Babcock, Surajit Chaudhuri, and Gautam Das. Dynamic sample selection for approximate query processing. In ACM [2290], pages 539–550. ISBN 1-58113-634-X. LCCN ????

Cheng:2003:EPQ

[1800] Reynold Cheng, Dmitri V. Kalashnikov, and Sunil Prabhakar. Evaluating probabilistic queries over imprecise data. In ACM [2290], pages 551–562. ISBN 1-58113-634-X. LCCN ????

Olston:2003:AFC

[1801] Chris Olston, Jing Jiang, and Jennifer Widom. Adaptive filters for continuous queries over distributed data streams. In ACM [2290], pages 563–574. ISBN 1-58113-634-X. LCCN ????

Aggarwal:2003:FDC

[1802] Charu C. Aggarwal. A framework for diagnosing changes in evolving data streams. In ACM [2290], pages 575– 586. ISBN 1-58113-634-X. LCCN ????

Kriegel:2003:USF

[1803] Hans-Peter Kriegel, Stefan Brecheisen, Peer Kröger, Martin Pfeifle, and Matthias Schubert. Using sets of feature vectors for similarity search on voxelized CAD objects. In ACM [2290], pages 587–598. ISBN 1-58113-634-X. LCCN ????

Kim:2003:QRF

[1804] Deok-Hwan Kim and Chin-Wan Chung. QCluster: relevance feedback using adaptive clustering for contentbased image retrieval. In ACM [2290], pages 599–610. ISBN 1-58113-634-X. LCCN ????

Chaudhuri:2003:RSX

[1805] Surajit Chaudhuri, Raghav Kaushik, and Jeffrey F. Naughton. On relational support for XML publishing: beyond sorting and tagging. In ACM [2290], pages 611–622. ISBN 1-58113-634-X. LCCN ????

DeHaan:2003:CXS

[1806] David DeHaan, David Toman, Mariano P. Consens, and M. Tamer Özsu. A comprehensive XQuery to SQL translation using dynamic interval encoding. In ACM [2290], pages 623–634. ISBN 1-58113-634-X. LCCN ????

Ives:2003:AII

[1807] Zachary Ives. Abstracts of invited industrial track presentations. In ACM [2290], page 635. ISBN 1-58113-634-X. LCCN ????

Padmanabhan:2003:MDC

[1808] Sriram Padmanabhan, Bishwaranjan Bhattacharjee, Tim Malkemus, Leslie Cranston, and Matthew Huras. Multidimensional clustering: a new data layout scheme in DB2. In ACM [2290], pages 637–641. ISBN 1-58113-634-X. LCCN ????

Goto:2003:IET

[1809] Koichi Goto and Yahiko Kambayashi.

Integration of electronic tickets and personal guide system for public transport using mobile terminals. In ACM [2290], pages 642–646. ISBN 1-58113-634-X. LCCN ????

Cranor:2003:GSD

[1810] Chuck Cranor, Theodore Johnson, Oliver Spataschek, and Vladislav Shkapenyuk. Gigascope: a stream database for network applications. In ACM [2290], pages 647–651. ISBN 1-58113-634-X. LCCN ????

Zuzarte:2003:WSE

[1811] Calisto Zuzarte, Hamid Pirahesh, Wenbin Ma, Qi Cheng, Linqi Liu, and Kwai Wong. WinMagic: subquery elimination using window aggregation. In ACM [2290], pages 652–656. ISBN 1-58113-634-X. LCCN ????

Shen:2003:CND

[1812] Jialie Shen, Anne H. H. Ngu, John Shepherd, Du Q. Huynh, and Quan Z. Sheng. CMVF: a novel dimension reduction scheme for efficient indexing in a large image database. In ACM [2290], page 657. ISBN 1-58113-634-X. LCCN ????

Lakshmanan:2003:SSO

[1813] Laks V. S. Lakshmanan, Jian Pei, and Yan Zhao. SOCQET: semantic OLAP with compressed cube and summarization. In ACM [2290], page 658. ISBN 1-58113-634-X. LCCN ????

Ooi:2003:PPP

[1814] Beng Chin Ooi, Kian-Lee Tan, Aoying Zhou, Chin Hong Goh, Yingguang Li, Chu Yee Liau, Bo Ling, Wee Siong Ng, Yanfeng Shu, Xiaoyu Wang, and Ming Zhang. PeerDB: peering into personal databases. In ACM [2290], page 659. ISBN 1-58113-634-X. LCCN ????

Liu:2003:GRI

[1815] David T. Liu, Michael J. Franklin, and Devesh Parekh. GridDB: a relational interface for the grid. In ACM [2290], page 660. ISBN 1-58113-634-X. LCCN ????

Larson:2003:TMT

[1816] Per-Åke Larson, Jonathan Goldstein, and Jingren Zhou. Transparent midtier database caching in SQL server. In ACM [2290], page 661. ISBN 1-58113-634-X. LCCN ????

Bornhovd:2003:DMT

[1817] Christof Bornhövd, Mehmet Altinel, Sailesh Krishnamurthy, C. Mohan, Hamid Pirahesh, and Berthold Reinwald. DBCache: middle-tier database caching for highly scalable e-business architectures. In ACM [2290], page 662. ISBN 1-58113-634-X. LCCN ????

Agichtein:2003:QBB

[1818] Eugene Agichtein and Luis Gravano.

QXtract: a building block for efficient information extraction from text databases. In ACM [2290], page 663. ISBN 1-58113-634-X. LCCN ????

Amer-Yahia:2003:PEA

[1819] Sihem Amer-Yahia, Mary Fernández, Divesh Srivastava, and Yu Xu. PIX: exact and approximate phrase matching in XML. In ACM [2290], page 664. ISBN 1-58113-634-X. LCCN ????

Arasu:2003:SSS

[1820] Arvind Arasu, Brian Babcock, Shivnath Babu, Mayur Datar, Keith Ito, Itaru Nishizawa, Justin Rosenstein, and Jennifer Widom. STREAM: the Stanford Stream Data Manager (demonstration description). In ACM [2290], page 665. ISBN 1-58113-634-X. LCCN ????

Abadi:2003:ADS

[1821] D. Abadi, D. Carney, U. Çetintemel, M. Cherniack, C. Convey, C. Erwin, E. Galvez, M. Hatoun, A. Maskey, A. Rasin, A. Singer, M. Stonebraker, N. Tatbul, Y. Xing, R. Yan, and S. Zdonik. Aurora: a data stream management system. In ACM [2290], page 666. ISBN 1-58113-634-X. LCCN ????

Deshpande:2003:IIS

[1822] Amol Deshpande, Suman Nath, Phillip B. Gibbons, and Srinivasan Seshan. IrisNet: Internet-scale resourceintensive sensor services. In ACM [2290], page 667. ISBN 1-58113-634-X. LCCN ????

Chandrasekaran:2003:TCD

[1823] Sirish Chandrasekaran, Owen Cooper, Amol Deshpande, Michael J. Franklin, Joseph M. Hellerstein, Wei Hong, Sailesh Krishnamurthy, Samuel R. Madden, Fred Reiss, and Mehul A. Shah. TelegraphCQ: continuous dataflow processing. In ACM [2290], page 668. ISBN 1-58113-634-X. LCCN ????

Cho:2003:LSE

[1824] SungRan Cho, Sihem Amer-Yahia, Laks V. S. Lakshmanan, and Divesh Srivastava. LockX: a system for efficiently querying secure XML. In ACM [2290], page 669. ISBN 1-58113-634-X. LCCN ????

Zhou:2003:TDC

[1825] Aoying Zhou, Qing Wang, Zhimao Guo, Xueqing Gong, Shihui Zheng, Hongwei Wu, Jianchang Xiao, Kun Yue, and Wenfei Fan. TREX: DTDconforming XML to XML transformations. In ACM [2290], page 670. ISBN 1-58113-634-X. LCCN ????

Zhang:2003:RMX

[1826] Xin Zhang, Katica Dimitrova, Ling Wang, Maged El Sayed, Brian Murphy, Bradford Pielech, Mukesh Mulchandani, Luping Ding, and Elke A. Rundensteiner. Rainbow: multi-XQuery optimization using materialized XML views. In ACM [2290], page 671. ISBN 1-58113-634-X. LCCN ????

Paparizos:2003:TNS

[1827] Stelios Paparizos, Shurug Al-Khalifa, Adriane Chapman, H. V. Jagadish, Laks V. S. Lakshmanan, Andrew Nierman, Jignesh M. Patel, Divesh Srivastava, Nuwee Wiwatwattana, Yuqing Wu, and Cong Yu. TIMBER: a native system for querying XML. In ACM [2290], page 672. ISBN 1-58113-634-X. LCCN ????

Bohannon:2003:RRL

[1828] Philip Bohannon, Xin (Luna) Dong, Sumit Ganguly, Henry F. Korth, Chengkai Li, P. P. S. Narayan, and Pradeep Shenoy. ROLEX: relational on-line exchange with XML. In ACM [2290], page 673. ISBN 1-58113-634-X. LCCN ????

Agrawal:2003:SWR

[1829] Rakesh Agrawal, Peter J. Haas, and Jerry Kiernan. A system for watermarking relational databases. In ACM [2290], page 674. ISBN 1-58113-634-X. LCCN ????

Zhu:2003:QHA

[1830] Yunyue Zhu, Dennis Shasha, and Xiaojian Zhao. Query by humming: in action with its technology revealed. In ACM [2290], page 675. ISBN 1-58113-634-X. LCCN ????

Sengar:2003:PRQ

[1831] Vibhuti S. Sengar and Jayant R. Haritsa. PLASTIC: reducing query optimization overheads through plan recycling. In ACM [2290], page 676. ISBN 1-58113-634-X. LCCN ????

Korn:2003:IVC

[1832] Flip Korn, S. Muthukrishnan, and Yunyue Zhu. IPSOFACTO: a visual correlation tool for aggregate network traffic data. In ACM [2290], page 677. ISBN 1-58113-634-X. LCCN ????

Gupta:2003:BMS

[1833] Amarnath Gupta, Bertram Ludäscher, and Maryann E. Martone. BIRN-M: a semantic mediator for solving realworld neuroscience problems. In ACM [2290], page 678. ISBN 1-58113-634-X. LCCN ????

Koudas:2003:PQN

[1834] Nick Koudas and Divesh Srivastava. Panel: querying networked databases. In ACM [2290], page 679. ISBN 1-58113-634-X. LCCN ????

Gray:2003:LR

[1835] Jim Gray, Hans Schek, Michael Stonebraker, and Jeff Ullman. The Lowell Report. In ACM [2290], page 680. ISBN 1-58113-634-X. LCCN ????

Johnson:2003:DQD

[1836] Theodore Johnson and Tamraparni Dasu. Data quality and data cleaning: an overview. In ACM [2290], page 681. ISBN 1-58113-634-X. LCCN ????

Chamberlin:2003:XQL

[1837] Don Chamberlin. XQuery: a query language for XML. In ACM [2290], page 682. ISBN 1-58113-634-X. LCCN ????

Jagatheesan:2003:DGM

[1838] Arun Jagatheesan and Arcot Rajasekar. Data grid management systems. In ACM [2290], page 683. ISBN 1-58113-634-X. LCCN ????

Gray:2004:NDR

[1839] Jim Gray. The next database revolution. In ACM [2291], pages 1–4. ISBN 1-58113-859-8. LCCN ????

Maurer:2004:RCD

[1840] Ueli Maurer. The role of cryptography in database security. In ACM [2291], pages 5–10. ISBN 1-58113-859-8. LCCN ????

Jain:2004:ASR

[1841] Ankur Jain, Edward Y. Chang, and Yuan-Fang Wang. Adaptive stream resource management using Kalman Filters. In ACM [2291], pages 11–22. ISBN 1-58113-859-8. LCCN ????

Wu:2004:OED

[1842] Huanmei Wu, Betty Salzberg, and Donghui Zhang. Online event-driven subsequence matching over financial data streams. In ACM [2291], pages 23–34. ISBN 1-58113-859-8. LCCN ????

Cormode:2004:HUS

[1843] Graham Cormode, Theodore Johnson, Flip Korn, S. Muthukrishnan, Oliver Spatscheck, and Divesh Srivastava. Holistic UDAFs at streaming speeds. In ACM [2291], pages 35–46. ISBN 1-58113-859-8. LCCN ????

Chen:2004:BEX

[1844] Yi Chen, Susan B. Davidson, and Yifeng Zheng. BLAS: an efficient XPath processing system. In ACM [2291], pages 47–58. ISBN 1-58113-859-8. LCCN ????

Jiang:2004:EPX

[1845] Haifeng Jiang, Hongjun Lu, and Wei Wang. Efficient processing of XML twig queries with OR-predicates. In ACM [2291], pages 59–70. ISBN 1-58113-859-8. LCCN ????

Paparizos:2004:TLC

[1846] Stelios Paparizos, Yuqing Wu, Laks V. S. Lakshmanan, and H. V. Jagadish. Tree logical classes for efficient evaluation of XQuery. In ACM [2291], pages 71–82. ISBN 1-58113-859-8. LCCN ????

Amer-Yahia:2004:FFS

[1847] Sihem Amer-Yahia, Laks V. S. Lakshmanan, and Shashank Pandit. FleX-Path: flexible structure and full-text querying for XML. In ACM [2291], pages 83–94. ISBN 1-58113-859-8. LCCN ????

Wu:2004:ICB

[1848] Wensheng Wu, Clement Yu, AnHai Doan, and Weiyi Meng. An interactive clustering-based approach to integrating source query interfaces on the deep Web. In ACM [2291], pages 95–106. ISBN 1-58113-859-8. LCCN ????

Zhang:2004:UWQ

[1849] Zhen Zhang, Bin He, and Kevin Chen-Chuan Chang. Understanding Web query interfaces: best-effort parsing with hidden syntax. In ACM [2291], pages 107–118. ISBN 1-58113-859-8. LCCN ????

Lerman:2004:USW

[1850] Kristina Lerman, Lise Getoor, Steven Minton, and Craig Knoblock. Using the structure of Web sites for automatic segmentation of tables. In ACM [2291], pages 119–130. ISBN 1-58113-859-8. LCCN ????

Vlachos:2004:ISP

[1851] Michail Vlachos, Chris Meek, Zografoula Vagena, and Dimitrios Gunopulos. Identifying similarities, periodicities and bursts for online search queries. In ACM [2291], pages 131–142. ISBN 1-58113-859-8. LCCN ????

Cong:2004:FFI

[1852] Gao Cong, Anthony K. H. Tung, Xin Xu, Feng Pan, and Jiong Yang. FARMER: finding interesting rule groups in microarray datasets. In ACM [2291], pages 143–154. ISBN 1-58113-859-8. LCCN ????

Cormode:2004:DRF

[1853] Graham Cormode, Flip Korn, S. Muthukrishnan, and Divesh Srivastava. Diamond in the rough: finding Hierarchical Heavy Hitters in multi-dimensional data. In ACM [2291], pages 155–166. ISBN 1-58113-859-8. LCCN ????

Chen:2004:CBL

[1854] Lei Chen, Zheng Huang, and Raghu Ramakrishnan. Cost-based labeling of groups of mass spectra. In ACM [2291], pages 167–178. ISBN 1-58113-859-8. LCCN ????

Bowman:2004:OQS

[1855] Ivan T. Bowman and Kenneth Salem.

Optimization of query streams using semantic prefetching. In ACM [2291], pages 179–190. ISBN 1-58113-859-8. LCCN ????

Zhou:2004:BDO

[1856] Jingren Zhou and Kenneth A. Ross.

Buffering database operations for enhanced instruction cache performance. In ACM [2291], pages 191–202. ISBN 1-58113-859-8. LCCN ????

Ilyas:2004:RAQ

[1857] Ihab F. Ilyas, Rahul Shah, Walid G. Aref, Jeffrey Scott Vitter, and Ahmed K. Elmagarmid. Rank-aware query optimization. In ACM [2291], pages 203–214. ISBN 1-58113-859-8. LCCN ????

Govindaraju:2004:FCD

[1858] Naga K. Govindaraju, Brandon Lloyd, Wei Wang, Ming Lin, and Dinesh Manocha. Fast computation of database operations using graphics processors. In ACM [2291], pages 215– 226. ISBN 1-58113-859-8. LCCN ????

Abiteboul:2004:LQE

[1859] Serge Abiteboul, Omar Benjelloun, Bogdan Cautis, Ioana Manolescu, Tova Milo, and Nicoleta Preda. Lazy query evaluation for Active XML. In ACM [2291], pages 227–238. ISBN 1-58113-859-8. LCCN ????

Bose:2004:DSM

[1860] Sujoe Bose and Leonidas Fegaras. Data stream management for historical XML data. In ACM [2291], pages 239–250. ISBN 1-58113-859-8. LCCN ????

Jagadish:2004:CXO

[1861] H. V. Jagadish, Laks V. S. Lakshmanan, Monica Scannapieco, Divesh Srivastava, and Nuwee Wiwatwattana. Colorful XML: one hierarchy isn’t enough. In ACM [2291], pages 251–262. ISBN 1-58113-859-8. LCCN ????

Polyzotis:2004:AXQ

[1862] Neoklis Polyzotis, Minos Garofalakis, and Yannis Ioannidis. Approximate XML query answers. In ACM [2291], pages 263–274. ISBN 1-58113-859-8. LCCN ????

Haas:2004:BLB

[1863] Peter J. Haas and Christian König. A bi-level Bernoulli scheme for database sampling. In ACM [2291], pages 275– 286. ISBN 1-58113-859-8. LCCN ????

Chaudhuri:2004:EUB

[1864] Surajit Chaudhuri, Gautam Das, and Utkarsh Srivastava. Effective use of block-level sampling in statistics estimation. In ACM [2291], pages 287–298. ISBN 1-58113-859-8. LCCN ????

Jermaine:2004:OMV

[1865] Christopher Jermaine, Abhijit Pol, and Subramanian Arumugam. Online maintenance of very large random samples. In ACM [2291], pages 299–310. ISBN 1-58113-859-8. LCCN ????

Bruno:2004:CSS

[1866] Nicolas Bruno and Surajit Chaudhuri.

Conditional selectivity for statistics on query expressions. In ACM [2291], pages 311–322. ISBN 1-58113-859-8. LCCN ????

Graefe:2004:TSI

[1867] Goetz Graefe and Michael Zwilling.

Transaction support for indexed summary views. In ACM [2291], pages 323– 334. ISBN 1-58113-859-8. LCCN ????

Yan:2004:GIF

[1868] Xifeng Yan, Philip S. Yu, and Jiawei Han. Graph indexing: a frequent structure-based approach. In ACM [2291], pages 335–346. ISBN 1-58113-859-8. LCCN ????

Arge:2004:PRT

[1869] Lars Arge, Mark de Berg, Herman J.

Haverkort, and Ke Yi. The Priority Rtree: a practically efficient and worstcase optimal R-tree. In ACM [2291], pages 347–358. ISBN 1-58113-859-8. LCCN ????

Agrawal:2004:IVH

[1870] Sanjay Agrawal, Vivek Narasayya, and Beverly Yang. Integrating vertical and horizontal partitioning into automated physical database design. In ACM [2291], pages 359–370. ISBN 1-58113-859-8. LCCN ????

Yu:2004:CBX

[1871] Cong Yu and Lucian Popa. Constraintbased XML query rewriting for data integration. In ACM [2291], pages 371– 382. ISBN 1-58113-859-8. LCCN ????

Dhamankar:2004:IDC

[1872] Robin Dhamankar, Yoonkyong Lee, AnHai Doan, Alon Halevy, and Pedro Domingos. iMAP: discovering complex semantic matches between database schemas. In ACM [2291], pages 383– 394. ISBN 1-58113-859-8. LCCN ????

Ives:2004:ASP

[1873] Zachary G. Ives, Alon Y. Halevy, and Daniel S. Weld. Adapting to source properties in processing data integration queries. In ACM [2291], pages 395– 406. ISBN 1-58113-859-8. LCCN ????

Babu:2004:AOP

[1874] Shivnath Babu, Rajeev Motwani, Kamesh Munagala, Itaru Nishizawa, and Jennifer Widom. Adaptive ordering of pipelined stream filters. In ACM [2291], pages 407–418. ISBN 1-58113-859-8. LCCN ????

Ayad:2004:SOC

[1875] Ahmed M. Ayad and Jeffrey F. Naughton. Static optimization of conjunctive queries with sliding windows over infinite streams. In ACM [2291], pages 419–430. ISBN 1-58113-859-8. LCCN ????

Zhu:2004:DPM

[1876] Yali Zhu, Elke A. Rundensteiner, and George T. Heineman. Dynamic plan migration for continuous queries over data streams. In ACM [2291], pages 431–442. ISBN 1-58113-859-8. LCCN ????

Yiu:2004:COS

[1877] Man Lung Yiu and Nikos Mamoulis.

Clustering objects on a spatial network. In ACM [2291], pages 443–454. ISBN 1-58113-859-8. LCCN ????

Bohm:2004:CCC

[1878] Christian Böhm, Karin Kailing, Peer Kröger, and Arthur Zimek. Computing clusters of correlation connected objects. In ACM [2291], pages 455–466. ISBN 1-58113-859-8. LCCN ????

Nassar:2004:IED

[1879] Samer Nassar, Jörg Sander, and Corrine Cheng. Incremental and effective data summarization for dynamic hierarchical clustering. In ACM [2291], pages 467–478. ISBN 1-58113-859-8. LCCN ????

Tian:2004:ISX

[1880] Feng Tian, Berthold Reinwald, Hamid Pirahesh, Tobias Mayr, and Jussi Myllymaki. Implementing a scalable XML publish/subscribe system using relational database systems. In ACM [2291], pages 479–490. ISBN 1-58113-859-8. LCCN ????

Yi:2004:IMX

[1881] Ke Yi, Hao He, Ioana Stanoi, and Jun Yang. Incremental maintenance of XML structural indexes. In ACM [2291], pages 491–502. ISBN 1-58113-859-8. LCCN ????

Bohannon:2004:IES

[1882] Philip Bohannon, Byron Choi, and Wenfei Fan. Incremental evaluation of schema-directed XML publishing. In ACM [2291], pages 503–514. ISBN 1-58113-859-8. LCCN ????

Bawa:2004:PVD

[1883] Mayank Bawa, Aristides Gionis, Hector Garcia-Molina, and Rajeev Motwani. The price of validity in dynamic networks. In ACM [2291], pages 515–526. ISBN 1-58113-859-8. LCCN ????

Deligiannakis:2004:CHI

[1884] Antonios Deligiannakis, Yannis Kotidis, and Nick Roussopoulos. Compressing historical information in sensor networks. In ACM [2291], pages 527–538. ISBN 1-58113-859-8. LCCN ????

Tatarinov:2004:EQR

[1885] Igor Tatarinov and Alon Halevy. Efficient query reformulation in peer data management systems. In ACM [2291], pages 539–550. ISBN 1-58113-859-8. LCCN ????

Rizvi:2004:EQR

[1886] Shariq Rizvi, Alberto Mendelzon, S. Sudarshan, and Prasan Roy. Extending query rewriting techniques for fine-grained access control. In ACM [2291], pages 551–562. ISBN 1-58113-859-8. LCCN ????

Agrawal:2004:OPE

[1887] Rakesh Agrawal, Jerry Kiernan, Ramakrishnan Srikant, and Yirong Xu. Order preserving encryption for numeric data. In ACM [2291], pages 563– 574. ISBN 1-58113-859-8. LCCN ????

Miklau:2004:FAI

[1888] Gerome Miklau and Dan Suciu. A formal analysis of information disclosure in data exchange. In ACM [2291], pages 575–586. ISBN 1-58113-859-8. LCCN ????

Fan:2004:SXQ

[1889] Wenfei Fan, Chee-Yong Chan, and Minos Garofalakis. Secure XML querying with security views. In ACM [2291], pages 587–598. ISBN 1-58113-859-8. LCCN ????

Cai:2004:IST

[1890] Yuhan Cai and Raymond Ng. Indexing spatio-temporal trajectories with Chebyshev polynomials. In ACM [2291], pages 599–610. ISBN 1-58113-859-8. LCCN ????

Tao:2004:PIM

[1891] Yufei Tao, Christos Faloutsos, Dimitris Papadias, and Bin Liu. Prediction and indexing of moving objects with unknown motion patterns. In ACM [2291], pages 611–622. ISBN 1-58113-859-8. LCCN ????

Mokbel:2004:SSI

[1892] Mohamed F. Mokbel, Xiaopeing Xiong, and Walid G. Aref. SINA: scalable incremental processing of continuous queries in spatio-temporal databases. In ACM [2291], pages 623–634. ISBN 1-58113-859-8. LCCN ????

Patel:2004:SEI

[1893] Jignesh M. Patel, Yun Chen, and V. Prasad Chakka. STRIPES: an efficient index for predicted trajectories. In ACM [2291], pages 635–646. ISBN 1-58113-859-8. LCCN ????

Ilyas:2004:CAD

[1894] Ihab F. Ilyas, Volker Markl, Peter Haas, Paul Brown, and Ashraf Aboulnaga. CORDS: automatic discovery of correlations and soft functional dependencies. In ACM [2291], pages 647–658. ISBN 1-58113-859-8. LCCN ????

Markl:2004:RQP

[1895] Volker Markl, Vijayshankar Raman, David Simmen, Guy Lohman, Hamid Pirahesh, and Miso Cilimdzic. Robust query processing through progressive optimization. In ACM [2291], pages 659–670. ISBN 1-58113-859-8. LCCN ????

Rao:2004:CAO

[1896] Jun Rao, Hamid Pirahesh, and Calisto Zuzarte. Canonical abstraction for outerjoin optimization. In ACM [2291], pages 671–682. ISBN 1-58113-859-8. LCCN ????

Enderle:2004:JID

[1897] Jost Enderle, Matthias Hampel, and Thomas Seidl. Joining interval data in relational databases. In ACM [2291], pages 683–694. ISBN 1-58113-859-8. LCCN ????

Das:2004:ATS

[1898] Abhinandan Das, Johannes Gehrke, and Mirek Riedewald. Approximation techniques for spatial data. In ACM [2291], pages 695–706. ISBN 1-58113-859-8. LCCN ????

Cohen:2004:SDA

[1899] Edith Cohen and Haim Kaplan.

Spatially-decaying aggregation over a network: model and algorithms. In ACM [2291], pages 707–718. ISBN 1-58113-859-8. LCCN ????

Hung:2004:TET

[1900] Edward Hung, Yu Deng, and V. S.

Subrahmanian. TOSS: an extension of TAX with ontologies and similarity queries. In ACM [2291], pages 719–730. ISBN 1-58113-859-8. LCCN ????

Andritsos:2004:ITT

[1901] Periklis Andritsos, Renée J. Miller, and Panayiotis Tsaparas. Informationtheoretic tools for mining database structure from large data sets. In ACM [2291], pages 731–742. ISBN 1-58113-859-8. LCCN ????

Sarawagi:2004:ESJ

[1902] Sunita Sarawagi and Alok Kirpal. Efficient set joins on similarity predicates. In ACM [2291], pages 743–754. ISBN 1-58113-859-8. LCCN ????

Chakrabarti:2004:ACQ

[1903] Kaushik Chakrabarti, Surajit Chaudhuri, and Seung won Hwang. Automatic categorization of query results. In ACM [2291], pages 755–766. ISBN 1-58113-859-8. LCCN ????

Ipeirotis:2004:WOS

[1904] Panagiotis G. Ipeirotis and Luis Gravano. When one sample is not enough: improving text database selection using shrinkage. In ACM [2291], pages 767–778. ISBN 1-58113-859-8. LCCN ????

Kaushik:2004:ISI

[1905] Raghav Kaushik, Rajasekar Krishnamurthy, Jeffrey F. Naughton, and Raghu Ramakrishnan. On the integration of structure indexes and inverted lists. In ACM [2291], pages 779–790. ISBN 1-58113-859-8. LCCN ????

Luo:2004:TPI

[1906] Gang Luo, Jeffrey F. Naughton, Curt J. Ellmann, and Michael W. Watzke. Toward a progress indicator for database queries. In ACM [2291], pages 791–802. ISBN 1-58113-859-8. LCCN ????

Chaudhuri:2004:EPE

[1907] Surajit Chaudhuri, Vivek Narasayya, and Ravishankar Ramamurthy. Estimating progress of execution for SQL queries. In ACM [2291], pages 803–814. ISBN 1-58113-859-8. LCCN ????

Guo:2004:RCC

[1908] Hongfei Guo, Per-Åke Larson, Raghu Ramakrishnan, and Jonathan Goldstein. Relaxed currency and consistency: how to say “good enough” in SQL. In ACM [2291], pages 815–826. ISBN 1-58113-859-8. LCCN ????

Shah:2004:HAF

[1909] Mehul A. Shah, Joseph M. Hellerstein, and Eric Brewer. Highly available, fault-tolerant, parallel dataflows. In ACM [2291], pages 827–838. ISBN 1-58113-859-8. LCCN ????

Gryz:2004:QSD

[1910] Jarek Gryz, Junjie Guo, Linqi Liu, and Calisto Zuzarte. Query sampling in DB2 Universal Database. In ACM [2291], pages 839–843. ISBN 1-58113-859-8. LCCN ????

Galindo-Legaria:2004:QPS

[1911] César A. Galindo-Legaria, Stefano Stefani, and Florian Waas. Query processing for SQL updates. In ACM [2291], pages 844–849. ISBN 1-58113-859-8. LCCN ????

Cruanes:2004:PSE

[1912] Thierry Cruanes, Benoit Dageville, and Bhaskar Ghosh. Parallel SQL execution in Oracle 10g. In ACM [2291], pages 850–854. ISBN 1-58113-859-8. LCCN ????

Gupta:2004:DDR

[1913] Abhinav Gupta, Sankar Subramanian, Srikanth Bellamkonda, Tolga Bozkaya, Nathan Folkert, Lei Sheng, and Andrew Witkowski. Data densification in a relational database system. In ACM [2291], pages 855–859. ISBN 1-58113-859-8. LCCN ????

Acheson:2004:HNR

[1914] Alazel Acheson, Mason Bendixen, José A. Blakeley, Peter Carlin, Ebru Ersan, Jun Fang, Xiaowei Jiang, Christian Kleinerman, Balaji Rathakrishnan, Gideon Schaller, Beysim Sezgin, Ramachandran Venkatesh, and Honggang Zhang. Hosting the .NET runtime in Microsoft SQL server. In ACM [2291], pages 860–865. ISBN 1-58113-859-8. LCCN ????

Ordonez:2004:VHP

[1915] Carlos Ordonez. Vertical and horizontal percentage aggregations. In ACM [2291], pages 866–871. ISBN 1-58113-859-8. LCCN ????

Little:2004:MWS

[1916] Mark Little. Models for Web Services transactions. In ACM [2291], page 872. ISBN 1-58113-859-8. LCCN ????

Agrawal:2004:ESI

[1917] Rakesh Agrawal, Dmitri Asonov, and Ramakrishnan Srikant. Enabling sovereign information sharing using Web Services. In ACM [2291], pages 873–877. ISBN 1-58113-859-8. LCCN ????

Mihic:2004:BDA

[1918] Matthew Mihic. Building dynamic application networks with Web Services. In ACM [2291], page 878. ISBN 1-58113-859-8. LCCN ????

Gudgin:2004:SRT

[1919] Martin Gudgin. Secure, reliable, transacted: innovation in Web Services architecture. In ACM [2291], pages 879– 880. ISBN 1-58113-859-8. LCCN ????

Kosugi:2004:SPQ

[1920] Naoko Kosugi, Yasushi Sakurai, and Masashi Morimoto. SoundCompass: a practical query-by-humming system; normalization of scalable and shiftable time-series data and effective subsequence generation. In ACM [2291], pages 881–886. ISBN 1-58113-859-8. LCCN ????

Bendsen:2004:MDB

[1921] Per Bendsen. Model-driven business UI based on maps. In ACM [2291], pages 887–891. ISBN 1-58113-859-8. LCCN ????

Dar:2004:DTD

[1922] Shaul Dar, Gil Hecht, and Eden Shochat. dbSwitch^{T M}: towards a database utility. In ACM [2291], pages 892–896. ISBN 1-58113-859-8. LCCN ????

Hall:2004:RPC

[1923] Dean E. Hall. Requirements and policy challenges in highly secure environments. In ACM [2291], pages 897–898. ISBN 1-58113-859-8. LCCN ????

Multari:2004:IAT

[1924] Nicholas J. Multari. Information assurance technical challenges. In ACM [2291], page 899. ISBN 1-58113-859-8. LCCN ????

Dill:2004:SOB

[1925] Marcus Dill, Achim Kraiss, Stefan Sigg, and Thomas Zurek. Service-oriented BI: towards tight integration of business intelligence into operational applications. In ACM [2291], page 900. ISBN 1-58113-859-8. LCCN ????

Carey:2004:XMX

[1926] Michael J. Carey. XML in the middle:

XQuery in the WebLogic Platform. In ACM [2291], pages 901–902. ISBN 1-58113-859-8. LCCN ????

ONeil:2004:OIF

[1927] Patrick O’Neil, Elizabeth O’Neil, Shankar Pal, Istvan Cseri, Gideon Schaller, and Nigel Westbury. ORD-PATHs: insert-friendly XML node labels. In ACM [2291], pages 903–908. ISBN 1-58113-859-8. LCCN ????

Brambilla:2004:DSW

[1928] Marco Brambilla, Stefano Ceri, Sara Comai, Marco Dario, Piero Fraternali, and Ioana Manolescu. Declarative specification of Web applications exploiting Web services and workflows.

In ACM [2291], pages 909–910. ISBN 1-58113-859-8. LCCN ????

Fernandez:2004:YHB

[1929] Mary Fernández, Nicola Onose, and Jérôme Siméon. Yoo-Hoo!: building a presence service with XQuery and WSDL. In ACM [2291], pages 911–912. ISBN 1-58113-859-8. LCCN ????

He:2004:KDD

[1930] Bin He, Zhen Zhang, and Kevin Chen-Chuan Chang. Knocking the door to the deep Web: integrating Web query interfaces. In ACM [2291], pages 913– 914. ISBN 1-58113-859-8. LCCN ????

Carreira:2004:EDD

[1931] Paulo Carreira and Helena Galhardas. Efficient development of data migration transformations. In ACM [2291], pages 915–916. ISBN 1-58113-859-8. LCCN ????

Borkar:2004:LDW

[1932] Vinayak Borkar. Liquid data for WebLogic: integrating enterprise data and services. In ACM [2291], pages 917– 918. ISBN 1-58113-859-8. LCCN ????

Cai:2004:MMA

[1933] Y. Dora Cai, David Clutter, Greg Pape, Jiawei Han, Michael Welge, and Loretta Auvil. MAIDS: mining alarming incidents from data streams. In ACM [2291], pages 919–920. ISBN 1-58113-859-8. LCCN ????

Qian:2004:FFE

[1934] Yu Qian, Gang Zhang, and Kang Zhang. FAÇADE: a fast and effective approach to the discovery of dense clusters in noisy spatial data. In ACM [2291], pages 921–922. ISBN 1-58113-859-8. LCCN ????

Serazi:2004:D

[1935] Masum Serazi, Vasily Malakhov, Dongmei Ren, Amal Perera, Imad Rahal, Weihua Wu, Qiang Ding, Fei Pan, and William Perrizo. DataMIME^{T M}. In ACM [2291], pages 923–924. ISBN 1-58113-859-8. LCCN ????

Kramer:2004:PPI

[1936] Jürgen Krämer and Bernhard Seeger.

PIPES: a public infrastructure for processing and exploring streams. In ACM [2291], pages 925–926. ISBN 1-58113-859-8. LCCN ????

Garg:2004:WCM

[1937] Shaveen Garg, Krithi Ramamritham, and Soumen Chakrabarti. Web-CAM: monitoring the dynamic Web to respond to continual queries. In ACM [2291], pages 927–928. ISBN 1-58113-859-8. LCCN ????

Balazinska:2004:LMH

[1938] Magdalena Balazinska, Hari Balakrishnan, and Michael Stonebraker. Load management and high availability in the Medusa distributed stream processing system. In ACM [2291], pages 929– 930. ISBN 1-58113-859-8. LCCN ????

Babu:2004:SAE

[1939] Shivnath Babu and Jennifer Widom.

StreaMon: an adaptive engine for stream query processing. In ACM [2291], pages 931–932. ISBN 1-58113-859-8. LCCN ????

Idreos:2004:PDE

[1940] Stratos Idreos, Manolis Koubarakis, and Christos Tryfonopoulos. P2P-DIET: an extensible P2P service that unifies ad-hoc and continuous querying in super-peer networks. In ACM [2291], pages 933–934. ISBN 1-58113-859-8. LCCN ????

Chun:2004:QIS

[1941] Brent Chun, Joseph M. Hellerstein, Ryan Huebsch, Shawn R. Jeffery, Boon Thau Loo, Sam Mardanbeigi, Timothy Roscoe, Sean Rhea, Scott Shenker, and Ion Stoica. Querying at Internet scale. In ACM [2291], pages 935–936. ISBN 1-58113-859-8. LCCN ????

Guo:2004:SRC

[1942] Hongfei Guo, Per-Åke Larson, Raghu Ramakrishnan, and Jonathan Goldstein. Support for relaxed currency and consistency constraints in MTCache. In ACM [2291], pages 937–938. ISBN 1-58113-859-8. LCCN ????

Crainiceanu:2004:IFP

[1943] Adina Crainiceanu, Prakash Linga, Ashwin Machanavajjhala, Johannes Gehrke, and Jayavel Shanmugasundaram. An indexing framework for peer-to-peer systems. In ACM [2291], pages 939–940. ISBN 1-58113-859-8. LCCN ????

Meng:2004:XII

[1944] Xiaofeng Meng, Yu Jiang, Yan Chen, and Haixun Wang. XSeq: an indexing infrastructure for tree pattern queries. In ACM [2291], pages 941–942. ISBN 1-58113-859-8. LCCN ????

Botev:2004:TBX

[1945] Chavdar Botev, Sihem Amer-Yahia, and Jayavel Shanmugasundaram. A TeXQuery-based XML full-text search engine. In ACM [2291], pages 943–944. ISBN 1-58113-859-8. LCCN ????

Fundulaki:2004:SYD

[1946] Irini Fundulaki and Arnaud Sahuguet.

“share your data, keep your secrets.”. In ACM [2291], pages 945–946. ISBN 1-58113-859-8. LCCN ????

Agrawal:2004:MHD

[1947] Rakesh Agrawal, Ameet Kini, Kristen LeFevre, Amy Wang, Yirong Xu, and Diana Zhou. Managing healthcare data hippocratically. In ACM [2291], pages 947–948. ISBN 1-58113-859-8. LCCN ????

Kumaran:2004:LMM

[1948] A. Kumaran and Jayant R. Haritsa.

LexEQUAL: multilexical matching operator in SQL. In ACM [2291], pages 949–950. ISBN 1-58113-859-8. LCCN ????

Huang:2004:IIT

[1949] B. Huang, Z. Huang, H. Li, D. Lin, H. Lu, and Y. Song. ITQS: an integrated transport query system. In ACM [2291], pages 951–952. ISBN 1-58113-859-8. LCCN ????

Bedathur:2004:BDH

[1950] Srikanta J. Bedathur, Abhijit Kadlag, and Jayant R. Haritsa. BODHI: a database habitat for bio-diversity information. In ACM [2291], pages 953–954. ISBN 1-58113-859-8. LCCN ????

Mehrotra:2004:CCA

[1951] Sharad Mehrotra, Carter Butts, Dmitri V. Kalashnikov, Nalini Venkatasubramanian, Kemal Altintas, Ram Hariharan, Haimin Lee, Yiming Ma, Amnon Myers, Jehan Wickramasuriya, Ron Eguchi, and Charles Huyck. CA-MAS: a citizen awareness system for crisis mitigation. In ACM [2291], pages 955–956. ISBN 1-58113-859-8. LCCN ????

Franklin:2004:RCR

[1952] Michael J. Franklin, Jennifer Widom, Anastassia Ailamaki, Philip A. Bernstein, David DeWitt, Alon Halevy, Zachary Ives, and Gerhard Weikum. Rethinking the conference reviewing process. In ACM [2291], page 957. ISBN 1-58113-859-8. LCCN ????

Hull:2004:TDC

[1953] Richard Hull and Jianwen Su. Tools for design of composite Web Services. In ACM [2291], pages 958–961. ISBN 1-58113-859-8. LCCN ????

Rosenthal:2004:SSD

[1954] Arnon Rosenthal and Marianne Winslett. Security of shared data in large systems: state of the art and research directions. In ACM [2291], pages 962–964. ISBN 1-58113-859-8. LCCN ????

Lerner:2004:FAT

[1955] Alberto Lerner, Dennis Shasha, Zhihua Wang, Xiaojian Zhao, and Yunyue Zhu. Fast algorithms for time series with applications to finance, physics, music, biology, and other suspects. In ACM [2291], pages 965–968. ISBN 1-58113-859-8. LCCN ????

Faloutsos:2004:IMS

[1956] Christos Faloutsos. Indexing and mining streams. In ACM [2291], page 969. ISBN 1-58113-859-8. LCCN ????

Lam:2005:CSP

[1957] Monica S. Lam, John Whaley, V. Benjamin Livshits, Michael C. Martin, Dzintars Avots, Michael Carbin, and Christopher Unkel. Context-sensitive program analysis as database queries. In ACM [2292], pages 1–12. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Arenas:2005:XDE

[1958] Marcelo Arenas and Leonid Libkin.

XML data exchange: consistency and query answering. In ACM [2292], pages 13–24. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Benedikt:2005:XSP

[1959] Michael Benedikt, Wenfei Fan, and Floris Geerts. XPath satisfiability in the presence of DTDs. In ACM [2292], pages 25–36. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Vansummeren:2005:DWD

[1960] Stijn Vansummeren. Deciding welldefinedness of XQuery fragments. In ACM [2292], pages 37–48. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Segoufin:2005:VQD

[1961] Luc Segoufin and Victor Vianu. Views and queries: determinacy and rewriting. In ACM [2292], pages 49–60. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Kolaitis:2005:SMD

[1962] Phokion G. Kolaitis. Schema mappings, data exchange, and metadata management. In ACM [2292], pages 61–75. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Leinders:2005:CDS

[1963] Dirk Leinders and Jan Van den Bussche. On the complexity of division and set joins in the relational algebra. In ACM [2292], pages 76–83. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Koch:2005:CNX

[1964] Christoph Koch. On the complexity of nonrecursive XQuery and functional query languages on complex values. In ACM [2292], pages 84–97. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Cohen:2005:IAC

[1965] Sara Cohen and Yehoshua Sagiv. An incremental algorithm for computing ranked full disjunctions. In ACM [2292], pages 98–107. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Abadi:2005:SAC

[1966] Martín Abadi and Bogdan Warinschi. Security analysis of cryptographically controlled access to XML documents. In ACM [2292], pages 108–117. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Kenthapadi:2005:SA

[1967] Krishnaram Kenthapadi, Nina Mishra, and Kobbi Nissim. Simulatable auditing. In ACM [2292], pages 118– 127. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Blum:2005:PPS

[1968] Avrim Blum, Cynthia Dwork, Frank McSherry, and Kobbi Nissim. Practical privacy: the SuLQ framework. In ACM [2292], pages 128–138. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Zhong:2005:PEA

[1969] Sheng Zhong, Zhiqiang Yang, and Rebecca N. Wright. Privacy-enhancing k-anonymization of customer data. In ACM [2292], pages 139–147. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Gottlob:2005:CCD

[1970] Georg Gottlob. Computing cores for data exchange: new algorithms and practical solutions. In ACM [2292], pages 148–159. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Fuxman:2005:PDE

[1971] Ariel Fuxman, Phokion G. Kolaitis, Renée J. Miller, and Wang-Chiew Tan. Peer data exchange. In ACM [2292], pages 160–171. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Nash:2005:CMG

[1972] Alan Nash, Philip A. Bernstein, and Sergey Melnik. Composition of mappings given by embedded dependencies. In ACM [2292], pages 172–183. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Fagin:2005:MSD

[1973] Ronald Fagin, R. Guha, Ravi Kumar, Jasmine Novak, D. Sivakumar, and Andrew Tomkins. Multi-structural databases. In ACM [2292], pages 184– 195. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Cheng:2005:DMM

[1974] David Cheng, Santosh Vempala, Ravi Kannan, and Grant Wang. A divideand-merge methodology for clustering. In ACM [2292], pages 196–205. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Fekete:2005:AIL

[1975] Alan Fekete. Allocating isolation levels to transactions. In ACM [2292], pages 206–215. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Bar-Yossef:2005:BQE

[1976] Ziv Bar-Yossef, Marcus Fontoura, and Vanja Josifovski. Buffering in query evaluation over XML streams. In ACM [2292], pages 216–227. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Dobra:2005:HRW

[1977] Alin Dobra. Histograms revisited: when are histograms the best approximation method for aggregates over joins? In ACM [2292], pages 228– 237. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Grohe:2005:LBS

[1978] Martin Grohe and Nicole Schweikardt. Lower bounds for sorting with few random accesses to external memory. In ACM [2292], pages 238–249. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Srivastava:2005:OPN

[1979] Utkarsh Srivastava, Kamesh Munagala, and Jennifer Widom. Operator placement for in-network stream query processing. In ACM [2292], pages 250– 258. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Ganguly:2005:JDA

[1980] Sumit Ganguly, Minos Garofalakis, Amit Kumar, and Rajeev Rastogi. Join-distinct aggregate estimation over update streams. In ACM [2292], pages 259–270. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Cormode:2005:SEM

[1981] Graham Cormode and S. Muthukrishnan. Space efficient mining of multigraph streams. In ACM [2292], pages 271–282. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Maneth:2005:XTC

[1982] S. Maneth, A. Berlea, T. Perst, and H. Seidl. XML type checking with macro tree transducers. In ACM [2292], pages 283–294. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Abiteboul:2005:RRA

[1983] Serge Abiteboul, Tova Milo, and Omar Benjelloun. Regular rewriting of active XML and unambiguity. In ACM [2292], pages 295–303. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Deutsch:2005:DSC

[1984] Alin Deutsch, Yannis Katsis, and Yannis Papakonstantinou. Determining source contribution in integration systems. In ACM [2292], pages 304– 315. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Gehrke:2005:MMP

[1985] Johannes Gehrke. Models and methods for privacy-preserving data publishing and analysis: invited tutorial. In ACM [2292], page 316. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Alon:2005:EAS

[1986] Noga Alon, Nick Duffield, Carsten Lund, and Mikkel Thorup. Estimating arbitrary subset sums with few probes. In ACM [2292], pages 317–325. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Sakurai:2005:FFS

[1987] Yasushi Sakurai, Masatoshi Yoshikawa, and Christos Faloutsos. FTW: fast similarity search under the time warping distance. In ACM [2292], pages 326–337. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Hershberger:2005:SCH

[1988] John Hershberger, Nisheeth Shrivastava, Subhash Suri, and Csaba D. Tóth. Space complexity of hierarchical heavy hitters in multi-dimensional data streams. In ACM [2292], pages 338–347. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Sayrafi:2005:DC

[1989] Bassem Sayrafi and Dirk Van Gucht. Differential constraints. In ACM [2292], pages 348–357. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Abiteboul:2005:DAD

[1990] Serge Abiteboul, Zoë Abrams, Stefan Haar, and Tova Milo. Diagnosis of asynchronous discrete event systems: datalog to the rescue! In ACM [2292], pages 358–367. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Li:2005:RRO

[1991] Haiquan Li, Jinyan Li, Limsoon Wong, Mengling Feng, and Yap-Peng Tan. Relative risk and odds ratio: a data mining perspective. In ACM [2292], pages 368–377. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

Halevy:2006:PDS

[1992] Alon Halevy, Michael Franklin, and David Maier. Principles of dataspace systems. In ACM [2293], pages 1–9. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Bojanczyk:2006:TVL

[1993] Mikolaj Bojańczyk, Claire David, Anca Muscholl, Thomas Schwentick, and Luc Segoufin. Two-variable logic on data trees and XML reasoning. In ACM [2293], pages 10–19. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Bender:2006:APM

[1994] Michael A. Bender and Haodong Hu.

An adaptive packed-memory array. In ACM [2293], pages 20–29. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Kolaitis:2006:CDE

[1995] Phokion G. Kolaitis, Jonathan Panttaja, and Wang-Chiew Tan. The complexity of data exchange. In ACM [2293], pages 30–39. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Gottlob:2006:DEC

[1996] Georg Gottlob and Alan Nash. Data exchange: computing cores in polynomial time. In ACM [2293], pages 40–49. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Fagin:2006:ISM

[1997] Ronald Fagin. Inverting schema mappings. In ACM [2293], pages 50– 59. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Libkin:2006:DEI

[1998] Leonid Libkin. Data exchange and incomplete information. In ACM [2293], pages 60–69. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Cohen:2006:EQC

[1999] Sara Cohen. Equivalence of queries combining set and bag-set semantics. In ACM [2293], pages 70–79. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Jayram:2006:CPR

[2000] T. S. Jayram, Phokion G. Kolaitis, and Erik Vee. The containment problem for Real conjunctive queries with inequalities. In ACM [2293], pages 80–89. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Deutsch:2006:VCD

[2001] Alin Deutsch, Liying Sui, Victor Vianu, and Dayou Zhou. Verification of communicating data-driven Web services.

In ACM [2293], pages 90–99. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Calders:2006:AWI

[2002] Toon Calders, Stijn Dekeyser, Jan Hidders, and Jan Paredaens. Analyzing workflows implied by instancedependent access rules. In ACM [2293], pages 100–109. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Halpern:2006:SKB

[2003] Joseph Y. Halpern. From statistical knowledge bases to degrees of belief: an overview. In ACM [2293], pages 110–113. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Kolahi:2006:RVD

[2004] Solmaz Kolahi and Leonid Libkin. On redundancy vs dependency preservation in normalization: an informationtheoretic study of 3NF. In ACM [2293], pages 114–123. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Gottlob:2006:TDD

[2005] Georg Gottlob, Reinhard Pichler, and Fang Wei. Tractable database design through bounded treewidth. In ACM [2293], pages 124–133. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Motwani:2006:EPP

[2006] Rajeev Motwani and Ying Xu. Evolution of page popularity under random Web graph models. In ACM [2293], pages 134–142. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Mishra:2006:PPS

[2007] Nina Mishra and Mark Sandler. Privacy via pseudorandom sketches. In ACM [2293], pages 143–152. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Aggarwal:2006:AAC

[2008] Gagan Aggarwal, Tomás Feder, Krishnaram Kenthapadi, Samir Khuller, Rina Panigrahy, Dilys Thomas, and An Zhu. Achieving anonymity via clustering. In ACM [2293], pages 153– 162. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Machanavajjhala:2006:ECP

[2009] Ashwin Machanavajjhala and Johannes Gehrke. On the efficiency of checking perfect privacy. In ACM [2293], pages 163–172. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Kimelfeld:2006:FAT

[2010] Benny Kimelfeld and Yehoshua Sagiv. Finding and approximating top-k answers in keyword proximity search. In ACM [2293], pages 173–182. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Yang:2006:EOS

[2011] Guizhen Yang, Michael Kifer, and Vinay K. Chaudhri. Efficiently ordering subgoals with access constraints. In ACM [2293], pages 183–192. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Condon:2006:FAT

[2012] Anne Condon, Amol Deshpande, Lisa Hellerstein, and Ning Wu. Flow algorithms for two pipelined filter ordering problems. In ACM [2293], pages 193–202. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Goel:2006:ARQ

[2013] Ashish Goel, Sudipto Guha, and Kamesh Munagala. Asking the right questions: model-driven optimization using probes. In ACM [2293], pages 203–212. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Koch:2006:PQT

[2014] Christoph Koch. Processing queries on tree-structured data efficiently. In ACM [2293], pages 213–224. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Pagh:2006:SCA

[2015] Anna Pagh and Rasmus Pagh. Scalable computation of acyclic joins. In ACM [2293], pages 225–232. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Bender:2006:COS

[2016] Michael A. Bender, Martin Farach-Colton, and Bradley C. Kuszmaul. Cache-oblivious string B-trees. In ACM [2293], pages 233–242. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Grohe:2006:RCL

[2017] Martin Grohe, André Hernich, and Nicole Schweikardt. Randomized computations on large data sets: tight lower bounds. In ACM [2293], pages 243–252. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Buriol:2006:CTD

[2018] Luciana S. Buriol, Gereon Frahling, Stefano Leonardi, Alberto Marchetti-Spaccamela, and Christian Sohler. Counting triangles in data streams. In ACM [2293], pages 253–262. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Cormode:2006:STE

[2019] Graham Cormode, Flip Korn, S. Muthukrishnan, and Divesh Srivastava. Spaceand time-efficient deterministic algorithms for biased quantiles over data streams. In ACM [2293], pages 263– 272. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Guha:2006:AQO

[2020] Sudipto Guha and Andrew McGregor. Approximate quantiles and the order of the stream. In ACM [2293], pages 273–279. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Ganguly:2006:DSS

[2021] Sumit Ganguly and Anirban Majumder. Deterministic k-set structure. In ACM [2293], pages 280–289. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Lee:2006:SME

[2022] L. K. Lee and H. F. Ting. A simpler and more efficient deterministic scheme for finding frequent items over sliding windows. In ACM [2293], pages 290–297. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Zhao:2006:FGI

[2023] Qi (George) Zhao, Mitsunori Ogihara, Haixun Wang, and Jun (Jim) Xu. Finding global icebergs over distributed data sets. In ACM [2293], pages 298– 307. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Kashyap:2006:EGB

[2024] Srinivas Kashyap, Supratim Deb, K. V. M. Naidu, Rajeev Rastogi, and Anand Srinivasan. Efficient gossipbased aggregate computation. In ACM [2293], pages 308–317. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Gyssens:2006:SCS

[2025] Marc Gyssens, Jan Paredaens, Dirk Van Gucht, and George H. L. Fletcher. Structural characterizations of the semantics of XPath as navigation tool on a document. In ACM [2293], pages 318–327. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

tenCate:2006:EXT

[2026] Balder ten Cate. The expressivity of XPath with transitive closure. In ACM [2293], pages 328–337. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Bohannon:2006:RLL

[2027] Aaron Bohannon, Benjamin C. Pierce, and Jeffrey A. Vaughan. Relational lenses: a language for updatable views. In ACM [2293], pages 338–347. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Gollapudi:2006:PC

[2028] Sreenivas Gollapudi, Ravi Kumar, and D. Sivakumar. Programmable clustering. In ACM [2293], pages 348– 354. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Franconi:2006:LRS

[2029] Enrico Franconi and Sergio Tessaris.

The logic of RDF and SPARQL: a tutorial. In ACM [2293], page 355. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Rosati:2006:DFC

[2030] Riccardo Rosati. On the decidability and finite controllability of query processing in databases with incomplete information. In ACM [2293], pages 356–365. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

Dalvi:2007:MPD

[2031] Nilesh Dalvi and Dan Suciu. Management of probabilistic data: foundations and challenges. In ACM [2294], pages 1–12. ISBN 1-59593-685-8. LCCN ????

Gottlob:2007:GHD

[2032] Georg Gottlob, Zoltan Miklos, and Thomas Schwentick. Generalized hypertree decompositions: NP-hardness and tractable variants. In ACM [2294], pages 13–22. ISBN 1-59593-685-8. LCCN ????

Marx:2007:QDV

[2033] Maarten Marx. Queries determined by views: pack your views. In ACM [2294], pages 23–30. ISBN 1-59593-685-8. LCCN ????

Green:2007:PS

[2034] Todd J. Green, Grigoris Karvounarakis, and Val Tannen. Provenance semirings. In ACM [2294], pages 31–40. ISBN 1-59593-685-8. LCCN ????

Schweikardt:2007:MML

[2035] Nicole Schweikardt. Machine models and lower bounds for query processing. In ACM [2294], pages 41–52. ISBN 1-59593-685-8. LCCN ????

Chakaravarthy:2007:DTE

[2036] Venkatesan T. Chakaravarthy, Vinayaka Pandit, Sambuddha Roy, Pranjal Awasthi, and Mukesh Mohania. Decision trees for entity identification: approximation algorithms and hardness results. In ACM [2294], pages 53–62. ISBN 1-59593-685-8. LCCN ????

Engelfriet:2007:XTT

[2037] Joost Engelfriet, Hendrik Jan Hoogeboom, and Bart Samwel. XML transformation by tree-walking transducers with invisible pebbles. In ACM [2294], pages 63–72. ISBN 1-59593-685-8. LCCN ????

tenCate:2007:CQC

[2038] Balder ten Cate and Carsten Lutz. The complexity of query containment in expressive fragments of XPath 2.0. In ACM [2294], pages 73–82. ISBN 1-59593-685-8. LCCN ????

Fan:2007:ECX

[2039] Wenfei Fan, Floris Geerts, and Frank Neven. Expressiveness and complexity of XML publishing transducers. In ACM [2294], pages 83–92. ISBN 1-59593-685-8. LCCN ????

Gemulla:2007:MBS

[2040] Rainer Gemulla, Wolfgang Lehner, and Peter J. Haas. Maintaining Bernoulli samples over evolving multisets. In ACM [2294], pages 93–102. ISBN 1-59593-685-8. LCCN ????

Chierichetti:2007:FNN

[2041] Flavio Chierichetti, Alessandro Panconesi, Prabhakar Raghavan, Mauro Sozio, Alessandro Tiberi, and Eli Upfal. Finding near neighbors through cluster pruning. In ACM [2294], pages 103–112. ISBN 1-59593-685-8. LCCN ????

Hernich:2007:CSD

[2042] Andre Hernich and Nicole Schweikardt. CWA-solutions for data exchange settings with target dependencies. In ACM [2294], pages 113–122. ISBN 1-59593-685-8. LCCN ????

Fagin:2007:QIS

[2043] Ronald Fagin, Phokion G. Kolaitis, Lucian Popa, and Wang-Chiew Tan. Quasi-inverses of schema mappings. In ACM [2294], pages 123–132. ISBN 1-59593-685-8. LCCN ????

DeGiacomo:2007:RDE

[2044] Giuseppe De Giacomo, Domenico Lembo, Maurizio Lenzerini, and Riccardo Rosati. On reconciling data exchange, data integration, and peer data management. In ACM [2294], pages 133–142. ISBN 1-59593-685-8. LCCN ????

VandenBussche:2007:CCD

[2045] Jan Van den Bussche, Dirk Van Gucht, and Stijn Vansummeren. A crash course on database queries. In ACM [2294], pages 143–154. ISBN 1-59593-685-8. LCCN ????

Kasneci:2007:CRA

[2046] Gjergji Kasneci and Thomas Schwentick. The complexity of reasoning about pattern-based XML schemas. In ACM [2294], pages 155–164. ISBN 1-59593-685-8. LCCN ????

Gottlob:2007:MDF

[2047] Georg Gottlob, Reinhard Pichler, and Fang Wei. Monadic datalog over finite structures with bounded treewidth. In ACM [2294], pages 165–174. ISBN 1-59593-685-8. LCCN ????

Machlin:2007:IBM

[2048] Rona Machlin. Index-based multidimensional array queries: safety and equivalence. In ACM [2294], pages 175– 184. ISBN 1-59593-685-8. LCCN ????

Badia:2007:NLP

[2049] Antonio Badia and Stijn Vansummeren. Non-linear prefixes in query languages. In ACM [2294], pages 185– 194. ISBN 1-59593-685-8. LCCN ????

Cautis:2007:RAX

[2050] Bogdan Cautis, Serge Abiteboul, and Tova Milo. Reasoning about XML update constraints. In ACM [2294], pages 195–204. ISBN 1-59593-685-8. LCCN ????

Filiot:2007:PTF

[2051] Emmanuel Filiot, Joachim Niehren, Jean-Marc Talbot, and Sophie Tison. Polynomial time fragments of XPath with variables. In ACM [2294], pages 205–214. ISBN 1-59593-685-8. LCCN ????

Munagala:2007:OCQ

[2052] Kamesh Munagala, Utkarsh Srivastava, and Jennifer Widom. Optimization of continuous queries with shared expensive filters. In ACM [2294], pages 215– 224. ISBN 1-59593-685-8. LCCN ????

Zhang:2007:VES

[2053] Linfeng Zhang and Yong Guan. Variance estimation over sliding windows. In ACM [2294], pages 225–232. ISBN 1-59593-685-8. LCCN ????

Alur:2007:MWT

[2054] Rajeev Alur. Marrying words and trees. In ACM [2294], pages 233–242. ISBN 1-59593-685-8. LCCN ????

Jayram:2007:ESA

[2055] T. S. Jayram, Andrew McGregor, S. Muthukrishnan, and Erik Vee. Estimating statistical aggregates on probabilistic data streams. In ACM [2294], pages 243–252. ISBN 1-59593-685-8. LCCN ????

Cohen:2007:SUD

[2056] Edith Cohen, Nick Duffield, Haim Kaplan, Carsten Lund, and Mikkel Thorup. Sketching unaggregated data streams for subpopulation-size queries. In ACM [2294], pages 253–262. ISBN 1-59593-685-8. LCCN ????

White:2007:WNE

[2057] Walker White, Mirek Riedewald, Johannes Gehrke, and Alan Demers. What is ‘next’ in event processing? In ACM [2294], pages 263–272. ISBN 1-59593-685-8. LCCN ????

Barak:2007:PAC

[2058] Boaz Barak, Kamalika Chaudhuri, Cynthia Dwork, Satyen Kale, Frank McSherry, and Kunal Talwar. Privacy, accuracy, and consistency too: a holistic solution to contingency table release. In ACM [2294], pages 273–282. ISBN 1-59593-685-8. LCCN ????

Senellart:2007:CMP

[2059] Pierre Senellart and Serge Abiteboul. On the complexity of managing probabilistic XML data. In ACM [2294], pages 283–292. ISBN 1-59593-685-8. LCCN ????

Dalvi:2007:DCQ

[2060] Nilesh Dalvi and Dan Suciu. The dichotomy of conjunctive queries on probabilistic structures. In ACM [2294], pages 293–302. ISBN 1-59593-685-8. LCCN ????

Kimelfeld:2007:MJP

[2061] Benny Kimelfeld and Yehoshua Sagiv.

Maximally joining probabilistic data. In ACM [2294], pages 303–312. ISBN 1-59593-685-8. LCCN ????

Buneman:2008:CD

[2062] Peter Buneman, James Cheney, Wang-Chiew Tan, and Stijn Vansummeren. Curated databases. In Lenzerini and Lembo [2295], pages 1–12. ISBN 1-60560-932-3. LCCN ????

Arenas:2008:RSM

[2063] Marcelo Arenas, Jorge Pérez, and Cristian Riveros. The recovery of a schema mapping: bringing exchanged data back. In Lenzerini and Lembo [2295], pages 13–22. ISBN 1-60560-932-3. LCCN ????

Senellart:2008:CDS

[2064] Pierre Senellart and Georg Gottlob.

On the complexity of deriving schema mappings from database instances. In Lenzerini and Lembo [2295], pages 23– 32. ISBN 1-60560-932-3. LCCN ????

Fagin:2008:TTS

[2065] Ronald Fagin, Phokion G. Kolaitis, Alan Nash, and Lucian Popa. Towards a theory of schema-mapping optimization. In Lenzerini and Lembo [2295], pages 33–42. ISBN 1-60560-932-3. LCCN ????

Schnaitter:2008:ERJ

[2066] Karl Schnaitter and Neoklis Polyzotis. Evaluating rank joins with optimal cost. In Lenzerini and Lembo [2295], pages 43–52. ISBN 1-60560-932-3. LCCN ????

Bojannzyk:2008:ECT

[2067] Mikolaj Bojanńzyk. Effective characterizations of tree logics. In Lenzerini and Lembo [2295], pages 53–66. ISBN 1-60560-932-3. LCCN ????

Sarma:2008:EPG

[2068] Atish Das Sarma, Sreenivas Gollapudi, and Rina Panigrahy. Estimating PageRank on graph streams. In Lenzerini and Lembo [2295], pages 69–78. ISBN 1-60560-932-3. LCCN ????

Liu:2008:GFA

[2069] Zhen Liu, Srinivasan Parthasarathy, Anand Ranganathan, and Hao Yang. A generic flow algorithm for shared filter ordering problems. In Lenzerini and Lembo [2295], pages 79–88. ISBN 1-60560-932-3. LCCN ????

Cormode:2008:TDA

[2070] Graham Cormode, Flip Korn, and Srikanta Tirthapura. Time-decaying aggregates in out-of-order streams. In Lenzerini and Lembo [2295], pages 89– 98. ISBN 1-60560-932-3. LCCN ????

Koch:2008:APE

[2071] Christoph Koch. Approximating predicates and expressive queries on probabilistic databases. In Lenzerini and Lembo [2295], pages 99–108. ISBN 1-60560-932-3. LCCN ????

Cohen:2008:ICP

[2072] Sara Cohen, Benny Kimelfeld, and Yehoshua Sagiv. Incorporating constraints in probabilistic XML. In Lenzerini and Lembo [2295], pages 109–118. ISBN 1-60560-932-3. LCCN ????

Jha:2008:QES

[2073] Abhay Jha, Vibhor Rastogi, and Dan Suciu. Query evaluation with soft-key constraints. In Lenzerini and Lembo [2295], pages 119–128. ISBN 1-60560-932-3. LCCN ????

Afrati:2008:AAQ

[2074] Foto Afrati and Phokion G. Kolaitis.

Answering aggregate queries in data exchange. In Lenzerini and Lembo [2295], pages 129–138. ISBN 1-60560-932-3. LCCN ????

Libkin:2008:DES

[2075] Leonid Libkin and Cristina Sirangelo.

Data exchange and schema mappings in open and closed worlds. In Lenzerini and Lembo [2295], pages 139–148. ISBN 1-60560-932-3. LCCN ????

Deutsch:2008:CR

[2076] Alin Deutsch, Alan Nash, and Jeff Remmel. The chase revisited. In Lenzerini and Lembo [2295], pages 149–158. ISBN 1-60560-932-3. LCCN ????

Fan:2008:DRI

[2077] Wenfei Fan. Dependencies revisited for improving data quality. In Lenzerini and Lembo [2295], pages 159–170. ISBN 1-60560-932-3. LCCN ????

Evfimievski:2008:EP

[2078] Alexandre Evfimievski, Ronald Fagin, and David P. Woodruff. Epistemic privacy. In Lenzerini and Lembo [2295], pages 171–180. ISBN 1-60560-932-3. LCCN ????

Ferragina:2008:SCS

[2079] Paolo Ferragina, Roberto Grossi, Ankur Gupta, Rahul Shah, and Jeffrey Scott Vitter. On searching compressed string collections cacheobliviously. In Lenzerini and Lembo [2295], pages 181–190. ISBN 1-60560-932-3. LCCN ????

Cormode:2008:AAC

[2080] Graham Cormode and Andrew Mc-Gregor. Approximation algorithms for clustering uncertain data. In Lenzerini and Lembo [2295], pages 191–200. ISBN 1-60560-932-3. LCCN ????

Anagnostopoulos:2008:AAC

[2081] Aris Anagnostopoulos, Anirban Dasgupta, and Ravi Kumar. Approximation algorithms for co-clustering. In Lenzerini and Lembo [2295], pages 201–210. ISBN 1-60560-932-3. LCCN ????

Gollapudi:2008:PTM

[2082] Sreenivas Gollapudi and Rina Panigrahy. The power of two min-hashes for similarity search among hierarchical data objects. In Lenzerini and Lembo [2295], pages 211–220. ISBN 1-60560-932-3. LCCN ????

Abiteboul:2008:SAA

[2083] Serge Abiteboul, Luc Segoufin, and Victor Vianu. Static analysis of active XML systems. In Lenzerini and Lembo [2295], pages 221–230. ISBN 1-60560-932-3. LCCN ????

Fan:2008:CCS

[2084] Wenfei Fan, Floris Geerts, Wouter Gelade, Frank Neven, and Antonella Poggi. Complexity and composition of synthesized web services. In Lenzerini and Lembo [2295], pages 231–240. ISBN 1-60560-932-3. LCCN ????

Bojanczyk:2008:XEL

[2085] Mikolaj Bojańczyk and Pawel Parys.

XPath evaluation in linear time. In Lenzerini and Lembo [2295], pages 241–250. ISBN 1-60560-932-3. LCCN ????

tenCate:2008:XTC

[2086] Balder ten Cate and Luc Segoufin.

XPath, transitive closure logic, and nested tree walking automata. In Lenzerini and Lembo [2295], pages 251–260. ISBN 1-60560-932-3. LCCN ????

Gardner:2008:LHR

[2087] Philippa A. Gardner, Gareth D. Smith, Mark J. Wheelhouse, and Uri D. Zarfaty. Local Hoare reasoning about DOM. In Lenzerini and Lembo [2295], pages 261–270. ISBN 1-60560-932-3. LCCN ????

Foster:2008:AXQ

[2088] J. Nathan Foster, Todd J. Green, and Val Tannen. Annotated XML: queries and provenance. In Lenzerini and Lembo [2295], pages 271–280. ISBN 1-60560-932-3. LCCN ????

Sozio:2008:NOD

[2089] Mauro Sozio, Thomas Neumann, and Gerhard Weikum. Near-optimal dynamic replication in unstructured peerto-peer networks. In Lenzerini and Lembo [2295], pages 281–290. ISBN 1-60560-932-3. LCCN ????

deMoor:2008:TID

[2090] Oege de Moor, Damien Sereni, Pavel Avgustinov, and Mathieu Verbaere.

Type inference for datalog and its application to query optimisation. In Lenzerini and Lembo [2295], pages 291–300. ISBN 1-60560-932-3. LCCN ????

Sharfman:2008:SSG

[2091] Izchak Sharfman, Assaf Schuster, and Daniel Keren. Shape sensitive geometric monitoring. In Lenzerini and Lembo [2295], pages 301–310. ISBN 1-60560-932-3. LCCN ????

Adler:2008:TWF

[2092] Isolde Adler. Tree-width and functional dependencies in databases. In Lenzerini and Lembo [2295], pages 311–320. ISBN 1-60560-932-3. LCCN ????

Dalvi:2009:OIT

[2093] Nilesh Dalvi, Ravi Kumar, Bo Pang, Raghu Ramakrishnan, Andrew Tomkins, Philip Bohannon, Sathiya Keerthi, and Srujana Merugu. Opening and invited talk: a web of concepts. In Paredaens and Su [2296], pages 1–12. ISBN 1-60558-553-X. LCCN ????

Marnette:2009:GSM

[2094] Bruno Marnette. Generalized schemamappings: from termination to tractability. In Paredaens and Su [2296], pages 13–22. ISBN 1-60558-553-X. LCCN ????

Fagin:2009:RDE

[2095] Ronald Fagin, Phokion G. Kolaitis, Lucian Popa, and Wang-Chiew Tan. Reverse data exchange: coping with nulls. In Paredaens and Su [2296], pages 23– 32. ISBN 1-60558-553-X. LCCN ????

Amano:2009:XSM

[2096] Shun’ichi Amano, Leonid Libkin, and Filip Murlak. XML schema mappings.

In Paredaens and Su [2296], pages 33– 42. ISBN 1-60558-553-X. LCCN ????

Beeri:2009:APA

[2097] Catriel Beeri, Phokion G. Kolaitis, and Christos H. Papadimitriou. The ACM PODS alberto O. mendelzon test-of-time-award 2009. In Paredaens and Su [2296], page 43. ISBN 1-60558-553-X. LCCN ????

Gottlob:2009:STB

[2098] Georg Gottlob, Stephanie Tien Lee, and Gregory J. Valiant. Size and treewidth bounds for conjunctive queries. In Paredaens and Su [2296], pages 45–54. ISBN 1-60558-553-X. LCCN ????

Parys:2009:XEL

[2099] Pawel Parys. XPath evaluation in linear time with polynomial combined complexity. In Paredaens and Su [2296], pages 55–64. ISBN 1-60558-553-X. LCCN ????

Libkin:2009:FMT

[2100] Leonid Libkin. The finite model theory toolbox of a database theoretician. In Paredaens and Su [2296], pages 65–76. ISBN 1-60558-553-X. LCCN ????

Cali:2009:GDB

[2101] Andrea Calì, Georg Gottlob, and Thomas Lukasiewicz. A general datalog-based framework for tractable query answering over ontologies. In Paredaens and Su [2296], pages 77–86. ISBN 1-60558-553-X. LCCN ????

Abiteboul:2009:SRQ

[2102] Serge Abiteboul, Pierre Bourhis, and Bogdan Marinoiu. Satisfiability and relevance for queries over active documents. In Paredaens and Su [2296], pages 87–96. ISBN 1-60558-553-X. LCCN ????

Fan:2009:RIC

[2103] Wenfei Fan and Floris Geerts. Relative information completeness. In Paredaens and Su [2296], pages 97–106. ISBN 1-60558-553-X. LCCN ????

Rastogi:2009:RPO

[2104] Vibhor Rastogi, Michael Hay, Gerome Miklau, and Dan Suciu. Relationship privacy: output perturbation for queries with joins. In Paredaens and Su [2296], pages 107–116. ISBN 1-60558-553-X. LCCN ????

Kirsch:2009:ERA

[2105] Adam Kirsch, Michael Mitzenmacher, Andrea Pietracaprina, Geppino Pucci, Eli Upfal, and Fabio Vandin. An efficient rigorous approach for identifying statistically significant frequent itemsets. In Paredaens and Su [2296], pages 117–126. ISBN 1-60558-553-X. LCCN ????

Chierichetti:2009:SC

[2106] Flavio Chierichetti, Ravi Kumar, and Sergei Vassilvitskii. Similarity caching. In Paredaens and Su [2296], pages 127– 136. ISBN 1-60558-553-X. LCCN ????

Agarwal:2009:IUD

[2107] Pankaj K. Agarwal, Siu-Wing Cheng, Yufei Tao, and Ke Yi. Indexing uncertain data. In Paredaens and Su [2296], pages 137–146. ISBN 1-60558-553-X. LCCN ????

Braverman:2009:OSS

[2108] Vladimir Braverman, Rafail Ostrovsky, and Carlo Zaniolo. Optimal sampling from sliding windows. In Paredaens and Su [2296], pages 147–156. ISBN 1-60558-553-X. LCCN ????

Berinde:2009:SOH

[2109] Radu Berinde, Graham Cormode, Piotr Indyk, and Martin J. Strauss. Space-optimal heavy hitters with strong error bounds. In Paredaens and Su [2296], pages 157–166. ISBN 1-60558-553-X. LCCN ????

Yi:2009:OTD

[2110] Ke Yi and Qin Zhang. Optimal tracking of distributed heavy hitters and quantiles. In Paredaens and Su [2296], pages 167–174. ISBN 1-60558-553-X. LCCN ????

Arge:2009:WCE

[2111] Lars Arge. Worst-case efficient range search indexing: invited tutorial. In Paredaens and Su [2296], pages 175– 176. ISBN 1-60558-553-X. LCCN ????

Pagh:2009:SIO

[2112] Rasmus Pagh and Srinivasa Rao Satti.

Secondary indexing in one dimension: beyond B-trees and bitmap indexes. In Paredaens and Su [2296], pages 177– 186. ISBN 1-60558-553-X. LCCN ????

Yi:2009:DIL

[2113] Ke Yi. Dynamic indexability and lower bounds for dynamic one-dimensional range query indexes. In Paredaens and Su [2296], pages 187–196. ISBN 1-60558-553-X. LCCN ????

Figueira:2009:SDX

[2114] Diego Figueira. Satisfiability of downward XPath with data equality tests. In Paredaens and Su [2296], pages 197– 206. ISBN 1-60558-553-X. LCCN ????

DeHaan:2009:ENQ

[2115] David DeHaan. Equivalence of nested queries with mixed semantics. In Paredaens and Su [2296], pages 207– 216. ISBN 1-60558-553-X. LCCN ????

Chirkova:2009:ESQ

[2116] Rada Chirkova and Michael R. Genesereth. Equivalence of SQL queries in presence of embedded dependencies. In Paredaens and Su [2296], pages 217– 226. ISBN 1-60558-553-X. LCCN ????

Cohen:2009:RTA

[2117] Sara Cohen, Benny Kimelfeld, and Yehoshua Sagiv. Running tree automata on probabilistic XML. In Paredaens and Su [2296], pages 227– 236. ISBN 1-60558-553-X. LCCN ????

Barcelo:2009:XII

[2118] Pablo Barceló, Leonid Libkin, Antonella Poggi, and Cristina Sirangelo. XML with incomplete information: models, properties, and query answering. In Paredaens and Su [2296], pages 237–246. ISBN 1-60558-553-X. LCCN ????

Abiteboul:2009:DXD

[2119] Serge Abiteboul, Georg Gottlob, and Marco Manna. Distributed XML design. In Paredaens and Su [2296], pages 247–258. ISBN 1-60558-553-X. LCCN ????

Li:2009:CAQ

[2120] Jian Li and Amol Deshpande. Consensus answers for queries over probabilistic databases. In Paredaens and Su [2296], pages 259–268. ISBN 1-60558-553-X. LCCN ????

Guha:2009:EEC

[2121] Sudipto Guha and Kamesh Munagala.

Exceeding expectations and clustering uncertain data. In Paredaens and Su [2296], pages 269–278. ISBN 1-60558-553-X. LCCN ????

Atallah:2009:CAS

[2122] Mikhail J. Atallah and Yinian Qi.

Computing all skyline probabilities for uncertain data. In Paredaens and Su [2296], pages 279–287. ISBN 1-60558-553-X. LCCN ????

Hellerstein:2010:DRE

[2123] Joseph M. Hellerstein. Datalog redux: experience and conjecture. In Van Gucht [2297], pages 1–2. ISBN 1-4503-0033-2. LCCN ????

Barcelo:2010:ELP

[2124] Pablo Barcelo, Carlos Hurtado, Leonid Libkin, and Peter Wood. Expressive languages for path queries over graph-structured data. In Van Gucht [2297], pages 3–14. ISBN 1-4503-0033-2. LCCN ????

Kimelfeld:2010:TMS

[2125] Benny Kimelfeld and Christopher Ré.

Transducing Markov sequences. In Van Gucht [2297], pages 15–26. ISBN 1-4503-0033-2. LCCN ????

Benedikt:2010:PHO

[2126] Michael Benedikt, Gabriele Puppis, and Huy Vu. Positive higher-order queries. In Van Gucht [2297], pages 27– 38. ISBN 1-4503-0033-2. LCCN ????

Su:2010:APA

[2127] Jianwen Su and Phokion G. Kolaitis.

The ACM PODS Alberto O. Mendelzon Test-of-Time-Award 2010. In Van Gucht [2297], pages 39–40. ISBN 1-4503-0033-2. LCCN ????

Kane:2010:OAD

[2128] Daniel M. Kane, Jelani Nelson, and David P. Woodruff. An optimal algorithm for the distinct elements problem. In Van Gucht [2297], pages 41–52. ISBN 1-4503-0033-2. LCCN ????

Re:2010:UCE

[2129] Christopher Ré and Dan Suciu. Understanding cardinality estimation using entropy maximization. In Van Gucht [2297], pages 53–64. ISBN 1-4503-0033-2. LCCN ????

Weikum:2010:IKH

[2130] Gerhard Weikum and Martin Theobald. From information to knowledge: harvesting entities and relationships from Web sources. In Van Gucht [2297], pages 65–76. ISBN 1-4503-0033-2. LCCN ????

Cormode:2010:OSD

[2131] Graham Cormode, S. Muthukrishnan, Ke Yi, and Qin Zhang. Optimal sampling from distributed streams. In Van Gucht [2297], pages 77–86. ISBN 1-4503-0033-2. LCCN ????

Koch:2010:IQE

[2132] Christoph Koch. Incremental query evaluation in a ring of databases. In Van Gucht [2297], pages 87–98. ISBN 1-4503-0033-2. LCCN ????

Nelson:2010:FMS

[2133] Jelani Nelson and David P. Woodruff. Fast Manhattan sketches in data streams. In Van Gucht [2297], pages 99–110. ISBN 1-4503-0033-2. LCCN ????

Meier:2010:SQO

[2134] Michael Meier, Michael Schmidt, Fang Wei, and Georg Lausen. Semantic query optimization in the presence of types. In Van Gucht [2297], pages 111– 122. ISBN 1-4503-0033-2. LCCN ????

Li:2010:OLC

[2135] Chao Li, Michael Hay, Vibhor Rastogi, Gerome Miklau, and Andrew McGregor. Optimizing linear counting queries under differential privacy. In Van Gucht [2297], pages 123–134. ISBN 1-4503-0033-2. LCCN ????

Gupte:2010:UOP

[2136] Mangesh Gupte and Mukund Sundararajan. Universally optimal privacy mechanisms for minimax agents. In Van Gucht [2297], pages 135–146. ISBN 1-4503-0033-2. LCCN ????

Kifer:2010:TAS

[2137] Daniel Kifer and Bing-Rong Lin. Towards an axiomatization of statistical privacy and utility. In Van Gucht [2297], pages 147–158. ISBN 1-4503-0033-2. LCCN ????

Jayram:2010:ICT

[2138] T. S. Jayram. Information complexity:

a tutorial. In Van Gucht [2297], pages 159–168. ISBN 1-4503-0033-2. LCCN ????

Fan:2010:CMT

[2139] Wenfei Fan and Floris Geerts. Capturing missing tuples and missing values. In Van Gucht [2297], pages 169–178. ISBN 1-4503-0033-2. LCCN ????

Wijsen:2010:FOE

[2140] Jef Wijsen. On the first-order expressibility of computing certain answers to conjunctive queries over uncertain databases. In Van Gucht [2297], pages 179–190. ISBN 1-4503-0033-2. LCCN ????

David:2010:CAX

[2141] Claire David, Leonid Libkin, and Filip Murlak. Certain answers for XML queries. In Van Gucht [2297], pages 191–202. ISBN 1-4503-0033-2. LCCN ????

Dalvi:2010:CQP

[2142] Nilesh Dalvi, Karl Schnaitter, and Dan Suciu. Computing query probability with incidence algebras. In Van Gucht [2297], pages 203–214. ISBN 1-4503-0033-2. LCCN ????

Deutch:2010:PFM

[2143] Daniel Deutch, Christoph Koch, and Tova Milo. On probabilistic fixpoint and Markov chain query languages. In Van Gucht [2297], pages 215–226. ISBN 1-4503-0033-2. LCCN ????

Arenas:2010:FSM

[2144] Marcelo Arenas, Jorge Pérez, Juan L. Reutter, and Cristian Riveros. Foundations of schema mapping management. In Van Gucht [2297], pages 227–238. ISBN 1-4503-0033-2. LCCN ????

Martens:2010:SDX

[2145] Wim Martens, Matthias Niewerth, and Thomas Schwentick. Schema design for XML repositories: complexity and tractability. In Van Gucht [2297], pages 239–250. ISBN 1-4503-0033-2. LCCN ????

Gelade:2010:SXS

[2146] Wouter Gelade, Tomasz Idziaszek, Wim Martens, and Frank Neven. Simplifying XML schema: single-type approximations of regular tree languages. In Van Gucht [2297], pages 251–260. ISBN 1-4503-0033-2. LCCN ????

Alexe:2010:CSM

[2147] Bogdan Alexe, Phokion G. Kolaitis, and Wang-Chiew Tan. Characterizing schema mappings via data examples. In Van Gucht [2297], pages 261–272. ISBN 1-4503-0033-2. LCCN ????

Fagin:2010:UQS

[2148] Ronald Fagin, Benny Kimelfeld, Yunyao Li, Sriram Raghavan, and Shivakumar Vaithyanathan. Understanding queries in a search database system. In Van Gucht [2297], pages 273– 284. ISBN 1-4503-0033-2. LCCN ????

Lemay:2010:LAT

[2149] Aurelien Lemay, Sebastian Maneth, and Joachim Niehren. A learning algorithm for top-down XML transformations. In Van Gucht [2297], pages 285– 296. ISBN 1-4503-0033-2. LCCN ????

Pagh:2010:COH

[2150] Rasmus Pagh, Zhewei Wei, Ke Yi, and Qin Zhang. Cache-oblivious hashing. In Van Gucht [2297], pages 297–304. ISBN 1-4503-0033-2. LCCN ????

Bender:2010:PGB

[2151] Michael A. Bender, Haodong Hu, and Bradley C. Kuszmaul. Performance guarantees for B-trees with differentsized atomic keys. In Van Gucht [2297], pages 305–316. ISBN 1-4503-0033-2. LCCN ????

Hartmann:2010:WDD

[2152] Sven Hartmann and Sebastian Link.

When data dependencies over SQL tables meet the logics of paradox and S-3. In Van Gucht [2297], pages 317–326. ISBN 1-4503-0033-2. LCCN ????

Greco:2010:PTP

[2153] Gianluigi Greco and Francesco Scarcello. The power of tree projections: local consistency, greedy algorithms, and larger islands of tractability. In Van Gucht [2297], pages 327–338. ISBN 1-4503-0033-2. LCCN ????

Deutch:2011:QBW

[2154] Daniel Deutch and Tova Milo. A quest for beauty and wealth (or, business processes for database researchers). In Lenzerini [2298], pages 1–12. ISBN 1-4503-0660-8. LCCN ????

URL http://dl.acm.org/citation.
cfm?id=1989284.

Cohen:2011:GMY

[2155] Edith Cohen and Haim Kaplan. Get the most out of your sample: optimal unbiased estimators using partial information. In Lenzerini [2298], pages 13– 24. ISBN 1-4503-0660-8. LCCN ????

URL http://dl.acm.org/citation.
cfm?id=1989284.

Sheng:2011:FID

[2156] Cheng Sheng and Yufei Tao. FIFO indexes for decomposable problems. In Lenzerini [2298], pages 25–35. ISBN 1-4503-0660-8. LCCN ????

URL http://dl.acm.org/citation.
cfm?id=1989284.

Mir:2011:PPA

[2157] Darakhshan Mir, S. Muthukrishnan, Aleksandar Nikolov, and Rebecca N. Wright. Pan-private algorithms via statistics on sketches. In Lenzerini [2298], pages 37–48. ISBN 1-4503-0660-

8. LCCN ???? URL http://dl.acm.
org/citation.cfm?id=1989284.

Jowhari:2011:TBS

[2158] Hossein Jowhari, Mert Sağlam, and Gábor Tardos. Tight bounds for L_{p} samplers, finding duplicates in streams, and related problems. In Lenzerini [2298], pages 49–58. ISBN 1-4503-0660-

8. LCCN ???? URL http://dl.acm.
org/citation.cfm?id=1989284.

Libkin:2011:IIC

[2159] Leonid Libkin. Incomplete information and certain answers in general data models. In Lenzerini [2298], pages 59– 70. ISBN 1-4503-0660-8. LCCN ????

URL http://dl.acm.org/citation.
cfm?id=1989284.

Fan:2011:DCD

[2160] Wenfei Fan, Floris Geerts, and Jef Wijsen. Determining the currency of data. In Lenzerini [2298], pages 71– 82. ISBN 1-4503-0660-8. LCCN ????

URL http://dl.acm.org/citation.
cfm?id=1989284.

Arenas:2011:DEB

[2161] Marcelo Arenas, Jorge Pérez, and Juan Reutter. Data exchange beyond complete data. In Lenzerini [2298], pages 83–94. ISBN 1-4503-0660-8. LCCN

????  URL http://dl.acm.org/
citation.cfm?id=1989284.

Orlandi:2011:SES

[2162] Alessio Orlandi and Rossano Venturini.

Space-efficient substring occurrence estimation. In Lenzerini [2298], pages 95– 106. ISBN 1-4503-0660-8. LCCN ????

URL http://dl.acm.org/citation.
cfm?id=1989284.

Sheng:2011:FSE

[2163] Cheng Sheng and Yufei Tao. On finding skylines in external memory. In Lenzerini [2298], pages 107–116. ISBN 1-4503-0660-8. LCCN ????

                                        URL http://dl.acm.org/citation.
                                        cfm?id=1989284.

Wei:2011:BSA

[2164] Zhewei Wei and Ke Yi. Beyond simple aggregates: indexing for summary queries. In Lenzerini [2298], pages 117– 128. ISBN 1-4503-0660-8. LCCN ????

                                        URL http://dl.acm.org/citation.
                                        cfm?id=1989284.

Sheng:2011:NRT

[2165] Cheng Sheng and Yufei Tao. New results on two-dimensional orthogonal range aggregation in external memory. In Lenzerini [2298], pages 129– 139. ISBN 1-4503-0660-8. LCCN ????

                                        URL http://dl.acm.org/citation.
                                        cfm?id=1989284.

Amsterdamer:2011:PM

[2166] Yael Amsterdamer, Daniel Deutch, Tova Milo, and Val Tannen. On provenance minimization. In Lenzerini [2298], pages 141–152. ISBN 1-4503-0660-8. LCCN ???? URL http://dl.

                                        acm.org/citation.cfm?id=1989284.

Amsterdamer:2011:PAQ

[2167] Yael Amsterdamer, Daniel Deutch, and Val Tannen. Provenance for aggregate queries. In Lenzerini [2298], pages 153– 164. ISBN 1-4503-0660-8. LCCN ????

                                        URL http://dl.acm.org/citation.
                                        cfm?id=1989284.

He:2011:CPP

[2168] Yeye He, Siddharth Barman, Di Wang, and Jeffrey F. Naughton. On the complexity of privacy-preserving complex event processing. In Lenzerini [2298], pages 165–174. ISBN 1-4503-0660-8.

LCCN ???? URL http://dl.acm.
org/citation.cfm?id=1989284.

Davidson:2011:PVM

[2169] Susan B. Davidson, Sanjeev Khanna, Tova Milo, Debmalya Panigrahi, and Sudeepa Roy. Provenance views for module privacy. In Lenzerini [2298], pages 175–186. ISBN 1-4503-0660-8.

LCCN ???? URL http://dl.acm.
org/citation.cfm?id=1989284.

Kimelfeld:2011:MCV

[2170] Benny Kimelfeld, Jan Vondrák, and Ryan Williams. Maximizing conjunctive views in deletion propagation. In Lenzerini [2298], pages 187– 198. ISBN 1-4503-0660-8. LCCN ????

URL http://dl.acm.org/citation.
cfm?id=1989284.

Barcelo:2011:QGP

[2171] Pablo Barceló, Leonid Libkin, and Juan L. Reutter. Querying graph patterns. In Lenzerini [2298], pages 199– 210. ISBN 1-4503-0660-8. LCCN ????

URL http://dl.acm.org/citation.
cfm?id=1989284.

Benedikt:2011:DRA

[2172] Michael Benedikt, Georg Gottlob, and Pierre Senellart. Determining relevance of accesses at runtime. In Lenzerini [2298], pages 211–222. ISBN 1-4503-0660-8. LCCN ???? URL http://dl.

acm.org/citation.cfm?id=1989284.

Koutris:2011:PEC

[2173] Paraschos Koutris and Dan Suciu. Parallel evaluation of conjunctive queries.

In Lenzerini [2298], pages 223–234. ISBN 1-4503-0660-8. LCCN ????

                                        URL http://dl.acm.org/citation.
                                        cfm?id=1989284.

Kimelfeld:2011:FMT

[2174] Benny Kimelfeld and Yehoshua Sagiv. Finding a minimal tree pattern under neighborhood constraints. In Lenzerini [2298], pages 235–246. ISBN 1-4503-0660-8. LCCN ???? URL http://dl.

                                        acm.org/citation.cfm?id=1989284.

Antonopoulos:2011:CTP

[2175] Timos Antonopoulos, Wim Martens, and Frank Neven. The complexity of text-preserving XML transformations. In Lenzerini [2298], pages 247– 258. ISBN 1-4503-0660-8. LCCN ????

                                        URL http://dl.acm.org/citation.
                                        cfm?id=1989284.

Bojanczyk:2011:EET

[2176] Mikolaj Bojanczyk and Diego Figueira. Efficient evaluation for a temporal logic on changing XML documents. In Lenzerini [2298], pages 259–270. ISBN 1-4503-0660-8. LCCN ????

                                        URL http://dl.acm.org/citation.
                                        cfm?id=1989284.

Fagin:2011:RRS

[2177] Ronald Fagin, Benny Kimelfeld, Yunyao Li, Sriram Raghavan, and Shivakumar Vaithyanathan. Rewrite rules for search database systems. In Lenzerini [2298], pages 271–282. ISBN 1-4503-0660-8. LCCN ???? URL http://dl.

                                        acm.org/citation.cfm?id=1989284.

Ameloot:2011:RTD

[2178] Tom J. Ameloot, Frank Neven, and Jan Van den Bussche. Relational transducers for declarative networking. In Lenzerini [2298], pages 283– 292. ISBN 1-4503-0660-8. LCCN ????

URL http://dl.acm.org/citation.
cfm?id=1989284.

Abiteboul:2011:RBL

[2179] Serge Abiteboul, Meghyn Bienvenu, Alban Galland, and Émilien Antoine. A rule-based language for web data management. In Lenzerini [2298], pages 293–304. ISBN 1-4503-0660-8. LCCN

????  URL http://dl.acm.org/
citation.cfm?id=1989284.

Arenas:2011:QSW

[2180] Marcelo Arenas and Jorge Pérez.

Querying Semantic Web data with SPARQL. In Lenzerini [2298], pages 305–316. ISBN 1-4503-0660-8. LCCN

????  URL http://dl.acm.org/
citation.cfm?id=1989284.

Muthukrishnan:2011:TDS

[2181] S. Muthukrishnan. Theory of data stream computing: where to go. In Lenzerini [2298], pages 317–319. ISBN 1-4503-0660-8. LCCN ????

URL http://dl.acm.org/citation.
cfm?id=1989284.

Chaudhuri:2012:WNH

[2182] Surajit Chaudhuri. What next?: a half-dozen data management research goals for big data and the cloud. In Krötzsch et al. [2299], pages 1–4. ISBN ???? LCCN ???? URL http://www.

sigmod.org/2012/.

Ahn:2012:GSS

[2183] Kook Jin Ahn, Sudipto Guha, and Andrew McGregor. Graph sketches: sparsification, spanners, and subgraphs. In Krötzsch et al. [2299], pages 5–14.

ISBN ???? LCCN ???? URL http:

                                        //www.sigmod.org/2012/.

Indyk:2012:ATH

[2184] Piotr Indyk, Reut Levi, and Ronitt Rubinfeld. Approximating and testing k-histogram distributions in sub-linear time. In Krötzsch et al. [2299], pages 15–22. ISBN ???? LCCN ???? URL

                                        http://www.sigmod.org/2012/.

Agarwal:2012:MS

[2185] Pankaj K. Agarwal, Graham Cormode, Zengfeng Huang, Jeff Phillips, Zhewei Wei, and Ke Yi. Mergeable summaries. In Krötzsch et al. [2299], pages 23–34. ISBN ???? LCCN ???? URL http://

                                        www.sigmod.org/2012/.

Hull:2012:APA

[2186] Richard Hull, Phokion G. Kolaitis, and Dirk Van Gucht. The ACM PODS Alberto O. Mendelzon Test-of-Time Award 2012. In Krötzsch et al. [2299], pages 35–36. ISBN ???? LCCN ????

                                        URL http://www.sigmod.org/2012/.

Ngo:2012:WCO

[2187] Hung Q. Ngo, Ely Porat, Christopher Ré, and Atri Rudra. Worst-case optimal join algorithms: [extended abstract]. In Krötzsch et al. [2299], pages 37–48. ISBN ???? LCCN ???? URL

                                        http://www.sigmod.org/2012/.

Groz:2012:DRE

[2188] Benotît Groz, Sebastian Maneth, and Slawek Staworko. Deterministic regular expressions in linear time. In Krötzsch et al. [2299], pages 49–60. ISBN ???? LCCN ???? URL http:

                                        //www.sigmod.org/2012/.

Pierce:2012:LFB

[2189] Benjamin C. Pierce. Linguistic foundations for bidirectional transformations: invited tutorial. In Krötzsch et al. [2299], pages 61–64. ISBN ???? LCCN

???? URL http://www.sigmod.org/
2012/.

Choromanski:2012:PDN

[2190] Krzysztof Choromanski and Tal Malkin. The power of the Dinur– Nissim algorithm: breaking privacy of statistical and graph databases. In Krötzsch et al. [2299], pages 65–76. ISBN ???? LCCN ???? URL http://

www.sigmod.org/2012/.

Kifer:2012:RCF

[2191] Daniel Kifer and Ashwin Machanavajjhala. A rigorous and customizable framework for privacy. In Krötzsch et al. [2299], pages 77–88. ISBN ???? LCCN ???? URL http://www.

sigmod.org/2012/.

Letelier:2012:SAO

[2192] Andrés Letelier, Jorge Pérez, Reinhard Pichler, and Sebastian Skritek. Static analysis and optimization of Semantic Web queries. In Krötzsch et al. [2299], pages 89–100. ISBN ???? LCCN ????

URL http://www.sigmod.org/2012/.

Losemann:2012:CEP

[2193] Katja Losemann and Wim Martens.

The complexity of evaluating path expressions in SPARQL. In Krötzsch et al. [2299], pages 101–112. ISBN ???? LCCN ???? URL http://www.

sigmod.org/2012/.

Nekrich:2012:SER

[2194] Yakov Nekrich. Space-efficient range reporting for categorical data. In Krötzsch et al. [2299], pages 113–120. ISBN ???? LCCN ???? URL http://

                                        www.sigmod.org/2012/.

Sheng:2012:DTK

[2195] Cheng Sheng and Yufei Tao. Dynamic top-k range reporting in external memory. In Krötzsch et al. [2299], pages 121–130. ISBN ???? LCCN ???? URL

                                        http://www.sigmod.org/2012/.

Tao:2012:IRS

[2196] Yufei Tao. Indexability of 2D range search revisited: constant redundancy and weak indivisibility. In Krötzsch et al. [2299], pages 131–142. ISBN ???? LCCN ???? URL http://www.

                                        sigmod.org/2012/.

Mahoney:2012:ACI

[2197] Michael W. Mahoney. Approximate computation and implicit regularization for very large-scale data analysis. In Krötzsch et al. [2299], pages 143– 154. ISBN ???? LCCN ???? URL

                                        http://www.sigmod.org/2012/.

Borodin:2012:MSD

[2198] Allan Borodin, Hyun Chul Lee, and Yuli Ye. Max-Sum diversification, monotone submodular functions and dynamic updates. In Krötzsch et al. [2299], pages 155–166. ISBN ????

                                        LCCN ????  URL http://www.
                                        sigmod.org/2012/.

Koutris:2012:QBD

[2199] Paraschos Koutris, Prasang Upadhyaya, Magdalena Balazinska, Bill Howe, and Dan Suciu. Query-based data pricing. In Krötzsch et al. [2299], pages 167–178. ISBN ???? LCCN ????

URL http://www.sigmod.org/2012/.

Fagin:2012:LTC

[2200] Ronald Fagin and Phokion G. Kolaitis.

Local transformations and conjunctivequery equivalence. In Krötzsch et al. [2299], pages 179–190. ISBN ????

LCCN ????  URL http://www.
sigmod.org/2012/.

Kimelfeld:2012:DCD

[2201] Benny Kimelfeld. A dichotomy in the complexity of deletion propagation with functional dependencies. In Krötzsch et al. [2299], pages 191–202. ISBN ???? LCCN ???? URL http://

www.sigmod.org/2012/.

Grossi:2012:WTM

[2202] Roberto Grossi and Giuseppe Ottaviano. The wavelet trie: maintaining an indexed sequence of strings in compressed space. In Krötzsch et al. [2299], pages 203–214. ISBN ???? LCCN ????

URL http://www.sigmod.org/2012/.

Xu:2012:OCP

[2203] Pan Xu and Srikanta Tirthapura. On the optimality of clustering properties of space filling curves. In Krötzsch et al. [2299], pages 215–224. ISBN ???? LCCN ???? URL http://www.

sigmod.org/2012/.

Agarwal:2012:NNS

[2204] Pankaj K. Agarwal, Alon Efrat, Swaminathan Sankararaman, and Wuzhou Zhang. Nearest-neighbor searching under uncertainty. In Krötzsch et al. [2299], pages 225–236.

ISBN ???? LCCN ???? URL http://

                                        www.sigmod.org/2012/.

Kostylev:2012:CAS

[2205] Egor V. Kostylev, Juan L. Reutter, and András Z. Salamon. Classification of annotation semirings over query containment. In Krötzsch et al. [2299], pages 237–248. ISBN ???? LCCN ????

                                        URL http://www.sigmod.org/2012/.

Barcelo:2012:EAC

[2206] Pablo Barceló, Leonid Libkin, and Miguel Romero. Efficient approximations of conjunctive queries. In Krötzsch et al. [2299], pages 249–260. ISBN ???? LCCN ???? URL http://

                                        www.sigmod.org/2012/.

Deng:2012:CPR

[2207] Ting Deng, Wenfei Fan, and Floris Geerts. On the complexity of package recommendation problems. In Krötzsch et al. [2299], pages 261–272. ISBN ???? LCCN ???? URL http://

                                        www.sigmod.org/2012/.

McGregor:2012:SEE

[2208] Andrew McGregor, A. Pavan, Srikanta Tirthapura, and David Woodruff. Space-efficient estimation of statistics over sub-sampled streams. In Krötzsch et al. [2299], pages 273–282. ISBN ???? LCCN ???? URL http://www.

                                        sigmod.org/2012/.

Tirthapura:2012:REA

[2209] Srikanta Tirthapura and David Woodruff. Rectangle-efficient aggregation in spatial data streams. In Krötzsch et al. [2299], pages 283–294. ISBN ????

                                        LCCN ????  URL http://www.
                                        sigmod.org/2012/.

Huang:2012:RAT

[2210] Zengfeng Huang, Ke Yi, and Qin Zhang. Randomized algorithms for tracking distributed count, frequencies, and ranks. In Krötzsch et al. [2299], pages 295–306. ISBN ???? LCCN ????

URL http://www.sigmod.org/2012/.

Liu:2012:CDC

[2211] Zhenming Liu, Bozidar Radunović, and Milan Vojnović. Continuous distributed counting for non-monotonic streams. In Krötzsch et al. [2299], pages 307–318. ISBN ???? LCCN ???? URL

http://www.sigmod.org/2012/.

Hull:2013:FDA

[2212] Richard Hull. Foundations of dataaware process analysis: a database theory perspective. In Hull and Fan [2300], pages 1–12. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL

http://dl.acm.org/citation.cfm?
id=2463664;   http://www.sigmod.
org/2013/.

Kimelfeld:2013:CMM

[2213] Benny Kimelfeld and Phokion G. Kolaitis. The complexity of mining maximal frequent subgraphs. In Hull and Fan [2300], pages 13–24. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN

????  URL http://dl.acm.org/
citation.cfm?id=2463664;   http:
//www.sigmod.org/2013/.

Gottlob:2013:DMD

[2214] Georg Gottlob. Deciding monotone duality and identifying frequent itemsets in quadratic logspace. In Hull and Fan [2300], pages 25–36. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN

????  URL http://dl.acm.org/
citation.cfm?id=2463664;   http:
//www.sigmod.org/2013/.

Fagin:2013:SFF

[2215] Ronald Fagin, Benny Kimelfeld, Frederick Reiss, and Stijn Vansummeren. Spanners: a formal framework for information extraction. In Hull and Fan [2300], pages 37–48. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN

????  URL http://dl.acm.org/
citation.cfm?id=2463664;   http:
//www.sigmod.org/2013/.

Abouzied:2013:LVQ

[2216] Azza Abouzied, Dana Angluin, Christos Papadimitriou, Joseph M. Hellerstein, and Avi Silberschatz. Learning and verifying quantified boolean queries by example. In Hull and Fan [2300], pages 49–60. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL

http://dl.acm.org/citation.cfm?
id=2463664;   http://www.sigmod.
org/2013/.

Benedikt:2013:APA

[2217] Michael Benedikt, Tova Milo, and Dirk Van Gucht. The ACM PODS Alberto O. Mendelzon test-of-time award 2013. In Hull and Fan [2300], pages 61–62. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL http://dl.

acm.org/citation.cfm?id=2463664;
 http://www.sigmod.org/2013/.

Bojanczyk:2013:VDD

[2218] Mikolaj Boja’nczyk, Luc Segoufin, and Szymon Toruńczyk. Verification of database-driven systems via amalgamation. In Hull and Fan [2300], pages 63–74. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL http://dl.

acm.org/citation.cfm?id=2463664;
 http://www.sigmod.org/2013/.

Gheerbrant:2013:WNE

[2219] Amélie Gheerbrant, Leonid Libkin, and Cristina Sirangelo. When is naive evaluation possible? In Hull and Fan [2300], pages 75–86. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL

http://dl.acm.org/citation.cfm?
id=2463664;   http://www.sigmod.
org/2013/.

Indyk:2013:SHH

[2220] Piotr Indyk. Sketching via hashing:

from heavy hitters to compressed sensing to Sparse Fourier Transform. In Hull and Fan [2300], pages 87–90. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN

????  URL http://dl.acm.org/
citation.cfm?id=2463664;   http:
//www.sigmod.org/2013/.

Abiteboul:2013:CDD

[2221] Serge Abiteboul and Victor Vianu.

Collaborative data-driven workflows: think global, act local. In Hull and Fan [2300], pages 91–102. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN

????  URL http://dl.acm.org/
citation.cfm?id=2463664;   http:
//www.sigmod.org/2013/.

Kejlberg-Rasmussen:2013:EPR

[2222] Casper Kejlberg-Rasmussen, Yufei Tao, Konstantinos Tsakalidis, Kostas Tsichlas, and Jeonghun Yoon. I/Oefficient planar range skyline and attrition priority queues. In Hull and Fan [2300], pages 103–114. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN

                                        ????  URL http://dl.acm.org/
                                        citation.cfm?id=2463664;   http:
                                        //www.sigmod.org/2013/.

Agarwal:2013:NNS

[2223] Pankaj K. Agarwal, Boris Aronov, Sariel Har-Peled, Jeff M. Phillips, Ke Yi, and Wuzhou Zhang. Nearest neighbor searching under uncertainty II. In Hull and Fan [2300], pages 115– 126. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL http://dl.

                                        acm.org/citation.cfm?id=2463664;
                                         http://www.sigmod.org/2013/.

Gogacz:2013:BFC

[2224] Tomasz Gogacz and Jerzy Marcinkowski. On the BDD/FC conjecture. In Hull and Fan [2300], pages 127–138. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN

                                        ????  URL http://dl.acm.org/
                                        citation.cfm?id=2463664;   http:
                                        //www.sigmod.org/2013/.

Ameloot:2013:EPU

[2225] Tom J. Ameloot, Jan Van den Bussche, and Emmanuel Waller. On the expressive power of update primitives. In Hull and Fan [2300], pages 139– 150. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL http://dl.

                                        acm.org/citation.cfm?id=2463664;
                                         http://www.sigmod.org/2013/.

Rudolph:2013:FCD

[2226] Sebastian Rudolph and Markus Krötzsch. Flag & check: data access with monadically defined queries. In Hull and Fan [2300], pages 151–162. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN

                                        ????  URL http://dl.acm.org/
                                        citation.cfm?id=2463664;   http:
                                        //www.sigmod.org/2013/.

Hariri:2013:VRD

[2227] Babak Bagheri Hariri, Diego Calvanese, Giuseppe De Giacomo, Alin Deutsch, and Marco Montali. Verification of relational data-centric dynamic systems with external services. In Hull and Fan [2300], pages 163– 174. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL http://dl.

acm.org/citation.cfm?id=2463664;
 http://www.sigmod.org/2013/.

Baeza:2013:QGD

[2228] Pablo Barceló Baeza. Querying graph databases. In Hull and Fan [2300], pages 175–188. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL

http://dl.acm.org/citation.cfm?
id=2463664;   http://www.sigmod.
org/2013/.

Wijsen:2013:CTF

[2229] Jef Wijsen. Charting the tractability frontier of certain conjunctive query answering. In Hull and Fan [2300], pages 189–200. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL

http://dl.acm.org/citation.cfm?
id=2463664;   http://www.sigmod.
org/2013/.

Libkin:2013:TRA

[2230] Leonid Libkin, Juan Reutter, and Domagoj Vrgoc. Trial for RDF: adapting graph query languages for RDF data. In Hull and Fan [2300], pages 201–212. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL http://dl.

acm.org/citation.cfm?id=2463664;
 http://www.sigmod.org/2013/.

Bienvenu:2013:OBD

[2231] Meghyn Bienvenu, Balder ten Cate, Carsten Lutz, and Frank Wolter. Ontology-based data access: a study through disjunctive datalog, CSP, and MMSNP. In Hull and Fan [2300], pages 213–224. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL http://dl.

                                        acm.org/citation.cfm?id=2463664;
                                         http://www.sigmod.org/2013/.

Hernich:2013:WFS

[2232] André Hernich, Clemens Kupke, Thomas Lukasiewicz, and Georg Gottlob. Well-founded semantics for extended datalog and ontological reasoning. In Hull and Fan [2300], pages 225–236. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL http://dl.

                                        acm.org/citation.cfm?id=2463664;
                                         http://www.sigmod.org/2013/.

Baeza:2013:SAG

[2233] Pablo Barceló Baeza, Miguel Romero, and Moshe Y. Vardi. Semantic acyclicity on graph databases. In Hull and Fan [2300], pages 237–248. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN

                                        ????  URL http://dl.acm.org/
                                        citation.cfm?id=2463664;   http:
                                        //www.sigmod.org/2013/.

Figueira:2013:XTA

[2234] Diego Figueira. On XPath with transitive axes and data tests. In Hull and Fan [2300], pages 249–260. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN

                                        ????  URL http://dl.acm.org/
                                        citation.cfm?id=2463664;   http:
                                        //www.sigmod.org/2013/.

Bagan:2013:TRS

[2235] Guillaume Bagan, Angela Bonifati, and Benoit Groz. A trichotomy for regular simple path queries on graphs. In Hull and Fan [2300], pages 261– 272. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL http://dl.

acm.org/citation.cfm?id=2463664;
http://www.sigmod.org/2013/.

Beame:2013:CSP

[2236] Paul Beame, Paraschos Koutris, and Dan Suciu. Communication steps for parallel query processing. In Hull and Fan [2300], pages 273–284. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN

????  URL http://dl.acm.org/
citation.cfm?id=2463664;   http:
//www.sigmod.org/2013/.

Wong:2013:DIE

[2237] Limsoon Wong. A dichotomy in the intensional expressive power of nested relational calculi augmented with aggregate functions and a powerset operator. In Hull and Fan [2300], pages 285–296. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL http://dl.

acm.org/citation.cfm?id=2463664;
http://www.sigmod.org/2013/.

Kazana:2013:EFO

[2238] Wojciech Kazana and Luc Segoufin.

Enumeration of first-order queries on classes of structures with bounded expansion. In Hull and Fan [2300], pages 297–308. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL http://dl.

acm.org/citation.cfm?id=2463664;
http://www.sigmod.org/2013/.

Chen:2013:FCC

[2239] Hubie Chen and Moritz Müller. The fine classification of conjunctive queries and parameterized logarithmic space complexity. In Hull and Fan [2300], pages 309–320. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN ???? URL

http://dl.acm.org/citation.cfm?
id=2463664;   http://www.sigmod.
org/2013/.

King:1975:ICM

[2240] W. F. King, editor. International Conference on Management of Data, San Jose, California, May 14, 15 and 16th, 1975. ACM Press, New York, NY 10036, USA, 1975. ISBN ???? LCCN ????

Rothnie:1976:ASI

[2241] James B. Rothnie, editor. ACM-SIGMOD International Conference on Management of Data: [proceedings], June 2, 3 and 4, 1976, Washington, DC. ACM Press, New York, NY 10036, USA, 1976. LCCN QA76.9.D3 I59 1976.

Smith:1977:ASI

[2242] Diane C. P. Smith, editor. ACM-SIGMOD International Conference on Management of Data, Toronto, Canada, August 3, 4, and 5, 1977: [proceedings]. ACM Press, New York, NY 10036, USA, 1977. LCCN QA76.9.D3 I59 1977.

Dale:1978:ASI

[2243] Nell Dale and Eugene Lowenthal, editors. ACM-SIGMOD International Conference on Management of Data, Austin, Texas, May 31, June 1–2, 1978. ACM Press, New York, NY 10036, USA, 1978. ISBN ???? LCCN ????

Bernstein:1979:ASI

[2244] Philip A. Bernstein, editor. ACM-Sigmod 1979 International Conference on Management of Data: proceedings, May 30–June 1, the 57 Park Plaza Hotel, Boston, Massachusetts. ACM Press, New York, NY 10036, USA, 1979. ISBN 0-89791-001-X. LCCN QA76.9.D3 I59 1979.

Chen:1980:ASI

[2245] Peter P. Chen and R. Clay Sprowls, editors. ACM-SIGMOD 1980 International Conference on Management of Data: proceedings, May 14–May 16, Miramar-Sheraton Hotel, Santa Monica, California. ACM Press, New York, NY 10036, USA, 1980. ISBN 0-89791-018-4. LCCN QA76.9.D3 I59 1980.

Lien:1982:ASI

[2246] Y. Edmund Lien, editor. ACM-SIGMOD 1981 International Conference on Management of Data: [proceedings] April 29–May 1, the University of Michigan, Ann Arbor, Michigan. ACM Press, New York, NY 10036, USA, 1982. ISBN 0-89791-040-0. LCCN QA76.9.D3 I59 1981. ACM order number 472810.

Schkolnick:1982:PIC

[2247] Mario Schkolnick, editor. Proceedings / International Conference on Management of Data, June 2–4, Orlando, Florida, 1982. ACM Press, New York, NY 10036, USA, 1982. ISBN 0-89791-073-7. LCCN QA 76.9 D3 I59 1982.

ACM:1982:PPA

[2248] ACM, editor. PODS ’82. Proceedings of the ACM Symposium on Principles of Database Systems: 29–31 March 1982, Marina del Ray Hotel, Los Angeles, California. ACM Press, New York, NY 10036, USA, 1982. ISBN 0-89791-070-2. LCCN QA76.9.D3 A33 1982.

DeWitt:1983:SPA

[2249] David J. DeWitt and Georges Gardarin, editors. SIGMOD 83: proceedings of annual meeting, database week, San Jose, May 23-26, 1983, volume 13(4) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 1983. ISBN 0-89791-104-0. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.13 no.4.

ACM:1983:PPS

[2250] ACM, editor. PODS ’83. Proceedings of the Second ACM SIGACT-SIGMOD Symposium on Principles of Database Systems: 21–23 March 1983, Colony Square Hotel, Atlanta, Georgia. ACM Press, New York, NY 10036, USA, 1983. ISBN 0-89791-097-4. LCCN QA76.9.D3 A15 1983. ACM order number 475830.

Yormark:1984:ASI

[2251] Beatrice Yormark, editor. ACM-Sigmod International Conference on Management of Data: Proceedings of annual meeting: Boston, MA, June 18–21, 1984, volume 14(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 1984. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I59 1984. ACM order number 472840.

ACM:1984:PPT

[2252] ACM, editor. PODS ’84. Proceedings of the Third ACM SIGACT-SIGMOD Symposium on Principles of Database Systems: April 2–4, 1984, Waterloo, Ontario, Canada. ACM Press, New York, NY 10036, USA, 1984. ISBN 0-89791-128-8. LCCN QA76.9.D3 A15 1984. ACM order number 475840.

ACM:1985:PPF

[2253] ACM, editor. PODS ’85. Proceedings of the Fourth ACM SIGACT-SIGMOD Symposium on Principles of Database Systems, March 25–27, 1985, Portland, Oregon. ACM Press, New York, NY 10036, USA, 1985. ISBN 0-89791-153-9. LCCN QA 76.9 D3 A296 1985.

Navathe:1985:PAS

[2254] Sham Navathe, editor. Proceedings of ACM-SIGMOD 1985 International Conference on Management of Data, May 28–31, 1985, LaMansion Hotel, Austin, Texas. ACM Press, New York, NY 10036, USA, 1985. ISBN 0-89791-160-1. LCCN QA 76.9 D3 I59 1985; QA1 .A87.

ACM:1986:PPF

[2255] ACM, editor. PODS ’86. Proceedings of the Fifth ACM SIGACT-SIGMOD Symposium on Principles of Database Systems, March 24–26, 1986, Cambridge, MA. ACM Press, New York, NY 10036, USA, 1986. ISBN 0-89791-179-2. LCCN QA76.9.D3 A296 1986. ACM order number 475860.

Zaniolo:1986:PAS

[2256] Carlo Zaniolo, editor. Proceedings of ACM SIGMOD ’86 International Conference on Management of Data, Washington, D.C. May 28–30, 1986. ACM Press, New York, NY 10036, USA, 1986. ISBN 0-89791-191-1. LCCN QA 76.9 D3 A18 1986.

ACM:1987:PPS

[2257] ACM, editor. PODS ’87. Proceedings of the Sixth ACM SIGACT-SIGMOD-SIGART Symposium on Principles of Database Systems, March 23–25, 1987, San Diego, California. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-89791-223-3. LCCN QA76.9.D3A296 1987. ACM Order Number 475870.

Dayal:1987:PAC

[2258] Umeshwar Dayal and Irv Traiger, editors. Proceedings of Association for Computing Machinery Special Interest Group on Management of Data 1987 annual conference, San Francisco, May 27–29, 1987. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-89791-236-5. LCCN QA 76.9 D3 P76 1987. ACM order number 472870.

ACM:1988:PAC

[2259] ACM, editor. Proceedings of Association for Computing Machinery Special Interest Group on Management of Data: 1988 Annual Conference, Chicago, Illinois, June 1–3. ACM Press, New York, NY 10036, USA, 1988. ISBN ???? LCCN ???? ACM order number 472880.

ACM:1988:PPS

[2260] ACM, editor. PODS ’88. Proceedings of the Seventh ACM SIGACT-SIGMOD-SIGART Symposium on Principles of Database Systems: March 21–23, 1988, Austin, Texas. ACM Press, New York, NY 10036, USA, 1988. ISBN 0-89791-263-2. LCCN QA76.9.D3 A15 1988.

ACM:1989:PPE

[2261] ACM, editor. PODS ’89. Proceedings of the Eighth ACM SIGACT-SIGMOD-SIGART Symposium on Principles of Database Systems, March 29–31, 1989, Philadelphia, PA. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-308-6. LCCN QA 76.9 D3 A26 1989.

Clifford:1989:PAS

[2262] James Clifford, Bruce Lindsay, and David Maier, editors. Proceedings of the 1989 ACM SIGMOD international conference on the management of data, Portland, Oregon, May 31– June 2, 1989. ACM Press, New York, NY 10036, USA, 1989. ISBN ???? LCCN ????

Garcia-Molina:1990:PAS

[2263] Hector Garcia-Molina and H. V. Jagadish, editors. Proceedings of the 1990 ACM SIGMOD International Conference on Management of Data, May 23–25, 1990, Atlantic City, NJ, volume 19(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, June 1990. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1 .A87.

ACM:1990:PPN

[2264] ACM, editor. PODS ’90. Proceedings of the Ninth ACM SIGACT-SIGMOD-SIGART Symposium on Principles of Database Systems: April 2–4, 1990, Nashville, Tennessee, volume 51(1) of Journal of Computer and Systems Sciences. ACM Press, New York, NY 10036, USA, 1990. ISBN 0-89791-352-3. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA 76.9 D3 A26 1990. A few papers from this conference were republished in 1995 in the Journal of Computer and Systems Sciences.

ACM:1991:PPT

[2265] ACM, editor. PODS ’91. Proceedings of the Tenth ACM SIGACT-SIGMOD-SOGART Symposium on Principles of Database Systems: May 29–31, 1991, Denver, Colorado, volume 51(2) of Journal of Computer and Systems Sciences. ACM Press, New York, NY 10036, USA, 1991. ISBN 0-89791-430-9. ISSN 0022-0000 (print), 1090-2724 (electronic). LCCN QA76.9.D3 A296 1991. ACM order number 475910.

Clifford:1998:ASP

[2266] James Clifford and Roger King, editors. ACM SIGMOD 91: proceedings of the 1991 ACM SIGMOD International Conference on Management of Data, Denver, Colorado, May 29–31, 1991, volume 20(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 1998. ISBN 0-89791-425-2. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.20 no.2 1991; QA1 .A87.

Kim:1992:DSN

[2267] Won Kim, Y. Kambayashi, and In Sup Paik, editors. Database systems for next-generation applications: principles and practice, volume 1 of Advanced Database Research and Development Series. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1992. ISBN 981-02-1315-8. LCCN QA76.9.D3 D3589 1992.

Stonebraker:1992:PAS

[2268] Michael Stonebraker, editor. Proceedings of the 1992 ACM SIGMOD International Conference on Management of Data, San Diego, California, June 2–5, 1992, volume 21(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-521-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ????

ACM:1992:PPE

[2269] ACM, editor. PODS ’92. Proceedings of the Eleventh ACM SIGACT-SIGMOD-SIGART Symposium on Principles of Database Systems, June 2–4, 1992, San Diego, CA. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-519-4 (paperback), 0-89791-520-8 (casebound). LCCN QA 76.9 D3 A26 1992. ACM order number 475920.

ACM:1993:PPT

[2270] ACM, editor. PODS ’93. Proceedings of the Twelfth ACM SIGACT-SIGMOD-SIGART Symposium on Principles of Database Systems: May 25–28, 1993, Washington, DC, volume 12 of Proceedings of the ACM SIGACT SIGMOD SIGART Symposium on Principles of Database Systems. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-593-3. LCCN QA 76.9 D3 A26 1993.

Buneman:1993:PAS

[2271] Peter Buneman and Sushil Jajodia, editors. Proceedings of the 1993 ACM SIGMOD International Conference on Management of Data, SIG-MOD ’93, Washington, DC, May 26– 28, 1993, volume 22(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.22 no.2 1993.

ACM:1994:PPT

[2272] ACM, editor. PODS ’94. Proceedings of the Thirteenth ACM SIGACT-SIGMOD-SIGART Symposium on Principles of Database Systems, May 24–26, 1994, Minneapolis, MN, volume 13. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-642-5. LCCN QA 76.9 D3 A26 1994.

Snodgrass:1994:PAS

[2273] Richard T. Snodgrass and Marianne Winslett, editors. Proceedings of the 1994 ACM SIGMOD International Conference on Management of Data / SIGMOD ’94, Minneapolis, Minnesota, May 24–27, 1994, volume 23(2) of SIG-MOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-639-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.23 no.2 1994.

ACM:1995:PPF

[2274] ACM, editor. PODS ’95. Proceedings of the Fourteenth ACM SIGACT-SIGMOD-SIGART Symposium on Principles of Database Systems, PODS 1995, San Jose, California, May 22– 25, 1995, volume 14. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-730-8. LCCN QA 76.9 D3 A26 1995.

ACM:1995:PAS

[2275] ACM, editor. Proceedings of the 1995 ACM SIGMOD International Conference on Management of Data: May 23–25, 1995, San Jose, California, volume 24(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-731-6. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1 .A87. ACM order number 472950.

ACM:1996:PPF

[2276] ACM, editor. PODS ’96. Proceedings of the Fifteenth ACM SIGACT-SIGMOD-SIGART Symposium on Principles of Database Systems, PODS 1996, Montréal, Canada, June 3–5, 1996, volume 15. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-781-2. LCCN QA 76.9 D3 A26 1996. ACM order number 475960.

Jagadish:1996:PAS

[2277] H. V. Jagadish and Inderpal Singh Mumick, editors. Proceedings of the 1996 ACM SIGMOD International Conference on Management of Data, Montreal, Quebec, Canada, June 4–6, 1996. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-794-4.

LCCN QA 76.9 D3 S53 v.25 no.2 1996; QA1 .A87.

Peckman:1997:PAS

[2278] Joan M. Peckman, editor. Proceedings, ACM SIGMOD International Conference on Management of Data: SIG-MOD 1997: May 13–15, 1997, Tucson, Arizona, USA, volume 26(2) of SIG-MOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-911-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1 .A87.

ACM:1997:PPS

[2279] ACM, editor. PODS ’97. Proceedings of the Sixteenth ACM SIG-SIGMOD-SIGART Symposium on Principles of Database Systems, May 12–14, 1997, Tucson, Arizona. ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-910-6. LCCN QA 76.9 D3 A26 1997.

Haas:1998:PAS

[2280] Laura Haas and Ashutosh Tiwary, editors. Proceedings of the 1998 ACM SIGMOD International Conference on Management of Data: June 1–4, 1998, Seattle, Washington, USA, volume 27(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 1998. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1 .A87.

ACM:1998:PPA

[2281] ACM, editor. PODS ’98. Proceedings of the ACM SIGACT–SIGMOD– SIGART Symposium on Principles of Database Systems, June 1–3, 1998, Seattle, Washington. ACM Press, New York, NY 10036, USA, 1998. ISBN 0-89791-996-3. LCCN QA76.9.D3 A296 1998. ACM order number 475980.

Delis:1999:PAS

[2282] Alex Delis, Christos Faloutsos, and Shahram Ghandeharizadeh, editors. Proceedings of the 1999 ACM SIGMOD International Conference on Management of Data: SIGMOD ’99, Philadelphia, PA, USA, June 1–3, 1999, volume 28(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 1999. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1 .A87.

ACM:1999:PEA

[2283] ACM, editor. Proceedings of the Eighteenth ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems: PODS 1999: Philadelphia, Pennsylvania, May 31– June 2, 1999. ACM Press, New York, NY 10036, USA, 1999. ISBN 1-58113-062-7. LCCN QA76.9.D3 A296 1999. ACM order number 475990.

Chen:2000:PAS

[2284] Weidong Chen, Jeffery Naughton, and Philip A. Bernstein, editors. Proceedings of the 2000 ACM SIGMOD International Conference on Management of Data: May 16–18, 2000, Dallas, Texas, volume 29(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 2000. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1 .A87.

ACM:2001:PTA

[2285] ACM, editor. Proceedings of the Twentieth ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems: PODS 2001: Santa Barbara, California, May 21–23, 2001, SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 2001. ISBN 1-58113-361-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A296 2001. ACM order number 475010.

Sellis:2001:PAS

[2286] Timos Sellis and Sharad Mehrotra, editors. Proceedings of the 2001 ACM SIGMOD International Conference on Management of Data 2001, Santa Barbara, California, United States, May 21–24, 2001. ACM Press, New York, NY 10036, USA, 2001. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Franklin:2002:PAS

[2287] Michael Franklin, Bongki Moon, and Anastassia Ailamaki, editors. Proceedings of the ACM SIGMOD International Conference on Management of Data, June 3–6, 2002, Madison, WI, USA. ACM Press, New York, NY 10036, USA, 2002. ISBN ???? LCCN ???? ACM order number 475020.

ACM:2002:PTF

[2288] ACM, editor. Proceedings of the Twenty-First ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems: PODS 2002: Madison, Wisconsin, June 3–5, 2002. ACM Press, New York, NY 10036, USA, 2002. ISBN 1-58113-507-6. LCCN QA76.9.D3 A296 2002. ACM order number 475021.

ACM:2003:PTS

[2289] ACM, editor. Proceedings of the Twenty-Second ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems: PODS 2003: San Diego, Calif., June 9–11, 2003. ACM Press, New York, NY 10036, USA, 2003. ISBN 1-58113-670-6. LCCN QA76.9.D3 A296 2003. ACM order number 475030.

ACM:2003:PAS

[2290] ACM, editor. Proceedings of the 2003 ACM SIGMOD International Conference on Management of Data 2003, San Diego, California, June 09–12, 2003. ACM Press, New York, NY 10036, USA, 2003. ISBN 1-58113-634-X. LCCN ????

ACM:2004:PAS

[2291] ACM, editor. Proceedings of the 2004 ACM SIGMOD International Conference on Management of Data 2004, Paris, France, June 13–18, 2004. ACM Press, New York, NY 10036, USA, 2004. ISBN 1-58113-859-8. LCCN ????

ACM:2005:PTF

[2292] ACM, editor. Proceedings of the Twenty-Fourth ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems: PODS 2005: Baltimore, Maryland, June 13-15, 2005. ACM Press, New York, NY 10036, USA, 2005. ISBN 1-59593-062-0. LCCN QA76.9.D3 A296 2005. ACM order number 475050.

ACM:2006:PTF

[2293] ACM, editor. Proceedings of the Twenty-Fifth ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems, Chicago, IL, USA June 26–28, 2006. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-318-2. LCCN QA76.9.D3 A296 2006.

ACM:2007:PTS

[2294] ACM, editor. Proceedings of the Twenty-Sixth ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems: PODS 2007, Beijing, China, June 11–13, 2007. ACM Press, New York, NY 10036, USA, 2007. ISBN 1-59593-685-8. LCCN ????

Lenzerini:2008:PTS

[2295] Maurizio Lenzerini and Domenico Lembo, editors. Proceedings of the Twenty-Seventh ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems: PODS’08, Vancouver, BC, Canada, June 9–11, 2008. ACM Press, New York, NY 10036, USA, 2008. ISBN 1-60560-932-3. LCCN ????

Paredaens:2009:PTE

[2296] Jan Paredaens and Jianwen Su, editors. Proceedings of the twenty-eighth ACM SIGMOD-SIGACT-SIGART symposium on Principles of database systems, PODS’09, Providence, Rhode Island, June 29–July 1, 2009. ACM Press, New York, NY 10036, USA, 2009. ISBN 1-60558-553-X. LCCN ????

VanGucht:2010:PPT

[2297] Dirk Van Gucht, editor. PODS’10: proceedings of the Twenty-Ninth ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems: Indianapolis, Indiana, USA, June 6–11, 2010. ACM Press, New York, NY 10036, USA, 2010. ISBN 1-4503-0033-2. LCCN ????

Lenzerini:2011:PPT

[2298] Maurizio Lenzerini, editor. PODS’11:

Proceedings of the thirtieth ACM SIGMOD-SIGACT-SIGART symposium on Principles of database systems: June 13–15, 2011, Athens, Greece. ACM Press, New York, NY 10036, USA, 2011. ISBN 1-4503-0660-

8. LCCN ???? URL http://dl.acm.
org/citation.cfm?id=1989284.

Krotzsch:2012:PPA

[2299] Markus Krötzsch, Maurizio Lenzerini, and Michael Benedikt, editors. PODS’12: Proceedings of the 31st ACM SIGMOD-SIGACT-SIGART symposium on Principles of database systems: May 20–24, 2012, Scottsdale, AZ, USA. ACM Press, New York, NY 10036, USA, 2012. ISBN ???? LCCN ???? URL http://www.

sigmod.org/2012/.

Hull:2013:SPC

[2300] Richard Hull and Wenfei Fan, editors. SIGMOD/PODS’13: compilation proceedings of the 2013 ACM Symposium on Principles of Database Systems, ACM SIGMOD international conference on management of data, and SIGMOD/PODS 2013 PhD symposium: June 22–27, 2013, New York, New York, USA. ACM Press, New York, NY 10036, USA, 2013. ISBN 1-4503-2066-X, 1-4503-2037-6. LCCN

????  URL http://dl.acm.org/

                                        citation.cfm?id=2463664;   http:
                                        //www.sigmod.org/2013/.